### v1

# Liftings of ideals in positive characteristic to those in characteristic zero: Higher dimension

Abstract.
We study a pair consisting of a smooth variety of arbitrary dimension over a field of positive characteristic and a multi-ideal with a real exponent. We prove that the set of log discrepancies for a fixed exponent is discrete. Additionally, we show that the set of log canonical thresholds (lcts) of multi-ideals on a smooth variety in positive characteristic is contained within the set of lcts of multi-ideals on a smooth variety over $\mathbb{C}$. As a result, we find that the accumulation points of log canonical thresholds are rational if all the exponents are rational. We also obtain ACC for the set of lcts of multi-ideals on a smooth varieties in positive characteristic. These findings generalize the author’s results presented in arXiv in 2024 to the higher-dimensional case.
Mathematical Subject Classification 2020: 14B05,14E18, 14J17
Key words: singularities in positive characteristic, jet schemes, minimal log discrepancy, log canonical threshold
The author is partially supported by Grant-In-Aid (c) 22K03428 of JSPS and by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.
1. Introduction
There are many tools available for studying singularities in characteristic zero. These include resolutions of singularities, Bertini’s theorem (which addresses generic smoothness), and vanishing theorems of Kodaira type. However, these tools do not necessarily hold in positive characteristic, which creates difficulties when studying singularities in that setting. In order to overcome these difficulties, we construct a “bridge” between positive characteristic and characteristic zero, allowing us to transport results from characteristic zero to positive characteristic. In general, such a strategy does not work. For example, the resolution of singularities in positive characteristic does not follow from that in characteristic zero. However, for our specific problems, this approach is effective.
Our focus is on the invariants of singularities, specifically log discrepancies, minimal log discrepancies, and log canonical thresholds. Our bridge illustrates the relationships between the set of invariants in positive characteristic and the set of those in characteristic zero. As a result, the properties of these sets in characteristic zero are inherited by their counterparts in positive characteristic.
The main theorem of this paper serves as a “bridge” between characteristic 00 and positive characteristic. In our previous paper [8], we addressed similar issues and obtained results for dimensions less than or equal to three. The present paper employs the same method and overcomes the difficulties that hindered progress in higher-dimensional cases. Concretely, in the paper [8], the following theorem was proved under the assumption that there exists a log resolution of the singularities by smooth blow-ups. Our main challenge in this paper is to prove the theorem without the assumption about log resolutions.
Theorem 1.1.
Let $N$ be an integer $\geq 2$, and $k$ an algebraically closed field of characteristic $p>0$. Let $A={\mathbb{A}}_{k}^{N}$ be the $N$-dimensional affine space over $k$ and 00 the origin of $A$. Let $E$ be a prime divisor over $(A,0)$, and ${\mathfrak{a}},{\mathfrak{a}}_{1},\ldots,{\mathfrak{a}}_{r}\subset k[x_{1},\ldots,x_{N}]$ non-zero ideals. Then, there exists a prime divisor $F_{\mathbb{C}}$ over the affine space $A_{\mathbb{C}}:={\mathbb{A}}_{\mathbb{C}}^{N}$ over ${\mathbb{C}}$ with the center at the origin ${\tilde{0}}$ satisfying the following:
  1. (i)
    $2(N-1)+k_{E}=k_{F_{\mathbb{C}}}$;
  2. (ii)
    there exists a lifting ${\widetilde{\mathfrak{a}}}\subset{\mathbb{C}}[x_{1},\ldots,x_{N}]$ of ${\mathfrak{a}}$ satisfying
    $v_{E}({\mathfrak{a}})=v_{F_{\mathbb{C}}}({\widetilde{\mathfrak{a}}});$
  3. (iii)
    there exist liftings ${\widetilde{\mathfrak{a}}}_{i}\subset{\mathbb{C}}[x_{1},\ldots,x_{N}]$ of ${\mathfrak{a}}_{i}$ $(i=1,\ldots,r)$ such that
    $v_{E}({\mathfrak{a}}_{i})=v_{F_{\mathbb{C}}}({\widetilde{\mathfrak{a}}}_{i})$
    and for every multi-ideal ${\mathfrak{a}}^{e}:={\mathfrak{a}}_{1}^{e_{1}}\cdots{\mathfrak{a}}_{r}^{e_{r}}$ $(e_{i}\in{\mathbb{R}}_{>0})$ on $A$, the lifting ${\widetilde{\mathfrak{a}}}^{e}={\widetilde{\mathfrak{a}}}_{1}^{e_{1}}\cdots{\widetilde{\mathfrak{a}}}_{r}^{e_{r}}$ of ${\mathfrak{a}}^{e}$ on $A_{\mathbb{C}}$ satisfies
    $2(N-1)+a(E;A,{\mathfrak{a}}^{e})=a(F_{\mathbb{C}};A_{\mathbb{C}},{\widetilde{\mathfrak{a}}}^{e}).$
We obtain the following corollaries as applications of the theorem. Here, note that we do not require the existence of log resolutions of the singularities, while we required it in [8].
Corollary 1.2.
Let $N$ be a positive integer and $0\in A$ a smooth closed point of an $N$-dimensional algebraic variety $A$ defined over an algebraically closed field $k$ of characteristic $p>0$. Let ${\mathfrak{a}}^{e}$ be a non-zero multi-ideal on $A$. Then, the following hold:
  1. (i)
    (discreteness of log discrepancies) For a fixed exponent $e\in{\mathbb{R}}_{>0}^{r}$ and a real number $a>0$, the set of log discrepancies
    $\Lambda_{N,e,a}:=\left\{a(E;A,{\mathfrak{a}}^{e})\leq a\ \middle|\begin{array}[]{l}\mbox{every}\ {\mathfrak{a}}\ \mbox{such\ that}\ (A,{\mathfrak{a}}^{e})\ \mbox{is\ log\ canonical\ at $0$},\\ E:\ \mbox{\ any \ prime divisor\ over }(A,0)\end{array}\right\}$
    is finite. In particular, for fixed $N$ and $e$, the set
[…]
    is finite. Also for a fixed ${\mathfrak{a}}^{e}$, the discreteness of log discrepancies implies the existence of a prime divisor computing ${\mathrm{mld}}(0,A,{\mathfrak{a}}^{e})$.
  2. (ii)
    (Inclusion of the sets of ${\operatorname{lct}}$’s ) Let $e\in{\mathbb{R}}_{>0}^{r}$ and
[…]
    $T_{N,e,{\mathbb{C}}}:=\{{\operatorname{lct}}({\tilde{0}};A_{\mathbb{C}},{\mathfrak{b}}^{e})\mid{\mathfrak{b}}^{e}:\mbox{non-zero\ multi-ideal\ on}\ A_{\mathbb{C}}\},$
    where ${\tilde{0}}\in A_{\mathbb{C}}$ is a smooth closed point of an $N$-dimensional variety $A_{\mathbb{C}}$ defined over ${\mathbb{C}}$. Then
    $T_{N,e,k}\subset T_{N,e,{\mathbb{C}}}.$
    In particular, when the exponents $e_{i}$’s in $e=(e_{1},\ldots,e_{r})$ are all rational, then ${\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})$ is a rational number, and an accumulation point of these log canonical thresholds is also a rational number.
  3. (iii)
    (ACC of ${\operatorname{lct}}$) Let $D\subset{\mathbb{R}}_{>0}$ be a DCC set. Then, ACC holds for the set
    $T_{N,k}(D):=\{{\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})\mid e=(e_{1},\ldots,e_{r}),\forall e_{i}\in D\}.$
For $N=1$, the statements of Corollary 1.2 are trivial, because a prime divisor over $(A,0)$ is just the divisor $\{0\}\subset A$, and $v_{\{0\}}({\mathfrak{a}}^{e}):=\sum_{i}e_{i}{\mathrm{ord}}({\mathfrak{a}}_{i})$ and $k_{\{0\}}+1=1$ hold for any base field. So, we may assume $N\geq 2$ and can apply Theorem 1.1.
Note that the statements in Corollary 1.2 are known to hold in characteristic 00. More specifically, statement (i) is proven by Kawakita in [10], the statement about the rationality in (ii) is established by T. De Fernex and M. Mustaţǎ in [3], and statement (iii) for a special case is demonstrated by T. De Fernex, L. Ein and M. Mustaţǎ in [2] and for the general case by C. Hacon, J. McKernan and C. Xu in [5].
These are excellent results, and the proofs are elaborated based on resolutions of singularities, Bertini’s theorem, and cohomology vanishing of the Kodaira type, which are not available in the positive characteristic case. Therefore, the same proofs cannot be applied in positive characteristic settings.
Our proof in positive characteristic involves transporting these results from characteristic 00 into the realm of positive characteristic, utilizing the bridge established in Theorem 1.1.
This paper is organized as follows: In the second section, we introduce basic notions on liftings of materials in positive characteristic to those in characteristic 00. The proofs of this part can be seen in [8].
In the third section we introduce log discrepancies, minimal log discrepancies, and log canonical thresholds which are birational invariants to measure the nature of singularities of pairs. The description of these invariants by the arc spaces will work to compare the invariants between in the realm of positive characteristic and in that of characteristic 00. In order to provide the reader with the basic notion for our study, most statements in the second and third sections overlap with those in the corresponding sections of the previous paper ([8]).
In the forth section, from a given prime divisor over the affine space in positive characteristic, we construct a prime divisor over the affine space defined over ${\mathbb{C}}$ such that the divisors have some common properties.
In the fifth section, we give the proof for the main theorem and the corollaries.
As there are still many open problems related to our topics, we should have concluded the paper by some conjectures related to the topics. However, in our previous paper ([8]), the last section was for some conjectures which are still open after this generalized version. So, we do not repeat conjectures here in this paper. The reader interested may refer to the sixth section of [8].
Acknowledgement. The author would like to thank Lawrence Ein, Kazuhiko Kurano, Mircea Mustaţǎ, Yuri Prokhorov, Kohsuke Shibata and Shoji Yokura for useful discussions and comments. The author expresses her heartfelt thanks to János Kollár for providing an important example and comments for the preliminary version of this paper.
Conventions
  1. (1)
    When we say just “a point” of a variety $A$, it is not necessarily a closed point. By abuse of notation, for a point $P\in A$ we write $\dim P:=\dim\overline{\{P\}}$. The codimension of $P$ is defined accordingly ; ${\mathrm{codim}}P:=\dim A-\dim P$.
  2. (2)
    A blow-up with the center by the defining ideal of the closure $\overline{\{P\}}$ is called “the blow-up at $P$” and we also call $P$ the center of the blow-up.
  3. (3)
    Let $L_{1}$ be a ring, $L_{2}$ an $L_{1}$-algebra and $A$ a scheme of finite type over $L_{1}$. We denote the scheme $A\times_{{\mathrm{Spec\ }}L_{1}}{\mathrm{Spec\ }}L_{2}$ by $A\otimes_{L_{1}}L_{2}$ to avoid the bulky expression.
  4. (4)
    For an integral domain $R$, we denote the field of fractions of $R$ by $Q(R)$.
2. Basics on liftings
In this paper we construct a “lifting” which is an items on a variety over ${\mathbb{C}}$ corresponding to that in positive characteristic. Basic notions about liftings are discussed in [7] and [8], and here we summarize some of them, which we will use in this paper.
Definition 2.1.
Let $R$ be an integral domain which is finitely generated $k$-algebra, where $k$ is a field of characteristic $p\geq 0$. A subring $S\subset R$ is called a skeleton of $R$, if $S$ is finitely generated ${{\mathbb{Z}}/(p)}$-subalgebra of $R$ such that $k[S]=R$.
Definition 2.2.
  1. (1)
    (compatible skeletons) Let $k$ be a field of characteristic $p>0$ and $K$ a field of characteristic 00. Let $R$ and ${\widetilde{R}}$ be finitely generated $k$-algebra and $K$-algebra, respectively. Let $S\subset R$ and ${\widetilde{S}}\subset{\widetilde{R}}$ be skeletons. If the skeletons satisfies
    ${\widetilde{S}}$ is ${\mathbb{Z}}$-flat at $p$, and
    $S={\widetilde{S}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)},$
    we call them compatible skeletons. We sometimes say that a skeleton ${\widetilde{S}}$ is a lifting of the skeleton $S$. We also call the canonical surjection $\Phi_{p}:{\widetilde{S}}\to S$ compatible skeletons. In this case we write
    ${\widetilde{R}}({\operatorname{mod}\ }p)=R.$
    If schemes ${\widetilde{A}}$ and $A$ are defined by ${\widetilde{R}}$ and $R$ (eg. as ${\mathrm{Spec\ }}$ or ${\mathrm{Proj\ }}$) which have compatible skeletons, we also write
    ${\widetilde{A}}({\operatorname{mod}\ }p)=A.$
    Here, we note that if ${\widetilde{R}}$ is an integral domain (therefor ${\widetilde{S}}$ is so), then ${\mathbb{Z}}$-flatness is automatic. In this case $S={\widetilde{S}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}$ is the only condition for ${\widetilde{S}}$ and $S$ to be compatible skeletons. (In the paper [8], we treat only in this case)
  2. (2)
    (lifting of elements) Let $\Phi_{p}:{\widetilde{S}}\to S$ be compatible skeletons of ${\widetilde{R}}$ and $R$. If ${\tilde{f}}\in{\widetilde{R}}$ and $f\in R$ satisfy ${\tilde{f}}\in{\widetilde{S}}$ and $f=\Phi_{p}({\tilde{f}})$, then we write ${\tilde{f}}({\operatorname{mod}\ }p)=f$ and call ${\tilde{f}}$ a lifting of $f$. If $\{{\tilde{f}}_{1},{\tilde{f}}_{2},\ldots,{\tilde{f}}_{n}\}\subset{\widetilde{S}}$ and $f_{i}=\Phi_{p}({\tilde{f}}_{i})$ for all $i=1,\ldots,n$, then we write $\{{\tilde{f}}_{1},{\tilde{f}}_{2},\ldots,{\tilde{f}}_{n}\}({\operatorname{mod}\ }p)=\{f_{1},\ldots,f_{n}\}$ and call $\{{\tilde{f}}_{1},{\tilde{f}}_{2},\ldots,{\tilde{f}}_{n}\}$ a lifting of $\{f_{1},\ldots,f_{n}\}$.
When we think of liftings of elements, we always use a common pair of compatible skeletons in a continuous discussion.
Example 2.3.
([8, Example 2.3])
(1) For any finite elements $a_{1},\ldots,a_{r}\in k$ there exist liftings ${\widetilde{a}}_{1},\ldots,{\widetilde{a}}_{r}\in{\mathbb{C}}$.
(2) Let $\Sigma\subset k$ and ${\widetilde{\Sigma}}\subset{\mathbb{C}}$ be compatible skeletons. Then $S:=\Sigma[x_{1},\ldots,x_{N}]\subset k[x_{1},\ldots,x_{N}]$ and ${\widetilde{S}}:={\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]\subset{\mathbb{C}}[x_{1},\ldots,x_{N}]$ are also compatible skeletons for ${\mathbb{C}}[x_{1},\ldots,x_{N}]({\operatorname{mod}\ }p)=k[x_{1},\ldots,x_{N}]$ in the canonical way. Explicitely, a polynomial $f\in S$ has a lifting ${\tilde{f}}\in{\widetilde{S}}$ in the following form:
$f=\sum_{i\in{\mathbb{Z}}^{N}_{\geq 0}}a_{i}{\mathbf{x}}^{i}\ \ \ ,\ \ \mbox{then}\ \ \ {\tilde{f}}=\sum_{i\in{\mathbb{Z}}^{N}_{\geq 0}}{\widetilde{a}}_{i}\mathbf{x}^{i},$
where ${\widetilde{a}}_{i}$ is a lifting of $a_{i}$.
Definition 2.4.
Let $R$ and ${\widetilde{R}}$ be finitely generated algebras over fields $k$ and $K$ of characteristic $p>0$ and of 0, respectively.
Assume ${\widetilde{R}}({\operatorname{mod}\ }p)=R$ holds by compatible skeletons $\Phi_{p}:{\widetilde{S}}\to S$, where $S\subset R$ and ${\widetilde{S}}\subset{\widetilde{R}}$.
For ideals ${\mathfrak{a}}_{R}\subset R$ and ${\widetilde{\mathfrak{a}}}_{\widetilde{R}}\subset{\widetilde{R}}$ we write ${\widetilde{\mathfrak{a}}}_{\widetilde{R}}({\operatorname{mod}\ }p)={\mathfrak{a}}_{R}$ and say ${\widetilde{\mathfrak{a}}}_{\widetilde{R}}$ is a lifting of ${\mathfrak{a}}_{R}$ if we can take a generator system of ${\widetilde{\mathfrak{a}}}_{\widetilde{R}}$ from ${\widetilde{S}}$ whose image by $\Phi_{p}$ generates ${\mathfrak{a}}_{R}\subset R$. It is equivalent to the following:
  1. (1)
    There are ideals ${\mathfrak{a}}\subset S$ and ${\widetilde{\mathfrak{a}}}\subset{\widetilde{S}}$ such that ${\widetilde{\mathfrak{a}}}_{\widetilde{R}}={\widetilde{\mathfrak{a}}}{\widetilde{R}}$, ${\mathfrak{a}}_{R}={\mathfrak{a}}R$, and
  2. (2)
    $\Phi_{p}({\widetilde{\mathfrak{a}}})={\mathfrak{a}}$.
In this case, we call ${\mathfrak{a}}$ a skeleton of ${\mathfrak{a}}_{R}$, and ${\widetilde{\mathfrak{a}}}$ a skeleton of ${\widetilde{\mathfrak{a}}}_{\widetilde{R}}$.
Remark 2.5.
According to the definition above, the unit ideal ${\widetilde{R}}$ itself is a lifting of any ideal ${\mathfrak{a}}_{R}\subset R$ generated by elements of the skeleton $S\subset R$. Because the pull-back ${\widetilde{\mathfrak{a}}}:=\Phi_{p}^{-1}({\mathfrak{a}})$ satisfies the conditions above and contains $p\in{\mathbb{Z}}\subset{\widetilde{R}}$, a unit element of ${\widetilde{R}}$.
This gives a trivial example that a lifting of a prime ideal in $R$ is not necessarily prime. But the following proposition shows that for rings of special type a prime ideal has at least one lifting which is prime in the “skeleton level” and gives a proper ideal in ${\widetilde{R}}$ by the extension. Here, we note that even when the extension is a proper ideal of ${\widetilde{R}}$, it is not necessarily prime. Throughout this paper, we work on such algebras $R$ and ${\widetilde{R}}$ as follows:
Proposition-Definition 2.6.
Let $k$ and $K$ be fields of characteristic $p>0$ and 00, respectively. Let $R$ and ${\widetilde{R}}$ be $k$-subalgebra of $k[x_{1},\ldots,x_{m}]$ and $K$-subalgebra of $K[x_{1},\ldots,x_{m}]$, respectively. Denote the maximal ideal $(x_{1},\ldots,x_{m})\subset k[x_{1},\ldots,x_{m}]$ by ${\mathfrak{m}}$ and the maximal ideal $(x_{1},\ldots,x_{m})\subset K[x_{1},\ldots,x_{m}]$ by ${\widetilde{\mathfrak{m}}}$. Let ${\mathfrak{p}}_{R}\subset R$ be a prime ideal contained in ${\mathfrak{m}}$. Let $S\subset R$ be a skeleton such that ${\mathfrak{p}}:={\mathfrak{p}}_{R}\cap S$ generates ${\mathfrak{p}}_{R}$ and ${\widetilde{S}}\subset{\widetilde{R}}$ a lifting skeleton of $S$ such that the canonical map
${\widetilde{S}}\cap{\widetilde{\mathfrak{m}}}\to S\cap{\mathfrak{m}}$
is surjective.
Then, the following hold:
  1. (1)
    There is at least one prime ideal ${\widetilde{{\mathfrak{p}}}}\subset{\widetilde{S}}\cap{\widetilde{\mathfrak{m}}}\subset{\widetilde{S}}$ such that
    $\Phi_{p}({\widetilde{{\mathfrak{p}}}})={\mathfrak{p}}.$
    In this case, we call ${\widetilde{{\mathfrak{p}}}}$ a prime lifting of ${\mathfrak{p}}$ in the skeleton level. When ${\mathfrak{p}}$ corresponds to a point $P$ in the scheme defined by $R$ (eg. as ${\mathrm{Spec\ }}$ or ${\mathrm{Proj\ }}$), then the point ${\widetilde{P}}$ corresponding to ${\widetilde{{\mathfrak{p}}}}$ in the scheme defined by ${\widetilde{S}}$ is called a lifting of the point $P$ in the skeleton level.
  2. (2)
    Let prime ideals ${\mathfrak{p}}^{\prime}_{R}\subset{\mathfrak{p}}_{R}\subset R$ be generated by ${\mathfrak{p}}^{\prime}:={\mathfrak{p}}^{\prime}_{R}\cap S$ and by ${\mathfrak{p}}:={\mathfrak{p}}_{R}\cap S$, respectively. Assume there is a prime lifting ${\widetilde{{\mathfrak{p}}}}^{\prime}\subset{\widetilde{S}}\cap{\widetilde{\mathfrak{m}}}$ of ${\mathfrak{p}}^{\prime}$ in the skeleton level. Then ${\mathfrak{p}}$ has a prime lifting ${\widetilde{{\mathfrak{p}}}}\subset{\widetilde{S}}$ in the skeleton level containing ${\widetilde{{\mathfrak{p}}}}^{\prime}$.
  3. (3)
    In both cases (1) and (2) above, the prime lifting ${\widetilde{{\mathfrak{p}}}}$ of ${\mathfrak{p}}$ satisfies
    $0pt{\widetilde{{\mathfrak{p}}}}_{\mathbb{Q}}=0pt{\mathfrak{p}},$
    where ${\widetilde{{\mathfrak{p}}}}_{\mathbb{Q}}={\widetilde{{\mathfrak{p}}}}\otimes_{\mathbb{Z}}{\mathbb{Q}}\subset{\widetilde{S}}\otimes_{\mathbb{Z}}{\mathbb{Q}}$.
Remark 2.7.
The proposition above also holds for graded algebras $R,{\widetilde{R}},S,{\widetilde{S}}$ and homogeneous prime ideals ${\mathfrak{p}},{\mathfrak{p}}_{i}$, ${\widetilde{{\mathfrak{p}}}},$ and ${\widetilde{{\mathfrak{p}}}}_{i}$. where the degree is given by a weighted degree of $x_{i}$’s.
In this paper, we work on algebras of the special type: i.e., a finitely generated $k$-subalgebra $R$ of a polynomial ring. On such an algebra $R$, properties on an ideal ${\mathfrak{a}}$ of $R$ are reduced to the properties of its skeleton ideal ${\mathfrak{a}}\cap S$ as follows:
Lemma 2.8 ([8, Lemma 2.8]).
Let $K$ be a field and $\Sigma$ a subring of $K$. Denote the field of fractions of $\Sigma$ by $Q$. Let $S\subset\Sigma[x_{1},\ldots,x_{m}]$ be a finitely generated $\Sigma$-subalgebra of the polynomial ring and ${\mathfrak{a}}\subset S$ an ideal which does not contain any nonzero element of $\Sigma$. Let
$Q[S]\subset Q[x_{1},\ldots,x_{m}]\ \ \mbox{ and}\ \ K[S]\subset K[x_{1},\ldots,x_{m}]$
be the $Q$- and $K$-subalgebras generated by $S$, respectively. Let
${\mathfrak{a}}_{Q}\subset Q[S]\ \ \mbox{ and}\ \ {\mathfrak{a}}_{K}\subset K[S]$
be the ideals generated by ${\mathfrak{a}}$. Then, it follows:
  1. (i)
    $0pt{\mathfrak{a}}=0pt{\mathfrak{a}}_{Q}=0pt{\mathfrak{a}}_{K}.$
  2. (ii)
    If ${\mathfrak{a}}$ is a prime ideal, then ${\mathfrak{a}}_{Q}$ is also prime, though ${\mathfrak{a}}_{K}$ is not necessarily prime. Moreover for an element $f\in S$ which is not in $\Sigma$, the following are equivalent
    1. (a)
      $f\in{\mathfrak{a}}^{n}\setminus{\mathfrak{a}}^{n+1}$,
    2. (b)
      $f\in{\mathfrak{a}}_{Q}^{n}\setminus{\mathfrak{a}}_{Q}^{n+1}$,
    3. (c)
      $f\in{\mathfrak{a}}_{K}^{n}\setminus{\mathfrak{a}}_{K}^{n+1}$,
  3. (iii)
    In particular, for ${\operatorname{char}}K=0$, assume ${\mathfrak{a}}_{Q}$ is a locally principal prime ideal and a minimal prime ideal ${\mathfrak{p}}$ of ${\mathfrak{a}}_{K}$ is locally principal, then, the following is also equivalent to (a)–(c) above:
    1. (d)
      $f\in{\mathfrak{p}}^{n}\setminus{\mathfrak{p}}^{n+1}$ for every minimal prime ${\mathfrak{p}}$ of ${\mathfrak{a}}_{K}$.
Proof.
The proofs for (i)–(iii) are given in [8, Lemma 2.8]. ∎
Remark 2.9.
Lemma 2.8 also hold for the case that $S\subset\Sigma[x_{1},\ldots,x_{N}]$ have a graded structure and ideals are homogeneous ideals, where the degree is given by a weighted degree of $x_{i}$’s.
Proposition 2.10 ([8, Lemma 2.10]).
Let $R\subset k[x_{1},\ldots,x_{N}]$ be a finitely generated $k$-subalgebra and ${\widetilde{R}}\subset{\mathbb{C}}[x_{1},\ldots,x_{N}]$ a finitely generated ${\mathbb{C}}$-algebra. Let ${\mathfrak{a}}\subset R$ be a non-zero proper ideal and ${\widetilde{\mathfrak{a}}}\subset{\widetilde{R}}$ its lifting. Then,
$0pt{\mathfrak{a}}\leq 0pt{\widetilde{\mathfrak{a}}}.$
3. log discrepancies for multi-ideals and log canonical thresholds of ideals
In this section we study a pair $(A,{\mathfrak{a}}^{e})$ consisting of a non-singular affine variety $A$ of dimension $N$ defined over a field of arbitrary characteristic and a “multi-ideal” ${\mathfrak{a}}^{e}={\mathfrak{a}}_{1}^{e_{1}}\cdots{\mathfrak{a}}_{r}^{e_{r}}$ on $A$, where ${\mathfrak{a}}_{i}$’s are non-zero coherent ideal sheaves on $A$ and $e$ is a combination of the exponents $e=(e_{1},\ldots,e_{r})\in{\mathbb{R}}_{>0}^{r}$.
Definition 3.1.
Let $A$ be a normal variety defined over a field $k$ and $P\in A$ a (not necessarily closed) point.
Let $\varphi_{1}:A_{1}\to A$ be a proper birational morphism with $A_{1}$ normal and $E_{1}\subset A_{1}$ an irreducible divisor. Let $\varphi_{2}:A_{2}\to A$ be another proper birational morphism from a normal $A_{2}$ with an irreducible divisor $E_{2}\subset A_{2}$. Define
$E_{1}\sim E_{2}$
if the birational map $\varphi_{2}^{-1}\circ\varphi_{1}:A_{1}\dasharrow A_{2}$ is a local isomorphism at the generic points of $E_{1}$ and $E_{2}$. Then, $\sim$ becomes a equivalence relation, and such an equivalence class is called a prime divisor over $A$. In this case we denote $E_{1}$ and $E_{2}$ by the same symbol, say $E$, and we say “$E$ appears on $A_{1}$ and on $A_{2}$”. (Strictly speaking, we should be talking about the corresponding valuation instead.) If $P$ is the generic point of $\varphi_{1}(E)$, then we call $E$ a prime divisor over $A$ with the center at $P$ or just a prime divisor over $(A,P)$.
Definition 3.2.
Let $(A,{\mathfrak{a}}^{e})$ be a pair as in the beginning of this section. The log discrepancy of such a pair at a prime divisor $E$ over $A$ is defined as
$a(E;A,{\mathfrak{a}}^{e}):=k_{E}+1-\sum_{i=1}^{r}e_{i}\cdot v_{E}({\mathfrak{a}}_{i}),$
where $k_{E}$ is the coefficient of the relative canonical divisor $K_{\overline{A}/A}$ at $E$.
Here, $\varphi:\overline{A}\to A$ is a birational morphism such that $E$ appears on a normal variety $\overline{A}$. The valuation of an ideal ${\mathfrak{a}}_{i}$ at $E$ is defined as follows:
$v_{E}({\mathfrak{a}}_{i}):=\min\{v_{E}(f)\mid f\in{\mathfrak{a}}_{i}\}$
We say that the pair $(A,{\mathfrak{a}}^{e})$ is log canonical at a (not necessarily closed) point $P\in A$ if
(1) $a(E;A,{\mathfrak{a}}^{e})\geq 0,$
holds for every exceptional prime divisor $E$ over $A$ with the center containing $P\in A$.
Definition 3.3.
Let $A$ be a non-singular variety, $P\in A$ a point and ${\mathfrak{a}}^{e}$ a non-zero multi-ideal on $A$. We define the log canonical threshold of the pair $(A,{\mathfrak{a}}^{e})$ at $P$ as follows:
${\operatorname{lct}}(P;A,{\mathfrak{a}}^{e})=\sup\{c\in{\mathbb{R}}_{>0}\mid(A,({\mathfrak{a}}^{e})^{c})\ \mbox{is\ log \ canonical\ at \ }P\}.$
Let $E$ be a prime divisor over $A$ whose center on $A$ is contained in the locus of ${\mathfrak{a}}$. Define
$z(E;A,{\mathfrak{a}}^{e}):=\frac{k_{E}+1}{v_{E}({\mathfrak{a}}^{e})},$
where, we define $v_{E}({\mathfrak{a}}^{e}):=\sum_{i}e_{i}\cdot v_{E}({\mathfrak{a}}_{i})$. Then, the following is well known:
${\operatorname{lct}}(P;A,{\mathfrak{a}}^{e})=\inf\{z(E;A,{\mathfrak{a}}^{e})\mid E\ \mbox{is\ a\ prime\ divisor\ over\ }A\ \mbox{with\ the \ center\ containing\ }P\}.$
Definition 3.4.
Let $(A,{\mathfrak{a}}^{e})$ and $P\in A$ be as in Definition 3.2. Then the minimal log discrepancy is defined as follows:
  1. (1)
    When $\dim A\geq 2$,
    ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e}):=\inf\{a(E;A,{\mathfrak{a}}^{e})\mid E:\ {\operatorname{prime\ divisor\ over}}\ (A,P)\}.$
  2. (2)
    When $\dim A=1$, define ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})$ by the same definitions as in (1) if the right hand side of the equality (1) is non-negative and otherwise define ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})=-\infty$.
Here, we remark that for every pair $(A,{\mathfrak{a}}^{e})$, either ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})\geq 0$ or ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})=-\infty$ holds.
Definition 3.5.
Let $A$, $N$, and ${\mathfrak{a}}^{e}$ and $P\in A$ be as in Definition 3.2. We say that a prime divisor $E$ over $(A,P)$ computes ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})$, if
$\left\{\begin{array}[]{ll}a(E;A,{\mathfrak{a}}^{e})={\mathrm{mld}}(P;A,{\mathfrak{a}}^{e}),&\mbox{when}\ \ {\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})\geq 0\\ a(E;A,{\mathfrak{a}}^{e})<0,&\mbox{when}\ \ {\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})=-\infty\\ \end{array}\right.$
Let $A$, $P$ and ${\mathfrak{a}}^{e}$ be as in Definition 3.3. Let $E$ be a prime divisor over $A$ with the center containing $P$. We say that a prime divisor $E$ computes ${\operatorname{lct}}(P;A,{\mathfrak{a}}^{e})$ if
$z(E;A,{\mathfrak{a}}^{e})={\operatorname{lct}}(P;A,{\mathfrak{a}}^{e}).$
Remark 3.6.
Let ${\mathfrak{m}}_{P}\subset{\mathcal{O}}_{A}$ be the ideal defining $P$. If there is a log resolution of $(A,{\mathfrak{m}}_{P}\cdot{\mathfrak{a}}^{e})$ in a neighborhood of $P$, or if $e$ is a set of rational numbers, then there exists a prime divisor computing ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})$. Otherwise, the existence of such a divisor was not known in general. We will show the existence of a prime divisor computing ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})$ for a smooth point of arbitrary dimension as a corollary of our main theorem (Corollary 1.2, (i)).
Similarly, if there is a log resolution of $(A,{\mathfrak{a}}^{e})$ in a neighborhood of $P$, then there exists a prime divisor over $A$ computing ${\operatorname{lct}}(P;A,{\mathfrak{a}}^{e})$. However, even in this case a prime divisor computing ${\operatorname{lct}}(P;A,{\mathfrak{a}}^{e})$ is not necessarily a prime divisor over $(A,0)$, while its center on $A$ just contains $P$.
Lemma 3.7 ([8, Lemma 3.12]).
Let $k$ be a perfect field of characteristic $p>0$ and $A:={\mathbb{A}}_{k}^{N}$, ${\widetilde{A}}:={\mathbb{A}}_{\mathbb{C}}^{N}$. Let ${\mathfrak{a}}^{e}$ be a multi-ideal on $A$ and ${\widetilde{\mathfrak{a}}}^{e}$ a multi-ideal on ${\widetilde{A}}$. We denote the origins in $A$ and ${\widetilde{A}}$ by 00 and ${\tilde{0}}$. Assume ${\widetilde{\mathfrak{a}}}_{i}({\operatorname{mod}\ }p)={\mathfrak{a}}_{i}$ for each $i$.
Then it follows
(2) ${\mathrm{mld}}(0;A,{\mathfrak{a}}^{e})\leq{\mathrm{mld}}({\tilde{0}};{\widetilde{A}},{\widetilde{\mathfrak{a}}}^{e}).$
(3) ${\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})\leq{\operatorname{lct}}({\tilde{0}};{\widetilde{A}},{\widetilde{\mathfrak{a}}}^{e})$
In [8, Lemma 3.12], the inequality of lct’s for “ideals” is stated; however, the proof works for “multi-ideals” as well.
4. Construction of a lifting of the prime divisor
As we mentioned just after the Corollary 1.2, the corollary holds for the dimension of the variety $N=1$. So, from now on, we assume $N\geq 2$.
Proposition-Definition 4.1.
(Zariski, see eg.[11, VI, 1.3]) Let $E$ be a prime divisor over a variety $A^{\prime}$ defined over a field $k$ and appear on $\psi:A^{\prime\prime}\to A^{\prime}$. Let $Q\in A^{\prime\prime}$ be the generic point of $E$ and $Q_{0}\in A^{\prime}$ the image of $Q$. Then, we have a sequence of varieties $A^{\prime}_{0},A^{\prime}_{1},\ldots,A^{\prime}_{n}$ and birational maps as follows:
  1. $(0)$
    $A^{\prime}_{0}=A^{\prime}$, $\psi^{\prime}_{0}=\psi$.
  2. $(i)$
    If $\psi^{\prime}_{i}:A^{\prime\prime}\dashrightarrow A^{\prime}_{i}$ is already defined, then let $Q_{i}=\psi_{i}(Q)$. Let $A^{\prime}_{i+1}=Bl_{Q_{i}}A^{\prime}_{i}$ and $\psi^{\prime}_{i+1}:A^{\prime\prime}\dashrightarrow A^{\prime}_{i}$ be the induced map.
  3. $(n)$
    Then, the final birational map $\psi^{\prime}_{n}:A^{\prime\prime}\dashrightarrow A^{\prime}_{n}$ is isomorphic at $Q$, i.e., $E$ appears on $A^{\prime}_{n}$.
Here, $Bl_{Q_{i}}A^{\prime}_{i}$ is the blow-up of $A^{\prime}_{i}$ at $Q_{i}$. Let $\varphi^{\prime}_{i}:A^{\prime}_{i}\to A^{\prime}_{i-1}$ be the blow-up morphism and $E_{i}\subset A^{\prime}_{i}$ be the exceptional divisor dominating $Q_{i}$. We call this sequence $A^{\prime}_{n}\to\cdots\to A^{\prime}_{1}\to A^{\prime}_{0}=A^{\prime}$ the Zariski sequence for $E$ over $A^{\prime}$.
Note that if $Q_{0}\in A^{\prime}$ is a non-singular point, then $Q_{i}$’s are all non-singular for $i=1,\ldots,n-1$. Indeed, this is proved inductively. As $A^{\prime}_{1}$ is the blow-up at a non-singular point $Q_{0}$, it follows that $A^{\prime}_{1}$ and $E_{1}$ are non-singular in a neighborhood of ${\psi^{\prime}_{1}}^{-1}(Q_{0})$. Since $Q_{1}\in{\psi^{\prime}_{1}}^{-1}(Q_{0})$, it follows that $Q_{1}$ is again a non-singular point. In the same way, we obtain that $Q_{i}$’s are all non-singular.
Remark 4.2 (see for example [12]).
Let $A_{n}\to A_{n-1}\to\cdots A_{1}\to A$ be a sequence of blow-ups with the center at subschemes over an affine smooth variety $A={\mathrm{Spec\ }}R_{0}$. Then, each $A_{i}$ is described as $A_{i}={\mathrm{Proj\ }}R_{i}$, where $R_{i}$ is a graded ring over $R_{0}$ as follows:
(4) $R_{i}=\oplus_{\ell\geq 0}R_{0}[I_{1}T_{1},\ldots,I_{i}T_{i}]_{(\ell g_{1},\ldots,\ell g_{i-1},\ell g_{i})},\ \ \ \mbox{for}\ \ g_{1}\gg g_{2}\gg\cdots\gg g_{i}\gg..\gg g_{n},$
where $T_{1},\ldots,T_{n}$ are indeterminates, and ${I}_{1}\subset R_{0},\ldots,I_{i}\subset R_{i-1},\ldots,I_{n}\subset R_{n-1}$ are the homogeneous ideals defining the center of each blow-up. We define the multi-degree of an element $rT_{1}^{e_{1}}\cdots T_{i}^{e_{i}}\in R_{i}$ ($r\in R_{0}$) as follows:
$\deg rT_{1}^{e_{1}}\cdots T_{i}^{e_{i}}=(e_{1},\cdots,e_{i}).$
By this degree, the ring $R_{0}[I_{1}T_{1},\ldots,I_{i}T_{i}]$ becomes a multi-graded ring and $R_{0}[I_{1}T_{1},\ldots,I_{i}T_{i}]_{(\ell g_{1},\ldots,\ell g_{i-1},\ell g_{i})}$ is its homogeneous part of degree $(\ell g_{1},\ldots,\ell g_{i-1},\ell g_{i})$. Then, $R_{i}$ is a “diagonal” sub-graded ring of $R_{0}[I_{1}T_{1},\ldots,I_{i}T_{i}]$. From now on, we regard $R_{i}$ as a graded ring with respect to $\ell\in{\mathbb{Z}}_{\geq 0}$ and a homogeneous ideal of $R_{i}$ is in this grading.
We will now present a fundamental lemma that will be used in the upcoming proposition:
Lemma 4.3.
Let $A$ be an irreducible Noetherian scheme and $Z\subset A$ a closed subscheme with locally a complete intersection in $A$. Let $\psi:A^{\prime}\to A$ be the blow-up by $Z$. Then, the following hold:
  1. (i)
    $\psi^{-1}(Z)\simeq{\mathbb{P}}(I_{Z}/I_{Z}^{2})$, where $I_{Z}$ is the defining ideal of $Z$ in $A$. In particular, the irreducible components of $\psi^{-1}(Z)$ and the irreducible components of $Z$ correspond bijectively.
  2. (ii)
    Assume $A$ is a normal Gorenstein variety over a field $k$, and for any irreducible component $Z^{\prime}$ of $Z$, the singular loci ${\mathrm{Sing}}A$ and ${\mathrm{Sing}}Z$ of $A$ and of $Z,$ respectively, satisfy:
    (5) $Z^{\prime}\not\subset({\mathrm{Sing}}A)\cup({\mathrm{Sing}}Z).$
    Then $\psi^{-1}(Z)$ is Gorenstein and $A^{\prime}$ is normal and Gorenstein.
Proof.
As the embedding $Z\subset A$ is locally a complete intersection, by [4, IV, 16.9.4], it follows
$\bigoplus_{n\geq 0}I_{Z}^{n}/I^{n+1}_{Z}\simeq S(I_{Z}/I_{Z}^{2}),$
where $I_{Z}/I_{Z}^{2}$ is a locally free ${\mathcal{O}}_{Z}$-modlue of rank $c={\mathrm{codim}}(Z,A)$. The description yields $\psi^{-1}(Z)\simeq{\mathbb{P}}(I_{Z}/I_{Z}^{2})$ as required in (i). On the other hand, as $\psi^{-1}(Z)$ is a projective bundle on $Z$, its irreducible component is in the form $\psi^{-1}(Z^{\prime})$ for an irreducible component $Z^{\prime}$ of $Z$, which yields the second statement of (i).
For (ii), we use (5) to see that $A^{\prime}$ is $R_{1}$. On the other hand, we have that $Z$ is Gorenstein by the assumption on $Z$ and $A$. Then, it follows $\psi^{-1}(Z)$ is locally isomorphic to $Z\times{\mathbb{P}}^{c-1}$, which implies $\psi^{-1}(Z)$ is Gorenstein. As $\psi^{-1}(Z)$ is locally a hypersurface of $A^{\prime}$, it follows $A^{\prime}$ is Gorenstein. Applying the Serre’s criterion to $A^{\prime}$, we obtain that $A^{\prime}$ is normal. ∎
Proposition 4.4.
Let $k$ be an algebraically closed field of characteristic $p\geq 0$ and $E$ a prime divisor over the affine space $A={\mathbb{A}}_{k}^{N}$ over $k$ with the center at the origin $0\in A$. Then, there exists a sequence:
$A_{n}\stackrel{{\scriptstyle\varphi_{n}}}{{\longrightarrow}}A_{n-1}\to\cdots A_{1}\stackrel{{\scriptstyle\varphi_{1}}}{{\longrightarrow}}A$
of blow-ups such that
  1. (i)
    $E$ appears on $A_{n}$,
  2. (ii)
    the first blow-up $\varphi_{1}$ is by the maximal ideal ${\mathfrak{m}}=(x_{1},\ldots,x_{N})\subset k[x_{1},\ldots,x_{N}]$,
  3. (iii)
    for each $i$, the morphism $\varphi_{i}:A_{i}\to A_{i-1}$ is the blow-up along a complete intersection subscheme $Z_{i}\subset A_{i-1}$. One irreducible components of $Z_{i}$ is the center of $E$ on $A_{i-1}$, while the other components, if exist, are not contained in the fiber over $0\in A(=:A_{0})$ on $A_{i-1}$,
  4. (iv)
    $A_{i}$’s are all normal Gorenstein varieties and smooth at the generic point of the center of $E$ and also smooth outside of the fiber over $0\in A$.
Proof.
Let $\varphi_{1}:A_{1}\to A$ be the blow-up by the maximal ideal $(x_{1},\ldots,x_{N})\subset R_{0}:=k[x_{1},\ldots,x_{N}]$, the ideal defining the origin $0\in A$ which is the center of $E$ on $A$. (automatically (ii) and (iii) are satisfied for $i=1$). As $A_{1}$ is smooth, (iv) is satisfied for $i=1$. We will construct $\varphi_{i}$ by induction in $i$ so that (iii) and (iv) are satisfied.
Let $i\geq 2$ and assume we constructed $A_{i-1}$ with the properties (iii) and (iv). As $A_{i-1}\to A$ is projective, we can write $A_{i-1}={\mathrm{Proj\ }}R_{i-1}$ and
$R_{i-1}=\oplus_{n=0}^{\infty}J_{n},$
where $R_{0}$-submodule $J_{n}\subset R_{i-2}$ is the degree $n$ part of $R_{i-1}$. By replacing $R_{i-1}$ by a suitable Veronese subring, we may assume that $R_{i-1}$ is generated by $J_{1}$. Let ${\mathfrak{p}}_{i}\subset R_{i-1}$ be the homogeneous prime ideal defining the center of $E$ in $A_{i-1}$.
Claim. We can take a subset $\{f_{1},\ldots,f_{t}\}\subset{\mathfrak{p}}_{i}$ such that
  1. (1)
    $V(f_{1},\ldots,f_{t})=V({\mathfrak{p}}_{i})$ and each $f_{j}$ has the common degree $d>0$ in the graded ring $R_{i-1}$,
  2. (2)
    the linear system $\Lambda$ defined by the $k$-vector subspace
    $\langle f_{1},\ldots,f_{t}\rangle\subset\Gamma(A_{i-1},{\mathcal{O}}_{A_{i-1}}(d))$
    is very ample on $A_{i-1}\setminus V({\mathfrak{p}}_{i})$. In prticular, $\Lambda$ is very ample on $A_{i-1}\setminus\phi_{i-1}^{-1}(0)$, where $\phi_{i-1}:A_{i-1}\to A$ is the composite $\varphi_{1}\circ\cdots\circ\varphi_{i-1}$ and $0\in A$ is the origin. Here, note that $V({\mathfrak{p}}_{i})\subset\phi_{i-1}^{-1}(0)$.
Let $\{h_{1},\ldots,h_{s}\}$ be a generator system of $J_{1}$ over $R_{0}$, then the linear system over $k$ defined by $\{h_{1},\ldots,h_{s}\}$ is relatively very ample over $A$. Then, the affine space $A$ has a very ample linear system defined by $y_{1},\ldots,y_{m}\in R_{0}$ such that the linear system defined by the following $k$-vector subspace
$\langle y_{\alpha}\cdot h_{\beta}\rangle_{\alpha,\beta}\subset\Gamma(A_{i-1},{\mathcal{O}}_{A_{i-1}}(1))$
is very ample on $A_{i-1}$. Denote $y_{\alpha}\cdot h_{\beta}$ by $X_{\alpha\beta}$.
To prove the Claim above, we start by taking any homogeneous generator system $\{f^{\prime}_{1},\ldots,f^{\prime}_{r}\}$ of ${\mathfrak{p}}_{i}\subset R_{i-1}$ with the degree $d_{1}\leq\cdots\leq d_{r}$, respectively. We will construct an appropriate system of elements with the properties required in (1) and (2). Here, $V({\mathfrak{p}}_{i})\subset A_{i-1}$ is the center of $E$ on $A_{i-1}$.
Let $d>d_{r}$. Define
$f_{i\gamma}:=f^{\prime}_{i}\cdot X_{\gamma}^{(d-d_{i})},$
where
$\left\{X_{\gamma}^{(d-d_{i})}\middle|\gamma=1,\ldots,\left(\begin{array}[]{c}ms+d-d_{i}-1\\ d-d_{i}\\ \end{array}\right)\right\}$
is the set of all monomials in variables $X_{\alpha\beta}$ of degree $d-d_{i}$. Then the set of elements
$\left\{f_{i\gamma}\middle|1\leq i\leq r,1\leq\gamma\leq\left(\begin{array}[]{c}ms+d-d_{i}-1\\ d-d_{i}\\ \end{array}\right)\right\}$
gives the closed subset $V(f_{i\gamma})=V({\mathfrak{p}}_{i})$. Note that all $f_{i\gamma}$ are of degree $d$. Thus (1) in the Claim holds. Let $\Lambda$ be the linear system defined by the $k$-vector subspace $\langle f_{i\gamma}\rangle\subset\Gamma(A_{i-1},{\mathcal{O}}_{A_{i-1}}(d))$. Then, $\Lambda$ has the base locus $V({\mathfrak{p}}_{i})$ and is very ample outside of $V({\mathfrak{p}}_{i})$. As $V({\mathfrak{p}}_{i})\subset\phi_{i-1}^{-1}(0)$, the linear system $\Lambda$ is very ample outside of $\phi_{i-1}^{-1}(0)$. Now, denote the $\{f_{i\gamma}\}$ just by $\{f_{i}\mid i=1,\ldots,t\}$ for simplicity of notation. Then the system $\{f_{i}\mid i=1,\ldots,t\}$ satisfies (2) required in the Claim.
Let $c=0pt{\mathfrak{p}}_{i}$ and take $c$ general linear combinations of $f_{i}$’s over $k$:
$\sum_{i=1}^{r}a_{ji}f_{i}\ (a_{ji}\in k\ \mbox{ for}\ j=1,\ldots,c)$
and rename them as $f_{1},\ldots,f_{c}$. Then, $\{f_{1},\ldots,f_{c}\}$ becomes a regular sequence in $R_{i-1}$. Let $I_{i}\subset R_{i-1}$ be the ideal generated by $\{f_{1},\ldots,f_{c}\}$ and $\mathcal{I}_{i-1}$ the ideal sheaf of ${\mathcal{O}}_{A_{i-1}}$ defined by $I_{i-1}$. Similarly, let $\mathcal{P}_{i}$ be the ideal sheaf of ${\mathcal{O}}_{A_{i-1}}$ defined by ${\mathfrak{p}}_{i}$. By the choice of $f_{1},\ldots,f_{c}$, these give a regular system of paremeters of ${\mathcal{O}}_{A_{i-1},P_{i}}$, where $P_{i}\in A_{i-1}$ is the generic point of $V({\mathfrak{p}}_{i})$. Therefore,
$\mathcal{I}_{i}{\mathcal{O}}_{A_{i-1},P_{i}}=\mathcal{P}_{i}{\mathcal{O}}_{A_{i-1},P_{i}}.$
Therefore, one of the irreducible components of $Z_{i}:=V(I_{i})\subset A_{i-1}$ is the center of $E$ on $A_{i-1}$.
As $Z_{i}\subset A_{i-1}$ is obtained by successive cutting $A_{i-1}$ with $c$ general hypersurfaces
$H_{i}:=V(f_{i})\ \ (i=1,\ldots,c)$
from the linear system $\Lambda$, we obtain the following:
  1. (a)
    $\dim({\mathrm{Sing}}A_{i-1})\cap Z_{i}<N-c$ and
  2. (b)
    $\dim(\phi_{i-1}^{-1}(0)\setminus V({\mathfrak{p}}_{i}))\cap Z_{i}<N-c$.
Indeed, first we have
$\dim({\mathrm{Sing}}A_{i-1}\setminus V({\mathfrak{p}}_{i}))\cap Z_{i}<N-c.$
Because $\Lambda$ is very ample on $A_{i-1}\setminus V({\mathfrak{p}}_{i})$, each cut by a general member of $\Lambda$ makes the dimension of the intersection with ${\mathrm{Sing}}A_{i-1}\setminus V({\mathfrak{p}}_{i})$ exactly 1 less. Next we have
$\dim({\mathrm{Sing}}A_{i-1}\cap V({\mathfrak{p}}_{i}))\cap Z_{i}<N-c,$
because $\dim V({\mathfrak{p}}_{i})=N-c$ and the generic point of $V({\mathfrak{p}}_{i})$ is smooth in $A_{i-1}$ by the induction hypothesis. Thus (a) follows, and (b) also follows in the similar way.
By (a), any irreducible component of $Z_{i}$ is not contained in ${\mathrm{Sing}}A_{i-1}$, since the dimension of every irreducible component of $Z_{i}$ is $N-c$. By (b), any irreducible component of $Z_{i}$, except for $V({\mathfrak{p}}_{i})$, is not contained in the fiber $\phi_{i-1}^{-1}(0)$. Then, $Z_{i}$ satisfies the condition (iii) of the proposition.
Moreover, we obtain that
$(A_{i-1}\setminus\phi_{i-1}^{-1}(0))\cap Z_{i}$
is a smooth scheme by the Bertini’s Theorem “for very ample linear system” $\Lambda$ on a smooth quasi-projective scheme $A_{i-1}\setminus\phi_{i-1}^{-1}(0)$. Note that the “very ample version” of the Bertini’s Theorem holds for any characteristic base field (see the proof of [6, II, Theorem 8.18]). A fortiori, every irreducible component of $Z_{i}$ is not contained in ${\mathrm{Sing}}Z_{i}$ nor in ${\mathrm{Sing}}A_{i-1}$. Now the condition (5) in Lemma 4.3 is satisfied. Therefore, if $\varphi_{i}:A_{i}\to A_{i-1}$ is the blow-up with the center $Z_{i}\subset A_{i-1}$, then, $A_{i}$ is also normal and Gorenstein by the lemma. As the blow-up $\varphi_{i}$ is the same as the blow-up at the center $V({\mathfrak{p}}_{i})\subset A_{i-1}$ of $E$ around the generic point of $V({\mathfrak{p}}_{i})$, the generic point of the center of $E$ on $A_{i}$ is also smooth (see the comment below Proposition-Definition 4.1).
As $A_{i-1}\setminus\phi_{i-1}^{-1}(0)$ and ($A_{i-1}\setminus\phi_{i-1}^{-1}(0))\cap Z_{i}$ are smooth, we also obtain the blown-up space
$A_{i}\setminus\phi_{i}^{-1}(0)$
of $A_{i-1}\setminus\phi_{i-1}^{-1}(0)$ by the center $(A_{i-1}\setminus\phi_{i-1}^{-1}(0))\cap Z_{i}$ is also smooth, which completes the of (iv).
As the sequence of the blow-ups is the same as the Zariski sequence for $E$ around the centers of $E$ on any $A_{i}$ by (iii), the prime divisor $E$ appears on $A_{n}$ for a certain $n\in{\mathbb{Z}}_{\geq 0}$. Hence, (i) follows. ∎
Definition 4.5.
For a prime divisor $E$ over $A$ with the center at a closed point $0\in A$, the sequence in Proposition 4.4 is called a normal Gorenstein Zariski sequence for $E$. This definition makes sense, as the sequence is the same as the Zariski sequence for $E$ around the generic point of the center of $E$ on each $A_{i}$ by the properties (i) and (iii) in Proposition 4.4.
Proposition 4.6.
Let $E$ be a prime divisor over $A={\mathbb{A}}_{k}^{N}$ for $N\geq 2$ and an algebraically closed field $k$ of characteristic $p>0$. Let
$A_{n}\stackrel{{\scriptstyle\varphi_{n}}}{{\longrightarrow}}A_{n-1}\to\cdots A_{1}\stackrel{{\scriptstyle\varphi_{1}}}{{\longrightarrow}}A$
be a normal Gorenstein Zariski sequence for $E$. We use the same notation as in Proposition 4.4. Let $P_{i}\in A_{i-1}$ be the generic point of the center of $E$ and $E_{i}\subset A_{i}$ the exceptional divisor mapped to $P_{i}\in A_{i-1}$ by the blow-up $\varphi_{i}$, and let $E^{\prime}_{i}\subset A_{i}$ be the other exceptional divisor for $\varphi_{i}$. Then, there exist a subfield $L\subset{\mathbb{C}}$ and a sequence:
${\widetilde{A}}_{n}\stackrel{{\scriptstyle\widetilde{\varphi}_{n}}}{{\longrightarrow}}{\widetilde{A}}_{n-1}\to\cdots{\widetilde{A}}_{1}\stackrel{{\scriptstyle\widetilde{\varphi}_{1}}}{{\longrightarrow}}{\widetilde{A}}={\mathbb{A}}_{L}^{N},$
satisfying the following:
  1. (i)
    ${\widetilde{A}}_{i}({\operatorname{mod}\ }p)=A_{i}$;
  2. (ii)
    $\widetilde{\varphi}_{i}$ is the blow-up by a lifting ${\widetilde{Z}}_{i}\subset{\widetilde{A}}_{i-1}$ of $Z_{i}$. One of the irreducible components of ${\widetilde{Z}}_{i}$ has the generic point ${\widetilde{P}}_{i}\in{\widetilde{A}}_{i}$ that corresponds to the prime lifting of $P_{i}\in A_{i-1}$;
  3. (iii)
    Let ${\widetilde{E}}_{i}\subset{\widetilde{A}}_{i}$ be the exceptional divisor mapped to ${\widetilde{P}}_{i}$ by the blow-up $\widetilde{\varphi}_{i}$, then, for $1\leq i\leq n$,
    $k_{{\widetilde{E}}_{i}}=k_{E_{i}}.$
Proof.
Denote $A={\mathrm{Spec\ }}R_{0}$, $A_{i}={\mathrm{Proj\ }}R_{i}$ $(i=1,\ldots,n)$, where $R_{0}=k[x_{1},\ldots,x_{N}]$, and
$R_{i}=\oplus_{\ell\geq 0}R_{0}[I_{1}T_{1},\ldots,I_{i}T_{i}]_{(\ell g_{1},\ldots,\ell g_{i-1},\ell g_{i})},\ \ \ \mbox{for}\ \ g_{1}\gg g_{2}\gg\cdots\gg g_{i}\gg..\gg g_{n},$
as in Remark 4.2. Here, $I_{i}\subset R_{i-1}$ is the defining ideal of $Z_{i}$, the center of the blow-up $\varphi_{i}:A_{i}\to A_{i-1}$. Let ${\mathfrak{p}}_{i}\subset R_{i-1}$ be the homogeneous prime ideal defining $P_{i}$, the center of $E$ on $A_{i-1}$. Then, $Z_{i}=V({\mathfrak{p}}_{i})\cup Z^{\prime}_{i}$, where $Z^{\prime}_{i}$ is the other components of $Z_{i}$ than $V({\mathfrak{p}}_{i})$. Note that $R_{i}$ is a finitely generated $k$-subalgebra
$R_{i}\subset k[x_{1},\ldots,x_{N},T_{1},\ldots,T_{i}].$
We define a skeleton
$S_{i}\subset\Sigma[x_{1},\ldots,x_{N},T_{1},\ldots,T_{i}],$
of $R_{i}$ under an appropriately chosen skeleton $\Sigma$ of $k$ for each $i=0,\ldots,n$ in the following way:
Step 0($p$). For an arbitrary chosen skeleton $\Sigma$ of $k$, we can define $S_{0}=\Sigma[x_{1},\ldots,x_{N}]$. Then $S_{0}\subset R_{0}$ is a skeleton of $R_{0}$. (Later, we will restrict the choice of $\Sigma$ such that it contains certain necessary elements in the following steps.)
Step 1($p$). As $I_{1}={\mathfrak{p}}_{1}={\mathfrak{m}}=(x_{1},\ldots,x_{N})\subset k[x_{1},\ldots,x_{N}]$, we can take
$\underline{I}_{1}=(x_{1},\ldots,x_{N})\subset\Sigma[x_{1},\ldots,x_{N}]$
for arbitrary chosen skeleton $\Sigma$ of $k$. If we denote
$S_{1}:=\oplus_{\ell\geq 0}S_{0}[\underline{I}_{1}T_{1}]_{(\ell g_{1})},$
then $S_{1}$ is a skeleton of $R_{1}$ for every choice of a skeleton $\Sigma$ of $k$. It is clear that
$\underline{{\mathfrak{p}}}_{1}\subset S_{0}\cap{\mathfrak{m}}_{0},$
where ${\mathfrak{m}}_{0}=(x_{1},\ldots,x_{N})\subset\Sigma[x_{1},\ldots,x_{N}]$.
Step 2($p$). Let ${\mathfrak{p}}_{2}$ be generated by $g_{2,1},\ldots,g_{2,r_{2}}\in R_{1}$ and $I_{2}$ be generated by the part $g_{2,1},\ldots,g_{2,c_{2}}$ ($c_{2}:=0ptI_{2}\leq r_{2}$) of them as in the discussion in Proposition 4.4.
As the ideal ${\mathfrak{p}}_{2}\subset R_{1}$ and also $I_{2}$ are generated by finite number of elements in
$R_{1}=\oplus_{\ell\geq 0}R_{0}[I_{1}T_{1}]_{\ell g_{1}}\subset k[x_{1},\ldots,x_{N},T_{1}],$
the coefficients of each monomial in $x_{i}$ and $x_{i}T_{1}$ of the generators of ${\mathfrak{p}}_{2}$ and also $I_{2}$ are finite. Concretely, let $g_{2,1},\ldots,g_{2,r_{2}}$ be generators of ${\mathfrak{p}}_{2}$. Then,
$g_{2,j}=G_{2,j}(x_{1},\ldots,x_{N},x_{1}T_{1},\ldots,x_{N}T_{1}),$
where $G_{2,j}$ is a polynomial $G_{2,j}\in k[x_{1},\ldots,x_{N},y_{1,1},\ldots,y_{N,1}]$. Then, as $G_{2,j}$ has only finite number of non-zero coefficients, we can take a skeleton $\Sigma$ of $k$ containing all those coefficients.
Then, for any skeleton $\Sigma$ of $k$ containing those coefficients, the generators ${\mathfrak{p}}_{2}$ and those of $I_{2}$ are contained in $S_{1}$. Let $\underline{{\mathfrak{p}}}_{2}:={\mathfrak{p}}_{2}\cap S_{1}$ and $\underline{I}_{2}:=I_{2}\cap S_{1}$, and define
$S_{2}:=\oplus_{\ell\geq 0}S_{0}\left[\underline{I}_{1}T_{1},\underline{I}_{2}T_{2}\right]_{(\ell g_{1},\ell g_{2})}\ \ \mbox{for}\ \ g_{1}\gg g_{2}.$
Then, $S_{2}$ is a skeleton of $R_{2}$. Note that the generators $g_{2,1},\ldots,g_{2,r_{2}}$ of $\underline{{\mathfrak{p}}}_{2}$ has no constant term in $x_{i}$’s and $x_{i}T_{1}$’s, i.e.,
$\underline{{\mathfrak{p}}}_{2}\subset S_{1}\cap{\mathfrak{m}}_{1},$
where ${\mathfrak{m}}_{1}=(x_{1},\ldots,x_{N},T_{1})\subset\Sigma[x_{1},\ldots,x_{N},T_{1}]$. This is because the center $P_{2}\in A_{1}$ is mapped to the origin in ${\mathbb{A}}_{k}^{N}$ and every $g_{2,j}$ has positive degree on $\ell$.
On the other hand, we note that the center of $E$ is on the exceptional divisor $E_{1}:=\varphi_{1}^{-1}(0)\subset B_{1}$. So we may assume that the defining ideal $I_{E_{1}}$ of $E_{1}$ satisfies
$I_{E_{1}}\subset\underline{{\mathfrak{p}}}_{2}.$
Step $\mathbf{i}$($p$). Assume we obtain a skeleton $S_{i-1}$ of $R_{i-1}$ over a skeleton $\Sigma\subset k$ which contains necessary elements and
$\underline{{\mathfrak{p}}}_{j}\subset S_{j-1}\cap{\mathfrak{m}}_{j-1}$
for any $j\leq i-1$. By using an appropriately enlarged skeleton $\Sigma^{\prime}\supset\Sigma$ of $k$, we replace $S_{i-1}$ by $\Sigma^{\prime}[S_{i-1}]$ so that we may assume that the prime ideal ${\mathfrak{p}}_{i}\subset R_{i-1}$ and $I_{i}\subset R_{i-1}$ are generated by elements of $S_{i-1}$. Indeed, let $g\in{\mathfrak{p}}_{i}\subset R_{i-1}$ be one of the generators of ${\mathfrak{p}}_{i}$. Then, there is a polynomial $G$ with the coefficients in $k$ such that
$g=G\left(x_{1},\ldots,x_{N},g_{1,1}T_{1},\ldots,g_{1,r_{1}}T_{1},\ldots,g_{i-1,1}T_{i-1},\ldots g_{i-1,r_{i-1}}T_{i-1}\right),$
where $g_{j,l}$’s are generators of $\underline{{\mathfrak{p}}}_{j}\subset S_{j-1}$, and the part $g_{j,l}$’s $(l=1,\ldots,c_{j}=0pt{\mathfrak{p}}_{j})$ form a generator system of $\underline{I}_{j}\subset S_{j-1}$ for each $11\leq j\leq i-$ 1. Let a skeleton $\Sigma^{\prime}\supset\Sigma$ contain all coefficients of $G$, then
$g\in\Sigma^{\prime}[S_{i-1}].$
Doing this procedure for all generators $g_{i,l}$ $(l=1,\ldots,r_{i})$ of ${\mathfrak{p}}_{i}\subset R_{i-1}$, we obtain a skeleton $\Sigma^{\prime}\supset\Sigma$ of $k$ such that
$g_{i,l}\in\Sigma^{\prime}[S_{i-1}],\ (l=1,\ldots,r_{i}).$
Thus, by replacing $S_{i-1}$ by $\Sigma^{\prime}[S_{i-1}]$ with a suitable $\Sigma^{\prime}$, the ideals ${\mathfrak{p}}_{i}\subset R_{i-1}$ and $I_{i}\subset R_{i-1}$ are generated by elements of $S_{i-1}$. Let $\underline{{\mathfrak{p}}}_{i}={\mathfrak{p}}_{i}\cap S_{i-1}$ and $\underline{I}_{i}=I_{i}\cap S_{i-1}.$ Let
$S_{i}=\oplus_{\ell\geq 0}S_{0}\left[\underline{I}_{1}T_{1},\ldots,\underline{I}_{i}T_{i}\right]_{(\ell g_{1},\ldots,\ell g_{i})}\ \ \mbox{for}\ \ g_{1}\gg\cdots\gg g_{i}.$
Then, $S_{i}$ is a skeleton of $R_{i}$ and we have
$\underline{{\mathfrak{p}}}_{i}\subset S_{i-1}\cap{\mathfrak{m}}_{i-1},$
where ${\mathfrak{m}}_{i-1}=(x_{1},\ldots,x_{N},T_{1},\ldots,T_{i-1})\subset\Sigma[x_{1},\ldots,x_{N},T_{1},\ldots,T_{i-1}]$.
By doing these procedures until $i=n$, we obtain skeletons $S_{i}$ of $R_{i}$ for all $0\leq i\leq n$. Let $B_{0}:={\mathrm{Spec\ }}S_{0}$ and $B_{i}:={\mathrm{Proj\ }}{S_{i}}$ ($i\geq 1$). Then the canonical sequence
$B_{n}\to B_{n-1}\to\ldots\to B_{1}\to B_{0}$
is recognized as the sequence of blow-ups by the ideals $\underline{I}_{i}$’s. The center of the blow-up $B_{i}\to B_{i-1}$ is denoted by $Z_{i}\subset B_{i-1}$ by the same symbol as $Z_{i}\subset A_{i-1}$. By the construction, we note that the irreducible component $V({\mathfrak{p}}_{i})$ of $Z_{i}$ is contained in $\phi_{i-1}^{-1}(0)$, but any other components are not contained in $\phi_{i-1}^{-1}(0)$.
Now we take any skeleton in $k$ with the necessary elements for each step. Fix now such a skeleton $\Sigma$ (Later, for the theorem, we may replace it by a larger one if necessary).
Next we define a characteristic 00 objects inductively on $i$. Let ${\widetilde{\Sigma}}\subset{\mathbb{C}}$ be a compatible skeleton of $\Sigma$ as in Example 2.3, (1), i.e., the equality
${\widetilde{\Sigma}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}=\Sigma$
holds.
Step 0(00). Let
${\widetilde{S}}_{0}:={\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]\subset{\mathbb{C}}[x_{1},\ldots,x_{N}].$
Then, it is a compatible skeleton of $S_{0}$. Denote ${\mathrm{Spec\ }}{\widetilde{S}}_{0}$ by ${\widetilde{B}}_{0}$.
Step 1(00). Let $\underline{\widetilde{I}}_{1}={\underline{\widetilde{\mathfrak{p}}}}_{1}:=(x_{1},\ldots,x_{N})\subset{\widetilde{S}}_{0}$. and
${\widetilde{S}}_{1}:=\oplus_{\ell\geq 0}{\widetilde{S}}_{0}[\underline{\widetilde{I}}_{1}T_{1}]_{\ell g_{1}}.$
Then, ${\widetilde{S}}_{1}$ is a lifting of the skeleton $S_{1}$ and ${\widetilde{B}}_{1}:={\mathrm{Proj\ }}{\widetilde{S}}_{1}$ is the blow-up of ${\widetilde{B}}_{0}$ by the ideal $\underline{\widetilde{I}}_{1}={\underline{\widetilde{\mathfrak{p}}}}_{1}$. Denote the exceptional divisor on ${\widetilde{B}}_{1}$ by ${\widetilde{E}}_{1}$.
Step 2 (00). As we saw in Step 2 ($p$), the ideal $\underline{{\mathfrak{p}}}_{2}\subset S_{1}\cap{\mathfrak{m}}_{1}$ giving the center of $E$ satisfies
$I_{E_{1}}\subset\underline{{\mathfrak{p}}}_{2},$
where $I_{E_{1}}$ is the defining ideal of $E_{1}\subset B_{1}$. On the other hand, in ${\widetilde{B}}_{1}$, we already have the exceptional divisor ${\widetilde{E}}_{1}$ mapped to ${\widetilde{P}}_{1}=0\in{\widetilde{B}}_{0}={\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]$ by ${\widetilde{\varphi}}_{1}$, that is also a lifting of $E_{1}$. Then, by (ii) of Proposition 2.6, there is a lifting ${\underline{\widetilde{\mathfrak{p}}}}_{2}\subset{\widetilde{S}}_{1}\cap{\widetilde{\mathfrak{m}}}_{1}$ of $\underline{{\mathfrak{p}}}_{2}$ such that
$I_{{\widetilde{E}}_{1}}\subset{\underline{\widetilde{\mathfrak{p}}}}_{2}.$
Now, by using suitable $c_{2}$ elements from the generators of ${\underline{\widetilde{\mathfrak{p}}}}_{2}$, we have a lifting $\underline{\widetilde{I}}_{2}$ of $\underline{I}_{2}$. Let ${\widetilde{Z}}_{2}\subset{\widetilde{B}}_{1}$ be the closed subscheme defined by $\underline{\widetilde{I}}_{2}$. The component $V({\underline{\widetilde{\mathfrak{p}}}}_{2})$ of ${\widetilde{Z}}_{2}$ is contained in ${\widetilde{E}}_{1}$, but every other irreducible component of ${\widetilde{Z}}_{2}$ is not contained in ${\widetilde{E}}_{1}$. Because if such an irreducible component of ${\widetilde{Z}}_{2}$ was contained in ${\widetilde{E}}_{1}$, a component of $Z_{2}\subset B_{1}$ other than $V({\mathfrak{p}}_{2})$ would be contained in $E_{1}$, a contradiction to the definition of $Z_{2}$. Let
${\widetilde{S}}_{2}:=\oplus_{\ell\geq 0}{\widetilde{S}}_{0}\left[\underline{\widetilde{I}}_{1}T_{1},\underline{\widetilde{I}}_{2}T_{2}\right]_{(\ell g_{1},\ell g_{2})}\ \ \mbox{for}\ \ g_{1}\gg g_{2}.$
Then, ${\widetilde{S}}_{2}$ gives the blow up ${\widetilde{\varphi}}_{2}:{\widetilde{B}}_{2}\to{\widetilde{B}}_{1}$ by the ideal $\underline{\widetilde{I}}_{2}$.
Let ${\widetilde{E}}_{2}\subset{\widetilde{B}}_{2}$ be the irreducible divisor mapped to ${\widetilde{P}}_{2}\in{\widetilde{B}}_{1}$ by ${\widetilde{\varphi}}_{2}$. Then, on $B_{2}$ and on ${\widetilde{B}}_{2}$ the configuration of the exceptional divisors $\{E_{1},E_{2}\}$ is the same as that of $\{{\widetilde{E}}_{1},{\widetilde{E}}_{2}\}$. Let ${\mathfrak{m}}_{2}:=(x_{1},\ldots,x_{N},T_{1},T_{2})\subset\Sigma[x_{1},\ldots,x_{N},T_{1},T_{2}]$, and ${\widetilde{\mathfrak{m}}}_{2}:=(x_{1},\ldots,x_{N},T_{1},T_{2})\subset{\widetilde{\Sigma}}[x_{1},\ldots,x_{N},T_{1},T_{2}]$. Then, it is clear that there is the canonical surjection:
${\widetilde{S}}_{2}\cap{\widetilde{\mathfrak{m}}}_{2}\to S_{2}\cap{\mathfrak{m}}_{2}.$
Step i (00). Let $i\geq 2$. For each $j\leq i-1$, assume we have constructed ${\widetilde{B}}_{j}={\mathrm{Proj\ }}{\widetilde{S}}_{j}$ such that ${\widetilde{S}}_{j}$ is a lifting of the skeleton $S_{j}$ and satisfying the following:
  1. (1)
    The center ${\widetilde{Z}}_{j}\subset{\widetilde{B}}_{j-1}$ is a complete intersection. There is a prime lifting ${\widetilde{P}}_{j}\in{\widetilde{B}}_{j-1}$ of $P_{j}\in B_{j-1}$, that gives an irreducible component of ${\widetilde{Z}}_{j}$.
  2. (2)
    We denoted the irreducible exceptional divisor on $B_{j}$ mapped to $P_{j}\in B_{j-1}$ by $E_{j}$. Now, we denote the irreducible exceptional divisor on ${\widetilde{B}}_{j}$ mapped to ${\widetilde{P}}_{j}\in{\widetilde{B}}_{j-1}$ by ${\widetilde{E}}_{j}$. We denoted the proper transform of $E_{l}$ $(l\leq j-1)$ on $B_{j-1}$ by the same symbol as $E_{l}$ $(\subset B_{l})$. We also denote the proper transform of ${\widetilde{E}}_{l}$ $(l\leq j-1)$ on ${\widetilde{B}}_{j-1}$ by the same symbol as ${\widetilde{E}}_{l}$ $(\subset{\widetilde{B}}_{l})$.
    Then $\{{\widetilde{E}}_{1},\ldots,{\widetilde{E}}_{j-1},{\widetilde{P}}_{j}\}$ is the lifting of $\{E_{1},\ldots,E_{j-1},P_{j}\}$ and the configuration of $\{E_{1},\ldots,E_{j-1}\}$ around $P_{j}$ is the same as that of $\{{\widetilde{E}}_{1},\ldots,{\widetilde{E}}_{j-1}\}$ around ${\widetilde{P}}_{j}$.
We will construct ${\widetilde{B}}_{i}={\mathrm{Proj\ }}{\widetilde{S}}_{i}$ as the blow-up of ${\widetilde{B}}_{i-1}$ by an appropriate center ${\widetilde{Z}}_{i}\subset{\widetilde{B}}_{i-1}$. By the induction hypothesis above, we have that the configuration of $\{E_{1},\ldots,E_{i-1}\}$ is the same as that of $\{{\widetilde{E}}_{1},\ldots,{\widetilde{E}}_{i-1}\}$.
If the center $P_{i}\in B_{i-1}$ of $E$ is contained in $E_{i_{1}},\ldots,E_{i_{s}}$, we can take a lifting $\widetilde{P}_{i}\in{\widetilde{B}}_{i-1}$ such that
$\widetilde{P}_{i}\in{\widetilde{E}}_{i_{1}}\cap\ldots\cap{\widetilde{E}}_{i_{s}},$
by Proposition 2.6, (2). To apply Proposition 2.6 (2), here we use the fact that $P_{i}$ is a non-singular point of $A_{i-1}$, and that $E_{i_{1}},\ldots,E_{i_{s}}$ intersect normally at $P_{i}$. This allows us to assume that the intersection $E_{i_{1}}\cap\ldots\cap E_{i_{s}}$ is irreducible in a neighborhood of $P_{i}$. Consequently, we may assume that the ideal $I_{E_{i_{1}}\cap\ldots\cap E_{i_{s}}}$ is prime and admits a prime lifting $I_{{\widetilde{E}}_{i_{1}}\cap\ldots\cap{\widetilde{E}}_{i_{s}}}$ at the skeleton level, so that Proposition 2.6 (2) can be applied.
Assume the prime ideal $\underline{{\mathfrak{p}}}_{i}\subset S_{i-1}$ corresponding to $P_{i}$ is generated by $g_{i1},\ldots g_{ir_{i}}$, and $\underline{I}_{i}\subset S_{i-1}$ is generated by $g_{i1},\ldots g_{ic_{i}}$ $(c_{i}\leq r_{i})$. Let ${\underline{\widetilde{\mathfrak{p}}}}_{i}\subset S_{i-1}$ be a prime lifting corresponding to ${\widetilde{P}}_{i}\in B_{i-1}$. Then, we can take liftings ${\tilde{g}}_{i1},\ldots{\tilde{g}}_{ir_{i}}\in{\widetilde{S}}_{i-1}$ of $g_{i1},\ldots g_{ir_{i}}$ generating ${\underline{\widetilde{\mathfrak{p}}}}_{i}$ corresponding to ${\widetilde{P}}_{i}\in{\widetilde{B}}_{i-1}$. Define
$\underline{\widetilde{I}}_{i}=({\tilde{g}}_{i1},\ldots{\tilde{g}}_{ic_{i}}).$
Then $\{{\tilde{g}}_{i1},\ldots{\tilde{g}}_{ic_{i}}\}$ is a regular sequence in ${\widetilde{S}}_{i-1}$, because $\{g_{i1},\ldots g_{ic_{i}}\}$ is a regular sequence in $S_{i-1}={\widetilde{S}}_{i-1}\otimes{{\mathbb{Z}}/(p)}$. Therefore it defines a complete intersection subscheme ${\widetilde{Z}}_{i}\subset{\widetilde{B}}_{i-1}$ in ${\widetilde{B}}_{i-1}$, that is a lifting of ${\widetilde{Z}}_{i}$ and has an irreducible component corresponding to ${\widetilde{P}}_{i}\in{\widetilde{B}}_{i-1}$.
Define
${\widetilde{S}}_{i}:=\oplus_{\ell\geq 0}{\widetilde{S}}_{0}\left[\underline{\widetilde{I}}_{1}T_{1},\ldots,\underline{\widetilde{I}}_{i}T_{i}\right]_{(\ell g_{1},\ldots,\ell g_{i})}.$
Then, ${\widetilde{S}}_{i}$ is a lifting of the skeleton $S_{i}$ and the canonical inclusion ${\widetilde{S}}_{i-1}\to{\widetilde{S}}_{i}$ gives the blow-up $\widetilde{\varphi}_{i}:{\widetilde{B}}_{i}\to{\widetilde{B}}_{i-1}$ by $\underline{\widetilde{I}}_{i}$ and obtain the exceptional divisor ${\widetilde{E}}_{i}\subset{\widetilde{B}}_{i}$ dominating ${\widetilde{P}}_{i}\in{\widetilde{E}}_{i-1}$. It also follows the surjection:
${\widetilde{S}}_{i}\cap{\widetilde{\mathfrak{m}}}_{i}\to S_{i}\cap{\mathfrak{m}}_{i},$
where ${\widetilde{\mathfrak{m}}}_{i}=(x_{1},\ldots,x_{N},T_{1},\ldots,T_{i})$.
By carrying out each step up to $n$, we obtain the sequence:
(6) ${\widetilde{B}}_{n}\to\ldots\to{\widetilde{B}}_{0}$
of blow-ups by the ideal $\underline{\widetilde{I}}_{i}\in{\widetilde{B}}_{i-1}$ such that
(7) ${\widetilde{P}}_{i}\in{\widetilde{E}}_{j}\ \ \mbox{iff}\ \ P_{i}\in E_{j}\ \ \mbox{ for}\ \ j\leq i-1.$
On the other hand, by (i) of Lemma 2.8, we have
(8) $0pt{\mathfrak{p}}_{i}=0pt\underline{{\mathfrak{p}}}_{i},\ \mbox{and}\ \ 0pt({\underline{\widetilde{\mathfrak{p}}}}_{i}\otimes_{\mathbb{Z}}{\mathbb{Q}})=0pt({\underline{\widetilde{\mathfrak{p}}}}_{i}\otimes_{\widetilde{\Sigma}}L),$
where, for the second equality, we note the equality of the fields of fractions
$Q({\widetilde{\Sigma}}\otimes_{\mathbb{Z}}{\mathbb{Q}})=Q({\widetilde{\Sigma}})=:L,$
and the fact that ${\underline{\widetilde{\mathfrak{p}}}}_{i}\otimes_{\mathbb{Z}}{\mathbb{Q}}$ does not contain any non-zero element of ${\widetilde{\Sigma}}\otimes_{\mathbb{Z}}{\mathbb{Q}}$.
By (iii) of Proposition 2.6 the prime lifting ${\underline{\widetilde{\mathfrak{p}}}}_{i}$ of $\underline{{\mathfrak{p}}}_{i}$ satisfies
(9) $0pt\underline{{\mathfrak{p}}}_{i}=0pt({\underline{\widetilde{\mathfrak{p}}}}_{i}\otimes_{\mathbb{Z}}{\mathbb{Q}}).$
By (8) and (9), it follows:
(10) $0pt{\mathfrak{p}}_{i}=0pt({\underline{\widetilde{\mathfrak{p}}}}_{i}\otimes_{\widetilde{\Sigma}}L),$
Let ${\widetilde{A}}_{i}:={\widetilde{B}}_{i}\otimes_{\widetilde{\Sigma}}L$, then ${\widetilde{A}}_{i}({\operatorname{mod}\ }p)=A_{i}$ with the compatible skeletons ${\widetilde{S}}_{i}$ and $S_{i}$, and it follows the sequence of blow-ups:
${\widetilde{A}}_{n}\stackrel{{\scriptstyle\widetilde{\varphi}_{n}}}{{\longrightarrow}}{\widetilde{A}}_{n-1}\to\cdots{\widetilde{A}}_{1}\stackrel{{\scriptstyle\widetilde{\varphi}_{1}}}{{\longrightarrow}}{\widetilde{A}}={\mathbb{A}}_{L}^{N}.$
We denote the exceptional divisor on ${\widetilde{A}}_{i}$ corresponding to ${\widetilde{E}}_{i}\subset{\widetilde{B}}_{i}$ by the same symbol ${\widetilde{E}}_{i}$. Then, (7) and (10) yields $k_{{\widetilde{E}}_{i}}=k_{E_{i}}$, because
$k_{E_{i}}=0pt{\mathfrak{p}}_{i}-1+\sum_{P_{i}\in E_{j}}k_{E_{j}}\ \ \mbox{and}$
$k_{{\widetilde{E}}_{i}}=0pt({\underline{\widetilde{\mathfrak{p}}}}_{i}\otimes_{\widetilde{\Sigma}}L)-1+\sum_{{\widetilde{P}}_{i}\in{\widetilde{E}}_{j}}k_{{\widetilde{E}}_{j}}.$
Here, we note that the point ${\widetilde{P}}_{i}\in{\widetilde{A}}_{i-1}$ corresponding to the lifting ${\underline{\widetilde{\mathfrak{p}}}}_{i}$ is non-singular, as the local ring ${\mathcal{O}}_{{\widetilde{A}}_{i-1},{\widetilde{P}}_{i}}$ has the maximal ideal generated by $c_{i}$ elements, where $c_{i}=0pt{\underline{\widetilde{\mathfrak{p}}}}_{i}=0pt\underline{{\mathfrak{p}}}_{i}$. Indeed, exactly $c_{i}$ elements from the generators of $\underline{{\mathfrak{p}}}_{i}$ give the regular system of parameters of the local ring ${\mathcal{O}}_{A_{i-1},P_{i}}$. Then, the lifting of these $c_{i}$ generators give the regular sequence defining the maximal ideal of the local ring ${\mathcal{O}}_{{\widetilde{A}}_{i-1},{\widetilde{P}}_{i}}$.
5. Proof of the Main Theorem and applications
We fix the notation used in this section. Let $k$ be an algebraically closed field of characteristic $p>0$. Let $A={\mathbb{A}}_{k}^{N}$ be the affine space over $k$ of dimension $N\geq 2$ and 00 the origin of $A$.
Let $E$ be a prime divisor over $(A,0)$ and
$A_{n}\stackrel{{\scriptstyle\varphi_{n}}}{{\longrightarrow}}A_{n-1}\to\cdots A_{1}\stackrel{{\scriptstyle\varphi_{1}}}{{\longrightarrow}}A$
be a normal Gorenstein Zariski sequence for $E$. Proposition 4.6 guarantees the existence of a reasonable lifting of the sequence:
${\widetilde{A}}_{n}\stackrel{{\scriptstyle\widetilde{\varphi}_{n}}}{{\longrightarrow}}{\widetilde{A}}_{n-1}\to\cdots{\widetilde{A}}_{1}\stackrel{{\scriptstyle\widetilde{\varphi}_{1}}}{{\longrightarrow}}{\widetilde{A}}={\mathbb{A}}_{L}^{N},$
In the following discussions, we will use the notation in Proposition 4.6.
As in [8], for a technical reason, we will blow-up two more times from $A_{n}$ and lift them, which will lead to the main statement in this paper.
Lemma 5.1.
Let $E$ be a prime divisor over $(A,0)$, and
$A_{n}\stackrel{{\scriptstyle\varphi_{n}}}{{\longrightarrow}}A_{n-1}\to\cdots A_{1}\stackrel{{\scriptstyle\varphi_{1}}}{{\longrightarrow}}A$
be a normal Gorenstein Zariski sequence for $E$, We use the same notation as in Proposition 4.4 and 4.6. Let
$A_{n+2}\stackrel{{\scriptstyle\varphi_{n+2}}}{{\longrightarrow}}A_{n+1}\stackrel{{\scriptstyle\varphi_{n+1}}}{{\longrightarrow}}A_{n}$
be the composite of two blow-ups at general closed points $P_{n+1}\in E_{n}\subset A_{n}$ and $P_{n+2}\in F_{1}\subset A_{n+1}$, where $F_{1}$ is the exceptional divisor for $\varphi_{n+1}$. Denote the exceptional divisor for $\varphi_{n+2}$ by $F_{2}$.
Then, by taking a skeletons $\Sigma\subset k$ sufficiently large, and letting ${\widetilde{\Sigma}}\subset{\mathbb{C}}$ be its lifting, one obtains liftings ${\widetilde{P}}_{n+1}\in{\widetilde{A}}_{n}$ and ${\widetilde{P}}_{n+2}\in{\widetilde{A}}_{n+1}$ of $P_{n}$ and $P_{n+1}$, respectively, and the sequence:
${\widetilde{A}}_{n+2}\stackrel{{\scriptstyle\widetilde{\varphi}_{n+2}}}{{\longrightarrow}}{\widetilde{A}}_{n+1}\stackrel{{\scriptstyle\widetilde{\varphi}_{n+1}}}{{\longrightarrow}}{\widetilde{A}}_{n},$
the composite of two blow-ups at ${\widetilde{P}}_{n+1}\in{\widetilde{E}}_{n}\subset{\widetilde{A}}_{n}$ and ${\widetilde{P}}_{n+2}\in{\widetilde{F}}_{1}\subset{\widetilde{A}}_{n+1}$, where ${\widetilde{F}}_{1}$ and ${\widetilde{F}}_{2}$ are the exceptional divisor of $\widetilde{\varphi}_{n+1}$ and $\widetilde{\varphi}_{n+2}$, respectively.
Then, we obtain
$k_{F_{1}}=k_{{\widetilde{F}}_{1}}=N-1+k_{E_{n}}=N-1+k_{{\widetilde{E}}_{n}}\ \mbox{and}$
$k_{F_{2}}=k_{{\widetilde{F}}_{2}}=2(N-1)+k_{E_{n}}=2(N-1)+k_{{\widetilde{E}}_{n}}.$
Proof.
Take a general closed point $P_{n+1}\in E_{n}\subset A_{n}$ such that
$P_{n+1}\not\in\bigcup_{j<n}(E_{j}\cup E^{\prime}_{j}),$
where we denote the proper transform of $E_{j}$ and $E^{\prime}_{j}$ on $A_{n}$ by the same symbols as $E_{j}$ and $E^{\prime}_{j}\subset A_{j}$. Here, we denote the fiber $\varphi_{j}^{-1}(Z_{j})=E_{j}+E^{\prime}_{j}$ as Cartier divisors on $A_{j}$.
Next, take a general closed point $P_{n+2}\in F_{1}\subset A_{n+1}$ such that
$P_{n+2}\not\in E_{n},$
where $E_{n}$ denotes the proper transform of $E_{n}$ on $A_{n+1}$ as well. Here, we note that $P_{n+1}$ and $P_{n+2}$ are $k$-valued point as $k$ is algebraically closed. So, we can realize their defining ideals in the skeletons $S_{n}$ and $S_{n+1}$ for sufficiently large $\Sigma\subset k$. Then, in the same way as in Proposition 4.4, we can construct a sequence
${\widetilde{A}}_{n+2}\stackrel{{\scriptstyle\widetilde{\varphi}_{n+2}}}{{\longrightarrow}}{\widetilde{A}}_{n+1}\stackrel{{\scriptstyle\widetilde{\varphi}_{n+1}}}{{\longrightarrow}}{\widetilde{A}}_{n},$
of blow-ups with the centers ${\widetilde{P}}_{n+1}\in{\widetilde{E}}_{n}$ and ${\widetilde{P}}_{n+2}\in{\widetilde{F}}_{1}$ of liftings of $P_{n+1}$ and $P_{n+2}$, respectively. Hence, by the choice of the points of blow-ups, we obtain the required equalities. ∎
Lemma 5.2.
Under the notation in the proof of Proposition 4.4 and Lemma 5.1, consider the ideals
$\underline{\widetilde{I}}_{1}={\widetilde{\mathfrak{m}}}=(x_{1},\ldots,x_{N})\subset{\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]\ \ \mbox{and}$
$\underline{I}_{1}={\mathfrak{m}}=(x_{1},\ldots,x_{N})\subset\Sigma[x_{1},\ldots,x_{N}].$
Then, the following hold for $21\leq i\leq n+$ 2:
  1. (i)
    ${\mathcal{O}}_{{\widetilde{B}}_{i}}/{\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{i}}$ is ${\mathbb{Z}}$-flat at $p\in{\mathbb{Z}}$ for every $t\in{\mathbb{N}}$, and
  2. (ii)
    ${\mathcal{O}}_{{\widetilde{B}}_{i}}/{\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{i}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}={\mathcal{O}}_{B_{i}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{B_{i}}$.
Proof.
For $21\leq i\leq n+$ 2, denote the proper transforms on $B_{i}$ of the prime divisors $E_{j}\subset B_{j}$ and ${\widetilde{E}}_{j}\subset{\widetilde{B}}_{j}$ $(i<j)$ by the same symbols $E_{j}$ and ${\widetilde{E}}_{j}$, as we did in the previous discussions. (For consistency of the notation, only in this proof, we denote $F_{i}$ and ${\widetilde{F}}_{i}$ $(i=1,2)$ by $E_{n+i}$ and ${\widetilde{E}}_{n+i}$, respectively.) As $Z_{i}$ ($i\geq 2$) is not necessarily irreducible, the irreducible component of the exceptional divisor of $\varphi_{i}:B_{i}\to B_{i-1}$ is not only $E_{i}$.
As $\widetilde{\varphi}_{1}$ is the blow-up by $\widetilde{\mathfrak{m}}=(x_{1},\ldots,x_{N})\subset{\widetilde{S}}_{0}$, the pull-back ${\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{1}}$ is an invertible ideal sheaf on ${\widetilde{B}}_{1}$. Then, for every $i\geq 1$ its pull-back ${\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{i}}$ on ${\widetilde{B}}_{i}$ is also an invertible ideal sheaf on ${\widetilde{B}}_{i}$. By the similar reason, ${\mathfrak{m}}^{t}{\mathcal{O}}_{B_{i}}$ is also an invertible ideal sheaf for every $i\geq 1$.
For the proof of (i), let ${\widetilde{D}}_{i}\subset{\widetilde{B}}_{i}$ be the subscheme defined by ${\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{i}}$, and $D_{i}\subset B_{i}$ the subscheme defined by ${\mathfrak{m}}^{t}{\mathcal{O}}_{B_{i}}$ for $21\leq i\leq n+$ 2. Then, ${\widetilde{D}}_{i}\subset{\widetilde{B}}_{i}$ is locally a hypersurface and its support is contained in the exceptional divisor of $\widetilde{\phi}_{i}:{\widetilde{B}}_{i}\to{\widetilde{B}}_{0}$. Since the exceptional divisor of $\widetilde{\phi}_{i}$ does not contain $B_{i}$, the closed subscheme ${\widetilde{D}}_{i}$ does not contain the fiber $B_{i}$ over $p$, which yields that ${\widetilde{D}}_{i}$ is flat over ${\mathbb{Z}}$ at $p$ (see, [14, Corollary, p.177]). This completes the proof of (i)
For the proof of (ii), we first claim that the supports of $D_{i}$ and ${\widetilde{D}}_{i}$ are expressed as follows:
(11) ${\mathrm{supp}}D_{i}=\bigcup_{j\leq i}E_{j},\ \mbox{and}\ \ {\mathrm{supp}}{\widetilde{D}}_{i}=\bigcup_{j\leq i}{\widetilde{E}}_{j}\ \ \mbox{in\ a\ neighborhood\ of}\ p.$
Here, we remark again that the irreducible components of the exceptional set of $\widetilde{\phi}_{i}$ are not only ${\widetilde{E}}_{1},\ldots,{\widetilde{E}}_{i}$, because each center of the blow-up ${\widetilde{\varphi}}_{j}$ $(j\leq i)$ is not necessarily irreducible. We will prove these equalities in (11) by induction on $i$.
For $i=1$ the equalities are trivial. So we assume $i\geq 2$ and the equalities hold for $i-1$. Now, assume there exists an irreducible component $C$ of $D_{i}$ such that
$C\not\subset\bigcup_{j\leq i}E_{j}.$
Then, $\varphi_{i}(C)\subset D_{i-1}$. Therefore, $\varphi_{i}(C)\subset Z_{i}$, because otherwise, $\varphi_{i}$ is isomorphic at the generic point of $C$ and by the hypothesis of the induction, we have $\varphi_{i}(C)\subset E_{j}\subset B_{i-1}$ for some $j<i$, which gives $C\subset E_{j}\subset B_{i}$, a contradiction to the assumption. Hence, we obtain
$\varphi_{i}(C)\subset Z_{i}\cap D_{i-1}\subset B_{i-1}\ \mbox{ and}\ \ \varphi_{i}(C)\not\subset V({\mathfrak{p}}_{i}).$
By the construction of $Z_{i}$ (see, Proposition 4.4), it follows
${\mathrm{codim}}(Z_{i}\cap D_{i-1}\setminus V(\underline{{\mathfrak{p}}}_{i}),B_{i-1})=c_{i}+1:=0pt{\mathfrak{p}}_{i}+1,$
which yields
${\mathrm{codim}}(C,B_{i})\geq{\mathrm{codim}}(\varphi_{i}^{-1}(\varphi_{i}(C)),B_{i})\geq{\mathrm{codim}}(\varphi_{i}(C),B_{i-1})-(c_{i}-1)$
$\geq{\mathrm{codim}}(Z_{i}\cap D_{i-1}\setminus V(\underline{{\mathfrak{p}}}_{i}),B_{i-1})-(c_{i}-1)=2,$
a contradiction to that $C$ is an irreducible component of a closed subscheme defined by a locally principal ideal on an irreducible reduced scheme $B_{i}$. Therefore we obtain
${\mathrm{supp}}D_{i}=\bigcup_{j\leq i}E_{j}.$
On the other hand, we have
${\mathrm{codim}}({\widetilde{Z}}_{i}\cap{\widetilde{D}}_{i-1}\setminus V({\underline{\widetilde{\mathfrak{p}}}}_{i}),{\widetilde{B}}_{i-1})\geq{\mathrm{codim}}(Z_{i}\cap D_{i-1}\setminus V(\underline{{\mathfrak{p}}}_{i}),B_{i-1})=c_{i}+1$
in a neighborhood of $p\in{\mathbb{Z}}$. Then, by the similar discussion as above, we obtain
${\mathrm{supp}}{\widetilde{D}}_{i}=\bigcup_{j\leq i}{\widetilde{E}}_{i}.$
In order to prove (ii), it is sufficient to show the equality
${\mathfrak{m}}^{t}{\mathcal{O}}_{B_{i}}={\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{i}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}$
of the ideal sheaves for every $11\leq i\leq n+$ 1. Note that it is equivalent to the equality of their extensions to the normal variety $B_{iQ}:=B_{i}\otimes_{\Sigma}Q(\Sigma)$, where $Q(\Sigma)$ is the field of fractions of $\Sigma$. As in Lemma 5.1, we denote ${\widetilde{A}}_{i}={\widetilde{B}}_{i}\otimes_{\widetilde{\Sigma}}Q({\widetilde{\Sigma}})$. Then, as $B_{iQ}$ and ${\widetilde{A}}_{i}$ are normal varieties, and from the claim proved above, the ideal sheaves have the following expressions:
${\mathfrak{m}}^{t}{\mathcal{O}}_{B_{iQ}}={\mathcal{O}}_{B_{iQ}}\left(-\sum_{j\leq i}a_{j}E_{j}\right),$
${\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{A}}_{i}}={\mathcal{O}}_{{\widetilde{A}}_{i}}\left(-\sum_{j\leq i}{\widetilde{a}}_{j}{\widetilde{E}}_{j}\right).$
As $\left({\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{i}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}\right)\otimes_{\Sigma}Q(\Sigma)={\mathcal{O}}_{B_{iQ}}(-\sum_{j\leq i}{\widetilde{a}}_{j}E_{j})$, for the proof of (ii), it is sufficient to show $a_{j}={\widetilde{a}}_{j}$ ($j\leq i$). We will prove this by induction on $i$.
For $i=1$, it is clear that $a_{1}={\widetilde{a}}_{1}$. Assume $i\geq 2$ and $a_{j}={\widetilde{a}}_{j}$ holds for $j\leq i-1$. The center $P_{i}\in B_{i-1Q}$ of the blow-up $\varphi_{i}$ and the center ${\widetilde{P}}_{i}\in{\widetilde{A}}_{i-1}$ of the blow-up ${\widetilde{\varphi}}_{i}$ satisfies
$P_{i}\in E_{j}$ if and only if ${\widetilde{P}}_{i}\in{\widetilde{E}}_{j}$ for $j\leq i$.
$a_{i}=\sum_{P_{i}\in E_{j}}a_{j}=\sum_{{\widetilde{P}}_{i}\in{\widetilde{E}}_{j}}{\widetilde{a}}_{j}={\widetilde{a}}_{i},$
where the first and the last equalities follows from the fact that $P_{i}\in A_{i-1}$ and ${\widetilde{P}}_{i}\in{\widetilde{A}}_{i-1}$ are non-singular (see the last part of the proof of Proposition 4.6). Hence, the statement holds for $i$. ∎
Theorem 5.3 (Theorem 1.1, (i), (ii)).
Let $N$ be an integer $\geq 2$, and $k$ an algebraically closed field of characteristic $p>0$. Let $A={\mathbb{A}}_{k}^{N}$ be the affine space over $k$ of dimension $N$, and 00 the origin of $A$. Let $E$ be a prime divisor over $A$ with the center at 00, and ${\mathfrak{a}}\subset k[x_{1},\ldots,x_{N}]$ a non-zero ideal. Then, there exists a prime divisor $F_{\mathbb{C}}$ over the affine space $A_{\mathbb{C}}={\mathbb{A}}_{\mathbb{C}}^{N}$ over ${\mathbb{C}}$ with the center at the origin ${\tilde{0}}$ satisfying the following:
  1. (i)
    $2(N-1)+k_{E}=k_{F_{\mathbb{C}}}$;
  2. (ii)
    there exists a lifting ${\widetilde{\mathfrak{a}}}\subset{\mathcal{O}}_{A_{\mathbb{C}}}$ of ${\mathfrak{a}}$ satisfying
    $v_{E}({\mathfrak{a}})=v_{F_{\mathbb{C}}}({\widetilde{\mathfrak{a}}}).$
Proof.
We may assume $N\geq 4$, because the theorem is proved for $N\leq 3$ by [8, Theorem 5.3]. We replace the sequence of non-singular blow-ups in [8, Proposition 4.2] by our normal Gorenstein Zariski sequence. By [8, Remark 5.8], only parts we need to verify are
  1. (1)
    Lemma 5.2 for $\widetilde{\phi}_{n+2}$, a lifting of $\phi_{n+2}$, which is our normal Gorenstein Zariski sequence plus two more blow-ups, and
  2. (2)
    $R^{1}\phi_{n+2}{\mathcal{O}}_{A_{n+2}}=0$.
Actually, the proof of (i) of the theorem follows immediately from Lemma 5.1 as in [8, Theorem 5.3]. The proof of (ii) in [8, Theorem 5.6] for $\dim\geq 4$ was divided into four statements; (a)–(d). These (a)–(d) were proved by using (1) and (2) for composite of smooth blow-ups in [8].
Now, in our case for normal Gorenstein Zariski sequence, Lemma 5.2 is proved above and the vanishing (2) holds for any proper birational morphism $A_{n+2}\to A$ from a normal variety $A_{n+2}$ to a smooth variety $A$ defined over a field of arbitrary characteristic ([9],[13]). Here, we note that $A_{n+2}$ is normal, since it is obtained by just two more blow-ups by smooth closed points from a normal Gorenstein variety $A_{n}$. ∎
Corollary 5.4 (Theorem 1.1, (iii)).
Let $k$ be an algebraically closed field of characteristic $p>0$, $A={\mathbb{A}}_{k}^{N}$ the affine space over $k$ of dimension $N\geq 2$ and 00 the origin of $A$. Let $E$ be a prime divisor over $(A,0)$ and ${\mathfrak{a}}_{1},\ldots,{\mathfrak{a}}_{r}\subset k[x_{1},\ldots,x_{N}]$ non-zero ideals. Then, there exists a prime divisor $F_{\mathbb{C}}$ over the affine space $A_{\mathbb{C}}={\mathbb{A}}_{\mathbb{C}}^{N}$ over ${\mathbb{C}}$ with the center at the origin ${\tilde{0}}$ satisfying the following:
There exist liftings ${\widetilde{\mathfrak{a}}}_{i}\subset{\mathcal{O}}_{A_{\mathbb{C}}}$ of ${\mathfrak{a}}_{i}$ $(i=1,\ldots,r)$ such that for every multi-ideal ${\mathfrak{a}}^{e}:={\mathfrak{a}}_{1}^{e_{1}}\cdots{\mathfrak{a}}_{r}^{e_{r}}$ $(e_{i}\in{\mathbb{R}}_{>0})$ on $A$, the lifting ${\widetilde{\mathfrak{a}}}^{e}={\widetilde{\mathfrak{a}}}_{1}^{e_{1}}\cdots{\widetilde{\mathfrak{a}}}_{r}^{e_{r}}$ of ${\mathfrak{a}}^{e}$ on $A_{\mathbb{C}}$ satisfies
$2(N-1)+a(E;A,{\mathfrak{a}}^{e})=a(F_{\mathbb{C}};A_{\mathbb{C}},{\widetilde{\mathfrak{a}}}^{e}).$
Proof.
By (i) and (ii) of the theorem, the equality
$2(N-1)+a(E;A,{\mathfrak{a}})=a(F_{\mathbb{C}};A_{\mathbb{C}},{\widetilde{\mathfrak{a}}})$
is obvious for an ideal. For multi-ideal case, as the number of ideals is finite in the multi-ideal, we can take common centers $P_{n+1}$ and $P_{n+2}$ general for all ideals. Then we obtain a lifting ${\widetilde{\mathfrak{a}}}_{i}$ of ${\mathfrak{a}}_{i}$ for each $i$ such that $v_{E}({\mathfrak{a}}_{i})=v_{F_{\mathbb{C}}}({\widetilde{\mathfrak{a}}}_{i})$, which completes the proof of our corollary. ∎
Until now, we assume the base space $A$ to be the affine space. However, we can apply the theorem also to a pair consisting of a non-singular variety and a multi-ideal with a real exponent. In such a case, we cannot define a lifting of a regular function around the point 00, but passing through the completion we can reduce the discussion to the case of the affine space.
Corollary 5.5 (Corollary 1.2, (i)).
Let $k$ be an algebraically closed field of characteristic $p>0$. Let $0\in A$ be a non-singular closed point of a variety $A$ of dimension $N$. Then, for fixed $N$ and an exponent $e\in{\mathbb{R}}_{>0}^{r}$, the set of log discrepancies
$\Lambda_{N,e}:=\left\{a(E;A,{\mathfrak{a}}^{e})\middle|\begin{array}[]{l}E:\mbox{prime\ divisor\ over\ }(A,0)\\ (A,{\mathfrak{a}}^{e}):\mbox{log\ canonical\ at\ }0\\ \end{array}\right\}$
[…]
The proof of Corollary 5.5 is given in the same way as [8, Corollary 5.10]. The following is obtained by applying Corollary 5.5 to a special case.
Corollary 5.6.
Let $k$ be an algebraically closed field of characteristic $p>0$. Let $0\in A$ be a non-singular closed point of a variety $A$ of dimension $N$, and ${\mathfrak{a}}^{e}$ a multi-ideal with a real exponent. Then, there exists a prime divisor over $A$ with the center at 00 computing
${\mathrm{mld}}(0;A,{\mathfrak{a}}^{e}).$
Proof.
If ${\mathrm{mld}}(0;A,{\mathfrak{a}}^{e})=-\infty$, then a prime divisor $E$ with $a(E;A,{\mathfrak{a}}^{e})<0$ computes the mld. If ${\mathrm{mld}}(0;A,{\mathfrak{a}}^{e})\geq 0$, the discreteness of log discrepancies in Corollary 5.5 gives the existence of a prime divisor computing the mld. ∎
The existence of a prime divisor computing ${\mathrm{mld}}$ in positive characteristic was already proved for a special case by K. Shibata ([15]).
Definition 5.7.
Let $k$ be a field and $0\in X$ a non-singular $k$-valued point of dimension $N\geq 2$. Fix an exponent $e\in{\mathbb{R}}_{>0}^{r}$. We define
$T_{N,e,k}:=\{{\operatorname{lct}}(0;X,{\mathfrak{a}}^{e})\mid{\mathfrak{a}}^{e}\ \mbox{ a\ multi-ideal\ with\ the\ exponent\ }e\mbox{ on}\ X\}$
the set of log canonical thresholds.
Corollary 5.8.
Let $k$ be an algebraically closed field $k$ of positive characteristic and $0\in A$ a non-singular closed point in a variety $A$ of dimension $N$. Fix an exponent $e\in{\mathbb{R}}_{>0}^{r}$. Then,
$T_{N,e,k}\subset T_{N,e,{\mathbb{C}}}.$
Proof.
As is seen in the proof of Corollary 5.19 in [8] we may assume that $A={\mathbb{A}}_{k}^{N}$ and 00 is the origin of ${\mathbb{A}}_{k}^{N}$. As the corollary is proved for $N\leq 3$ in [8] we may assume that $N\geq 4$. For simplicity, we denote ${\sum e_{i}v_{E}({\mathfrak{a}}_{i})}$ by $v_{E}({\mathfrak{a}}^{e})$.
In [3], they prove that $T_{N-1,e,{\mathbb{C}}}\subset T_{N,e,{\mathbb{C}}}$ and the accumulation points of $T_{N,e,{\mathbb{C}}}$ belong to $T_{N-1,e,{\mathbb{C}}}$. Actually, they stated it only for $e=(1)\in{\mathbb{R}}_{>0}$. However, their proof also works for every fixed $e\in{\mathbb{R}}^{r}_{>0}$. Now we look at a positive characteristic pair $(A,{\mathfrak{a}}^{e})$ and let $c={\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})$. If $c\in T_{N-1,e,{\mathbb{C}}}$, then $c\in T_{N,e,{\mathbb{C}}}$, which completes the proof. Assume $c\not\in T_{N-1,e,{\mathbb{C}}}$, then by the result of [3], $c$ is not an accumulation point of $T_{N,e,{\mathbb{C}}}$. Therefore, we obtain $\epsilon>0$ such that:
$(c,c+\epsilon)\cap T_{N,{\mathbb{C}}}=\emptyset.$
We claim that there is a prime divisor $E$ over $(A,0)$ such that
(12) $z(E;A,{\mathfrak{a}}^{e})=\frac{k_{E}+1}{v_{E}({\mathfrak{a}}^{e})}<c+\frac{\epsilon}{2},\ \ \mbox{and}$
(13) $\frac{2(N-1)}{v_{E}({\mathfrak{a}}^{e})}<\frac{\epsilon}{2}.$
Indeed, by the definition of ${\operatorname{lct}}$, it is clear that there exists a prime divisor $E^{\prime}$ over $A$ with the center “containing 00” and satisfies the inequality (12):
$z(E^{\prime};A,{\mathfrak{a}}^{e})=\frac{k_{E^{\prime}}+1}{v_{E^{\prime}}({\mathfrak{a}}^{e})}<c+\frac{\epsilon}{2}.$
We need to find a prime divisor “with the center at 00” satisfying the inequality (12). If the center of $E^{\prime}$ is not $\{0\}$, then by suitable blow-ups, we obtain a normal variety $\overline{A}$ on which $E^{\prime}$ appears, and there is an exceptional divisor $E^{\prime}_{1}$ on $\overline{A}$ with the center at 00 such that
$E^{\prime}\cap E^{\prime}_{1}\neq\emptyset.$
We may assume that $\overline{A}$ is smooth around $E^{\prime}\cap E^{\prime}_{1}$ outside a closed subset of codimension $4$, and $E^{\prime}$ and $E^{\prime}_{1}$ intersect normally there. Because even over a field of positive characteristic, there is a log resolution for dimension $\leq 3$ ([1, Proposition 4.2]).
If $z(E^{\prime}_{1};A,{\mathfrak{a}}^{e})\geq c+{\epsilon}/{2}$, then take the blow-up at the intersection $E^{\prime}\cap E^{\prime}_{1}$ to obtain a new prime divisor $E^{\prime}_{2}$ dominating $E^{\prime}\cap E^{\prime}_{1}$ such that
$z(E^{\prime};A,{\mathfrak{a}}^{e})<z(E^{\prime}_{2};A,{\mathfrak{a}}^{e})<z(E^{\prime}_{1};A,{\mathfrak{a}}^{e}).$
Next, blow-up at $E^{\prime}\cap E^{\prime}_{2}$ to obtain $E^{\prime}_{3}$. Continue this process, then we have a sequence of prime divisors $\{E^{\prime}_{i}\}$ over $(A,0)$ such that
$\lim_{i\to\infty}z(E^{\prime}_{i};A,{\mathfrak{a}}^{e})=z(E^{\prime};A,{\mathfrak{a}}^{e}).$
This gives a prime divisor $E^{\prime}_{n}$ over $(A,0)$ satisfying (12). The inequality (13) in the claim holds when one takes sufficiently many times of blow-ups.
For $E$ and ${\mathfrak{a}}^{e}$, take a prime divisor $F_{\mathbb{C}}$ and a lifting ${\widetilde{\mathfrak{a}}}^{e}$ on ${\mathbb{A}}_{\mathbb{C}}^{N}$ as in Theorem 1.1. Then we obtain the inequalities by using Lemma 3.7:
$c+\epsilon>\frac{k_{E}+1+2(N-1)}{v_{E}({\mathfrak{a}}^{e})}=\frac{k_{F_{\mathbb{C}}}+1}{v_{F_{\mathbb{C}}}({\widetilde{\mathfrak{a}}}^{e})}\geq{\operatorname{lct}}({\tilde{0}};{\mathbb{A}}_{\mathbb{C}}^{N},{\widetilde{\mathfrak{a}}}^{e})\geq{\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})=c.$
By the definition of $\epsilon$, the last inequality must be the equality:
${\operatorname{lct}}({\tilde{0}};{\mathbb{A}}_{\mathbb{C}}^{N},{\widetilde{\mathfrak{a}}}^{e})={\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})=c,$
which yields ${\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})\in T_{N,e,{\mathbb{C}}}$.
Corollary 5.9 (Corollary 1.2,(ii),(iii)).
Assume $0\in A$ is a non-singular closed point of a variety of dimension $N$ over an algebraically closed field. Then the following hold:
  1. (ii)
    (Rationality of ${\operatorname{lct}}$’s and their accumulation points) For a rational exponent $e$ and the multi-ideal ${\mathfrak{a}}^{e}$ on $A$, the log canonical threshold ${\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})$ is a rational number and an accumulation point of the log canonical thresholds is also a rational number.
  2. (iii)
    (ACC of ${\operatorname{lct}}$) Let $D\subset{\mathbb{R}}_{>0}$ be a DCC set. Then, ACC holds for the set
    $T_{N,k}(D):=\{{\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})\mid e=(e_{1},\ldots,e_{r}),\forall e_{i}\in D\}.$
Proof.
By Corollary 5.8, for rational $e$, every value $c={\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})$ belongs to $T_{N,e,{\mathbb{C}}}\subset{\mathbb{Q}}$, and $T_{N,e,{\mathbb{C}}}$ is known to be closed ([3]), therefore (ii) follows immediately. As the set $T_{N,{\mathbb{C}}}(D)$ satisfies ACC by [2] for a special case and by [5] for the general case, its subset $T_{N,k}(D)$ satisfies ACC as well. ∎