### v2

# Liftings of ideals in positive characteristic to those in characteristic zero

Abstract.
We study a pair consisting of a smooth variety of arbitrary dimension over a field of positive characteristic and a multi-ideal with a real exponent. We prove that the set of log discrepancies for a fixed exponent is discrete. Additionally, we show that the set of log canonical thresholds (lcts) of multi-ideals on a smooth variety in positive characteristic is contained within the set of lcts of multi-ideals on a smooth variety over $\mathbb{C}$. As a result, we find that the accumulation points of log canonical thresholds are rational if all the exponents are rational. We also obtain ACC for the set of lcts of multi-ideals on a smooth varieties in positive characteristic.
Mathematical Subject Classification 2020: 14B05,14E18, 14J17
Key words: singularities in positive characteristic, jet schemes, minimal log discrepancy, log canonical threshold
The author is partially supported by Grant-In-Aid (c) 22K03428 of JSPS and by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.
1. Introduction
There are many tools available for studying singularities in characteristic zero. These include resolutions of singularities, Bertini’s theorem (which addresses generic smoothness), and vanishing theorems of Kodaira type. However, these tools do not necessarily hold in positive characteristic, which creates difficulties when studying singularities in that setting. In order to overcome these difficulties, we construct a “bridge” between positive characteristic and characteristic zero, allowing us to transport results from characteristic zero to positive characteristic. In general, such a strategy does not work. For example, the resolution of singularities in positive characteristic does not follow from that in characteristic zero. However, for our specific problems, this approach is effective.
Our focus is on the invariants of singularities, specifically log discrepancies, minimal log discrepancies, and log canonical thresholds. Our bridge illustrates the relationships between the set of invariants in positive characteristic and the set of those in characteristic zero. As a result, the properties of these sets in characteristic zero are inherited by their counterparts in positive characteristic.
The main theorem of this paper serves as a “bridge” between characteristic 0 and positive characteristic.
Theorem 1.1.
Let $N$ be an integer $\geq 2$, and $k$ an algebraically closed field of characteristic $p>0$. Let $A={\mathbb{A}}_{k}^{N}$ be the $N$-dimensional affine space over $k$ and 0 the origin of $A$. Let $E$ be a prime divisor over $(A,0)$, and ${\mathfrak{a}},{\mathfrak{a}}_{1},\ldots,{\mathfrak{a}}_{r}\subset k[x_{1},\ldots,x_{N}]$ non-zero ideals. Then, there exists a prime divisor $F_{\mathbb{C}}$ over the affine space $A_{\mathbb{C}}:={\mathbb{A}}_{\mathbb{C}}^{N}$ over $\mathbb{C}$} with the center at the origin $\tilde{0}$} satisfying the following:
  1. (i)
    $2(N-1)+k_{E}=k_{F_{\mathbb{C}}}$;
  2. (ii)
    there exists a lifting ${\widetilde{\mathfrak{a}}}\subset{\mathbb{C}}[x_{1},\ldots,x_{N}]$ of $\mathfrak{a}$} satisfying
[…]
  3. (iii)
    there exist liftings ${\widetilde{\mathfrak{a}}}_{i}\subset{\mathbb{C}}[x_{1},\ldots,x_{N}]$ of ${\mathfrak{a}}_{i}$ $(i=1,\ldots,r)$ such that
    $v_{E}({\mathfrak{a}}_{i})=v_{F_{\mathbb{C}}}({\widetilde{\mathfrak{a}}}_{i})$
    and for every multi-ideal ${\mathfrak{a}}^{e}:={\mathfrak{a}}_{1}^{e_{1}}\cdots{\mathfrak{a}}_{r}^{e_{r}}$ $(e_{i}\in{\mathbb{R}}_{>0})$ on $A$, the lifting ${\widetilde{\mathfrak{a}}}^{e}={\widetilde{\mathfrak{a}}}_{1}^{e_{1}}\cdots{\widetilde{\mathfrak{a}}}_{r}^{e_{r}}$ of ${\mathfrak{a}}^{e}$ on $A_{\mathbb{C}}$ satisfies
    $2(N-1)+a(E;A,{\mathfrak{a}}^{e})=a(F_{\mathbb{C}};A_{\mathbb{C}},{\widetilde{\mathfrak{a}}}^{e}).$
The constant term 2(N1)2(N-1) may look mysterious to the reader. This arises from two more blow-ups of the variety on which $E$ appears. The reason why we do this is briefly explained before Lemma 5.1.
We obtain the following corollaries as applications of the theorem.
Corollary 1.2.
Let $N$ be a positive integer and $0\in A$ a smooth closed point of an $N$-dimensional algebraic variety $A$ defined over an algebraically closed field $k$ of characteristic $p>0$. Let ${\mathfrak{a}}^{e}$ be a non-zero multi-ideal on $A$. Then, the following hold:
  1. (i)
    (discreteness of log discrepancies) For a fixed exponent $e\in{\mathbb{R}}_{>0}^{r}$ and a real number $a>0$, the set of log discrepancies
[…]
    is finite. In particular, for fixed $N$ and $e$, the set
    $\{{\mathrm{mld}}(0,A,{\mathfrak{a}}^{e})\mid{\mathfrak{a}}^{e}\ \mbox{non-zero\ multi-ideal}\ \}$
    is finite. Also for a fixed ${\mathfrak{a}}^{e}$, the discreteness of log discrepancies implies the existence of a prime divisor computing ${\mathrm{mld}}(0,A,{\mathfrak{a}}^{e})$.
  2. (ii)
    (Inclusion of the sets of $\operatorname{lct}$}’s ) Let $e\in{\mathbb{R}}_{>0}^{r}$ and
    $T_{N,e,k}:=\{{\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})\mid{\mathfrak{a}}^{e}:\mbox{non-zero\ multi-ideal\ on \ A\ }\},\mbox{and}$
    $T_{N,e,{\mathbb{C}}}:=\{{\operatorname{lct}}({\tilde{0}};A_{\mathbb{C}},{\mathfrak{b}}^{e})\mid{\mathfrak{b}}^{e}:\mbox{non-zero\ multi-ideal\ on}\ A_{\mathbb{C}}\},$
    where ${\tilde{0}}\in A_{\mathbb{C}}$ is a smooth closed point of an $N$-dimensional variety $A_{\mathbb{C}}$ defined over $\mathbb{C}$}. Then
    $T_{N,e,k}\subset T_{N,e,{\mathbb{C}}}.$
    In particular, when the exponents $e_{i}$’s in $e=(e_{1},\ldots,e_{r})$ are all rational, then ${\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})$ is a rational number, and an accumulation point of these log canonical thresholds is also a rational number.
  3. (iii)
    (ACC of $\operatorname{lct}$}) Let $D\subset{\mathbb{R}}_{>0}$ be a DCC set. Then, ACC holds for the set
    $T_{N,k}(D):=\{{\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})\mid e=(e_{1},\ldots,e_{r}),\forall e_{i}\in D\}.$
For $N=1$, the statements of Corollary 1.2 are trivial, because a prime divisor over $(A,0)$ is just the divisor $\{0\}\subset A$, and $v_{\{0\}}({\mathfrak{a}}^{e}):=\sum_{i}e_{i}{\mathrm{ord}}({\mathfrak{a}}_{i})$ and $k_{\{0\}}+1=1$ hold for any base field. So, we may assume $N\geq 2$ and can apply Theorem 1.1.
Note that the statements in Corollary 1.2 are known to hold in characteristic 0. More specifically, statement (i) is proven by Kawakita in [12], the statement about the rationality in (ii) is established by T. De Fernex and M. Mustaţǎ in [4], and statement (iii) for a special case is demonstrated by T. De Fernex, L. Ein and M. Mustaţǎ in [3] and for the general case by C. Hacon, J. McKernan and C. Xu in [7].
These are excellent results, and the proofs are elaborated based on resolutions of singularities, Bertini’s theorem, and cohomology vanishing of the Kodaira type, which are not available in the positive characteristic case. Therefore, the same proofs cannot be applied in positive characteristic settings.
Our proof in positive characteristic involves transporting these results from characteristic 0 into the realm of positive characteristic, utilizing the bridge established in Theorem 1.1.
This paper is organized as follows: In the second section, we introduce basic notions on liftings of materials in positive characteristic to those in characteristic 0.
In the third section we introduce log discrepancies, minimal log discrepancies, and log canonical thresholds which are birational invariants to measure the nature of singularities of pairs. The description of these invariants by the arc spaces will work to compare the invariants between in the realm of positive characteristic and in that of characteristic 0.
In the forth section, from a given prime divisor over the affine space in positive characteristic, we construct a prime divisor over the affine space defined over $\mathbb{C}$} such that the divisors have some common properties.
In the fifth section, we give the proof for the main theorem and the corollaries.
As there are still many open problems related to our topics, we conclude the paper by some conjectures related to the topics in the sixth section.
Acknowledgement. The author would like to thank Lawrence Ein, Kazuhiko Kurano, Mircea Mustaţǎ, Yuri Prokhorov, Kohsuke Shibata and Shoji Yokura for useful discussions and comments. The author expresses her heartfelt thanks to János Kollár for providing critical examples and comments for the preliminary version of this paper.
Conventions
  1. (1)
    When we say just “a point” of a variety $A$, it is not necessarily a closed point. By abuse of notation, for a point $P\in A$ we write $\dim P:=\dim\overline{\{P\}}$. The codimension of $P$ is defined accordingly ; ${\mathrm{codim}}P:=\dim A-\dim P$.
  2. (2)
    A blow-up with the center by the defining ideal of the closure $\overline{\{P\}}$ is called “the blow-up at PP” and we also call $P$ the center of the blow-up.
  3. (3)
    Let $L_{1}$ be a ring, $L_{2}$ an $L_{1}$-algebra and $A$ a scheme of finite type over $L_{1}$. We denote the scheme $A\times_{{\mathrm{Spec\ }}L_{1}}{\mathrm{Spec\ }}L_{2}$ by $A\otimes_{L_{1}}L_{2}$ to avoid the bulky expression.
  4. (4)
    For an integral domain $R$, we denote the field of fractions of $R$ by $Q(R)$.
  5. (5)
    A prime ideal ${\mathfrak{p}}\subset R$ of a Noetherian integral domain is called a regular prime ideal if the local ring associated to $\mathfrak{p}$} is a regular local ring. Here, note that if $R$ is a graded ring and $\mathfrak{p}$} is a homogeneous prime ideal the associated local ring is $R_{({\mathfrak{p}})}$ by the symbol in [8, Definition, p.116].
  6. (6)
    Let ${\mathfrak{p}}\subset R$ be a regular prime ideal. A set of elements $\{f_{1},\ldots,f_{c}\}\subset{\mathfrak{p}}$ is called a regular system of parameters (RSP for short) of $\mathfrak{p}$} if it gives an RSP in the local ring associated to $\mathfrak{p}$}. If $R$ is graded and $\mathfrak{p}$} is homogeneous, the set gives an RSP in $R_{({\mathfrak{p}})}$.
2. Basics on liftings
In this paper we construct a “lifting” which is an item on a variety over $\mathbb{C}$} corresponding to that in positive characteristic. Basic notions about liftings are discussed in [10], and here we summarize some of them, which we will use in this paper.
Definition 2.1.
Let $R$ be an integral domain which is finitely generated $k$-algebra, where $k$ is a field of characteristic $p\geq 0$. A subring $S\subset R$ is called a skeleton of $R$, if $S$ is finitely generated ${\mathbb{Z}}/(p)$}-subalgebra of $R$ such that $k[S]=R$.
Definition 2.2.
  1. (1)
    (compatible skeletons) Let $k$ be a field of characteristic $p>0$ and $K$ a field of characteristic 0. Let $R$ and $\widetilde{R}$} be finitely generated $k$-algebra and $K$-algebra, respectively. Let $S\subset R$ and ${\widetilde{S}}\subset{\widetilde{R}}$ be skeletons. If the skeletons satisfies
    $\widetilde{S}$} is $\mathbb{Z}$}-flat at $p$, and
    $S={\widetilde{S}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)},$
    we call them compatible skeletons. We sometimes say that a skeleton $\widetilde{S}$} is a lifting of the skeleton $S$. We also call the canonical surjection $\Phi_{p}:{\widetilde{S}}\to S$ compatible skeletons. In this case we write
    ${\widetilde{R}}({\operatorname{mod}\ }p)=R.$
    If schemes $\widetilde{A}$} and $A$ are defined by $\widetilde{R}$} and $R$ (eg. as $\mathrm{Spec\ }$} or $\mathrm{Proj\ }$}) which have compatible skeletons, we also write
    ${\widetilde{A}}({\operatorname{mod}\ }p)=A.$
    Here, we note that if $\widetilde{R}$} is an integral domain (therefor, so is $\widetilde{S}$}), then $\mathbb{Z}$}-flatness is automatic. In this case $S={\widetilde{S}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}$ is the only condition for $\widetilde{S}$} and $S$ to be compatible skeletons.
[…]
When we think of liftings of elements, we always use a common pair of compatible skeletons in a continuous discussion.
Example 2.3.
The following are examples of liftings. The first one is a trivial example and the second and third are more practical and will play basic roles in this paper. Assume that $k$ is a field of characteristic $p>0$.
(1) We have ${\mathbb{C}}({\operatorname{mod}\ }p)=k$ with compatible skeletons $S={{\mathbb{Z}}/(p)}$ and $\widetilde{S}$$\mathbb{Z}$}.
(2) For any finite elements $a_{1},\ldots,a_{r}\in k$ there exist liftings ${\widetilde{a}}_{1},\ldots,{\widetilde{a}}_{r}\in{\mathbb{C}}$.
Indeed compatible skeletons $\Sigma\subset k$ and $\widetilde{\Sigma}\subset{\mathbb{C}}$ giving these liftings are constructed in [10, Proposition 2.3]. However, since this statement is fundamental to the paper, we restate the proof here.
Proof.
Considering the subring ${{\mathbb{Z}}/(p)}[a_{1},\ldots,a_{n}]\subset k$, we obtain canonical surjections:
$R:={\mathbb{Z}}[Y_{1},\ldots,Y_{n}]\stackrel{{\scriptstyle\psi}}{{\longrightarrow}}S:={{\mathbb{Z}}/(p)}[Y_{1},\ldots,Y_{n}]\stackrel{{\scriptstyle\varphi}}{{\longrightarrow}}{{\mathbb{Z}}/(p)}[a_{1},\ldots,a_{n}]$
with $Y_{i}\mapsto a_{i}$. Let $P:=Ker\varphi\subset S$ and $Q:=Ker\varphi\circ\psi\subset R$, then these are prime ideals in regular rings. Therefore $R_{Q}$ and $S_{P}$ are also regular local rings. Hence we obtain
  1. (1)
    $f_{1},\ldots,f_{c}\ (\in P)$ which form a regular system of parameters of $S_{P}$,
  2. (2)
    ${\tilde{f}}_{1},\ldots,{\tilde{f}}_{c},p\ (\in Q)$ which form a regular system of parameters of $R_{Q}$ and
  3. (3)
    $\psi({\tilde{f}}_{i})=f_{i}$ for $i=1,\ldots,c$.
Then, $R_{Q}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})R_{Q}$ is also a regular local ring, in particular it is an integral domain. By considering the homomorphism ${\mathbb{Z}}\to R_{Q}$ and the regular sequence $\{f_{1},\ldots,f_{c}\}\subset S_{P}=R_{Q}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}$, we obtain that $R_{Q}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})R_{Q}$ is flat over $\mathbb{Z}$} by [16, Corollary, p.177]. In particular, the homomorphism ${\mathbb{Z}}\to R_{Q}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})R_{Q}$ is injective, which implies the ring $R_{Q}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})R_{Q}$ is of characteristic 0. Note that the algebras ${\widetilde{\Sigma}}^{\prime}:=R_{Q}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})R_{Q}$ and $\Sigma^{\prime}:=S_{P}/(f_{1},\ldots,f_{c})S_{P}$ satisfies
${\widetilde{\Sigma}}^{\prime}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}=\Sigma^{\prime}={{\mathbb{Z}}/(p)}(a_{1},\ldots,a_{n}).$
However, these $\widetilde{\Sigma}$$^{\prime}$ and $\Sigma^{\prime}$ are not appropriate for the skeletons, because these are not finitely generated as $\mathbb{Z}$}-algebra and ${\mathbb{Z}}/(p)$}-algebra, respectively. So we replace $\widetilde{\Sigma}$$^{\prime}$ and $\Sigma^{\prime}$ by smaller ones.
As $PS_{P}=(f_{1},\ldots,f_{c})S_{P}$, there exists $h\in S\setminus P$ such that $PS_{h}=(f_{1},\ldots,f_{c})S_{h}$. Take ${\tilde{h}}\in R\setminus Q$ such that $\psi({\tilde{h}})=h$ and let
$\widetilde{\Sigma}:=R_{\tilde{h}}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})R_{\tilde{h}}$.
This, as well, is an integral domain and of characteristic 0. Now noting that $S/P={{\mathbb{Z}}/(p)}[a_{1},\ldots,a_{c}]$, we obtain
$\Sigma:=\widetilde{\Sigma}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}=\frac{R_{\tilde{h}}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})R_{\tilde{h}}}{p({R_{\tilde{h}}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})R_{\tilde{h}}})}=S_{h}/(f_{1},\ldots,f_{c})S_{h}$
$(S/P)_{h}\subset{{\mathbb{Z}}/(p)}(a_{1},\ldots,a_{c})\subset k.$
[…]
$\phi:K_{0}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\phi(K_{0})\subset{\mathbb{C}}.$
Then we obtain $\{\widetilde{a}_{1},\ldots,\widetilde{a}_{n}\}\subset{\mathbb{C}}$ and $\{\widetilde{a}_{1},\ldots,\widetilde{a}_{n}\}({\operatorname{mod}\ }p)=\{a_{1},\ldots,a_{n}\}$. As is seen, $\widetilde{\Sigma}$$\subset{\mathbb{C}}$ and $\Sigma\otimes{{\mathbb{Z}}/(p)}$ are compatible skeletons. ∎
(3) Let $\Sigma\subset k$ and $\widetilde{\Sigma}$$\subset{\mathbb{C}}$ be compatible skeletons. Then $S:=\Sigma[x_{1},\ldots,x_{N}]\subset k[x_{1},\ldots,x_{N}]$ and ${\widetilde{S}}:={\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]\subset{\mathbb{C}}[x_{1},\ldots,x_{N}]$ are also compatible skeletons for ${\mathbb{C}}[x_{1},\ldots,x_{N}]({\operatorname{mod}\ }p)=k[x_{1},\ldots,x_{N}]$ in the canonical way. Explicitly, a polynomial $f\in S$ has a lifting ${\tilde{f}}\in{\widetilde{S}}$ in the following form:
$f=\sum_{i\in{\mathbb{Z}}^{N}_{\geq 0}}a_{i}{\mathbb{x}}^{i}\ \ \ ,\ \ \mbox{then}\ \ \ {\tilde{f}}=\sum_{i\in{\mathbb{Z}}^{N}_{\geq 0}}{\widetilde{a}}_{i}\mathbb{x}^{i},$
where ${\widetilde{a}}_{i}$ is a lifting of $a_{i}$.
Definition 2.4.
Let $R$ and $\widetilde{R}$} be finitely generated algebras over fields $k$ and $K$ of characteristic $p>0$ and of 0, respectively.
Assume ${\widetilde{R}}({\operatorname{mod}\ }p)=R$ holds by compatible skeletons $\Phi_{p}:{\widetilde{S}}\to S$, where $S\subset R$ and ${\widetilde{S}}\subset{\widetilde{R}}$.
For ideals ${\mathfrak{a}}_{R}\subset R$ and ${\widetilde{\mathfrak{a}}}_{\widetilde{R}}\subset{\widetilde{R}}$ we write ${\widetilde{\mathfrak{a}}}_{\widetilde{R}}({\operatorname{mod}\ }p)={\mathfrak{a}}_{R}$ and say ${\widetilde{\mathfrak{a}}}_{\widetilde{R}}$ is a lifting of ${\mathfrak{a}}_{R}$ if we can take a generator system of ${\widetilde{\mathfrak{a}}}_{\widetilde{R}}$ from $\widetilde{S}$} whose image by $\Phi_{p}$ generates ${\mathfrak{a}}_{R}\subset R$. It is equivalent to the following:
  1. (1)
    There are ideals ${\mathfrak{a}}\subset S$ and ${\widetilde{\mathfrak{a}}}\subset{\widetilde{S}}$ such that ${\widetilde{\mathfrak{a}}}_{\widetilde{R}}={\widetilde{\mathfrak{a}}}{\widetilde{R}}$, ${\mathfrak{a}}_{R}={\mathfrak{a}}R$, and
  2. (2)
    $\Phi_{p}({\widetilde{\mathfrak{a}}})={\mathfrak{a}}$.
In this case, we call $\mathfrak{a}$} a skeleton of ${\mathfrak{a}}_{R}$, and $\widetilde{\mathfrak{a}}$} a skeleton of ${\widetilde{\mathfrak{a}}}_{\widetilde{R}}$.
Remark 2.5.
According to the definition above, the unit ideal $\widetilde{R}$} itself is a lifting of any ideal ${\mathfrak{a}}_{R}\subset R$ generated by elements of the skeleton $S\subset R$. Because the pull-back ${\widetilde{\mathfrak{a}}}:=\Phi_{p}^{-1}({\mathfrak{a}})$ satisfies the conditions above and contains $p\in{\mathbb{Z}}\subset{\widetilde{R}}$, a unit element of $\widetilde{R}$}.
This gives a trivial example that a lifting of a prime ideal in $R$ is not necessarily prime. For a prime ideal, we define a “prime lifting ” and “weak prime lifting” which are again a prime ideals in the skeleton rings.
Definition 2.6.
Let $S\subset R$ and ${\widetilde{S}}\subset{\widetilde{R}}$ be compatible skeletons for finitely generated algebras $R$ and $\widetilde{R}$} over fields $k$ and $K$ of characteristic $p>0$ and of 0, respectively. Let ${\mathfrak{p}}\subset S$ and ${\widetilde{{\mathfrak{p}}}}\subset{\widetilde{S}}$ be prime ideals.
  1. (i)
    we call $\widetilde{{\mathfrak{p}}}$} a prime lifting of $\mathfrak{p}$} in the skeleton level if $\Phi_{p}({\widetilde{{\mathfrak{p}}}})={\mathfrak{p}}$ and $0pt{\mathfrak{p}}=0pt{\widetilde{{\mathfrak{p}}}}$.
  2. (ii)
[…]
Remark 2.7.
The following are equivalent:
  1. (1)
    A prime ideal ${\widetilde{{\mathfrak{p}}}}\subset{\widetilde{S}}$ is a weak prime lifting of ${\mathfrak{p}}\subset S$;
  2. (2)
    There exist an element $g\in S$ and its lifting ${\tilde{g}}\in{\widetilde{S}}$ such that ${\widetilde{{\mathfrak{p}}}}_{{\tilde{g}}}$ is a prime lifting of ${\mathfrak{p}}_{g}$;
  3. (3)
    The inclusions ${\widetilde{{\mathfrak{p}}}}\subset\Phi_{p}^{-1}({\mathfrak{p}})\subset{\widetilde{S}}$ holds with the equality $0pt{\mathfrak{p}}=0pt{\widetilde{{\mathfrak{p}}}}$. And the canonical homomorphism ${\widetilde{S}}_{\Phi_{p}^{-1}({\mathfrak{p}})}\to S_{\mathfrak{p}}$ induces the surjection
    ${\widetilde{{\mathfrak{p}}}}{\widetilde{S}}_{\Phi_{p}^{-1}({\mathfrak{p}})}\to{\mathfrak{p}}S_{\mathfrak{p}}.$
Here we omit the term “in the skeleton level” when it is clear that we are working in the skeletons.
A prime lifting in the skeleton level does not exist in general. But the following proposition shows that for rings of special type a regular homogeneous prime ideal has at least one weak prime lifting in the skeleton level and gives a proper ideal in $\widetilde{R}$} by the extension. Here, we note that even when the extension is a proper ideal of $\widetilde{R}$}, it is not necessarily prime in $\widetilde{R}$}. In this paper, we work on such algebras $R$ and $\widetilde{R}$} as follows:
Proposition 2.8.
Let $k$ and $K$ be fields of characteristic $p>0$ and 0, respectively. Let $\Sigma\subset k$ and ${\widetilde{\Sigma}}\subset K$ be compatible skeletons. By fixing degree $\deg x_{i}:=w_{i}>0$, we take four polynomial rings and their graded subalgebras:
$\begin{array}[]{cccccc}\Sigma[x_{1},\ldots,x_{m}]&\subset&k[x_{1},\ldots,x_{m}],&\ \ \ \ {\widetilde{\Sigma}}[x_{1},\ldots,x_{m}]&\subset&K[x_{1},\ldots,x_{m}]\\ \bigcup&&\bigcup&\ \ \ \ \bigcup&&\bigcup\\ S&\subset&R,&\ \ \ \ {\widetilde{S}}&\subset&{\widetilde{R}}\\ \end{array}$
Here, we assume that $R$ is a $k$-subalgebra of $k[x_{1},\ldots,x_{m}]$, $\widetilde{R}$} is a $K$-subalgebra of $K[x_{1},\ldots,x_{m}]$, and $S\subset R$ and ${\widetilde{S}}\subset{\widetilde{R}}$ are compatible skeletons and the homomorphism
$\Phi_{p}:{\widetilde{S}}\to S$
is the restriction of the canonical map ${\widetilde{\Sigma}}[x_{1},\ldots,x_{m}]\to\Sigma[x_{1},\ldots,x_{m}]$ induced from $\Phi_{p}:{\widetilde{\Sigma}}\to\Sigma$.
Let ${\mathfrak{p}}_{R}\subset R$ be a regular homogeneous prime ideal generated by ${\mathfrak{p}}:={\mathfrak{p}}_{R}\cap S$. Then, there is a weak homogeneous prime lifting
${\widetilde{{\mathfrak{p}}}}\subset{\widetilde{S}}_{+}:=\oplus_{i>0}{\widetilde{S}}_{i}$
of $\mathfrak{p}$} in the skeleton level.
Proof.
First, let $S_{+}:=\oplus_{i>0}S_{i}$ and $R_{+}:=\oplus_{i>0}R_{i}$, then ${\mathfrak{p}}\subset S_{+}$, because ${\mathfrak{p}}_{R}\subset R_{+}$. Take an RSP $\{f_{1},\ldots,f_{c}\}\subset S_{+}$ of $\mathfrak{p}$}. By the condition of the proposition, $\Phi_{p}:{\widetilde{S}}\to S$ is a surjective graded homomorphism, therefore there exists a lifting $\{{\tilde{f}}_{1},\ldots,{\tilde{f}}_{c}\}\subset{\widetilde{S}}_{+}$ of $\{f_{1},\ldots,f_{c}\}$, where ${\tilde{f}}_{i}$ is homogeneous with $\deg{\tilde{f}}_{i}=\deg f_{i}$ for every $i$. The subset $\{{\tilde{f}}_{1},\ldots,{\tilde{f}}_{c}\}\subset\Phi_{p}^{-1}({\mathfrak{p}})\subset{\widetilde{S}}$ forms a regular sequence in ${\widetilde{S}}_{\Phi_{p}^{-1}({\mathfrak{p}})}$, and also the ring
$W:={\widetilde{S}}_{\Phi_{p}^{-1}({\mathfrak{p}})}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})$
is $\mathbb{Z}$}-flat at $p$ by [16, Corollary p.177]. Here $p\in W$ is a prime element by the equality
$W\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}=S_{\mathfrak{p}}/(f_{1},\ldots,f_{c})=S_{\mathfrak{p}}/{\mathfrak{p}}S_{\mathfrak{p}},$
where the last ring is an integral domain. By the primeness of $p\in W$ and the flatness of $W$ over $\mathbb{Z}$} we obtain that $W$ is an integral domain. This implies that $({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})\subset{\widetilde{S}}_{\Phi_{p}^{-1}({\mathfrak{p}})}$ is a prime ideal. This yields a primary component of $({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})\subset{\widetilde{S}}$ contained in $\Phi_{p}^{-1}({\mathfrak{p}})$} is unique and coincides with its radical, which we denote by $\widetilde{{\mathfrak{p}}}$}. Then we obtain the canonical homomorphism
${\widetilde{{\mathfrak{p}}}}{\widetilde{S}}_{\Phi_{p}^{-1}({\mathfrak{p}})}=({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c}){\widetilde{S}}_{\Phi_{p}^{-1}({\mathfrak{p}})}\to(f_{1},\ldots,f_{c})S_{\mathfrak{p}}={\mathfrak{p}}S_{\mathfrak{p}}$
is surjective. As $({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})$ is regular sequence, $0pt{\widetilde{{\mathfrak{p}}}}=c=0pt{\mathfrak{p}}$, we otain that $\widetilde{{\mathfrak{p}}}$} is a weak prime lifting of $\mathfrak{p}$} in the skeleton level.
For the proof of ${\widetilde{{\mathfrak{p}}}}\subset{\widetilde{S}}_{+}$, assume contrary. As the associated prime $\widetilde{{\mathfrak{p}}}$} of the homogeneous ideal $({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})$ is also homogeneous, our assumption ${\widetilde{{\mathfrak{p}}}}\not\subset{\widetilde{S}}_{+}$ implies that $\widetilde{{\mathfrak{p}}}$} contains a degree 0 element $s\in{\widetilde{\Sigma}}$. As $\Phi_{p}({\widetilde{{\mathfrak{p}}}})\subset{\mathfrak{p}}\subset S_{+}$, the element $s$ is divided by $p\in{\mathbb{Z}}$. By primeness of $\widetilde{{\mathfrak{p}}}$}, we may assume that
$p\in{\widetilde{{\mathfrak{p}}}}$.
Then the ideal $\widetilde{{\mathfrak{p}}}$} would contain a regular sequence $\{{\tilde{f}}_{1},\ldots,{\tilde{f}}_{c},p\}$ of length $c+1$, which is a contradiction to the fact that $\widetilde{{\mathfrak{p}}}$} is an associated prime of the ideal generated by a regular sequence of length $c$. ∎
Definition 2.9.
In the proof of Proposition 2.8, one can see that for a given homogeneous RSP $\{f_{1},\ldots,f_{c}\}\subset{\mathfrak{p}}$ and its lifting $\{{\tilde{f}}_{1},\ldots,{\tilde{f}}_{c}\}$, one obtain the unique weak prime lifting $\widetilde{{\mathfrak{p}}}$} containing $\{{\tilde{f}}_{1},\ldots,{\tilde{f}}_{c}\}$. We call such $\widetilde{{\mathfrak{p}}}$} the weak prime lifting with respect to $\{f_{1},\ldots,f_{c}\}$ and $\{{\tilde{f}}_{1},\ldots,{\tilde{f}}_{c}\}$.
Lemma 2.10.
Let $R$, $S$, $\widetilde{S}$}, and ${\mathfrak{p}}\subset S$ be as in Proposition 2.8. If ${\widetilde{{\mathfrak{p}}}}\subset{\widetilde{S}}_{+}$ is a weak prime lifting of $\mathfrak{p}$}. Then, for any RSP $\{f_{1},\ldots,f_{c}\}\subset{\mathfrak{p}}$, there is a lifting $\{{\tilde{f}}_{1},\ldots,{\tilde{f}}_{c}\}\subset{\widetilde{S}}$ such that $\widetilde{{\mathfrak{p}}}$} is a weak prime lifting of $\mathfrak{p}$} with respect to them.
Proof.
As $\widetilde{{\mathfrak{p}}}$} is a weak prime lifting of $\mathfrak{p}$}, it follows $0pt{\widetilde{{\mathfrak{p}}}}=c$ and there is a surjection
${\widetilde{{\mathfrak{p}}}}{\widetilde{S}}_{\Phi_{p}^{-1}({\mathfrak{p}})}\to{\mathfrak{p}}S_{\mathfrak{p}}.$
By the equality ${\mathfrak{p}}S_{\mathfrak{p}}=(f_{1},\ldots,f_{c})S_{\mathfrak{p}}$ and the above surjection, we obtain a set of homogeneous elements $\{{\tilde{f}}_{1},\ldots,{\tilde{f}}_{c}\}\subset{\widetilde{S}}_{+}$ corresponding to $\{f_{1},\ldots,f_{c}\}$. Then, by [16, Corollary, p.177], $\{{\tilde{f}}_{1},\ldots,{\tilde{f}}_{c}\}$ is a regular sequence in ${\widetilde{S}}_{\Phi_{p}^{-1}({\mathfrak{p}})}$ and $W:={\widetilde{S}}_{\Phi_{p}^{-1}({\mathfrak{p}})}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})$ is $\mathbb{Z}$}-flat at $p$. As in the proof of Proposition 2.8, $({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})\subset{\widetilde{S}}_{\Phi_{p}^{-1}({\mathfrak{p}})}$ is a prime ideal with height $c$ contained in $\widetilde{{\mathfrak{p}}}$} which yields $({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})={\widetilde{{\mathfrak{p}}}}$ in ${\widetilde{S}}_{\Phi_{p}^{-1}({\mathfrak{p}})}$. Therefore, it follows that $\widetilde{{\mathfrak{p}}}$} is a weak prime lifting of $\mathfrak{p}$} with respect to $\{f_{1},\ldots,f_{c}\}$ and $\{{\tilde{f}}_{1},\ldots,{\tilde{f}}_{c}\}$. ∎
Proposition 2.11.
Let $R$, $S$ and $\widetilde{S}$} be as in the Proposition 2.8. Let two regular prime ideals
$I_{R}\subset{\mathfrak{p}}_{R}\subset R$
generated by $I:=I_{R}\cap S$ and by ${\mathfrak{p}}:={\mathfrak{p}}_{R}\cap S$, respectively. Assume there is a weak prime lifting ${\widetilde{I}}\subset{\widetilde{S}}_{+}$ of $I$ in the skeleton level. Then $\mathfrak{p}$} has a weak prime lifting ${\widetilde{{\mathfrak{p}}}}\subset{\widetilde{S}}_{+}$ such that
${\widetilde{I}}\subset{\widetilde{{\mathfrak{p}}}}$.
More precisely, for any homogeneous RSP $\{f_{1},\ldots,f_{c},f_{c+1},\ldots,f_{e}\}$ of $\mathfrak{p}$} such that $\{f_{1},\ldots,f_{c}\}$ is an RSP of $I$, and a lifting $\{{\tilde{f}}_{1},\ldots,{\tilde{f}}_{c}\}$ of $\{f_{1},\ldots,f_{c}\}$ which gives a weak prime lifting $\widetilde{I}$} of $I$, $\{f_{c+1},\ldots,f_{e}\}$, there is a prime lifting ${\widetilde{{\mathfrak{p}}}}\subset{\widetilde{S}}\cap{\widetilde{\mathfrak{m}}}$ of $\mathfrak{p}$} with respect to the extended systems $\{f_{1},\ldots,f_{c},f_{c+1},\ldots,f_{e}\}$ and $\{{\tilde{f}}_{1},\ldots,{\tilde{f}}_{c},\}$, which yields ${\widetilde{I}}\subset{\widetilde{{\mathfrak{p}}}}$. Here, we can choose any homogeneous lifting $\{{\tilde{f}}_{c+1},\ldots,{\tilde{f}}_{e}\}\subset{\widetilde{S}}_{+}$ of $f_{c+1},\ldots,f_{e}\}$.
Proof.
By Lemma 2.10, we can take an RSP $\{{\tilde{f}}_{1},\ldots,{\tilde{f}}_{c}\}\subset{\widetilde{S}}_{+}$ of $\widetilde{I}$}, which is a lifting of $\{f_{1},\ldots,f_{c}\}\subset I$. Take any lifting $\{{\tilde{f}}_{c+1},\ldots,{\tilde{f}}_{e}\}\subset{\widetilde{S}}_{+}$ of $\{f_{c+1},\ldots,f_{e}\}$. Then, the extended system $\{{\tilde{f}}_{1},\ldots,{\tilde{f}}_{c},\ldots,{\tilde{f}}_{e}\}$ is a lifting of $\{f_{1},\ldots,f_{c},\ldots f_{e}\}$ and they give a weak prime lifting $\widetilde{{\mathfrak{p}}}$} of $\mathfrak{p}$} with the required properties. ∎
Remark 2.12.
Let $R$ be a finitely generated graded $k$-algebra with $R_{0}=k$. Assume $R_{1}$ generates $R$ and let $X={\mathrm{Proj\ }}R$. Let $P\in X$ be a point and ${\mathfrak{p}}\subset R$ the homogeneous prime ideal defining $\overline{\{P\}}$. Then, $P\in X$ is a regular point if and only if the local ring $R_{\mathfrak{p}}$ is a regular local ring. In this case, a set $\{f_{1},\ldots,f_{c}\}\subset R$ of homogeneous elements is an RSP of $R_{\mathfrak{p}}$ if and only if $\{f_{1}/g^{e_{1}},\ldots,f_{c}/g^{e_{c}}\}$ is an RSP of $R_{({\mathfrak{p}})}$, where $e_{i}=\deg f_{i}$ $(i=1,\ldots,c)$ and $g\in R_{1}$ is not contained in $\mathfrak{p}$}.
In this paper, we work on algebras of the special type: i.e., a finitely generated $k$-subalgebra $R$ of a polynomial ring. On such an algebra $R$, properties on an ideal $\mathfrak{a}$} of $R$ are reduced to the properties of its skeleton ideal ${\mathfrak{a}}\cap S$ as follows:
Lemma 2.13.
Let $K$ be a field and $\Sigma$ a subring of $K$. Denote the field of fractions of $\Sigma$ by $Q$. Let $S\subset\Sigma[x_{1},\ldots,x_{m}]$ be a finitely generated $\Sigma$-subalgebra of the polynomial ring and ${\mathfrak{a}}\subset S$ an ideal which does not contain any nonzero element of $\Sigma$. Let
$Q[S]\subset Q[x_{1},\ldots,x_{m}]\ \ \mbox{ and}\ \ K[S]\subset K[x_{1},\ldots,x_{m}]$
be the $Q$- and $K$-subalgebras generated by $S$, respectively. Let
${\mathfrak{a}}_{Q}\subset Q[S]\ \ \mbox{ and}\ \ {\mathfrak{a}}_{K}\subset K[S]$
be the ideals generated by $\mathfrak{a}$}. Then, it follows:
  1. (i)
    $0pt{\mathfrak{a}}=0pt{\mathfrak{a}}_{Q}=0pt{\mathfrak{a}}_{K}.$
  2. (ii)
    If $\mathfrak{a}$} is a prime ideal, then ${\mathfrak{a}}_{Q}$ is also prime, though ${\mathfrak{a}}_{K}$ is not necessarily prime. Moreover for an element $f\in S$ which is not in $\Sigma$, the following are equivalent
    1. (a)
      $f\in{\mathfrak{a}}^{n}\setminus{\mathfrak{a}}^{n+1}$,
    2. (b)
      $f\in{\mathfrak{a}}_{Q}^{n}\setminus{\mathfrak{a}}_{Q}^{n+1}$,
    3. (c)
      $f\in{\mathfrak{a}}_{K}^{n}\setminus{\mathfrak{a}}_{K}^{n+1}$,
  3. (iii)
    In particular, for ${\operatorname{char}}K=0$, assume ${\mathfrak{a}}_{Q}$ is a locally principal prime ideal and a minimal prime ideal $\mathfrak{p}$} of ${\mathfrak{a}}_{K}$ is locally principal, then, the following is also equivalent to (a)–(c) above:
    1. (d)
      $f\in{\mathfrak{p}}^{n}\setminus{\mathfrak{p}}^{n+1}$ for every minimal prime $\mathfrak{p}$} of ${\mathfrak{a}}_{K}$.
Proof.
As $Q\hookrightarrow K$ is faithfully flat, it follows $Q[S]\hookrightarrow K[S]=K\otimes_{Q}Q[S]$ is also faithfully flat, in particular it is flat. Then, by [8, III, Corollary 9.6] (or algebraically [16, Theorem 15.1]), we have
$0pt{\mathfrak{a}}_{Q}=0pt({\mathfrak{a}}_{Q}K[S])=0pt{\mathfrak{a}}_{K}.$
This gives the second equality in (i).
Since for every $g\in Q[S]$ there exists an element $r\in\Sigma$ such that $rg\in S$, we have the following bijective correspondence:
$\left\{\begin{array}[]{l}\mbox{prime ideals\ of\ }S\\ \mbox{without\ nonzero\ elements\ in\ }\Sigma\\ \end{array}\right\}$ $\leftrightarrow$ $\left\{\begin{array}[]{l}\mbox{prime\ ideals\ of\ }\\ Q[S]\\ \end{array}\right\}$
by associating ${\mathfrak{p}}\subset S$ to ${\mathfrak{p}}_{Q}\subset Q[S]$ in the forward direction and ${\mathfrak{q}}\subset Q[S]$ to ${\mathfrak{q}}\cap S\subset S$ in the backward direction. Here, the inclusion relations are preserved under the correspondence, therefore the height is preserved. Hence, we have the bijective correspondence between minimal associated primes of $\mathfrak{a}$} and those of ${\mathfrak{a}}_{Q}$. This completes the proof of (i). As for the proof of (ii), the equivalence of (a) and (b) follows from that $f\in{\mathfrak{a}}^{n}$ if and only if $f\in{\mathfrak{a}}_{Q}^{n}$. By the faithful flatness of $Q\hookrightarrow K$ it follows ${\mathfrak{a}}_{K}^{n}\cap Q[S]={\mathfrak{a}}_{Q}^{n}$. Then (b) holds if and only if $f\in{\mathfrak{a}}_{K}^{n}\setminus{\mathfrak{a}}_{K}^{n+1}$, which yields (c).
Now, let $\mathfrak{p}$} be a minimal prime of ${\mathfrak{a}}_{K}$ and generated locally by $g\in K[S]$. Here, “generated locally” means generated in the localized ring by $\mathfrak{p}$}. We note that
${\mathfrak{p}}\cap Q[S]={\mathfrak{a}}_{Q},$
indeed, the inclusion $\supset$ is trivial and the equality follows from the equality of the heights of prime ideals of the both-hand-side. Then, $K[S]_{{\mathfrak{p}}}$ dominates $Q[S]_{{\mathfrak{a}}_{Q}}$. Therefore, a local generator $g^{\prime}\in Q[S]$ of ${\mathfrak{a}}_{Q}$ is expressed as $g^{\prime}=g^{e}\cdot u$ with a unit $u$ in the local ring $K[S]_{{\mathfrak{p}}}$. Noting that ${\operatorname{char}}K=0$, the extension $Q\subset K$ is separable and therefore $e=1$. Hence, $f\in{\mathfrak{a}}_{Q}^{n}\setminus{\mathfrak{a}}_{Q}^{n+1}$ if and only if $f\in{\mathfrak{p}}^{n}\setminus{\mathfrak{p}}^{n+1}$. This shows the equivalence of (c) and (d). ∎
Remark 2.14.
Lemma 2.13 also hold for the case that $S\subset\Sigma[x_{1},\ldots,x_{N}]$ have a graded structure and ideals are homogeneous ideals, where the degree is given by a weighted degree of $x_{i}$’s.
Proposition 2.15.
Let $R\subset k[x_{1},\ldots,x_{N}]$ be a finitely generated $k$-subalgebra and ${\widetilde{R}}\subset{\mathbb{C}}[x_{1},\ldots,x_{N}]$ a finitely generated $\mathbb{C}$}-algebra. Let ${\mathfrak{a}}\subset R$ be a non-zero proper ideal and ${\widetilde{\mathfrak{a}}}\subset{\widetilde{R}}$ its lifting. Then,
$0pt{\mathfrak{a}}\leq 0pt{\widetilde{\mathfrak{a}}}$.
Proof.
Let $S\subset R$ and ${\widetilde{S}}\subset{\widetilde{R}}$ be the compatible skeletons giving the lifting $\widetilde{\mathfrak{a}}$} of $\mathfrak{a}$}. Then, as $\widetilde{S}$} is integral domain containing $\mathbb{Z}$}, it is flat over $\mathbb{Z}$}. Denote the skeleton ideal of $\widetilde{\mathfrak{a}}$} by $\underline{{\widetilde{\mathfrak{a}}}}$ and that of $\mathfrak{a}$} by $\underline{{\mathfrak{a}}}$. Then, their extensions $\underline{{\widetilde{\mathfrak{a}}}}_{\mathbb{Q}}\subset{\widetilde{S}}\otimes_{\mathbb{Z}}{\mathbb{Q}}$ and $\underline{{\mathfrak{a}}}=\underline{{\widetilde{\mathfrak{a}}}}_{{\mathbb{Z}}/(p)}:=\Phi_{p}(\underline{{\widetilde{\mathfrak{a}}}})\subset{\widetilde{S}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}$ have the relation:
$0pt\underline{{\widetilde{\mathfrak{a}}}}_{\mathbb{Q}}\geq 0pt\underline{{\widetilde{\mathfrak{a}}}}_{{\mathbb{Z}}/(p)}=0pt\underline{{\mathfrak{a}}}.$
Then, by Lemma 2.13, (i), this inequality implies
$0pt{\widetilde{\mathfrak{a}}}\geq 0pt{\mathfrak{a}}.$
3. log discrepancies for multi-ideals and log canonical thresholds of ideals
In this section we study a pair $(A,{\mathfrak{a}}^{e})$ consisting of a non-singular affine variety $A$ of dimension $N$ defined over a field of arbitrary characteristic and a “multi-ideal” ${\mathfrak{a}}^{e}={\mathfrak{a}}_{1}^{e_{1}}\cdots{\mathfrak{a}}_{r}^{e_{r}}$ on $A$, where ${\mathfrak{a}}_{i}$’s are non-zero coherent ideal sheaves on $A$ and $e$ is a combination of the exponents $e=(e_{1},\ldots,e_{r})\in{\mathbb{R}}_{>0}^{r}$.
Definition 3.1.
Let $A$ be a normal variety defined over a field $k$ and $P\in A$ a (not necessarily closed) point.
Let $\varphi_{1}:A_{1}\to A$ be a proper birational morphism with $A_{1}$ normal and $E_{1}\subset A_{1}$ an irreducible divisor. Let $\varphi_{2}:A_{2}\to A$ be another proper birational morphism from a normal $A_{2}$ with an irreducible divisor $E_{2}\subset A_{2}$. Define
$E_{1}\sim E_{2}$
if the birational map $\varphi_{2}^{-1}\circ\varphi_{1}:A_{1}\dasharrow A_{2}$ is a local isomorphism at the generic points of $E_{1}$ and $E_{2}$. Then, $\sim$ becomes a equivalence relation, and such an equivalence class is called a prime divisor over $A$. In this case we denote $E_{1}$ and $E_{2}$ by the same symbol, say $E$, and we say “$E$ appears on $A_{1}$ and on $A_{2}$”. (Strictly speaking, we should be talking about the corresponding valuation instead.) If $P$ is the generic point of $\varphi_{1}(E)$, then we call $E$ a prime divisor over $A$ with the center at $P$ or just a prime divisor over $(A,P)$.
Definition 3.2.
Let $(A,{\mathfrak{a}}^{e})$ be a pair as in the beginning of this section. The log discrepancy of such a pair at a prime divisor $E$ over $A$ is defined as
$a(E;A,{\mathfrak{a}}^{e}):=k_{E}+1-\sum_{i=1}^{r}e_{i}\cdot v_{E}({\mathfrak{a}}_{i}),$
where $k_{E}$ is the coefficient of the relative canonical divisor $K_{{A}^{\prime}/A}$ at $E$.
Here, $\varphi:{A^{\prime}}\to A$ is a birational morphism such that $E$ appears on a normal variety $A^{\prime}$}. The valuation of an ideal ${\mathfrak{a}}_{i}$ at $E$ is defined as follows:
$v_{E}({\mathfrak{a}}_{i}):=\min\{v_{E}(f)\mid f\in{\mathfrak{a}}_{i}\}$
We say that the pair $(A,{\mathfrak{a}}^{e})$ is log canonical at a (not necessarily closed) point $P\in A$ if
(1) $a(E;A,{\mathfrak{a}}^{e})\geq 0,$
holds for every exceptional prime divisor $E$ over $A$ with the center containing $P\in A$.
Definition 3.3.
Let $A$ be a non-singular variety, $P\in A$ a point and ${\mathfrak{a}}^{e}$ a non-zero multi-ideal on $A$. We define the log canonical threshold of the pair $(A,{\mathfrak{a}}^{e})$ at $P$ as follows:
${\operatorname{lct}}(P;A,{\mathfrak{a}}^{e})=\sup\{c\in{\mathbb{R}}_{>0}\mid(A,({\mathfrak{a}}^{e})^{c})\ \mbox{is\ log \ canonical\ at \ }P\}.$
Let $E$ be a prime divisor over $A$ whose center on $A$ is contained in the locus of $\mathfrak{a}$}. Define
$z(E;A,{\mathfrak{a}}^{e}):=\frac{k_{E}+1}{v_{E}({\mathfrak{a}}^{e})},$
where, we define $v_{E}({\mathfrak{a}}^{e}):=\sum_{i}e_{i}\cdot v_{E}({\mathfrak{a}}_{i})$. Then, the following is well known:
[…]
Definition 3.4.
Let $(A,{\mathfrak{a}}^{e})$ and $P\in A$ be as in Definition 3.2. Then the minimal log discrepancy is defined as follows:
  1. (1)
    When $\dim A\geq 2$,
    ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e}):=\inf\{a(E;A,{\mathfrak{a}}^{e})\mid E:\ {\operatorname{prime\ divisor\ over}}\ (A,P)\}.$
  2. (2)
    When $\dim A=1$, define ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})$ by the same definitions as in (1) if the right hand side of the equality (1) is non-negative and otherwise define ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})=$-$\infty$.
Here, we remark that for every pair $(A,{\mathfrak{a}}^{e})$, either ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})\geq 0$ or ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})=$-$\infty$ holds.
Definition 3.5.
Let $A$, NN, and ${\mathfrak{a}}^{e}$ and $P\in A$ be as in Definition 3.2. We say that a prime divisor $E$ over $(A,P)$ computes ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})$, if
[…]
Let $A$, $P$ and ${\mathfrak{a}}^{e}$ be as in Definition 3.3. Let $E$ be a prime divisor over $A$ with the center containing $P$. We say that a prime divisor $E$ computes ${\operatorname{lct}}(P;A,{\mathfrak{a}}^{e})$ if
$z(E;A,{\mathfrak{a}}^{e})={\operatorname{lct}}(P;A,{\mathfrak{a}}^{e}).$
Remark 3.6.
Let ${\mathfrak{m}}_{P}\subset{\mathcal{O}}_{A}$ be the ideal defining $P$. If there is a log resolution of $(A,{\mathfrak{m}}_{P}\cdot{\mathfrak{a}}^{e})$ in a neighborhood of $P$, or if $e$ is a set of rational numbers, then there exists a prime divisor computing ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})$. Otherwise, the existence of such a divisor was not known in general. We will show the existence of a prime divisor computing ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})$ for a smooth point of arbitrary dimension as a corollary of our main theorem (Corollary 1.2, (i)).
Similarly, if there is a log resolution of $(A,{\mathfrak{a}}^{e})$ in a neighborhood of $P$, then there exists a prime divisor over $A$ computing ${\operatorname{lct}}(P;A,{\mathfrak{a}}^{e})$. However, even in this case a prime divisor computing ${\operatorname{lct}}(P;A,{\mathfrak{a}}^{e})$ is not necessarily a prime divisor over $(A,0)$, while its center on $A$ just contains $P$.
Definition 3.7.
Let $X$ be a scheme of finite type over a field $k$ and $k^{\prime}\supset k$ a field extension. For $m\in{\mathbb{Z}}_{\geq 0}$ a $k$-morphism ${\mathrm{Spec\ }}k^{\prime}[t]/(t^{m+1})\to X$ is called an $m$-jet of $X$ and $k$-morphism ${\mathrm{Spec\ }}k^{\prime}[[t]]\to X$ is called an arc of $X$.
Let $X_{m}$ be the space of $m$-jets of $X$. It is well known that $X_{m}$ has a scheme structure of finite type over $k$. If $X$ is an affine variety, then there exists the projective limit
$X_{\infty}:=\lim_{\overleftarrow{m}}X_{m}$
and it is called the space of arcs or the space of $\infty$-jet of $X$.
Definition 3.8.
Denote the canonical truncation morphisms induced from $k[[t]]\to k[t]/(t^{m+1})$ and $k[t]/(t^{m+1})\to k$ by $\psi_{m}:X_{\infty}\to X_{m}$ and $\pi_{m}:X_{m}\to X$, respectively. In particular we denote the morphism $\psi_{0}=\pi_{\infty}:X_{\infty}\to X$ by $\pi$.
Definition 3.9.
Let $\mathfrak{a}$} be a non-zero ideal on a variety $X$. We define the subsets “contact loci of $\mathfrak{a}$}” in the space of arcs as follows:
${\mathrm{Cont}}^{\geq m}({\mathfrak{a}})=\{\gamma\in X_{\infty}\mid{\mathrm{ord}}_{\gamma}({\mathfrak{a}}):={\mathrm{ord}}_{t}\gamma^{*}({\mathfrak{a}})\geq m\},$
where $\gamma^{*}:{\mathcal{O}}_{X}\to k^{\prime}[[t]]$ is the homomorphism of rings corresponding to $\gamma$.
Example 3.10.
Let $Z$ be a closed subscheme of affine $N$-space $A:={\mathbb{A}}_{k}^{N}\\ ={\mathrm{Spec\ }}k[x_{1},\ldots,x_{N}]$ defined over a field $k$. Let ${\mathfrak{a}}\subset k[x_{1},\ldots,x_{N}]$ be the defining ideal of $Z$ in $A$. Assume $\mathfrak{a}$} is generated by $f_{1},\ldots,f_{r}\in k[x_{1},\ldots,x_{N}]$. We define polynomials
$F_{i}^{(j)}\in k\left[x_{\ell}^{(q)}\middle|1\leq\ell\leq N,0\leq q\leq j\right]$
so that
$f_{i}\left(\sum_{q\geq 0}x_{1}^{(q)}t^{q},\sum_{q\geq 0}x_{2}^{(q)}t^{q},\ldots,\sum_{q\geq 0}x_{N}^{(q)}t^{q}\right)=F_{i}^{(0)}+F_{i}^{(1)}t+\cdots+F_{i}^{(j)}t^{j}+\cdots.$
Then the contact locus ${\mathrm{Cont}}^{\geq m}{\mathfrak{a}}\subset A_{\infty}$ is defined by the ideal of $k[x_{\ell}^{(q)}\mid 1\leq\ell\leq N,0\leq q]$ generated by
$F_{i}^{(j)}\ \ (i=1,\ldots,r,j=0,1,\ldots,m-1).$
Here, we note that if all coefficients of $f_{i}$’s are in a subring $\Sigma\subset k$, then all coefficients of $F_{i}^{(j)}$’s are also in $\Sigma$. (This fact will be used in the proof of Lemma 3.12.)
The fiber $\pi_{m}^{-1}(0)$ of the origin $0\in{\mathbb{A}}_{k}^{N}=A$ by the truncation morphism $\pi_{m}:A_{m}\to A$ is defined by the ideal
$(x_{1}^{(0)},\ldots,x_{N}^{(0)})\subset k[x_{\ell}^{(q)}\mid 1\leq\ell\leq N,0\leq q\leq m].$
The following Proposition 3.11 and Lemma 3.12 require perfectness of the base field $k$. As our base field $k$ is algebraic closed, we can apply these formula.
Proposition 3.11.
Let $A$ be a non-singular variety defined over a perfect field, $P\in A$ a closed point and $e=(e_{1},\ldots,e_{r})\in{\mathbb{R}}_{>0}^{r}$. Let ${\mathfrak{a}}^{e}$ be a multi-ideal ${\mathfrak{a}}_{1}^{e_{1}}\cdots{\mathfrak{a}}_{r}^{e_{r}}$ on $A$. Then, it follows:
(2) ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})=\inf_{m\in{\mathbb{Z}}_{\geq 0}^{r}}\left\{{\mathrm{codim}}\left(\bigcap_{i}({\mathrm{Cont}}^{\geq m_{i}}({\mathfrak{a}}_{i})\cap\pi^{-1}(P),A_{\infty}\right)-\sum_{i=1}^{r}m_{i}e_{i}\right\}.$
(3) ${\operatorname{lct}}(P;A,{\mathfrak{a}}^{e})=\inf_{m\in{\mathbb{Z}}_{\geq 0}}\left\{\frac{{\mathrm{codim}}_{P}\left(\bigcap_{i}{\mathrm{Cont}}^{\geq m_{i}}({\mathfrak{a}}_{i}),A_{\infty}\right)}{\sum_{i}m_{i}e_{i}}\right\},$
where ${\mathrm{codim}}_{P}({\mathrm{Cont}}^{\geq m}({\mathfrak{a}}),A_{\infty})$ is the minimum of codimensions of the irreducible components $T$ of ${\mathrm{Cont}}^{\geq m}({\mathfrak{a}})$ in $A_{\infty}$ such that $\pi(T)\cap\{P\}\neq\emptyset$.
Proof.
The equality (2) is obtained by [5] for characteristic 0 and by [9] for arbitrary characteristic. The equality (3) is obtained by [17] for characteristic 0 and by [20] for arbitrary characteristic. ∎
Lemma 3.12.
[…]
(4) ${\mathrm{mld}}(0;A,{\mathfrak{a}}^{e})\leq{\mathrm{mld}}({\tilde{0}};{\widetilde{A}},{\widetilde{\mathfrak{a}}}^{e}),$
(5) ${\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})\leq{\operatorname{lct}}({\tilde{0}};{\widetilde{A}},{\widetilde{\mathfrak{a}}}^{e}).$
Proof.
By the assumption, there are compatible skeletons $\widetilde{\Sigma}\to\Sigma$ with $\widetilde{\Sigma}\subset{\mathbb{C}}$ and $\Sigma\subset k$ such that the coefficients of the generators of ${\mathfrak{a}}_{i}$ and of ${\widetilde{\mathfrak{a}}}_{i}$ are in $\Sigma$ and in $\widetilde{\Sigma}$, respectively.
First, we prove the following inequalities which are essential for (4) and (5):
[…]
Let ${\mathfrak{a}}_{i}$ be generated by $f_{ih}\in\Sigma[x_{1},\ldots,x_{N}]$ and ${\widetilde{\mathfrak{a}}}_{i}$ by their liftings ${\tilde{f}}_{ih}\in{\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]$ $(i=1,\ldots,r,h=1,\ldots,s_{i})$, then the defining ideal $I$ of
$\bigcap_{i=1}^{r}{\mathrm{Cont}}^{\geq m_{i}}({\mathfrak{a}}_{i})\cap\pi^{-1}(0)$
in $A_{\infty}$ is generated by $x_{\ell}^{(0)}(\ell=1,\ldots,N)$ and
$F_{ih}^{(j)}\in\Sigma\left[x_{\ell}^{(q)}\middle|\ell=1,\ldots,N,q\geq 0\right]\ \ (i=1,\ldots,r,h=1,\ldots,s_{i},j\leq m_{i}-1).$
Then, we can see that the liftings $x_{\ell}^{(0)}(\ell=1,\ldots,N)$ and
$\widetilde{F}_{ih}^{(j)}\in\widetilde{\Sigma}\left[x_{\ell}^{(q)}\middle|\ell=1,\ldots,N,q\geq 0\right]\ \ (i=1,\ldots,r,h=1,\ldots,s_{i},j\leq m_{i}-1)$
generate the ideal $\widetilde{I}$ defining
[…]
in $\widetilde{A}$$_{\infty}$. Here, $\widetilde{F}_{ih}^{(j)}$ is obtained from $\widetilde{f}_{ih}^{(j)}$ as in Example 3.10. So, we note that $\widetilde{F}_{ih}^{(j)}$ is a lifting of $F_{ih}^{(j)}$. By Lemma 2.13 and Proposition 2.15, it follows that $0ptI\leq 0pt\widetilde{I}$, which yields (6). This completes the proof of (4), by using (2) in Proposition 3.11. Here, we note that Proposition 2.15 states for finitely generated $k$-algebra and $\mathbb{C}$}-algebra, while our statement is for infinitely generated algebras. However as the monomials appearing in the generators and the number of generators of $I$ and $\widetilde{I}$ are finite, we can work in the same way a in Proposition 2.15. Concretely, we may assume for $m\gg 0$, the ideal $I\subset k\left[x_{\ell}^{(q)}\middle|\ell=1,\ldots,N,q\geq 0\right]$ is generated by elements of $k\left[x_{\ell}^{(q)}\middle|\ell=1,\ldots,N,0\leq q\leq m\right]$. Let
$I_{m}\subset k\left[x_{\ell}^{(q)}\middle|\ell=1,\ldots,N,0\leq q\leq m\right]$
be the ideal generated by the generators of $I$. In the same way, we obtain the ideal
${\widetilde{I}}_{m}\subset{\mathbb{C}}\left[x_{\ell}^{(q)}\middle|\ell=1,\ldots,N,0\leq q\leq m\right]$
generating $\widetilde{I}$}. As ${\widetilde{I}}_{m}$ is a lifting of $I_{m}$, applying Proposition 2.15, we have
$0ptI_{m}\leq 0pt{\widetilde{I}}_{m},$
which yields
$0ptI\leq 0pt{\widetilde{I}}$,
where we use $0ptI=0ptI_{m}$ and $0pt{\widetilde{I}}=0pt{\widetilde{I}}_{m}$. The inequality (4) follows from this using the formula in Proposition 3.11.
For (5), we need one more step. We note that
${\operatorname{lct}}(0;A,{\mathfrak{a}})=\inf_{m}\frac{{\mathrm{codim}}_{0}(\bigcap_{i}{\mathrm{Cont}}^{\geq m_{i}}({\mathfrak{a}}_{i}),A_{\infty})}{\sum_{i}m_{i}e_{i}}\leq\inf_{m}\frac{{\mathrm{codim}}(\bigcap_{i}{\mathrm{Cont}}^{\geq m_{i}}({\mathfrak{a}}_{i})\cap\pi^{-1}(0),A_{\infty})}{\sum_{i}m_{i}e_{i}},$
because an irreducible component of $\bigcap_{i}{\mathrm{Cont}}^{\geq m_{i}}({\mathfrak{a}}_{i})\cap\pi^{-1}(0)$ is contained in an irreducible component $T$ of $\bigcap_{i}{\mathrm{Cont}}^{\geq m_{i}}({\mathfrak{a}}_{i})$ with $\pi(T)\cap\{0\}\neq\emptyset$. Then, by (6), it follows
$\inf_{m}\frac{{\mathrm{codim}}(\bigcap_{i}{\mathrm{Cont}}^{\geq m_{i}}({\mathfrak{a}}_{i})\cap\pi^{-1}(0),A_{\infty})}{\sum_{i}m_{i}e_{i}}\leq\inf_{m}\frac{{\mathrm{codim}}(\bigcap_{i}{\mathrm{Cont}}^{\geq m_{i}}({\widetilde{\mathfrak{a}}}_{i})\cap\pi^{-1}({\tilde{0}}),{\widetilde{A}}_{\infty})}{\sum_{i}m_{i}e_{i}}$
On the other hand, in characteristic 0, by [4, Lemma 2.6], we have the equality
${\operatorname{lct}}({\tilde{0}};{\widetilde{A}},{\widetilde{\mathfrak{a}}}^{e})=\inf_{m}\frac{{\mathrm{codim}}(\bigcap_{i}{\mathrm{Cont}}^{\geq m_{i}}({\widetilde{\mathfrak{a}}}_{i})\cap\pi^{-1}({\tilde{0}}),{\widetilde{A}}_{\infty})}{\sum_{i}m_{i}e_{i}},$
which yields (5).
4. Construction of a weak lifting of the prime divisor
As we mentioned just after the Corollary 1.2, the corollary holds for the dimension of the variety $N=1$. So, from now on, we assume $N\geq 2$.
Proposition-Definition 4.1.
(Zariski, see eg.[14, VI, 1.3]) Let $E$ be a prime divisor over a variety $A^{\prime}$ defined over a field $k$ and appear on $\psi:A^{\prime\prime}\to A^{\prime}$. Let $Q\in A^{\prime\prime}$ be the generic point of $E$ and $Q_{0}\in A^{\prime}$ the image of $Q$. Then, we have a sequence of varieties $A^{\prime}_{0},A^{\prime}_{1},\ldots,A^{\prime}_{n}$ and birational maps as follows:
  1. $(0)$
    $A^{\prime}_{0}=A^{\prime}$, $\psi^{\prime}_{0}=\psi$.
  2. $(i)$
    If $\psi^{\prime}_{i}:A^{\prime\prime}\dashrightarrow A^{\prime}_{i}$ is already defined, then let $Q_{i}=\psi_{i}(Q)$. Let $A^{\prime}_{i+1}=Bl_{Q_{i}}A^{\prime}_{i}$ and $\psi^{\prime}_{i+1}:A^{\prime\prime}\dashrightarrow A^{\prime}_{i}$ be the induced map.
  3. $(n)$
    Then, the final birational map $\psi^{\prime}_{n}:A^{\prime\prime}\dashrightarrow A^{\prime}_{n}$ is isomorphic at $Q$, i.e., $E$ appears on $A^{\prime}_{n}$.
Here, $Bl_{Q_{i}}A^{\prime}_{i}$ is the blow-up of $A^{\prime}_{i}$ at $Q_{i}$. Let $\varphi^{\prime}_{i}:A^{\prime}_{i}\to A^{\prime}_{i-1}$ be the blow-up morphism and $E_{i}\subset A^{\prime}_{i}$ be the exceptional divisor dominating $Q_{i}$. We call this sequence $A^{\prime}_{n}\to\cdots\to A^{\prime}_{1}\to A^{\prime}_{0}=A^{\prime}$ the Zariski sequence for $E$ over $A^{\prime}$.
Note that if $Q_{0}\in A^{\prime}$ is a non-singular point, then $Q_{i}$’s are all non-singular for $i=1,\ldots,n-1$. Indeed, this is proved inductively. As $A^{\prime}_{1}$ is the blow-up at a non-singular point $Q_{0}$, it follows that $A^{\prime}_{1}$ and $E_{1}$ are non-singular in a neighborhood of ${\psi^{\prime}_{1}}^{-1}(Q_{0})$. Since $Q_{1}\in{\psi^{\prime}_{1}}^{-1}(Q_{0})$, it follows that $Q_{1}$ is again a non-singular point. In the same way, we obtain that $Q_{i}$’s are all non-singular.
Remark 4.2 (see for example [15]).
Let $A_{n}\to A_{n-1}\to\cdots A_{1}\to A$ be a sequence of blow-ups with the center at subschemes over the projective space $A={\mathbb{P}}_{k}^{N}$. Then, each $A_{i}$ is described as $A_{i}={\mathrm{Proj\ }}R_{i}$, where $R_{i}$ is a graded ring with its degree 0 part $(R_{i})_{0}=k$ as follows:
(7) $R_{i}=\oplus_{\ell\geq 0}k[x_{0},x_{1},\ldots,x_{N},I_{1}T_{1},\ldots,I_{i}T_{i}]_{(\ell g_{0},\ell g_{1},\ldots,\ell g_{i-1},\ell g_{i})},$
$\mbox{for}\ \ g_{0}\gg g_{1}\gg\cdots\gg g_{i}\gg..\gg g_{n},$
where $x_{i}$ and $T_{j}$ are indeterminates for all $i$ and $j$, and ${I}_{1}\subset R_{0},\ldots,I_{i}\subset R_{i-1},\ldots,I_{n}\subset R_{n-1}$ are the homogeneous ideals defining the center of each blow-up. We define the multi-degree of an element $r\left(\prod_{j=0}^{N}x_{j}^{a_{j}}\right)T_{1}^{e_{1}}\cdots T_{i}^{e_{i}}\in R_{i}$ $(r\in k,\ e_{0}:=\sum_{j=0}^{N}a_{j})$ as follows:
$\deg r\left(\prod_{j=0}^{N}x_{j}^{a_{j}}\right)T_{1}^{e_{1}}\cdots T_{i}^{e_{i}}=(e_{0},\cdots,e_{i}).$
By this degree, the ring $R_{0}[I_{1}T_{1},\ldots,I_{i}T_{i}]$ becomes a multi-graded ring and $R_{0}[I_{1}T_{1},\ldots,I_{i}T_{i}]_{(\ell g_{0},\ldots,\ell g_{i-1},\ell g_{i})}$ is its homogeneous part of degree $(\ell g_{0},\ldots,\ell g_{i-1},\ell g_{i})$. Then, $R_{i}$ is a “diagonal” sub-graded ring of $R_{0}[I_{1}T_{1},\ldots,I_{i}T_{i}]$. From now on, we regard $R_{i}$ as a graded ring with respect to $\ell\in{\mathbb{Z}}_{\geq 0}$ and a homogeneous ideal of $R_{i}$ is in this grading.
We will now present a fundamental lemma that will be used in the upcoming proposition:
Lemma 4.3.
Let $A$ be an irreducible Noetherian scheme and $Z\subset A$ a closed subscheme with locally a complete intersection in $A$. Let $\psi:A^{\prime}\to A$ be the blow-up by $Z$. Then, the following hold:
  1. (i)
    $\psi^{-1}(Z)\simeq{\mathbb{P}}(I_{Z}/I_{Z}^{2})$, where $I_{Z}$ is the defining ideal of $Z$ in $A$. In particular, the irreducible components of $\psi^{-1}(Z)$ and the irreducible components of $Z$ correspond bijectively.
  2. (ii)
    Assume $A$ is a normal Gorenstein variety over a field $k$, and for any irreducible component $Z^{\prime}$ of $Z$, the singular loci ${\mathrm{Sing}}A$ and ${\mathrm{Sing}}Z$ of $A$ and of $Z,$ respectively, satisfy:
    (8) $Z^{\prime}\not\subset({\mathrm{Sing}}A)\cup({\mathrm{Sing}}Z).$
    Then $\psi^{-1}(Z)$ is Gorenstein and $A^{\prime}$ is normal and Gorenstein.
Proof.
As the embedding $Z\subset A$ is locally a complete intersection, by [6, IV, 16.9.4], it follows
$\bigoplus_{n\geq 0}I_{Z}^{n}/I^{n+1}_{Z}\simeq S(I_{Z}/I_{Z}^{2}),$
where $I_{Z}/I_{Z}^{2}$ is a locally free ${\mathcal{O}}_{Z}$-modlue of rank $c={\mathrm{codim}}(Z,A)$. The description yields $\psi^{-1}(Z)\simeq{\mathbb{P}}(I_{Z}/I_{Z}^{2})$ as required in (i). On the other hand, as $\psi^{-1}(Z)$ is a projective bundle on $Z$, its irreducible component is in the form $\psi^{-1}(Z^{\prime})$ for an irreducible component $Z^{\prime}$ of $Z$, which yields the second statement of (i).
For (ii), we use (8) to see that $A^{\prime}$ satisfies $R_{1}$ condition. On the other hand, we have that $Z$ is Gorenstein by the assumption on $Z$ and $A$. Then, it follows $\psi^{-1}(Z)$ is locally isomorphic to $Z\times{\mathbb{P}}^{c-1}$, which implies $\psi^{-1}(Z)$ is Gorenstein. As $\psi^{-1}(Z)$ is locally a hypersurface of $A^{\prime}$, it follows $A^{\prime}$ is Gorenstein. Applying the Serre’s criterion to $A^{\prime}$, we obtain that $A^{\prime}$ is normal. ∎
For later use, we prove the following vanishing theorem for a blow-up as in Lemma 4.3.
Lemma 4.4.
Let $A$ be a normal variety over a field $k$ and $Z\subset A$ a closed subscheme locally a complete intersection in $A$. Let $\psi:A^{\prime}\to A$ be the blow-up by the center $Z$. Then, it follows:
[…]
Proof.
The equality $\psi_{*}{\mathcal{O}}_{A^{\prime}}={\mathcal{O}}_{A}$ follows immediately from normality of $A$. For the vanishing $R^{i}\psi_{*}{\mathcal{O}}_{A^{\prime}}=0$ $(i>0)$, we remark the following facts:
  1. (1)
    $R^{i}\psi_{*}{\mathcal{O}}_{E}=0$ $(i>0)$ for $E:=\psi^{-1}(Z)={\mathbb{P}}(I_{Z}/I_{Z}^{2})$,
  2. (2)
    $R^{i}\psi_{*}(\mathcal{I}^{j}\otimes{\mathcal{O}}_{E})=0$ $(i,j>0)$ for the defining ideal sheaf $\mathcal{I}$ of $E$ in $A^{\prime}$,
  3. (3)
    $R^{i}\psi_{*}\mathcal{I}^{n}=0$ for $n\gg 0$.
(1) follows from that $E\to Z$ is locally isomorphic to the canonical projection $Z\times{\mathbb{P}}^{c-1}\to Z$, where $c={\mathrm{codim}}(Z,A)$. (2) follows from the same reason and $\mathcal{I}$ is relatively very ample on $E$. (3) follows from the Serre’s vanishing for the relatively ample invertible sheaf $\mathcal{I}$ on $A^{\prime}$.
Now consider the $R^{i}\psi_{*}$ of the exact sequences:
$0\to\mathcal{I}^{j+1}\to\mathcal{I}^{j}\to\mathcal{I}^{j}\otimes{\mathcal{O}}_{E}\to 0$
for $j=0,\ldots,n-1$. Then, by (1)–(3), we obtain $R^{i}\psi_{*}{\mathcal{O}}_{A^{\prime}}=0$ $(i>0)$.
Corollary 4.5.
Let $A_{0}$ be a normal Gorenstein variety over a field $k$ and $\psi$ be the composite
$A_{n}\stackrel{{\scriptstyle\varphi_{n}}}{{\longrightarrow}}A_{n-1}\to\cdots A_{1}\stackrel{{\scriptstyle\varphi_{1}}}{{\longrightarrow}}A_{0}$
of blow-ups $\varphi_{i}$ by the center of locally a complete intersection in $A_{i-1}$ satisfying the condition of (ii) of Lemma 4.3. Then, it follows
$R^{i}\psi_{*}{\mathcal{O}}_{A_{n}}=\left\{\begin{array}[]{ll}0&i>1\\ {\mathcal{O}}_{A_{0}}&i=0\\ \end{array}\right.$.
Proof.
As it holds if $i=1$ by Lemma 4.4, we prove the corollary by induction on $n$. Assume $n\geq 2$ and for $\psi^{\prime}:=\varphi_{1}\circ\cdots\circ\varphi_{n-1}$,
$R^{i}\psi^{\prime}_{*}{\mathcal{O}}_{A_{n-1}}=\left\{\begin{array}[]{ll}0&i>1\\ {\mathcal{O}}_{A_{0}}&i=0\\ \end{array}\right.$.
Consider the Leray’s spectral sequence.
$E_{2}^{p,q}=R^{p}\psi^{\prime}_{*}R^{q}{\varphi_{n}}_{*}{\mathcal{O}}_{A_{n}}\Rightarrow E^{p+q}=R^{p+q}(\psi^{\prime}\circ\varphi_{n})_{*}{\mathcal{O}}_{A_{n}}$
As we know that $A_{n-1}$ is normal Gorenstein by Lemma 4.3, it follows $(\varphi_{n})_{*}{\mathcal{O}}_{A_{n}}={\mathcal{O}}_{A_{n-1}}$. Then, we have $E_{2}^{p,q}=0$ for $(p,q)\neq(0,0)$. Therefore $E^{i}=0$ for $i>0$ and $E^{0}={\mathcal{O}}_{A_{0}}$ which give the required equalities.
Proposition 4.6.
Let $k$ be an algebraically closed field of characteristic $p\geq 0$ and $E$ a prime divisor over the affine space $A={\mathbb{A}}_{k}^{N}$ over $k$ with the center at the origin $0\in A$. Then, there exists a sequence:
$A_{n}\stackrel{{\scriptstyle\varphi_{n}}}{{\longrightarrow}}A_{n-1}\to\cdots A_{1}\stackrel{{\scriptstyle\varphi_{1}}}{{\longrightarrow}}A$
of blow-ups such that
  1. (i)
    $E$ appears on $A_{n}$,
  2. (ii)
    the first blow-up $\varphi_{1}$ is by the maximal ideal ${\mathfrak{m}}=(x_{1},\ldots,x_{N})\subset k[x_{1},\ldots,x_{N}]$,
  3. (iii)
    for each $i$, the morphism $\varphi_{i}:A_{i}\to A_{i-1}$ is the blow-up along a closed subscheme $Z_{i}\subset A_{i-1}$ of locally a complete intersection. One of the irreducible components of $Z_{i}$ is the center of $E$ on $A_{i-1}$, while the other components, if exist, are not contained in the fiber over $0\in A(=:A_{0})$ on $A_{i-1}$,
  4. (iv)
    $A_{i}$’s are all normal Gorenstein varieties and smooth at the generic point of the center of $E$ and also smooth outside of the fiber over $0\in A$.
Proof.
We need the blow-up sequence only over $A={\mathbb{A}}_{k}^{N}$, however by a technical reason we compactify it to $\overline{A}={\mathbb{P}}_{k}^{N}$ with the identification
$A\subset{\overline{A}},\ \ \ (x_{1},\ldots,x_{N})\mapsto(1:x_{1}:\cdots:x_{N})$
and will construct the sequence over $\overline{A}$}. Then, we can restrict it over $A$ to obtain the required sequence of the proposition.
Let $\varphi_{1}:{\overline{A}}_{1}\to{\overline{A}}$ be the blow-up by $(x_{1},\ldots,x_{N})\subset{\overline{R}}_{0}:=k[x_{0},x_{1},\cdots,x_{N}]$, the ideal defining $0:=(1:0:\ldots:0)\in A\subset{\overline{A}}$ which is the center of $E$ on $A$. (automatically (ii) and (iii) are satisfied for $i=1$). As ${\overline{A}}_{1}$ is smooth, (iv) is satisfied for $i=1$. We will construct $\varphi_{i}$ by induction on $i$ so that (iii) and (iv) are satisfied.
Let $i\geq 2$ and assume we constructed ${\overline{A}}_{i-1}$ with the properties (iii) and (iv). As ${\overline{A}}_{i-1}\to{\overline{A}}$ is projective and $\overline{A}$} is projective, we can write ${\overline{A}}_{i-1}={\mathrm{Proj\ }}{\overline{R}}_{i-1}$ and
${\overline{R}}_{i-1}=\oplus_{n=0}^{\infty}J_{n},$
where $J_{n}$ is the degree $n$ part of the graded ring ${\overline{R}}_{i-1}$ and $J_{0}=k$. By replacing ${\overline{R}}_{i-1}$ by a suitable Veronese subring, we may assume that ${\overline{R}}_{i-1}$ is generated by $J_{1}$. Let ${\mathfrak{p}}_{i}\subset{\overline{R}}_{i-1}$ be the homogeneous prime ideal defining the center of $E$ in ${\overline{A}}_{i-1}$.
Claim. We can take a subset $\{f_{1},\ldots,f_{t}\}\subset{\mathfrak{p}}_{i}$ such that
  1. (1)
    $V(f_{1},\ldots,f_{t})=V({\mathfrak{p}}_{i})$ and each $f_{j}$ has the common degree $d>0$ in the graded ring ${\overline{R}}_{i-1}$,
  2. (2)
    the linear system $\Lambda$ defined by the $k$-vector subspace
    $\langle f_{1},\ldots,f_{t}\rangle\subset\Gamma({\overline{A}}_{i-1},{\mathcal{O}}_{{\overline{A}}_{i-1}}(d))$
    is very ample on ${\overline{A}}_{i-1}\setminus V({\mathfrak{p}}_{i})$. In prticular, $\Lambda$ is very ample on ${\overline{A}}_{i-1}\setminus\phi_{i-1}^{-1}(0)$, where $\phi_{i-1}:{\overline{A}}_{i-1}\to{\overline{A}}$ is the composite $\varphi_{1}\circ\cdots\circ\varphi_{i-1}$ and $0\in{\overline{A}}$ is the origin of $A$. Here, note that $V({\mathfrak{p}}_{i})\subset\phi_{i-1}^{-1}(0)$.
Let $\{h_{1},\ldots,h_{s}\}$ be a generator system of $J_{1}$ over $k$, then the linear system over $k$ defined by $\{h_{1},\ldots,h_{s}\}$ is very ample on ${\overline{A}}_{i-1}$. To prove the Claim above, we start by taking any homogeneous generator system $\{f^{\prime}_{1},\ldots,f^{\prime}_{r}\}$ of ${\mathfrak{p}}_{i}\subset{\overline{R}}_{i-1}$ with the degree $d_{1}\leq\cdots\leq d_{r}$, respectively. We will construct an appropriate system of generators with the properties claimed in (1) and (2). Here, $V({\mathfrak{p}}_{i})\subset{\overline{A}}_{i-1}$ is the center of $E$ on ${\overline{A}}_{i-1}$.
Let $d>d_{r}$. Define
$f_{i\gamma}:=f^{\prime}_{i}\cdot X_{\gamma}^{(d-d_{i})},$
where
$\left\{X_{\gamma}^{(d-d_{i})}\middle|\gamma=1,\ldots,\left(\begin{array}[]{c}ms+d-d_{i}-1\\ d-d_{i}\\ \end{array}\right)\right\}$
is the set of all monomials in variables $h_{j}$’s of degree $d-d_{i}$. Then the set of elements
$\left\{f_{i\gamma}\middle|1\leq i\leq r,1\leq\gamma\leq\left(\begin{array}[]{c}ms+d-d_{i}-1\\ d-d_{i}\\ \end{array}\right)\right\}$
gives the closed subset $V(f_{i\gamma})=V({\mathfrak{p}}_{i})$. Note that all $f_{i\gamma}$ are of degree $d$. Thus (1) in the Claim holds. Let $\Lambda$ be the linear system defined by the $k$-vector subspace $\langle f_{i\gamma}\rangle\subset\Gamma({\overline{A}}_{i-1},{\mathcal{O}}_{{\overline{A}}_{i-1}}(d))$. Then, $\Lambda$ has the base locus $V({\mathfrak{p}}_{i})$ and is very ample outside of $V({\mathfrak{p}}_{i})$. As $V({\mathfrak{p}}_{i})\subset\phi_{i-1}^{-1}(0)$, the linear system $\Lambda$ is very ample outside of $\phi_{i-1}^{-1}(0)$. Now, denote the $\{f_{i\gamma}\}$ just by $\{f_{i}\mid i=1,\ldots,t\}$ for simplicity of notation. Then the system $\{f_{i}\mid i=1,\ldots,t\}$ satisfies (2) required in the Claim.
Now, let $c=0pt{\mathfrak{p}}_{i}$ and take $c$ general linear combinations of $f_{i}$’s over $k$:
$\sum_{i=1}^{t}a_{ji}f_{i}\ (a_{ji}\in k\ \mbox{ for}\ j=1,\ldots,c)$
and rename them as $f_{1},\ldots,f_{c}$. Then, $\{f_{1},\ldots,f_{c}\}$ becomes a regular sequence in $R_{i-1}$. Let $I_{i}\subset{\overline{R}}_{i-1}$ be the ideal generated by $\{f_{1},\ldots,f_{c}\}$ and $\mathcal{I}_{i-1}$ the ideal sheaf of ${\mathcal{O}}_{{\overline{A}}_{i-1}}$ defined by $I_{i-1}$. Similarly, let $\mathcal{P}_{i}$ be the ideal sheaf of ${\mathcal{O}}_{{\overline{A}}_{i-1}}$ defined by ${\mathfrak{p}}_{i}$. By the choice of $f_{1},\ldots,f_{c}$, these give a regular system of paremeters of ${\mathcal{O}}_{{\overline{A}}_{i-1},P_{i}}$, where $P_{i}\in{\overline{A}}_{i-1}$ is the generic point of $V({\mathfrak{p}}_{i})$. Therefore,
$\mathcal{I}_{i}{\mathcal{O}}_{{\overline{A}}_{i-1},P_{i}}=\mathcal{P}_{i}{\mathcal{O}}_{{\overline{A}}_{i-1},P_{i}}.$
Therefore, one of the irreducible components of $Z_{i}:=V(I_{i})\subset{\overline{A}}_{i-1}$ is the center of $E$ on ${\overline{A}}_{i-1}$.
As $Z_{i}\subset{\overline{A}}_{i-1}$ is obtained by successive cutting ${\overline{A}}_{i-1}$ with $c$ general hypersurfaces
$H_{i}:=V(f_{i})\ \ (i=1,\ldots,c)$
from the linear system $\Lambda$, we obtain the following:
  1. (a)
    $\dim({\mathrm{Sing}}{\overline{A}}_{i-1})\cap Z_{i}<N-c$ and
  2. (b)
    $\dim(\phi_{i-1}^{-1}(0)\setminus V({\mathfrak{p}}_{i}))\cap Z_{i}<N-c$.
Indeed, first we have
$\dim({\mathrm{Sing}}{\overline{A}}_{i-1}\setminus V({\mathfrak{p}}_{i}))\cap Z_{i}<N-c.$
Because $\Lambda$ is very ample on ${\overline{A}}_{i-1}\setminus V({\mathfrak{p}}_{i})$, each cut by a general member of $\Lambda$ makes the dimension of the intersection with ${\mathrm{Sing}}{\overline{A}}_{i-1}\setminus V({\mathfrak{p}}_{i})$ exactly 1 less. Next we have
$\dim({\mathrm{Sing}}{\overline{A}}_{i-1}\cap V({\mathfrak{p}}_{i}))\cap Z_{i}<N-c,$
because $\dim V({\mathfrak{p}}_{i})=N-c$ and the generic point of $V({\mathfrak{p}}_{i})$ is smooth in ${\overline{A}}_{i-1}$. Thus (a) follows, and (b) also follows in the similar way.
By (a), any irreducible component of $Z_{i}$ is not contained in ${\mathrm{Sing}}{\overline{A}}_{i-1}$, since the dimension of every irreducible component of $Z_{i}$ is NcN-c. By (b), any irreducible component of $Z_{i}$, except for $V({\mathfrak{p}}_{i})$, is not contained in the fiber $\phi_{i-1}^{-1}(0)$. Then, $Z_{i}$ satisfies the condition (iii) of the proposition.
Moreover, we obtain that
$({\overline{A}}_{i-1}\setminus\phi_{i-1}^{-1}(0))\cap Z_{i}$
is a smooth scheme by the Bertini’s Theorem “for very ample linear system” $\Lambda$ on a smooth quasi-projective scheme ${\overline{A}}_{i-1}\setminus\phi_{i-1}^{-1}(0)$. Note that the “very ample version” of the Bertini’s Theorem holds for any characteristic base field (see the proof of [8, II, Theorem 8.18]). A fortiori, every irreducible component of $Z_{i}$ is not contained in ${\mathrm{Sing}}Z_{i}$ nor in ${\mathrm{Sing}}{\overline{A}}_{i-1}$. Now the condition (8) in Lemma 4.3 is satisfied. Therefore, if $\varphi_{i}:{\overline{A}}_{i}\to{\overline{A}}_{i-1}$ is the blow-up with the center $Z_{i}\subset{\overline{A}}_{i-1}$, then, ${\overline{A}}_{i}$ is also normal and Gorenstein by the lemma. As the blow-up $\varphi_{i}$ is the same as the blow-up at the center $V({\mathfrak{p}}_{i})\subset{\overline{A}}_{i-1}$ of $E$ around the generic point of $V({\mathfrak{p}}_{i})$, the generic point of the center of $E$ on $A_{i}$ is also smooth (see the comment below Proposition-Definition 4.1).
As ${\overline{A}}_{i-1}\setminus\phi_{i-1}^{-1}(0)$ and (${\overline{A}}_{i-1}\setminus\phi_{i-1}^{-1}(0))\cap Z_{i}$ are smooth, we also obtain the blown-up space
${\overline{A}}_{i}\setminus\phi_{i}^{-1}(0)$
of ${\overline{A}}_{i-1}\setminus\phi_{i-1}^{-1}(0)$ by the center $({\overline{A}}_{i-1}\setminus\phi_{i-1}^{-1}(0))\cap Z_{i}$ is also smooth, which completes the proof of (iv).
As the sequence of the blow-ups is the same as the Zariski sequence for $E$ around the centers of $E$ on any ${\overline{A}}_{i}$ by (iii), the prime divisor $E$ appears on ${\overline{A}}_{n}$ for a certain $n\in{\mathbb{Z}}_{\geq 0}$. Hence, (i) follows. ∎
Definition 4.7.
For a prime divisor $E$ over $A$ with the center at a closed point $0\in A$, the sequence obtained satisfying the conditions (i)–(iv) in Proposition 4.6 is called a normal Gorenstein Zariski sequence for $E$. This name makes sense, as the sequence is the same as the Zariski sequence for $E$ around the generic point of the center of $E$ on each $A_{i}$ by the properties (i) and (iii) in Proposition 4.6.
Proposition 4.8.
Let $E$ be a prime divisor over $A={\mathbb{A}}_{k}^{N}$ for $N\geq 2$ and an algebraically closed field $k$ of characteristic $p>0$. Let
$A_{n}\stackrel{{\scriptstyle\varphi_{n}}}{{\longrightarrow}}A_{n-1}\to\cdots A_{1}\stackrel{{\scriptstyle\varphi_{1}}}{{\longrightarrow}}A$
be a normal Gorenstein Zariski sequence for $E$ constructed in Proposition 4.6. We use the same notation as in Proposition 4.6. Let $P_{i}\in A_{i-1}$ be the generic point of the center of $E$ and $E_{i}\subset A_{i}$ the exceptional divisor mapped to $P_{i}\in A_{i-1}$ by the blow-up $\varphi_{i}$, and let $E^{\prime}_{i}\subset A_{i}$ be the other exceptional divisor for $\varphi_{i}$. Then, there exist a subfield $L\subset{\mathbb{C}}$ and a sequence:
${\widetilde{A}}_{n}\stackrel{{\scriptstyle\widetilde{\varphi}_{n}}}{{\longrightarrow}}{\widetilde{A}}_{n-1}\to\cdots{\widetilde{A}}_{1}\stackrel{{\scriptstyle\widetilde{\varphi}_{1}}}{{\longrightarrow}}{\widetilde{A}}={\mathbb{A}}_{L}^{N},$
satisfying the following:
  1. (i)
    ${\widetilde{A}}_{i}({\operatorname{mod}\ }p)=A_{i}$;
  2. (ii)
    $\widetilde{\varphi}_{i}$ is the blow-up by a lifting ${\widetilde{Z}}_{i}\subset{\widetilde{A}}_{i-1}$ of $Z_{i}$. One of the irreducible components of ${\widetilde{Z}}_{i}$ has the generic point ${\widetilde{P}}_{i}\in{\widetilde{A}}_{i}$ that corresponds to a weak prime lifting of $P_{i}\in A_{i-1}$;
  3. (iii)
    Let ${\widetilde{E}}_{i}\subset{\widetilde{A}}_{i}$ be the exceptional divisor mapped to ${\widetilde{P}}_{i}$ by the blow-up $\widetilde{\varphi}_{i}$, then, for $1\leq i\leq n$,
    $k_{{\widetilde{E}}_{i}}=k_{E_{i}}.$
Proof.
As in the proof of Proposition 4.6, we may assume that $A={\mathbb{P}}_{k}^{N}$ and Denote $A={\mathrm{Proj\ }}R_{0}$, $A_{i}={\mathrm{Proj\ }}R_{i}$ $(i=1,\ldots,n)$, where $R_{0}=k[x_{1},\ldots,x_{N}]$, and
$R_{i}=\oplus_{\ell\geq 0}R_{0}[I_{1}T_{1},\ldots,I_{i}T_{i}]_{(\ell g_{0},\ldots,\ell g_{i-1},\ell g_{i})},\ \ \ \mbox{for}\ \ g_{0}\gg g_{1}\gg\cdots\gg g_{i}\gg..\gg g_{n},$
as in Remark 4.2. Here, $I_{i}\subset R_{i-1}$ is the defining ideal of $Z_{i}$, the center of the blow-up $\varphi_{i}:A_{i}\to A_{i-1}$. Let ${\mathfrak{p}}_{i}\subset R_{i-1}$ be the homogeneous prime ideal defining $P_{i}$, the center of $E$ on $A_{i-1}$. Then, $Z_{i}=V({\mathfrak{p}}_{i})\cup Z^{\prime}_{i}$, where $Z^{\prime}_{i}$ is the other components of $Z_{i}$ than $V({\mathfrak{p}}_{i})$. Note that $R_{i}$ is a finitely generated $k$-subalgebra
$R_{i}\subset k[x_{0},\ldots,x_{N},T_{1},\ldots,T_{i}].$
We define a skeleton
$S_{i}\subset\Sigma[x_{0},\ldots,x_{N},T_{1},\ldots,T_{i}],$
of $R_{i}$ under an appropriately chosen skeleton $\Sigma$ of $k$ for each $i=0,\ldots,n$ in the following way:
Step 0($p$). For an arbitrary chosen skeleton $\Sigma$ of $k$, we can define $S_{0}=\Sigma[x_{1},\ldots,x_{N}]$. Then $S_{0}\subset R_{0}$ is a skeleton of $R_{0}$. (Later, we will restrict the choice of $\Sigma$ such that it contains certain necessary elements in the following steps.)
Step 1($p$). As $I_{1}={\mathfrak{p}}_{1}={\mathfrak{m}}=(x_{1},\ldots,x_{N})\subset k[x_{0},x_{1},\ldots,x_{N}]$, we can take
$\underline{I}_{1}=\underline{{\mathfrak{p}}}_{1}=(x_{1},\ldots,x_{N})\subset\Sigma[x_{0},x_{1},\ldots,x_{N}]$
for arbitrary chosen skeleton $\Sigma$ of $k$. If we denote
$S_{1}:=\oplus_{\ell\geq 0}S_{0}[\underline{I}_{1}T_{1}]_{(\ell g_{0},\ell g_{1})},$
then $S_{1}$ is a skeleton of $R_{1}$ for every choice of a skeleton $\Sigma$ of $k$. It is clear that
$\underline{{\mathfrak{p}}}_{1}\subset(S_{0})_{+}.$
Step 2($p$). Let ${\mathfrak{p}}_{2}$ be generated by $g_{2,1},\ldots,g_{2,r_{2}}\in R_{1}$ and $I_{2}$ be generated by the part $g_{2,1},\ldots,g_{2,c_{2}}$ ($c_{2}:=0ptI_{2}\leq r_{2}$) of them as in the discussion in Proposition 4.6.
As the ideal ${\mathfrak{p}}_{2}\subset R_{1}$ and also $I_{2}$ are generated by finite number of elements in
$R_{1}=\oplus_{\ell\geq 0}R_{0}[I_{1}T_{1}]_{(\ell g_{0},\ell g_{1})}\subset k[x_{1},\ldots,x_{N},T_{1}],$
the coefficients of each monomial in $x_{i}$ and $x_{i}T_{1}$ of the generators of ${\mathfrak{p}}_{2}$ and also $I_{2}$ are finite. Concretely, let $g_{2,1},\ldots,g_{2,r_{2}}$ be generators of ${\mathfrak{p}}_{2}$. Then,
$g_{2,j}=G_{2,j}(x_{0},\ldots,x_{N},x_{1}T_{1},\ldots,x_{N}T_{1}),$
where $G_{2,j}$ is a polynomial $G_{2,j}\in k[x_{0},\ldots,x_{N},y_{1,1},\ldots,y_{N,1}]$. Then, as $G_{2,j}$ has only finite number of non-zero coefficients, we can take a skeleton $\Sigma$ of $k$ containing all those coefficients.
Then, for any skeleton $\Sigma$ of $k$ containing those coefficients, the generators ${\mathfrak{p}}_{2}$ and those of $I_{2}$ are contained in $S_{1}$. Let $\underline{{\mathfrak{p}}}_{2}:={\mathfrak{p}}_{2}\cap S_{1}$ and $\underline{I}_{2}:=I_{2}\cap S_{1}$, and define
$S_{2}:=\oplus_{\ell\geq 0}S_{0}\left[\underline{I}_{1}T_{1},\underline{I}_{2}T_{2}\right]_{(\ell g_{0},\ell g_{1},\ell g_{2})}\ \ \mbox{for}\ \ g_{0}\gg g_{1}\gg g_{2}.$
Then, $S_{2}$ is a skeleton of $R_{2}$. Note that the generators $g_{2,1},\ldots,g_{2,r_{2}}$ of $\underline{{\mathfrak{p}}}_{2}$ has no constant term in $x_{i}$’s and $x_{i}T_{1}$’s, i.e.,
$\underline{{\mathfrak{p}}}_{2}\subset(S_{1})_{+}.$
On the other hand, we note that the center of $E$ is on the exceptional divisor $E_{1}:=\varphi_{1}^{-1}(0)\subset B_{1}$. So we may assume that the defining ideal $I_{E_{1}}$ of $E_{1}$ satisfies
$I_{E_{1}}\subset\underline{{\mathfrak{p}}}_{2}.$
Step $\mathbf{i}$($p$). Assume we obtain a skeleton $S_{i-1}$ of $R_{i-1}$ over a skeleton $\Sigma\subset k$ which contains necessary elements and
$\underline{{\mathfrak{p}}}_{j}\subset(S_{j-1})_{+}$
for any $j\leq i-1$. By using an appropriately enlarged skeleton $\Sigma^{\prime}\supset\Sigma$ of $k$, we replace $S_{i-1}$ by $\Sigma^{\prime}[S_{i-1}]$ so that we may assume that the prime ideal ${\mathfrak{p}}_{i}\subset R_{i-1}$ and $I_{i}\subset R_{i-1}$ are generated by elements of $S_{i-1}$. Indeed, let $g\in{\mathfrak{p}}_{i}\subset R_{i-1}$ be one of the generators of ${\mathfrak{p}}_{i}$. Then, there is a polynomial $G$ with the coefficients in $k$ such that
$g=G\left(x_{0},\ldots,x_{N},g_{1,1}T_{1},\ldots,g_{1,r_{1}}T_{1},\ldots,g_{i-1,1}T_{i-1},\ldots g_{i-1,r_{i-1}}T_{i-1}\right),$
where $g_{j,l}$’s are generators of $\underline{{\mathfrak{p}}}_{j}\subset S_{j-1}$, and the part $g_{j,l}$’s $(l=1,\ldots,c_{j}=0pt{\mathfrak{p}}_{j})$ form a generator system of $\underline{I}_{j}\subset S_{j-1}$ for each $1\leq j\leq i-1$. Let a skeleton $\Sigma^{\prime}\supset\Sigma$ contain all coefficients of $G$, then
$g\in\Sigma^{\prime}[S_{i-1}].$
Doing this procedure for all generators $g_{i,l}$ $(l=1,\ldots,r_{i})$ of ${\mathfrak{p}}_{i}\subset R_{i-1}$, we obtain a skeleton $\Sigma^{\prime}\supset\Sigma$ of $k$ such that
$g_{i,l}\in\Sigma^{\prime}[S_{i-1}],\ (l=1,\ldots,r_{i}).$
Thus, by replacing $S_{i-1}$ by $\Sigma^{\prime}[S_{i-1}]$ with a suitable $\Sigma^{\prime}$, the ideals ${\mathfrak{p}}_{i}\subset R_{i-1}$ and $I_{i}\subset R_{i-1}$ are generated by elements of $S_{i-1}$. Let $\underline{{\mathfrak{p}}}_{i}={\mathfrak{p}}_{i}\cap S_{i-1}$ and $\underline{I}_{i}=I_{i}\cap S_{i-1}.$ Let
$S_{i}=\oplus_{\ell\geq 0}S_{0}\left[\underline{I}_{1}T_{1},\ldots,\underline{I}_{i}T_{i}\right]_{(\ell g_{0},\ldots,\ell g_{i})}\ \ \mbox{for}\ \ g_{0}\gg\cdots\gg g_{i}.$
Then, $S_{i}$ is a skeleton of $R_{i}$ and we have
$\underline{{\mathfrak{p}}}_{i}\subset(S_{i-1})_{+}.$
By doing these procedures until $i=n$, we obtain skeletons $S_{i}$ of $R_{i}$ for all $0\leq i\leq n$. Let $B_{0}:={\mathrm{Proj\ }}S_{0}$ and $B_{i}:={\mathrm{Proj\ }}{S_{i}}$ ($i\geq 1$). Then the canonical sequence
$B_{n}\to B_{n-1}\to\ldots\to B_{1}\to B_{0}$
is recognized as the sequence of blow-ups by the ideals $\underline{I}_{i}$’s. The center of the blow-up $B_{i}\to B_{i-1}$ is denoted by $Z_{i}\subset B_{i-1}$ by the same symbol as $Z_{i}\subset A_{i-1}$. By the construction, we note that the irreducible component $V(\underline{{\mathfrak{p}}}_{i})$ of $Z_{i}$ is contained in $\phi_{i-1}^{-1}(0)$, but any other components are not contained in $\phi_{i-1}^{-1}(0)$. Here, $\phi_{i-1}:B_{i-1}\to B_{0}$ is the composite $\varphi_{1}\circ\cdots\varphi_{i-1}$.
Now we take any skeleton in $k$ with the necessary elements for each step. Fix now such a skeleton $\Sigma$ (Later, for the theorem, we may replace it by a larger one if necessary).
Next we define a characteristic 0 objects inductively on $i$. Let $\widetilde{\Sigma}$$\subset{\mathbb{C}}$ be a compatible skeleton of $\Sigma$ as in Example 2.3, (2), i.e., the equality
${\widetilde{\Sigma}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}=\Sigma$
holds.
Step 0(0). Let
${\widetilde{S}}_{0}:={\widetilde{\Sigma}}[x_{0},\ldots,x_{N}]\subset{\mathbb{C}}[x_{0},\ldots,x_{N}].$
Then, it is a compatible skeleton of $S_{0}$. Denote ${\mathrm{Proj\ }}{\widetilde{S}}_{0}$ by ${\widetilde{B}}_{0}$.
Step 1(0). Let $\underline{\widetilde{I}}_{1}={\underline{\widetilde{\mathfrak{p}}}}_{1}:=(x_{1},\ldots,x_{N})\subset{\widetilde{S}}_{0}$. and
${\widetilde{S}}_{1}:=\oplus_{\ell\geq 0}{\widetilde{S}}_{0}[\underline{\widetilde{I}}_{1}T_{1}]_{(\ell g_{0},\ell g_{1})}.$
Then, ${\widetilde{S}}_{1}$ is a lifting of the skeleton $S_{1}$ and ${\widetilde{B}}_{1}:={\mathrm{Proj\ }}{\widetilde{S}}_{1}$ is the blow-up of ${\widetilde{B}}_{0}$ by the ideal $\underline{\widetilde{I}}_{1}={\underline{\widetilde{\mathfrak{p}}}}_{1}$. Denote the exceptional divisor on ${\widetilde{B}}_{1}$ by ${\widetilde{E}}_{1}$. Then, ${\widetilde{E}}_{1}$ is a lifting of $E_{1}$, because its center ${\widetilde{P}}_{1}$ is a lifting of the center $P_{1}$ of $E_{1}$.
Step 2 (0). As we saw in Step 2 ($p$), the ideal $\underline{{\mathfrak{p}}}_{2}\subset(S_{1})_{+}$ giving the center of $E$ on $B_{1}$ satisfies
(9) $I_{E_{1}}\subset\underline{{\mathfrak{p}}}_{2},$
where $I_{E_{1}}$ is the defining ideal of $E_{1}\subset B_{1}$. On the other hand, in ${\widetilde{B}}_{1}$, we already have the exceptional divisor ${\widetilde{E}}_{1}$ mapped to ${\widetilde{P}}_{1}=0\in{\widetilde{B}}_{0}={\mathrm{Proj\ }}{\widetilde{S}}_{0}$ by ${\widetilde{\varphi}}_{1}$, that is also a lifting of $E_{1}$. On the other hand, let a system $\{g_{2,1},\ldots g_{2,c_{2}}\}\subset\underline{{\mathfrak{p}}}_{2}\subset S_{1}$ be generators of $\underline{I}_{2}$ defining $Z_{2}$, the center of the blow-up $\varphi_{2}$ as in Step 2($p$). This system was taken so that it gives an RSP of $\underline{{\mathfrak{p}}}_{2}$ in Step 2($p$). For each
$g_{2,i}=G_{2,i}(x_{0},\ldots,x_{N},x_{1}T_{1},\ldots,x_{N}T_{1})\in\Sigma[x_{0},\ldots,x_{N},x_{1}T_{1},\ldots,x_{N}T_{1}],$
we have a lifting
${\tilde{g}}_{2,i}=\widetilde{G}_{2,i}(x_{0},\ldots,x_{N},x_{1}T_{1},\ldots,x_{N}T_{1})\in{\widetilde{\Sigma}}[x_{0},\ldots,x_{N},x_{1}T_{1},\ldots,x_{N}T_{1}],$
by lifting the coefficient of $G_{2,i}$ at each monomial in $x_{i}$ $(0\leq i\leq N)$ and $x_{j}T_{1}$ $(1\leq j\leq N)$ such that $\deg{\tilde{g}}_{2,i}=\deg g_{2,i}$. We may assume $\{g_{2,1}\}$ gives an RSP of $I_{E_{1}}$. Then, apply Proposition 2.11, we obtain a lifting $\{{\tilde{g}}_{2,1},\ldots{\tilde{g}}_{2,c_{2}}\}\subset({\widetilde{S}}_{1})_{+}$ of $\{g_{2,1},\ldots g_{2,c_{2}}\}\subset\underline{{\mathfrak{p}}}_{2}\subset S_{1}$, which gives a weak prime lifting ${\underline{\widetilde{\mathfrak{p}}}}_{2}\subset({\widetilde{S}}_{1})_{+}$ with
(10) $I_{{\widetilde{E}}_{1}}\subset{\underline{\widetilde{\mathfrak{p}}}}_{2}.$
Let $\underline{\widetilde{I}}_{2}=({\tilde{g}}_{2,1},\ldots{\tilde{g}}_{2,c_{2}})\subset{\widetilde{S}}_{1}$, then it is a lifting of $\underline{I}_{2}\subset S_{1}$. Let ${\widetilde{Z}}_{2}\subset{\widetilde{B}}_{1}$ be the closed subscheme defined by $\underline{\widetilde{I}}_{2}$. The component $V({\underline{\widetilde{\mathfrak{p}}}}_{2})$ of ${\widetilde{Z}}_{2}$ is contained in ${\widetilde{E}}_{1}$, but every other irreducible component of ${\widetilde{Z}}_{2}$ is not contained in ${\widetilde{E}}_{1}$. Because if such an irreducible component of ${\widetilde{Z}}_{2}$ was contained in ${\widetilde{E}}_{1}$, a component of $Z_{2}\subset B_{1}$ other than $V(\underline{{\mathfrak{p}}}_{2})$ would be contained in $E_{1}$, a contradiction to the definition of $Z_{2}$. Let
${\widetilde{S}}_{2}:=\oplus_{\ell\geq 0}{\widetilde{S}}_{0}\left[\underline{\widetilde{I}}_{1}T_{1},\underline{\widetilde{I}}_{2}T_{2}\right]_{(\ell g_{0},\ell g_{1},\ell g_{2})}\ \ \mbox{for}\ \ g_{0}\gg g_{1}\gg g_{2}.$
Then, ${\widetilde{S}}_{2}$ is a lifting of $S_{2}$ and gives the blow up ${\widetilde{\varphi}}_{2}:{\widetilde{B}}_{2}\to{\widetilde{B}}_{1}$ by the ideal $\underline{\widetilde{I}}_{2}$.
Let ${\widetilde{E}}_{2}\subset{\widetilde{B}}_{2}$ be the irreducible divisor mapped to ${\widetilde{P}}_{2}\in{\widetilde{B}}_{1}$ by ${\widetilde{\varphi}}_{2}$. Then, by (9) and (10), the configuration of the exceptional divisors $\{E_{1},E_{2}\}$ on $B_{2}$ is the same as that of $\{{\widetilde{E}}_{1},{\widetilde{E}}_{2}\}$ on ${\widetilde{B}}_{2}$. Here, note that $E_{1}$ and $E_{2}$ cross normally at the generic point of the intersection, and so do ${\widetilde{E}}_{1}$ and ${\widetilde{E}}_{2}$. Note that $\Phi_{p}:{\widetilde{S}}_{2}\to S_{2}$ is a surjective graded homomorphism. Hence, $({\widetilde{S}}_{2})_{+}\to(S_{2})_{+}$ is surjective.
Step i (0). Let $i\geq 2$. For each $j\leq i-1$, assume we have constructed ${\widetilde{B}}_{j}={\mathrm{Proj\ }}{\widetilde{S}}_{j}$ such that ${\widetilde{S}}_{j}$ is a lifting of the skeleton $S_{j}$ and satisfying the following:
  1. (1)
    The center ${\widetilde{Z}}_{j}\subset{\widetilde{B}}_{j-1}$ is generated by a regular sequence which gives RSP at $P_{j}$, the generic point of $V({\mathfrak{p}}_{j})$. There is a weak prime lifting ${\widetilde{P}}_{j}\in{\widetilde{B}}_{j-1}$ of $P_{j}\in B_{j-1}$, that gives an irreducible component of ${\widetilde{Z}}_{j}$.
  2. (2)
    Let $E_{l}$ be the irreducible exceptional divisor on $B_{l}$ mapped to $P_{l}\in B_{l-1}$. We also denote the proper transform of $E_{l}$ $(l\leq j-1)$ on $B_{j-1}$ by the same symbol as $E_{l}$ $(\subset B_{l})$.
    Now, in the same way, let ${\widetilde{E}}_{l}$ be the irreducible exceptional divisor on ${\widetilde{B}}_{j}$ mapped to ${\widetilde{P}}_{j}\in{\widetilde{B}}_{j-1}$. We also denote the proper transform of ${\widetilde{E}}_{l}$ $(l\leq j-1)$ on ${\widetilde{B}}_{j-1}$ by the same symbol as ${\widetilde{E}}_{l}$ $(\subset{\widetilde{B}}_{l})$.
    Then $\{{\widetilde{E}}_{1},\ldots,{\widetilde{E}}_{j-1},{\widetilde{P}}_{j}\}$ is weak prime liftings of $\{E_{1},\ldots,E_{j-1},P_{j}\}$ with the same configuration around ${\widetilde{P}}_{j}$ and $P_{j}$. That means each ${\widetilde{E}}_{l}$ $(1\leq l\leq)$ or ${\widetilde{P}}_{j}$ is a weak prime lifting of $E_{l}$ or $P_{j}$ and if $E_{l_{1}}\cap\ldots\cap E_{l_{s}}\ni P_{j}$, then ${\widetilde{E}}_{l_{1}}\cap\ldots\cap{\widetilde{E}}_{l_{s}}$ contains ${\widetilde{P}}_{j}$ and it is a weak prime lifting of $E_{l_{1}}\cap\ldots\cap E_{l_{s}}$.
Under these induction hypothesis for $j\leq i-1$ we will construct ${\widetilde{B}}_{i}={\mathrm{Proj\ }}{\widetilde{S}}_{i}$ as the blow-up of ${\widetilde{B}}_{i-1}$ by an appropriate center ${\widetilde{Z}}_{i}\subset{\widetilde{B}}_{i-1}$ with the properties (1) and (2) for ${\widetilde{B}}_{i}$.
As the point $P_{i}$ is a regular point of $A_{i-1}$, and that $E_{i_{1}},\ldots,E_{i_{s}}$ intersect normally at $P_{i}$, we may assume that the intersection $E_{i_{1}}\cap\ldots\cap E_{i_{s}}\subset B_{i-1}$ is irreducible in a neighborhood of $P_{i}$. (We are using the same symbols $P_{i}$ for a point of $A_{i-1}$ and of $B_{i-1}$, as well as $E_{l}$ for a divisor on $A_{i}$ and on $B_{i}$.)
Let $\{g_{i,1},\ldots g_{i,c_{i}}\}$ give an RSP of $\underline{{\mathfrak{p}}}_{i}\subset S_{i-1}$ and generate $\underline{I}_{i}\subset S_{i-1}$ that defines the center $Z_{i}$ of the blow-up $\varphi_{i}:B_{i}\to B_{i-1}$. Note that as the prime ideal $I_{E_{i_{1}}\cap\ldots\cap E_{i_{s}}}$ is contained in $\underline{{\mathfrak{p}}}_{i}$ and $I_{E_{i_{1}}\cap\ldots\cap E_{i_{s}}}$ has a weak prime lifting $I_{{\widetilde{E}}_{i_{1}},\ldots,{\widetilde{E}}_{i_{s}}}$, we can apply Proposition 2.11. Then, there ia a lifting $\{{\tilde{g}}_{i,1},\ldots{\tilde{g}}_{i,c_{i}}\}\subset{\widetilde{S}}_{i-1}$ of $\{g_{i,1},\ldots g_{i,c_{i}}\}$ which gives an RSP of a prime ideal $\underline{{\widetilde{{\mathfrak{p}}}}}_{i}\subset({\widetilde{S}}_{i-1})_{+}$ such that
$I_{{\widetilde{E}}_{i_{1}}\cap\ldots\cap{\widetilde{E}}_{i_{s}}}\subset\underline{{\widetilde{{\mathfrak{p}}}}}_{i}.$
Define
$\underline{\widetilde{I}}_{i}=({\tilde{g}}_{i1},\ldots{\tilde{g}}_{ic_{i}}).$
Then $\{{\tilde{g}}_{i1},\ldots{\tilde{g}}_{ic_{i}}\}$ is a regular sequence in ${\widetilde{S}}_{i-1}$, because $\{g_{i1},\ldots g_{ic_{i}}\}$ is a regular sequence in $S_{i-1}={\widetilde{S}}_{i-1}\otimes{{\mathbb{Z}}/(p)}$. Therefore it defines a complete intersection subscheme ${\widetilde{Z}}_{i}\subset{\widetilde{B}}_{i-1}$ in ${\widetilde{B}}_{i-1}$, that is a lifting of $Z_{i}$ and has an irreducible component corresponding to ${\widetilde{P}}_{i}\in{\widetilde{B}}_{i-1}$.
Define
${\widetilde{S}}_{i}:=\oplus_{\ell\geq 0}{\widetilde{S}}_{0}\left[\underline{\widetilde{I}}_{1}T_{1},\ldots,\underline{\widetilde{I}}_{i}T_{i}\right]_{(\ell g_{0},\ldots,\ell g_{i})}.$
Then, ${\widetilde{S}}_{i}$ is a lifting of the skeleton $S_{i}$ and the canonical inclusion ${\widetilde{S}}_{i-1}\hookrightarrow{\widetilde{S}}_{i}$ gives the blow-up $\widetilde{\varphi}_{i}:{\widetilde{B}}_{i}\to{\widetilde{B}}_{i-1}$ by $\underline{\widetilde{I}}_{i}$ and obtain the exceptional divisor ${\widetilde{E}}_{i}\subset{\widetilde{B}}_{i}$ dominating ${\widetilde{P}}_{i}\in{\widetilde{E}}_{i-1}$.
Geometrically, if the center $P_{i}\in B_{i-1}$ of $E$ is contained in $E_{i_{1}},\ldots,E_{i_{s}}$, we have a weak prime lifting $\widetilde{P}_{i}\in{\widetilde{B}}_{i-1}$ such that
$\widetilde{P}_{i}\in{\widetilde{E}}_{i_{1}}\cap\ldots\cap{\widetilde{E}}_{i_{s}}.$
By carrying out each step up to $n$, we obtain the sequence:
(11) ${\widetilde{B}}_{n}\to\ldots\to{\widetilde{B}}_{0}$
of blow-ups by the ideal $\underline{\widetilde{I}}_{i}\in{\widetilde{B}}_{i-1}$ such that
(12) ${\widetilde{P}}_{i}\in{\widetilde{E}}_{j}\ \ \mbox{iff}\ \ P_{i}\in E_{j}\ \ \mbox{ for}\ \ j\leq i-1.$
On the other hand, by (i) of Lemma 2.13, we have
(13) $0pt{\mathfrak{p}}_{i}=0pt\underline{{\mathfrak{p}}}_{i},\ \mbox{and}\ \ 0pt{\underline{\widetilde{\mathfrak{p}}}}_{i}=0pt({\underline{\widetilde{\mathfrak{p}}}}_{i}\otimes_{\widetilde{\Sigma}}L),$
[…]
As ${\underline{\widetilde{\mathfrak{p}}}}_{i}$ is a weak prime lifting of $\underline{{\mathfrak{p}}}_{i}$, we obtain:
(14) $0pt\underline{{\mathfrak{p}}}_{i}=0pt{\underline{\widetilde{\mathfrak{p}}}}_{i}.$
By (13) and (14), it follows:
(15) $0pt{\mathfrak{p}}_{i}=0pt({\underline{\widetilde{\mathfrak{p}}}}_{i}\otimes_{\widetilde{\Sigma}}L),$
Let ${\widetilde{A}}_{i}:={\widetilde{B}}_{i}\otimes_{\widetilde{\Sigma}}L$, then ${\widetilde{A}}_{i}({\operatorname{mod}\ }p)=A_{i}$ with the compatible skeletons ${\widetilde{S}}_{i}$ and $S_{i}$, and it follows the sequence of blow-ups:
${\widetilde{A}}_{n}\stackrel{{\scriptstyle\widetilde{\varphi}_{n}}}{{\longrightarrow}}{\widetilde{A}}_{n-1}\to\cdots{\widetilde{A}}_{1}\stackrel{{\scriptstyle\widetilde{\varphi}_{1}}}{{\longrightarrow}}{\widetilde{A}}={\mathbb{A}}_{L}^{N}.$
We denote the exceptional divisor on ${\widetilde{A}}_{i}$ corresponding to ${\widetilde{E}}_{i}\subset{\widetilde{B}}_{i}$ by the same symbol ${\widetilde{E}}_{i}$. Then, (12) and (15) yields $k_{{\widetilde{E}}_{i}}=k_{E_{i}}$, because
$k_{E_{i}}=0pt{\mathfrak{p}}_{i}-1+\sum_{P_{i}\in E_{j}}k_{E_{j}}\ \ \mbox{and}$
$k_{{\widetilde{E}}_{i}}=0pt({\underline{\widetilde{\mathfrak{p}}}}_{i}\otimes_{\widetilde{\Sigma}}L)-1+\sum_{{\widetilde{P}}_{i}\in{\widetilde{E}}_{j}}k_{{\widetilde{E}}_{j}}.$
Here, we note that the point ${\widetilde{P}}_{i}\in{\widetilde{A}}_{i-1}$ corresponding to the lifting ${\underline{\widetilde{\mathfrak{p}}}}_{i}$ is non-singular, as the local ring ${\mathcal{O}}_{{\widetilde{A}}_{i-1},{\widetilde{P}}_{i}}$ has the maximal ideal generated by $c_{i}$ elements, where $c_{i}=0pt{\underline{\widetilde{\mathfrak{p}}}}_{i}=0pt\underline{{\mathfrak{p}}}_{i}$. Indeed, exactly $c_{i}$ elements from the generators of $\underline{{\mathfrak{p}}}_{i}$ give the regular system of parameters of the local ring ${\mathcal{O}}_{A_{i-1},P_{i}}$. Then, the lifting of these $c_{i}$ generators give the regular sequence defining the maximal ideal of the local ring ${\mathcal{O}}_{{\widetilde{A}}_{i-1},{\widetilde{P}}_{i}}$. ∎
Corollary 4.9.
Under the notation of Proposition 4.8, let ${\widetilde{C}}=V({\widetilde{{\mathfrak{p}}}}_{i})$ be given by the weak prime lifting ${\widetilde{{\mathfrak{p}}}}_{i}$ of ${\mathfrak{p}}_{i}$ constructed in the proof of Proposition 4.8. Let $C:=V({\mathfrak{p}}_{i})$. Then, scheme theoretically
${\widetilde{C}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}=C\cup C^{\prime},$
where $C^{\prime}$ is either empty or the union of embedded components in $C$.
Proof.
Denote the point in $B_{0}$ defined by $(x_{1},\ldots,x_{N})\subset S_{0}=\Sigma[x_{1},\ldots,x_{N}]$ by 0, and the point in ${\widetilde{B}}_{0}$ defined by $(x_{1},\ldots,x_{N})\subset{\widetilde{S}}_{0}={\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]$ by $\widetilde{0}$. Then, set theoretically we have an inclusion
${\widetilde{\phi}_{i-1}^{-1}(\widetilde{0})}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}\subset\phi_{i-1}^{-1}(0),$
[…]
${\widetilde{C}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}=C\cup C^{\prime}$
for some $C^{\prime}\subset Z_{i}\cap\phi_{i-1}^{-1}(0)$. Here, by the construction of $Z_{i}$ ((iii) in Proposition 4.6), we know
${\mathrm{codim}}(Z_{i}\cap(\phi_{i-1}^{-1}(0)\setminus C),B_{i-1})>{\mathrm{codim}}(C,B_{i-1}).$
Therefore, if $C^{\prime}\not\subset C$ set theoretically, then ${\mathrm{codim}}({C^{\prime}},B_{i-1})>{\mathrm{codim}}(C,B_{i-1})$ and $C\cup C^{\prime}$ is the fiber of $(p)\in{\mathrm{Spec\ }}{\mathbb{Z}}$ by the flat morphism ${\widetilde{C}}\to{\mathrm{Spec\ }}{\mathbb{Z}}$, which is a contradiction. This shows that $C^{\prime}\subset C$ set theoretically.
Remark 4.10.
In the above proposition, we focus on the exceptional divisor $E_{i}$ mapped to $P_{i}$ by $\varphi_{i}$ and ${\widetilde{E}}_{i}$ mapped to ${\widetilde{P}}_{i}$ by ${\widetilde{\varphi}}_{i}$, respectively. However we should not forget that there are some other irreducible components of $\varphi_{i}^{-1}(Z_{i})$ and ${\widetilde{\varphi}}_{i}^{-1}({\widetilde{Z}}_{i})$.
Remark 4.11.
Under the notation of Proposition 4.6, the reader may expect that for an ideal ${\mathfrak{a}}\subset k[x_{1},\ldots,x_{N}]$, there exists a lifting ${\widetilde{\mathfrak{a}}}\subset L[x_{1},\ldots,x_{N}]$ such that
[…]
It is a reasonable and preferable expectation. In some simple cases, we can verify it (see Proposition 4.12 below). However, at present, we cannot construct a lifting $\widetilde{\mathfrak{a}}$} satisfying (16) for $E_{n}$ and ${\widetilde{E}}_{n}$. A compromised statement is our main theorem (Theorem 1.1) which will be proved in the next section. The following are simple examples that the equality (16) holds for an appropriate lifting $\widetilde{\mathfrak{a}}$} of $\mathfrak{a}$}.
Proposition 4.12.
Under the notation of the statement in Proposition 4.6, the following hold:
  1. (i)
[…]
  2. (ii)
    Let $E_{2}\subset A_{2}$ be a prime divisor obtained by the blow-up at a $k$-valued closed point $P_{2}\in E_{1}\subset A_{1}$ and ${\mathfrak{a}}\subset k[x_{1},\ldots,x_{N}]$ a nonzero ideal. Then, the following hold:
    1. (a)
[…]
    2. (b)
      By choosing compatible skeletons $\Sigma\subset k$ and $\widetilde{\Sigma}$$\subset{\mathbb{C}}$, construct the skeleton schemes $B_{i}$ and ${\widetilde{B}}_{i}$ for $i=1,2$ as in the proof of Proposition 4.6. Denote the canonical morphism $B_{2}\to{\mathbb{A}}_{\Sigma}^{N}$ and ${\widetilde{B}}_{2}\to{\mathbb{A}}_{\widetilde{\Sigma}}^{N}$ by $\psi$ and $\widetilde{\psi}$, respectively. Then, the canonical map
      $\widetilde{\psi}_{*}({\mathcal{O}}_{{\widetilde{B}}_{2}}(-m{\widetilde{E}}_{2}))\to\psi_{*}({\mathcal{O}}_{B_{2}}(-mE_{2}))$
      is surjective for every $m\in{\mathbb{Z}}_{\geq 0}$.
    3. (c)
      For every pair $m<t$ of positive integers the canonical map is surjective:
      $\widetilde{\psi}_{*}\left(({\mathcal{O}}_{{\widetilde{B}}_{2}}(-m{\widetilde{E}}_{2}))/{\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{2}}\right)\to\psi_{*}\left(({\mathcal{O}}_{B_{2}}(-mE_{2}))/{\mathfrak{m}}^{t}{\mathcal{O}}_{B_{2}}\right),$
      where ${\mathfrak{m}}=(x_{1},\ldots,x_{N})\subset\Sigma[x_{1},\ldots,x_{N}]$ and ${\widetilde{\mathfrak{m}}}=(x_{1},\ldots,x_{N})\subset{\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]$.
Proof.
[…]
For (i), take appropriate compatible skeletons $\Sigma$ and $\widetilde{\Sigma}$} so that $f\in\Sigma[x_{1},\ldots,x_{N}]$. For
$f=\sum_{i\in{\mathbb{Z}}^{N}_{\geq 0}}a_{i}{\mathbb{x}}^{i}\ \ \ ,\ \ \mbox{take}\ \ \ {\tilde{f}}=\sum_{i\in{\mathbb{Z}}^{N}_{\geq 0}}{\widetilde{a}}_{i}\mathbb{x}^{i},$
where ${\widetilde{a}}_{i}\in{\widetilde{\Sigma}}$ is a lifting of $a_{i}\in\Sigma$ such that ${\widetilde{a}}_{i}=0$ if $a_{i}=0$. Then we can see that ${\mathrm{ord}}_{\mathbb{x}}f={\mathrm{ord}}_{\mathbb{x}}{\tilde{f}}$, which implies
[…]
For (ii), we may assume that locally the following hold as $P_{2}$ is a $k$-valued point:
$(B_{1},P_{2})=\left({\mathrm{Spec\ }}\Sigma\left[x_{1},\frac{x_{2}}{x_{1}},\ldots,\frac{x_{N}}{x_{1}}\right],\left(x_{1},\frac{x_{2}}{x_{1}},\ldots,\frac{x_{N}}{x_{1}}\right)\right)$
and
$({\widetilde{B}}_{1},{\widetilde{P}}_{2})=\left({\mathrm{Spec\ }}{\widetilde{\Sigma}}\left[x_{1},\frac{x_{2}}{x_{1}},\ldots,\frac{x_{N}}{x_{1}}\right],\left(x_{1},\frac{x_{2}}{x_{1}},\ldots,\frac{x_{N}}{x_{1}}\right)\right),$
for an appropriate compatible skeletons $\Sigma\subset k$ and $\widetilde{\Sigma}$$\subset{\mathbb{C}}$.
For $f\in\Sigma[x_{1},\ldots,x_{N}]$, if ${\mathrm{ord}}_{\mathbb{x}}f=m_{1}$, then $f$ factors as follows:
[…]
on $B_{1}$. If
$g=\sum_{e}a_{e}x_{1}^{e_{1}}\left(\frac{x_{2}}{x_{1}}\right)^{e_{2}}\ldots\left(\frac{x_{N}}{x_{1}}\right)^{e_{N}},$
where $a_{e}\in\Sigma$ and $e=(e_{1},\ldots,e_{N})$, define
${\tilde{g}}=\sum_{e}{\widetilde{a}}_{e}x_{1}^{e_{1}}\left(\frac{x_{2}}{x_{1}}\right)^{e_{2}}\ldots\left(\frac{x_{N}}{x_{1}}\right)^{e_{N}},$
by using a lifting ${\widetilde{a}}_{e}\in{\widetilde{\Sigma}}$ of $a_{e}$ such that ${\widetilde{a}}_{e}=0$ if $a_{e}=0$. Then, ${\mathrm{mult}}_{P_{2}}g={\mathrm{mult}}_{{\widetilde{P}}_{2}}{\tilde{g}}$, which we denote by $m_{2}$. Therefore, the lifting ${\tilde{f}}:=x_{1}^{m_{1}}{\tilde{g}}$ of $f$ satisfies
[…]
which yields (ii), (a).
For the proof of (ii),(b), note that for every polynomial $f\in\Sigma[x_{1},\ldots,x_{N}]$ with $v_{E_{2}}(f)=m$ ($m\in{\mathbb{Z}}_{\geq 0}$) has a lifting $\tilde{f}$} such that $v_{{\widetilde{E}}_{2}}({\tilde{f}})=m$ as is constructed above. This yields that the modulo $p$ reduction $\widetilde{\psi}_{*}({\mathcal{O}}_{{\widetilde{B}}_{2}}(-m{\widetilde{E}}_{2}))\to\psi_{*}({\mathcal{O}}_{B_{2}}(-mE_{2}))$ is surjective for $m\in{\mathbb{Z}}_{\geq 0}$.
For the proof of (ii), (c), in the construction $\tilde{f}$} from $f$, by the expression of $g$ and $\tilde{g}$}, we have
$\deg_{x}f=\max\{e_{1}+m_{1}-(e_{2}+\cdots+e_{N})\mid a_{e}\neq 0\}=\deg_{x}{\tilde{f}}$.
[…]
5. Proof of the Main Theorem and applications
We fix the notation used in this section. Let $k$ be an algebraically closed field of characteristic $p>0$. Let $A={\mathbb{A}}_{k}^{N}$ be the affine space over $k$ of dimension $N\geq 2$ and 0 the origin of $A$.
Let $E$ be a prime divisor over $(A,0)$ and
$A_{n}\stackrel{{\scriptstyle\varphi_{n}}}{{\longrightarrow}}A_{n-1}\to\cdots A_{1}\stackrel{{\scriptstyle\varphi_{1}}}{{\longrightarrow}}A$
be a normal Gorenstein Zariski sequence for $E$. Proposition 4.8 guarantees the existence of a reasonable lifting of the sequence:
${\widetilde{A}}_{n}\stackrel{{\scriptstyle\widetilde{\varphi}_{n}}}{{\longrightarrow}}{\widetilde{A}}_{n-1}\to\cdots{\widetilde{A}}_{1}\stackrel{{\scriptstyle\widetilde{\varphi}_{1}}}{{\longrightarrow}}{\widetilde{A}}={\mathbb{A}}_{L}^{N},$
For a technical reason, we will blow up two more times from $A_{n}$ and lift them to the blow-ups from ${\widetilde{A}}_{n}$, which will lead to the main statement in this paper. Here, focusing at the part of the extra two blow-ups, we see that an ideal in positive characteristic is lifted to that in characteristic 0 with the preferable property (see, Proposition 4.12, (ii)). We will reduce our problem into the two blow-ups case in Proposition 4.12, (ii). In the following discussions, we will use the notation in Proposition 4.8.
Lemma 5.1.
Let $E$ be a prime divisor over $(A,0)$, and
$A_{n}\stackrel{{\scriptstyle\varphi_{n}}}{{\longrightarrow}}A_{n-1}\to\cdots A_{1}\stackrel{{\scriptstyle\varphi_{1}}}{{\longrightarrow}}A$
be a normal Gorenstein Zariski sequence for $E$, We use the same notation as in Proposition 4.8. Let
$A_{n+2}\stackrel{{\scriptstyle\varphi_{n+2}}}{{\longrightarrow}}A_{n+1}\stackrel{{\scriptstyle\varphi_{n+1}}}{{\longrightarrow}}A_{n}$
be the composite of two blow-ups at general closed points $P_{n+1}\in E_{n}\subset A_{n}$ and $P_{n+2}\in F_{1}\subset A_{n+1}$, where $F_{1}$ is the exceptional divisor for $\varphi_{n+1}$. Denote the exceptional divisor for $\varphi_{n+2}$ by $F_{2}$.
Then, by taking a skeletons $\Sigma\subset k$ sufficiently large, and letting $\widetilde{\Sigma}$$\subset{\mathbb{C}}$ be its lifting, one obtains liftings ${\widetilde{P}}_{n+1}\in{\widetilde{A}}_{n}$ and ${\widetilde{P}}_{n+2}\in{\widetilde{A}}_{n+1}$ of $P_{n}$ and $P_{n+1}$, respectively, and the sequence:
${\widetilde{A}}_{n+2}\stackrel{{\scriptstyle\widetilde{\varphi}_{n+2}}}{{\longrightarrow}}{\widetilde{A}}_{n+1}\stackrel{{\scriptstyle\widetilde{\varphi}_{n+1}}}{{\longrightarrow}}{\widetilde{A}}_{n},$
the composite of two blow-ups at ${\widetilde{P}}_{n+1}\in{\widetilde{E}}_{n}\subset{\widetilde{A}}_{n}$ and ${\widetilde{P}}_{n+2}\in{\widetilde{F}}_{1}\subset{\widetilde{A}}_{n+1}$, where ${\widetilde{F}}_{1}$ and ${\widetilde{F}}_{2}$ are the exceptional divisor of $\widetilde{\varphi}_{n+1}$ and $\widetilde{\varphi}_{n+2}$, respectively.
Then, we obtain
$k_{F_{1}}=k_{{\widetilde{F}}_{1}}=N-1+k_{E_{n}}=N-1+k_{{\widetilde{E}}_{n}}\ \mbox{and}$
$k_{F_{2}}=k_{{\widetilde{F}}_{2}}=2(N-1)+k_{E_{n}}=2(N-1)+k_{{\widetilde{E}}_{n}}.$
Proof.
Take a general closed point $P_{n+1}\in E_{n}\subset A_{n}$ such that
$P_{n+1}\not\in\bigcup_{j<n}(E_{j}\cup E^{\prime}_{j}),$
where we denote the proper transform of $E_{j}$ and $E^{\prime}_{j}$ on $A_{n}$ by the same symbols as $E_{j}$ and $E^{\prime}_{j}\subset A_{j}$. Here, we denote the fiber $\varphi_{j}^{-1}(Z_{j})=E_{j}+E^{\prime}_{j}$ as Cartier divisors on $A_{j}$.
Next, take a general closed point $P_{n+2}\in F_{1}\subset A_{n+1}$ such that
$P_{n+2}\not\in E_{n},$
where $E_{n}$ denotes the proper transform of $E_{n}$ on $A_{n+1}$ as well. Here, we note that $P_{n+1}$ and $P_{n+2}$ are $k$-valued point as $k$ is algebraically closed. So, we can realize their defining ideals in the skeletons $S_{n}$ and $S_{n+1}$ for sufficiently large $\Sigma\subset k$. Then, in the same way as in Proposition 4.6, we can construct a sequence
${\widetilde{A}}_{n+2}\stackrel{{\scriptstyle\widetilde{\varphi}_{n+2}}}{{\longrightarrow}}{\widetilde{A}}_{n+1}\stackrel{{\scriptstyle\widetilde{\varphi}_{n+1}}}{{\longrightarrow}}{\widetilde{A}}_{n},$
of blow-ups with the centers ${\widetilde{P}}_{n+1}\in{\widetilde{E}}_{n}$ and ${\widetilde{P}}_{n+2}\in{\widetilde{F}}_{1}$ of liftings of $P_{n+1}$ and $P_{n+2}$, respectively. Hence, by the choice of the points of blow-ups, we obtain the required equalities. ∎
Lemma 5.2.
Under the notation as in the beginning of this section and in Lemma 5.1, consider the ideals
$\underline{\widetilde{I}}_{1}={\widetilde{\mathfrak{m}}}=(x_{1},\ldots,x_{N})\subset{\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]={\widetilde{S}}_{0}\ \ \mbox{and}$
$\underline{I}_{1}={\mathfrak{m}}=(x_{1},\ldots,x_{N})\subset\Sigma[x_{1},\ldots,x_{N}]=S_{0}.$
Then, the following hold for $1\leq i\leq n+2$:
  1. (i)
    ${\mathcal{O}}_{{\widetilde{B}}_{i}}/{\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{i}}$ is $\mathbb{Z}$}-flat at $p\in{\mathbb{Z}}$ for every $t\in{\mathbb{N}}$, and
  2. (ii)
    ${\mathcal{O}}_{{\widetilde{B}}_{i}}/{\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{i}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}={\mathcal{O}}_{B_{i}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{B_{i}}$.
Proof.
For $1\leq i\leq n+2$, denote the proper transforms on $B_{i}$ of the prime divisors $E_{j}\subset B_{j}$ and ${\widetilde{E}}_{j}\subset{\widetilde{B}}_{j}$ $(i<j)$ by the same symbols $E_{j}$ and ${\widetilde{E}}_{j}$, as we did in the previous discussions. (For consistency of the notation, only in this proof, we denote $F_{i}$ and ${\widetilde{F}}_{i}$ $(i=1,2)$ by $E_{n+i}$ and ${\widetilde{E}}_{n+i}$, respectively.) As $Z_{i}$ ($i\geq 2$) is not necessarily irreducible, the irreducible component of the exceptional divisor of $\varphi_{i}:B_{i}\to B_{i-1}$ is not only $E_{i}$.
As $\widetilde{\varphi}_{1}$ is the blow-up by $\widetilde{\mathfrak{m}}=(x_{1},\ldots,x_{N})\subset{\widetilde{S}}_{0}$, the pull-back ${\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{1}}$ is an invertible ideal sheaf on ${\widetilde{B}}_{1}$. Then, for every $i\geq 1$ its pull-back ${\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{i}}$ on ${\widetilde{B}}_{i}$ is also an invertible ideal sheaf on ${\widetilde{B}}_{i}$. By the similar reason, ${\mathfrak{m}}^{t}{\mathcal{O}}_{B_{i}}$ is also an invertible ideal sheaf for every $i\geq 1$.
For the proof of (i), let ${\widetilde{D}}_{i}\subset{\widetilde{B}}_{i}$ be the subscheme defined by ${\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{i}}$, and $D_{i}\subset B_{i}$ the subscheme defined by ${\mathfrak{m}}^{t}{\mathcal{O}}_{B_{i}}$ for $1\leq i\leq n+2$. Then, ${\widetilde{D}}_{i}\subset{\widetilde{B}}_{i}$ is locally a hypersurface and its support is contained in the exceptional divisor of $\widetilde{\phi}_{i}:{\widetilde{B}}_{i}\to{\widetilde{B}}_{0}$. Since the exceptional divisor of $\widetilde{\phi}_{i}$ does not contain $B_{i}$, the closed subscheme ${\widetilde{D}}_{i}$ does not contain the fiber $B_{i}$ over $p$, which yields that ${\widetilde{D}}_{i}$ is flat over $\mathbb{Z}$} at $p$ (see, [16, Corollary, p.177]). This completes the proof of (i).
For the proof of (ii), we first claim that the supports of $D_{i}$ and ${\widetilde{D}}_{i}$ are expressed as follows:
(17) ${\mathrm{supp}}D_{i}=\bigcup_{j\leq i}E_{j},\ \mbox{and}\ \ {\mathrm{supp}}{\widetilde{D}}_{i}=\bigcup_{j\leq i}{\widetilde{E}}_{j}\ \ \mbox{in\ a\ neighborhood\ of}\ p.$
Here, we remark again that the irreducible components of the exceptional set of $\widetilde{\phi}_{i}$ are not only ${\widetilde{E}}_{1},\ldots,{\widetilde{E}}_{i}$, because each center of the blow-up ${\widetilde{\varphi}}_{j}$ $(j\leq i)$ is not necessarily irreducible. We will prove the equalities in (17) by induction on $i$.
For $i=1$ the equalities are trivial. So we assume $i\geq 2$ and the equalities hold for i1i-1. Now, assume there exists an irreducible component $C$ of $D_{i}$ such that
$C\not\subset\bigcup_{j\leq i}E_{j}.$
Then, obviously $\varphi_{i}(C)\subset D_{i-1}$. Therefore, $\varphi_{i}(C)\subset Z_{i}$, because otherwise, $\varphi_{i}$ is isomorphic at the generic point of $C$ and by the hypothesis of the induction, we have $\varphi_{i}(C)\subset E_{j}\subset B_{i-1}$ for some $j<i$, which gives $C\subset E_{j}\subset B_{i}$, a contradiction to the assumption. Hence, we obtain
$\varphi_{i}(C)\subset Z_{i}\cap D_{i-1}\subset B_{i-1}\ \mbox{ and}\ \ \varphi_{i}(C)\not\subset V({\mathfrak{p}}_{i}).$
By the construction of $Z_{i}$ (see, Proposition 4.6), it follows
${\mathrm{codim}}(Z_{i}\cap D_{i-1}\setminus V(\underline{{\mathfrak{p}}}_{i}),B_{i-1})=c_{i}+1:=0pt{\mathfrak{p}}_{i}+1,$
which yields
${\mathrm{codim}}(C,B_{i})\geq{\mathrm{codim}}(\varphi_{i}^{-1}(\varphi_{i}(C)),B_{i})\geq{\mathrm{codim}}(\varphi_{i}(C),B_{i-1})-(c_{i}-1)$
$\geq{\mathrm{codim}}(Z_{i}\cap D_{i-1}\setminus V(\underline{{\mathfrak{p}}}_{i}),B_{i-1})-(c_{i}-1)=2,$
a contradiction to that $C$ is an irreducible component of a closed subscheme defined by a locally principal ideal on an irreducible reduced scheme $B_{i}$. Therefore we obtain
${\mathrm{supp}}D_{i}=\bigcup_{j\leq i}E_{j}.$
On the other hand, we have
${\mathrm{codim}}({\widetilde{Z}}_{i}\cap{\widetilde{D}}_{i-1}\setminus V({\underline{\widetilde{\mathfrak{p}}}}_{i}),{\widetilde{B}}_{i-1})\geq{\mathrm{codim}}(Z_{i}\cap D_{i-1}\setminus V(\underline{{\mathfrak{p}}}_{i}),B_{i-1})=c_{i}+1$
in a neighborhood of $p\in{\mathbb{Z}}$. Then, by the similar discussion as above, we obtain
${\mathrm{supp}}{\widetilde{D}}_{i}=\bigcup_{j\leq i}{\widetilde{E}}_{i}.$
In order to prove (ii), it is sufficient to show the equality
${\mathfrak{m}}^{t}{\mathcal{O}}_{B_{i}}={\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{i}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}$
[…]
${\mathfrak{m}}^{t}{\mathcal{O}}_{B_{iQ}}={\mathcal{O}}_{B_{iQ}}\left(-\sum_{j\leq i}a_{j}E_{j}\right),$
${\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{A}}_{i}}={\mathcal{O}}_{{\widetilde{A}}_{i}}\left(-\sum_{j\leq i}{\widetilde{a}}_{j}{\widetilde{E}}_{j}\right).$
As $\left({\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{i}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}\right)\otimes_{\Sigma}Q(\Sigma)={\mathcal{O}}_{B_{iQ}}(-\sum_{j\leq i}{\widetilde{a}}_{j}E_{j})$, for the proof of (ii), it is sufficient to show $a_{j}={\widetilde{a}}_{j}$ ($j\leq i$). We will prove this by induction on $i$.
For $i=1$, it is clear that $a_{1}={\widetilde{a}}_{1}$. Assume $i\geq 2$ and $a_{j}={\widetilde{a}}_{j}$ holds for $j\leq i-1$. By the construction, the center $P_{i}\in B_{i-1Q}$ of the blow-up $\varphi_{i}$ and the center ${\widetilde{P}}_{i}\in{\widetilde{A}}_{i-1}$ of the blow-up ${\widetilde{\varphi}}_{i}$ satisfies
$P_{i}\in E_{j}$ if and only if ${\widetilde{P}}_{i}\in{\widetilde{E}}_{j}$ for $j\leq i$.
$a_{i}=\sum_{P_{i}\in E_{j}}a_{j}=\sum_{{\widetilde{P}}_{i}\in{\widetilde{E}}_{j}}{\widetilde{a}}_{j}={\widetilde{a}}_{i},$
where the first and the last equalities follows from the fact that $P_{i}\in A_{i-1}$ and ${\widetilde{P}}_{i}\in{\widetilde{A}}_{i-1}$ are non-singular (see the last part of the proof of Proposition 4.8). Hence, the statement holds for $i$. ∎
Theorem 5.3 (Theorem 1.1, (i), (ii)).
Let $N$ be an integer $\geq 2$, and $k$ an algebraically closed field of characteristic $p>0$. Let $A={\mathbb{A}}_{k}^{N}$ be the affine space over $k$ of dimension $N$, and 0 the origin of $A$. Let $E$ be a prime divisor over $A$ with the center at 0, and ${\mathfrak{a}}\subset k[x_{1},\ldots,x_{N}]$ a non-zero ideal. Then, there exists a prime divisor $F_{\mathbb{C}}$ over the affine space $A_{\mathbb{C}}={\mathbb{A}}_{\mathbb{C}}^{N}$ over $\mathbb{C}$} with the center at the origin $\tilde{0}$} satisfying the following:
  1. (i)
    $2(N-1)+k_{E}=k_{F_{\mathbb{C}}}$;
  2. (ii)
    there exists a lifting ${\widetilde{\mathfrak{a}}}\subset{\mathcal{O}}_{A_{\mathbb{C}}}$ of $\mathfrak{a}$} satisfying
[…]
Proof.
[…]
Let $F_{\mathbb{C}}$ be one of the irreducible components of ${\widetilde{F}}_{2}\times_{{\mathbb{A}}_{L}^{N}}{\mathbb{C}}$ on ${\widetilde{A}}_{n+2}\times_{{\mathbb{A}}_{L}^{N}}{\mathbb{C}}$. Then, as $k_{{\widetilde{F}}_{2}}=k_{F_{\mathbb{C}}}$, the statement (i) follows from Lemma 5.1.
We prove the statements (ii) for $N\geq 4$ first, and then, applying Lemma 5.4, we will obtain the results for $N\leq 3$.
Now, let us assume $N\geq 4$. First remember the definition of the blow-up centers. In Lemma 5.1, we took the closed point $P_{n+1}\in A_{n}$ general. We may assume
$P_{n+1}\not\in Z({\mathfrak{a}}_{A_{n}})\cup\bigcup_{j<n}E_{j},$
where ${\mathfrak{a}}_{A_{n}}$ is the weak transform of the ideal ${\mathfrak{a}}\subset R_{0}$ on $A_{n}$ and $E_{j}$ $(j<n)$ denotes the proper transform of $E_{j}$ on $A_{n}$. (For the definition of weak transform, see, eg. [13]. The important property we use is only that the locus of the weak transform does not contain the whole $E_{n}$.) We take $P_{n+2}\in A_{n+1}$ as in Lemma 5.1.
By the choice of $P_{n+1}$ and $P_{n+2}$, it follows
$v_{E_{n}}({\mathfrak{a}})=v_{F_{2}}({\mathfrak{a}}).$
Denote the divisors on $B_{n}$ and on $B_{n+j}$ $(j=1,2)$ corresponding to $E_{n}$ and $F_{j}$ by the same symbols, and the ideal in the skeleton $S_{0}$ generating $\mathfrak{a}$} is denoted by $\underline{{\mathfrak{a}}}$. Then, by Lemma 2.13, (ii), we have
$v_{E_{n}}(\underline{{\mathfrak{a}}})=v_{F_{2}}(\underline{{\mathfrak{a}}}).$
Therefore, it is sufficient to show that for a nonzero ideal ${\mathfrak{a}}\subset k[x_{1},\ldots,x_{N}]$ there exist compatible skeletons $\Sigma\subset k$ and $\widetilde{\Sigma}$$\subset{\mathbb{C}}$ such that $\mathfrak{a}$} is generated by $\underline{{\mathfrak{a}}}\subset\Sigma[x_{1},\ldots,x_{N}]$ and a lifting ${\widetilde{\mathfrak{a}}}\subset{\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]$ of $\underline{{\mathfrak{a}}}$ satisfies
[…]
We will prove (18) for the case $N\geq 4$ first, and then will prove for the case $N\leq 3$.
[Proof of (ii) for the case $N\geq 4$]
Assume $N\geq 4$. Consider the commutative diagram
$\begin{array}[]{ccc}B_{n+2}&\stackrel{{\scriptstyle\varphi}}{{\longrightarrow}}&B_{0}\\ \downarrow&&\downarrow\\ {\widetilde{B}}_{n+2}&\stackrel{{\scriptstyle{\widetilde{\varphi}}}}{{\longrightarrow}}&{\widetilde{B}}_{0},\end{array}$
where the vertical arrows are the canonical closed immersions induced from the surjection ${\widetilde{\Sigma}}\to\Sigma$.
Note that $v_{F_{2}}(\underline{{\mathfrak{a}}})\geq m$ if and only if $\underline{{\mathfrak{a}}}\subset\varphi_{*}({\mathcal{O}}_{B_{n+2}}(-mF_{2}))$. Then, to prove (18), it is sufficient to prove that there exists a lifting ${\widetilde{\mathfrak{a}}}\subset{\widetilde{S}}_{0}={\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]$ of $\underline{{\mathfrak{a}}}$ such that
${\widetilde{\mathfrak{a}}}\subset{\widetilde{\varphi}}_{*}({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2}))\ \ \mbox{for}\ \ m=v_{F_{2}}(\underline{{\mathfrak{a}}}),$
because $v_{{\widetilde{F}}_{2}}({\widetilde{\mathfrak{a}}})\leq v_{F_{2}}(\underline{{\mathfrak{a}}})$ is clear for any lifting $\widetilde{\mathfrak{a}}$} of $\underline{{\mathfrak{a}}}$. Therefore, we have only to prove the surjectivity of the canonical map:
(19) ${\widetilde{\varphi}}_{*}({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2}))\to\varphi_{*}({\mathcal{O}}_{B_{n+2}}(-mF_{2}))$
for every $m\in{\mathbb{N}}$. As the monomials of degree >m>m do not affect the valuation $m=v_{F_{2}}(f)$ of the polynomial $f\in\Sigma[x_{1},\ldots,x_{N}]$, we can neglect higher degree term of the polynomial. Hence, it is sufficient to show the surjectivity of
(20) ${\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2}))/{\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{n+2}}\right)\stackrel{{\scriptstyle\alpha}}{{\longrightarrow}}\varphi_{*}\left({\mathcal{O}}_{B_{n+2}}(-mF_{2}))/{\mathfrak{m}}^{t}{\mathcal{O}}_{B_{n+2}}\right),$
for $t>m$. Now we reduce the problem into the problem about coherent sheaves with the support on the exceptional divisors. Here, denote ${\widetilde{\mathfrak{m}}}=(x_{1},\ldots,x_{N})\subset{\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]$ and ${\mathfrak{m}}=(x_{1},\ldots,x_{N})\subset\Sigma[x_{1},\ldots,x_{N}]$. For simplicity of the notation, we denote
${\mathcal{F}}:={\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2}))/{\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{n+2}},$
then, $\mathcal{F}$} is $\mathbb{Z}$}-flat and the following holds
${\mathcal{O}}_{B_{n+2}}(-mF_{2}))/{\mathfrak{m}}^{t}{\mathcal{O}}_{B_{n+2}}={\mathcal{F}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)},$
by Lemma 5.2.
As the map $\alpha$ in (20) is a part of the following long exact sequence:
[…]
it is sufficient to show that the last map is injective, i.e.,
[…]
Fix an integer $t$ such that $t>m$ and factor the locally principal ideals ${\mathfrak{m}}^{t}{\mathcal{O}}_{B_{n+2}}$ and ${\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{n+2}}$ into the following forms:
${\mathfrak{m}}^{t}{\mathcal{O}}_{B_{n+2}}=I\cdot I_{1}^{r_{1}}\cdot I_{2}^{r_{2}},$
${\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{n+2}}={\widetilde{I}}\cdot{\widetilde{I}}_{1}^{r_{1}}\cdot{\widetilde{I}}_{2}^{r_{2}},$
where for $i=1,2$, the ideals $I_{i}$ and ${\widetilde{I}}_{i}$ are the defining ideals of the proper transforms of the exceptional divisors $F_{i}$ and ${\widetilde{F}}_{i}$ on $B_{n+2}$ and ${\widetilde{B}}_{n+2}$, respectively. We denote the remaining factors by $I$ and $\widetilde{I}$}, respectively. Here, by the construction of blow-ups, ${\widetilde{I}},{\widetilde{I}}_{1}$ and ${\widetilde{I}}_{2}$ are liftings of $I,I_{1}$ and $I_{2}$, and
$r_{1}=r_{2}=r,$
where $r$ is the exponent of the defining ideal of $E_{n}$ in $I$, which is the same as that of ${\widetilde{E}}_{n}$ in $\widetilde{I}$} by Lemma 5.2. Consider the following exact sequence of Mayer-Vietoris type:
0
$\downarrow$
${\mathcal{F}}={\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/({\widetilde{I}}\cdot{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r})$
(21) $\downarrow$
$({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}})\oplus({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r})$
$\downarrow$
${\mathcal{O}}_{{\widetilde{B}}_{n+2}}/({\widetilde{I}}+{\widetilde{I}}_{1}^{r})$
$\downarrow$
$0.$
From this exact sequence, we obtain the long exact sequence of direct image sheaves. First consider the case $m=0$ as follows:
0
$\downarrow$
${\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\cdot{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)$
$\iota_{0}\downarrow$
${\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\right)\oplus{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)$
$\pi_{0}\downarrow$
${\widetilde{\varphi}}_{*}({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/({\widetilde{I}}+{\widetilde{I}}_{1}^{r}))$
(22) $\downarrow$
$R^{1}{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\cdot{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)\$
$\downarrow$
$R^{1}{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\right)\oplus R^{1}{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)$
$\downarrow$
$R^{1}{\widetilde{\varphi}}_{*}{\mathcal{O}}_{{\widetilde{B}}_{n+2}}/({\widetilde{I}}+I_{1}^{r}).$
Here, we will show the following:
  1. (a)
[…]
  2. (b)
    $R^{1}{\widetilde{\varphi}}_{*}({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\cdot{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r})=0.$
  3. (c)
    $R^{1}{\widetilde{\varphi}}_{*}({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}+{\widetilde{I}}_{1}^{r})=0$.
  4. (d)
    $R^{1}{\widetilde{\varphi}}_{*}({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}})=R^{1}{\widetilde{\varphi}}_{*}({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r})=0.$
[Proof of (a)] We will prove the following canonical map is isomorphic
[…]
We study of ${\widetilde{B}}_{n+1}$ instead of ${\widetilde{B}}_{n+2}$. Denote ${\widetilde{\mathfrak{m}}}{\mathcal{O}}_{{\widetilde{B}}_{n+1}}={{{{\widetilde{J}}}_{1}}}^{r}\cdot{\widetilde{I}}$, where ${\widetilde{J}}_{1}$ is the defining ideal of the latest exceptional divisor ${\widetilde{F}}_{1}\subset{\widetilde{B}}_{n+1}$. By the definition of the center ${\widetilde{P}}_{n+2}\in{\widetilde{B}}_{n+1}$ of blow-up ${\widetilde{\varphi}}_{n+2}$, we have the isomorphism
${\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\simeq{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{I}}$.
Here, we use the same symbol $\widetilde{I}$} to describe the defining ideals of the image of $Z({\widetilde{I}})\subset{\widetilde{B}}_{n+2}$ in ${\widetilde{B}}_{n+1}$. As
${\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\cdot{{\widetilde{I}}_{1}}^{r}\cdot{{\widetilde{I}}_{2}}^{r}\right)={\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]/(x_{1},\ldots,x_{N})^{t}={\widetilde{\psi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{I}}\cdot{{\widetilde{J}}_{1}}^{r}\right),$
it is sufficient to show
[…]
where ${\widetilde{\psi}}:{\widetilde{B}}_{n+1}\to{\widetilde{B}}_{0}$ is the composite ${\widetilde{\varphi}}_{1}\circ\cdots\circ{\widetilde{\varphi}}_{n+1}$. Considering the exact sequence:
$0\to{\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{{\widetilde{J}}_{1}}^{r}\to{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{I}}\cdot{{\widetilde{J}}_{1}}^{r}\to{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{I}}\to 0,$
we will see the following vanishing for $i=0,1$
(25) $R^{i}{\widetilde{\psi}}_{*}({\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{{\widetilde{J}}_{1}}^{r})=0,$
which yields the isomorphism (24).
Indeed, for $j\leq r-1$, noting that
(26) ${\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{F}}_{1}}={\widetilde{I}}_{{\widetilde{E}}_{n}}^{r}\otimes{\mathcal{O}}_{{\widetilde{F}}_{1}}={\mathcal{O}}_{{\mathbb{P}}_{\widetilde{\Sigma}}^{N-1}}(-r)\ \ \mbox{and}\ \ {\widetilde{I}}\otimes{\widetilde{J}}_{1}^{j}\otimes{\mathcal{O}}_{{\widetilde{F}}_{1}}={\mathcal{O}}_{{\mathbb{P}}_{\widetilde{\Sigma}}^{N-1}}(j-r),$
[…]
$0\to{\widetilde{I}}\otimes{\widetilde{J}}^{j}_{1}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}\to{\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/({\widetilde{J}}_{1}^{j+1})\to{\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/({\widetilde{J}}_{1}^{j})\to 0,$
we obtain the long exact sequence of the relative cohomologies
$\to R^{i}{\widetilde{\psi}}_{*}({\widetilde{I}}\otimes{\widetilde{J}}^{j}_{1}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1})\to R^{i}{\widetilde{\psi}}_{*}({\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/({\widetilde{J}}_{1}^{j+1}))\to R^{i}{\widetilde{\psi}}_{*}({\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/({\widetilde{J}}_{1}^{j}))\to.$
Assume i<N1i<N-1, apply the vanishings of $R^{i}{\widetilde{\psi}}$ for the two sheaves in (26) with $j=1$ first, then by induction on $j$ up to j=r1j=r-1, we obtain :
$R^{i}{\widetilde{\psi}}_{*}({\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{{\widetilde{J}}_{1}}^{r})=0,$
in particular, it holds for $i=0,1$ as we assume $N\geq 4$. This completes the proof of (24), and therefore the proof of $(a)$ as well. Here, note that (25) also holds for $i=2$ if $N>3$ (this comment is for later use in the proof of (c)).
[Proof of (b)]First we show for every $t\in{\mathbb{N}}$
(27) $R^{1}\varphi_{*}({\mathcal{O}}_{A_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{n+2}})=0,$
where we denote the maximal ideal $(x_{1},\ldots,x_{N})\subset k[x_{1},\ldots,x_{N}]$ also by $\mathfrak{m}$}. Factor the morphism $\varphi:A_{n+2}\to A$ by $\psi:A_{n+2}\to A_{1}$ and $\varphi_{1}:A_{1}\to A$. By Corollary 4.5 it follows:
(28) $R^{i}\psi_{*}{\mathcal{O}}_{A_{n+2}}=\left\{\begin{array}[]{cc}0&i>0\\ {\mathcal{O}}_{A_{1}}&i=0.\end{array}\right$.
Therefore, we have
(29) $R^{i}\psi_{*}{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{n+2}}=R^{i}\psi_{*}(\psi^{*}({\mathfrak{m}}^{t}{\mathcal{O}}_{A_{1}}))=(R^{i}\psi_{*}{\mathcal{O}}_{A_{n+2}})\otimes{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{1}}=\left\{\begin{array}[]{cc}0&i>0\\ {\mathfrak{m}}^{t}{\mathcal{O}}_{A_{1}}&i=0,\end{array}\right$.
where, as ${\mathfrak{m}}^{t}{\mathcal{O}}_{A_{1}}$ is invertible, we can apply the projection formula to obtain the second equality. Considering the exact sequence
$0\to{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{n+2}}\to{\mathcal{O}}_{A_{n+2}}\to{\mathcal{O}}_{A_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{n+2}}\to 0,$
by using (28) and (29) we obtain
(30) $R^{i}\psi_{*}({\mathcal{O}}_{A_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{n+2}})=\left\{\begin{array}[]{cc}0&i>0\\ {\mathcal{O}}_{A_{1}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{1}}&i=0.\end{array}\right$.
Now, take the Leray’s spectral sequence:
$E_{2}^{p,q}=R^{p}\varphi_{1*}R^{q}\psi_{*}({\mathcal{O}}_{A_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{n+2}})\Rightarrow E^{p+q}=R^{p+q}\varphi_{*}({\mathcal{O}}_{A_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{n+2}}).$
As $E_{2}^{1,0}=R^{1}\varphi_{1*}({\mathcal{O}}_{A_{1}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{1}})=0$ and $E_{2}^{0,1}=\varphi_{1*}R^{1}\psi_{*}({\mathcal{O}}_{A_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{n+2}})=0$, the canonical exact sequence $0\to E_{2}^{1,0}\to E^{1}\to E_{2}^{0,1}$ gives $E^{1}=0$, i.e.
$R^{1}\varphi_{*}({\mathcal{O}}_{A_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{n+2}})=0,$
which is first required. This cohomology is the faithfully flat base change of
$R^{1}\varphi_{*}(({\mathcal{O}}_{B_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{B_{n+2}})\otimes_{\Sigma}Q(\Sigma))$
by the field extension $Q(\Sigma)\hookrightarrow k$, where $Q(\Sigma)$ is the field of fractions of $\Sigma$. Therefore,
(31) $R^{1}\varphi_{*}(({\mathcal{O}}_{B_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{B_{n+2}})\otimes_{\Sigma}Q(\Sigma))=0.$
Now, we denote
${\mathcal{G}}:={\mathcal{O}}_{B_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{B_{n+2}}\ \mbox{and}$
${\widetilde{{\mathcal{G}}}}:={\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{n+2}}.$
[…]
(32) $(R^{1}\varphi_{*}{\mathcal{G}})\otimes_{\Sigma}Q(\Sigma)\simeq R^{1}\varphi_{*}({\mathcal{G}}\otimes_{\Sigma}Q(\Sigma))=0,$
where the vanishing follows from (31). On the other hand, the following exact sequence
$\to R^{1}{\widetilde{\varphi}}_{*}{\widetilde{{\mathcal{G}}}}\stackrel{{\scriptstyle\times p}}{{\longrightarrow}}R^{1}{\widetilde{\varphi}}_{*}{\widetilde{{\mathcal{G}}}}\to R^{1}\varphi_{*}{\mathcal{G}}\to\cdots$
gives the inclusion $R^{1}{\widetilde{\varphi}}_{*}{\widetilde{{\mathcal{G}}}}\otimes_{\widetilde{\Sigma}}\Sigma\hookrightarrow R^{1}\varphi_{*}{\mathcal{G}}$ from the cokernel of the map $\times p$. By applying the tensor $\otimes_{\Sigma}Q(\Sigma)$ on the inclusion, we obtain
$R^{1}{\widetilde{\varphi}}_{*}{\widetilde{{\mathcal{G}}}}\otimes_{\widetilde{\Sigma}}Q(\Sigma)\hookrightarrow R^{1}\varphi_{*}{\mathcal{G}}\otimes_{\Sigma}Q(\Sigma).$
By the vanishing (32), we obtain $R^{1}{\widetilde{\varphi}}_{*}{\widetilde{{\mathcal{G}}}}\otimes_{\widetilde{\Sigma}}Q(\Sigma)=0$, which yields
$R^{1}{\widetilde{\varphi}}_{*}{\widetilde{{\mathcal{G}}}}=0\ \ \mbox{around}\ \ \ p{\widetilde{\Sigma}}\in{\mathrm{Spec\ }}{\widetilde{\Sigma}}$.
[Proof of (c)] The support of ${\mathcal{O}}_{{\widetilde{B}}_{n+2}}/({\widetilde{I}}+{\widetilde{I}}_{1}^{r})$ is $C:={\widetilde{E}}_{n}\cap{\widetilde{F}}_{1}\simeq{\mathbb{P}}_{\widetilde{\Sigma}}^{N-2}$. As ${\widetilde{B}}_{n+2}$ is isomorphic to ${\widetilde{B}}_{n+1}$ around $C$, we will study ${\widetilde{B}}_{n+1}$ instead of ${\widetilde{B}}_{n+2}$ as we did in the proof of (a). We use the same symbols as in the proof of (a). Consider the following exact sequence:
$0\to{\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}^{r}\to{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}^{r}\to{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/({\widetilde{I}}+{\widetilde{J}}_{1}^{r})\to 0.$
As we noted at the end of the proof of (a), it follows:
(33) $R^{2}{\widetilde{\psi}}_{*}({\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}^{r})=0.$
On the other hand, we can prove
(34) $R^{1}{\widetilde{\psi}}_{*}{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}^{r}=0,$
by using $R^{1}{\widetilde{\psi}}_{*}({\widetilde{J}}_{1}^{j}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1})=H^{1}\left({\mathbb{P}}^{N-1}_{\widetilde{\Sigma}},{\mathcal{O}}_{{\mathbb{P}}^{N-1}_{\widetilde{\Sigma}}}(j)\right)=0$ $(j\geq 0)$ and the exact sequences for $j>0$:
$0\to{\widetilde{J}}_{1}^{j}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}\to{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}^{j+1}\to{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}^{j}\to 0.$
Here, we use $N\geq 4$ for the vanishing $H^{1}\left({\mathbb{P}}^{N-1}_{\widetilde{\Sigma}},{\mathcal{O}}_{{\mathbb{P}}^{N-1}_{\widetilde{\Sigma}}}(j)\right)=0$. Now in the exact sequence:
$R^{1}{\widetilde{\psi}}_{*}{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}^{r}\to R^{1}{\widetilde{\psi}}_{*}{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/({\widetilde{I}}+{\widetilde{J}}_{1}^{r})\to R^{2}{\widetilde{\psi}}_{*}({\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}^{r})$
the first and the third terms are 0 by (33) and (34), therefore it follows
$R^{1}{\widetilde{\psi}}_{*}{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/({\widetilde{I}}+{\widetilde{J}}_{1}^{r})=0,$
which implies the vanishing required in (c).
[Proof of (d)] Due to the sequence (22), the statement (d) follows automatically from (b) and (c).
Now, we are on the last step to show $\mathbb{Z}$}-flatness of $R^{1}{\widetilde{\varphi}}_{*}{\mathcal{F}}$ at $p$ for $N\geq 4$. We can use (a)–(d). For $m>0$, by taking direct image sheaves of the exact sequence (21), we obtain the following long exact sequence:
0
$\downarrow$
${\widetilde{\varphi}}_{*}{\mathcal{F}}={\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/{\widetilde{I}}\cdot{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)$
$\iota_{m}\downarrow$
${\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\right)\oplus{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)$
$\pi_{m}\downarrow$
${\widetilde{\varphi}}_{*}({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/({\widetilde{I}}+{\widetilde{I}}_{1}^{r}))$
(35) $\downarrow$
$R^{1}{\widetilde{\varphi}}_{*}{\mathcal{F}}=R^{1}{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/{\widetilde{I}}\cdot{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)$
$\downarrow$
$R^{1}{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\right)\oplus R^{1}{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)$
$\downarrow$
$R^{1}{\widetilde{\varphi}}_{*}{\mathcal{O}}_{{\widetilde{B}}_{n+2}}/({\widetilde{I}}+{\widetilde{I}}_{1}^{r}).$
Applying (a) – (d) to the exact sequence (35), we obtain
$R^{1}{\widetilde{\varphi}}_{*}{\mathcal{O}}_{{\widetilde{B}}_{n+2}}/({\widetilde{I}}+{\widetilde{I}}_{1}^{r})=0,\ \mbox{and}$
$R^{1}{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\right)=0.$
Then, we have the exact sequence:
$0\to{\rm Coker\ }\pi_{m}\to R^{1}{\widetilde{\varphi}}_{*}{\mathcal{F}}\to R^{1}{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)\to 0,$
where the last term is $\mathbb{Z}$}-flat at $p$, because our last two blow-ups are the same as the ones in Proposition 4.12, (ii) and the surjection in (c) of Proposition 4.12, (ii) gives the $\mathbb{Z}$}-flatness of the last term at $p$. Hence, in order to prove that $R^{1}{\widetilde{\varphi}}_{*}{\mathcal{F}}$ is $\mathbb{Z}$}-flat at $p$, it is sufficient to prove that ${\rm Coker\ }\pi_{m}$ is $\mathbb{Z}$}-flat at $p$.
[…]
${\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\right)\to{\widetilde{\varphi}}_{*}({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/({\widetilde{I}}+{\widetilde{I}}_{1}^{r}))$
is the zero map. Here, noting that ${\widetilde{F}}_{2}$ and the co-support of $\widetilde{I}$} and are disjoint, we have
(36) ${\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\right)\to{\widetilde{\varphi}}_{*}({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/({\widetilde{I}}+{\widetilde{I}}_{1}^{r}))$
is the zero map.
On the other hand, (a) and (b) yield the isomorphism
(37) ${\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)\simeq{\widetilde{\varphi}}_{*}({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/({\widetilde{I}}+{\widetilde{I}}_{1}^{r}))$
Considering (36) and (37) in the exact sequence (35), we have ${\rm Coker\ }\pi_{m}={\rm Coker\ }\beta$, where
$\beta:{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)\to{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right).$
As the map $\beta$ appears in the following long exact sequence:
$0\to{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)\stackrel{{\scriptstyle\beta}}{{\longrightarrow}}{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)\to{\widetilde{\varphi}}_{*}{\mathcal{O}}_{m{\widetilde{F}}_{2}},$
${\rm Coker\ }\beta$ is contained in the $\mathbb{Z}$}-flat module ${\widetilde{\varphi}}_{*}{\mathcal{O}}_{m{\widetilde{F}}_{2}}$. Therefore, ${\rm Coker\ }\beta$ is also $\mathbb{Z}$}-flat at $p$ as required. [QED for $N\geq 4$]
[Proof of (ii) for the case $N\leq 3$]
As we proved (ii) of Theorem 5.3 for $N\geq 4$, the following lemma provides the proof of (ii) for $N\leq 3$. ∎
Lemma 5.4.
Let $N\in{\mathbb{Z}}_{\geq 2}$ be the dimension of the affine space $A={\mathbb{A}}_{k}^{N}$. If the Theorem 5.3 is proved for the $(N+1)$-dimensional affine space by the method of the proof as above, then the theorem also holds for $N$-dimensional affine space.
Proof.
We use the notation in the proof of Theorem 5.3 for $N\geq 4$ shown above. Let $E$ be a prime divisor over $(A,0)$, where $A={\mathbb{A}}_{k}^{N}$ and $0\in A$ is the origin. In the proof of Theorem 5.3, we constructed sequence of schemes of compatible skeletons for any dimension $N$:
[…]
where the prime divisor $E=E_{n}$ appears on $A_{n}$ whose skeleton is $B_{n}$. Let $P_{i}\in B_{i-1}$ be the center of $E$ and also give an irreducible component of the center $Z_{i}\subset B_{i-1}$ of the blow-up $B_{i}\to B_{i-1}$. On the other hand, ${\widetilde{P}}_{i}\in{\widetilde{B}}_{i-1}$ is the weak prime lifting of $P_{i}$ and gives an irreducible component of the center ${\widetilde{Z}}_{i}\subset{\widetilde{B}}_{i-1}$ of the blow-up ${\widetilde{B}}_{i}\to{\widetilde{B}}_{i-1}$. Here, ${\widetilde{Z}}_{i}$ is a lifting of $Z_{i}$. Then, we already proved the statement (i) for arbitrary dimension $N$. Therefore, the problem is (ii). In order to reduce the discussion to the $(N+1)$-dimensional case, embed our affine spaces into higher dimensional affine spaces as follows:
$B_{0}={\mathbb{A}}_{\Sigma}^{N}={\mathrm{Spec\ }}\Sigma[x_{1},\ldots,x_{N}]\subset\overline{B}_{0}:={\mathbb{A}}_{\Sigma}^{N+1}={\mathrm{Spec\ }}\Sigma[x_{1},\ldots,x_{N},x_{N+1}]$
${\widetilde{B}}_{0}={\mathbb{A}}_{\widetilde{\Sigma}}^{N}={\mathrm{Spec\ }}{\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]\subset\overline{{\widetilde{B}}}_{0}:={\mathbb{A}}_{\widetilde{\Sigma}}^{N+1}={\mathrm{Spec\ }}{\widetilde{\Sigma}}[x_{1},\ldots,x_{N},x_{N+1}],$
as the hyperplanes defined by $x_{N+1}=0$. For $i=1,2,\ldots,n+2$, take the blow-ups
$\overline{B}_{i}\to\overline{B}_{i-1}\ \mbox{ with\ the\ center\ by\ }Z_{i}\subset B_{i-1}\subset\overline{B}_{i-1}\ \mbox{and}$
$\overline{{\widetilde{B}}}_{i}\to\overline{{\widetilde{B}}}_{i-1}\ \mbox{ with\ the\ center\ by\ }{\widetilde{Z}}_{i}\subset{\widetilde{B}}_{i-1}\subset\overline{{\widetilde{B}}}_{i-1}.$
Here, we note that $B_{i}$ and ${\widetilde{B}}_{i}$ are isomorphic to the proper transforms of $B_{0}$ and ${\widetilde{B}}_{0}$ in $\overline{B}_{i}$ and in $\overline{{\widetilde{B}}}_{i}$, respectively. So we identify $B_{i}$ and ${\widetilde{B}}_{i}$ with the closed subschemes in $\overline{B}_{i}$ and in $\overline{{\widetilde{B}}}_{i}$, respectively. As $\overline{{\widetilde{B}}}_{0}$ and $\overline{B}_{0}$ are compatible skeletons, and ${\widetilde{Z}}_{1}={\widetilde{P}}_{1}=0\in\overline{{\widetilde{B}}}_{0}$ is a lifting of $Z_{1}=P_{1}=0\in\overline{B}_{0}$, it follows that $\overline{{\widetilde{B}}}_{1}$ and $\overline{B}_{1}$ are compatible skeletons.
For $i\geq 2$ the center ${\widetilde{Z}}_{i}\subset{\widetilde{B}}_{i-1}\subset\overline{{\widetilde{B}}}_{i-1}$ is a lifting of $Z_{i}\subset B_{i-1}\subset\overline{{\widetilde{B}}}_{i-1}$. Then, inductively it follows that $\overline{{\widetilde{B}}}_{i}$ and $\overline{B}_{i}$ are compatible skeletons, and ${\widetilde{P}}_{i+1}\in\overline{{\widetilde{B}}}_{i}$ is a weak prime lifting of $P_{i+1}\in\overline{B}_{i}$ for all $0\leq i\leq n+1$.
Here, we note that
$\overline{B}_{n+2}\to\overline{B}_{n+1}\to\overline{B}_{n}\to\ldots\to\overline{B}_{1}\to\overline{B}_{0}={\mathbb{A}}_{\Sigma}^{N+1}$
is a normal Gorenstein Zariski sequence, i.e. it satisfies (i)–(iv) of Proposition 4.6. And
$\overline{{\widetilde{B}}}_{n+2}\to\overline{{\widetilde{B}}}_{n+1}\to\overline{{\widetilde{B}}}_{n}\to\ldots\to\overline{{\widetilde{B}}}_{1}\to\overline{{\widetilde{B}}}_{0}={\mathbb{A}}_{\widetilde{\Sigma}}^{N+1}$
is its lifting.
Let $E_{i}\subset B_{i}$ and $\overline{E}_{i}\subset\overline{B}_{i}$ be the exceptional divisors dominating $P_{i}\in B_{i-1}$ and $P_{i}\in\overline{B}_{i-1}$, respectively. Let ${\widetilde{E}}_{i}\subset{\widetilde{B}}_{i}$ and $\overline{{\widetilde{E}}}_{i}\subset\overline{{\widetilde{B}}}_{i}$ be the exceptional divisors dominating ${\widetilde{P}}_{i}$. (For $i=n+1,n+2$, we use character $F$ instead of $E$ as we did before.)
Let ${\mathfrak{a}}\subset k[x_{1},\ldots,x_{N}]$ be a non-zero ideal. We may assume that $\mathfrak{a}$} is generated by elements of $\Sigma[x_{1},\ldots,x_{N}]$ by taking $\Sigma\subset k$ sufficiently large. Denote the ideal ${\mathfrak{a}}\cap\Sigma[x_{1},\ldots,x_{N}]$ by the same symbol $\mathfrak{a}$}. Let
$\overline{{\mathfrak{a}}}:={\mathfrak{a}}[x_{N+1}]\subset\Sigma[x_{1},\ldots,x_{N},x_{N+1}].$
Then, applying the theorem to $\overline{{\widetilde{E}}}_{n}$ and $\overline{{\mathfrak{a}}}$, there exists a lifting $\widetilde{\overline{{\mathfrak{a}}}}\subset{\widetilde{\Sigma}}[x_{1},\ldots,x_{N},x_{N+1}]$ of $\overline{{\mathfrak{a}}}$ such that
(38) $v_{\overline{{\widetilde{F}}}_{2}}(\widetilde{\overline{{\mathfrak{a}}}})=v_{\overline{F}_{2}}({\overline{{\mathfrak{a}}}})=v_{\overline{E}_{n}}(\overline{{\mathfrak{a}}}).$
Define ${\mathfrak{a}}^{\prime}\subset{\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]$ to be the image of $\widetilde{\overline{{\mathfrak{a}}}}\subset{\widetilde{\Sigma}}[x_{1},\ldots,x_{N},x_{N+1}]$ by the canonical projection. Then, $\mathfrak{a}$$^{\prime}$ is a lifting of $\mathfrak{a}$}. By the definition of the ideals it follows
(39) $v_{{\widetilde{F}}_{2}}({\mathfrak{a}}^{\prime})=v_{\overline{{\widetilde{F}}}_{2}}({\mathfrak{a}}^{\prime}[x_{N+1}])\geq v_{\overline{{\widetilde{F}}}_{2}}(\widetilde{\overline{{\mathfrak{a}}}})=v_{\overline{F}_{2}}({\overline{{\mathfrak{a}}}})=v_{F_{2}}({\mathfrak{a}}),$
where the inequality in the middle follows from the minimality of ${\mathfrak{a}}^{\prime}[x_{N+1}]$ among the ${\widetilde{\Sigma}}[x_{1},\ldots,x_{N+1}]$-ideals whose images in ${\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]$ coincide with $\mathfrak{a}$$^{\prime}$. The inequality (39) gives $v_{{\widetilde{F}}_{2}}({\mathfrak{a}}^{\prime})\geq v_{F_{2}}({\mathfrak{a}}),$ while the opposite inequality holds in general. Now, we obtain that the lifting $\mathfrak{a}$$^{\prime}$ of $\mathfrak{a}$} satisfies
$v_{{\widetilde{F}}_{2}}({\mathfrak{a}}^{\prime})=v_{F_{2}}({\mathfrak{a}})=v_{E_{n}}({\mathfrak{a}}),$
as desired. ∎
Corollary 5.5 (Theorem 1.1, (iii)).
Let $k$ be an algebraically closed field of characteristic $p>0$, $A={\mathbb{A}}_{k}^{N}$ the affine space over $k$ of dimension $N\geq 2$ and 0 the origin of $A$. Let $E$ be a prime divisor over $(A,0)$ and ${\mathfrak{a}}_{1},\ldots,{\mathfrak{a}}_{r}\subset k[x_{1},\ldots,x_{N}]$ non-zero ideals. Then, there exists a prime divisor $F_{\mathbb{C}}$ over the affine space $A_{\mathbb{C}}={\mathbb{A}}_{\mathbb{C}}^{N}$ over $\mathbb{C}$} with the center at the origin $\tilde{0}$} satisfying the following:
There exist liftings ${\widetilde{\mathfrak{a}}}_{i}\subset{\mathcal{O}}_{A_{\mathbb{C}}}$ of ${\mathfrak{a}}_{i}$ $(i=1,\ldots,r)$ such that for every multi-ideal ${\mathfrak{a}}^{e}:={\mathfrak{a}}_{1}^{e_{1}}\cdots{\mathfrak{a}}_{r}^{e_{r}}$ $(e_{i}\in{\mathbb{R}}_{>0})$ on $A$, the lifting ${\widetilde{\mathfrak{a}}}^{e}={\widetilde{\mathfrak{a}}}_{1}^{e_{1}}\cdots{\widetilde{\mathfrak{a}}}_{r}^{e_{r}}$ of ${\mathfrak{a}}^{e}$ on $A_{\mathbb{C}}$ satisfies
$2(N-1)+a(E;A,{\mathfrak{a}}^{e})=a(F_{\mathbb{C}};A_{\mathbb{C}},{\widetilde{\mathfrak{a}}}^{e}).$
Proof.
By (i) and (ii) of the theorem, the equality
[…]
is obvious for an ideal $\mathfrak{a}$}. For multi-ideal case, as the number of ideals is finite in the multi-ideal, we can take common centers $P_{n+1}$ and $P_{n+2}$ general for all ideals. Then we obtain a lifting ${\widetilde{\mathfrak{a}}}_{i}$ of ${\mathfrak{a}}_{i}$ for each $i$ such that $v_{E}({\mathfrak{a}}_{i})=v_{F_{\mathbb{C}}}({\widetilde{\mathfrak{a}}}_{i})$, which completes the proof of our corollary. ∎
Until now, we assume the base space $A$ to be the affine space. However, we can apply the theorem also to a pair consisting of a non-singular variety and a multi-ideal with a real exponent. In such a case, we cannot define a lifting of a regular function around the point 0, but passing through the completion we can reduce the discussion to the case of the affine space.
Corollary 5.6 (Corollary 1.2, (i)).
Let $k$ be an algebraically closed field of characteristic $p>0$. Let $0\in A$ be a non-singular closed point of a variety $A$ of dimension $N$. Then, for fixed $N$ and an exponent $e\in{\mathbb{R}}_{>0}^{r}$, the set of log discrepancies
$\Lambda_{N,e}:=\left\{a(E;A,{\mathfrak{a}}^{e})\middle|\begin{array}[]{l}E:\mbox{prime\ divisor\ over\ }(A,0)\\ (A,{\mathfrak{a}}^{e}):\mbox{log\ canonical\ at\ }0\\ \end{array}\right\}$
is discrete in the sense of Corollary 1.2, (i). In particular, for a fixed $e$, the set $\{{\mathrm{mld}}(0,A,{\mathfrak{a}}^{e})\mid{\mathfrak{a}}:\ \mbox{non-zero\ ideal}\ \}$ is finite.
Proof.
[…]
$v_{E}(f)=v_{E}(f_{<i}).$
Note that $f_{<i}\in k[x_{1},\ldots,x_{N}]$. For a given ideal $\mathfrak{a}$} on $A$, take its generators and truncate them at degree $i\gg 0$, then we obtain an ideal ${\mathfrak{a}}^{\prime}\subset k[x_{1},\ldots,x_{N}]$. Now, by the discussion above, we obtain $v_{E}({\mathfrak{a}})=v_{E}({\mathfrak{a}}^{\prime})$ which reduces the discussion to the case of a pair $({\mathbb{A}}_{k}^{N},0)$.
From now on, we assume that $A$ is the affine space of dimension $N$ over a field of positive characteristic. By Theorem 1.1, (iii), for a fixed exponent $e\in{\mathbb{R}}_{>0}^{r}$, the set of numbers of the following form:
$\{2(N-1)+a(E;A,{\mathfrak{a}}^{e})\mid(A,{\mathfrak{a}}^{e})\ \mbox{is\ log\ canonical\ at\ }0,\ E:\ \mbox{prime\ divisor \ over \ }(A,0)\}$
is contained in
$\{a(E^{\prime};A_{\mathbb{C}},{\mathfrak{b}}^{e})\mid(A_{\mathbb{C}},{\mathfrak{b}}^{e})\ \mbox{is\ log\ canonical\ at\ }0,\ E^{\prime}:\ \mbox{prime\ divisor \ over \ }(A_{\mathbb{C}},0)\}$
which is known to be discrete by [12]. Here, we use that a lifting ${\widetilde{\mathfrak{a}}}^{e}$ of log canonical multi-ideal ${\mathfrak{a}}^{e}$ is also log canonical by Lemma 3.12. Hence, we obtain that the set of log discrepancies of log canonical multi-ideals with a fixed exponent in positive characteristic is also discrete.
As the minimal log discrepancy is either -$\infty$ or in the interval $[0,N]$ the discreteness of log discrepancies implies the finiteness of the set of the minimal log discrepancies. ∎
Corollary 5.7.
Let $k$ be an algebraically closed field of characteristic $p>0$. Let $0\in A$ be a non-singular closed point of a variety $A$ of dimension $N$, and ${\mathfrak{a}}^{e}$ a multi-ideal with a real exponent. Then, there exists a prime divisor over $A$ with the center at 0 computing
${\mathrm{mld}}(0;A,{\mathfrak{a}}^{e}).$
Proof.
If ${\mathrm{mld}}(0;A,{\mathfrak{a}}^{e})=$-$\infty$, then a prime divisor $E$ with $a(E;A,{\mathfrak{a}}^{e})<0$ computes the mld. If ${\mathrm{mld}}(0;A,{\mathfrak{a}}^{e})\geq 0$, the discreteness of log discrepancies in Corollary 5.6 gives the existence of a prime divisor computing the mld. ∎
The existence of a prime divisor computing $\mathrm{mld}$} in positive characteristic was already proved for a special case by K. Shibata ([18]).
Definition 5.8.
Let $k$ be a field and $0\in X$ a non-singular $k$-valued point of dimension $N\geq 2$. Fix an exponent $e\in{\mathbb{R}}_{>0}^{r}$. We define
$T_{N,e,k}:=\{{\operatorname{lct}}(0;X,{\mathfrak{a}}^{e})\mid{\mathfrak{a}}^{e}\ \mbox{ a\ multi-ideal\ with\ the\ exponent\ }e\mbox{ on}\ X\}$
the set of log canonical thresholds.
Corollary 5.9.
Let $k$ be an algebraically closed field $k$ of positive characteristic and $0\in A$ a non-singular closed point in a variety $A$ of dimension $N$. Fix an exponent $e\in{\mathbb{R}}_{>0}^{r}$. Then,
$T_{N,e,k}\subset T_{N,e,{\mathbb{C}}}.$
Proof.
As is seen in the proof of Corollary 5.6 we may assume that $A={\mathbb{A}}_{k}^{N}$ and 0 is the origin of ${\mathbb{A}}_{k}^{N}$. For simplicity, we denote $\sum e_{i}v_{E}({\mathfrak{a}}_{i})$} by $v_{E}({\mathfrak{a}}^{e})$.
In [4], they prove that $T_{N-1,e,{\mathbb{C}}}\subset T_{N,e,{\mathbb{C}}}$ and the accumulation points of $T_{N,e,{\mathbb{C}}}$ belong to $T_{N-1,e,{\mathbb{C}}}$. Actually, they stated it only for $e=(1)\in{\mathbb{R}}_{>0}$. However, their proof also works for every fixed $e\in{\mathbb{R}}^{r}_{>0}$. Now we look at a positive characteristic pair $(A,{\mathfrak{a}}^{e})$ and let $c={\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})$. If $c\in T_{N-1,e,{\mathbb{C}}}$, then $c\in T_{N,e,{\mathbb{C}}}$, which completes the proof. Assume $c\not\in T_{N-1,e,{\mathbb{C}}}$, then by the result of [4], $c$ is not an accumulation point of $T_{N,e,{\mathbb{C}}}$. Therefore, we obtain $\epsilon>0$ such that:
$(c,c+\epsilon)\cap T_{N,{\mathbb{C}}}=\emptyset.$
We claim that there is a prime divisor $E$ over $(A,0)$ such that
(40) $z(E;A,{\mathfrak{a}}^{e})=\frac{k_{E}+1}{v_{E}({\mathfrak{a}}^{e})}<c+\frac{\epsilon}{2},\ \ \mbox{and}$
(41) $\frac{2(N-1)}{v_{E}({\mathfrak{a}}^{e})}<\frac{\epsilon}{2}.$
Indeed, by the definition of $\operatorname{lct}$}, it is clear that there exists a prime divisor $E^{\prime}$ over $A$ with the center “containing 0” and satisfies the inequality (40):
$z(E^{\prime};A,{\mathfrak{a}}^{e})=\frac{k_{E^{\prime}}+1}{v_{E^{\prime}}({\mathfrak{a}}^{e})}<c+\frac{\epsilon}{2}.$
We need to find a prime divisor “with the center at 0” satisfying the inequality (40). If the center of $E^{\prime}$ is not $\{0\}$, then by suitable blow-ups, we obtain a normal variety $\overline{A}$ on which $E^{\prime}$ appears, and there is an exceptional divisor $E^{\prime}_{1}$ on $\overline{A}$ with the center at 0 such that
$E^{\prime}\cap E^{\prime}_{1}\neq\emptyset.$
We may assume that $\overline{A}$ is smooth around $E^{\prime}\cap E^{\prime}_{1}$ outside a closed subset of codimension $4$, and $E^{\prime}$ and $E^{\prime}_{1}$ intersect normally there. Because even over a field of positive characteristic, there is a log resolution for dimension $\leq 3$ ([1],[2]).
If $z(E^{\prime}_{1};A,{\mathfrak{a}}^{e})\geq c+{\epsilon}/{2}$, then take the blow-up at the intersection $E^{\prime}\cap E^{\prime}_{1}$ to obtain a new prime divisor $E^{\prime}_{2}$ dominating $E^{\prime}\cap E^{\prime}_{1}$ such that
$z(E^{\prime};A,{\mathfrak{a}}^{e})<z(E^{\prime}_{2};A,{\mathfrak{a}}^{e})<z(E^{\prime}_{1};A,{\mathfrak{a}}^{e}).$
Next, blow-up at $E^{\prime}\cap E^{\prime}_{2}$ to obtain $E^{\prime}_{3}$. Continue this process, then we have a sequence of prime divisors $\{E^{\prime}_{i}\}$ over $(A,0)$ such that
$\lim_{i\to\infty}z(E^{\prime}_{i};A,{\mathfrak{a}}^{e})=z(E^{\prime};A,{\mathfrak{a}}^{e}).$
This gives a prime divisor $E^{\prime}_{n}$ over $(A,0)$ satisfying (40). The inequality (41) in the claim holds when one takes sufficiently many times of blow-ups.
For $E$ and ${\mathfrak{a}}^{e}$, take a prime divisor $F_{\mathbb{C}}$ and a lifting ${\widetilde{\mathfrak{a}}}^{e}$ on ${\mathbb{A}}_{\mathbb{C}}^{N}$ as in Theorem 1.1. Then we obtain the inequalities by using Lemma 3.12:
$c+\epsilon>\frac{k_{E}+1+2(N-1)}{v_{E}({\mathfrak{a}}^{e})}=\frac{k_{F_{\mathbb{C}}}+1}{v_{F_{\mathbb{C}}}({\widetilde{\mathfrak{a}}}^{e})}\geq{\operatorname{lct}}({\tilde{0}};{\mathbb{A}}_{\mathbb{C}}^{N},{\widetilde{\mathfrak{a}}}^{e})\geq{\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})=c.$
By the definition of $\epsilon$, the last inequality must be the equality:
${\operatorname{lct}}({\tilde{0}};{\mathbb{A}}_{\mathbb{C}}^{N},{\widetilde{\mathfrak{a}}}^{e})={\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})=c,$
which yields ${\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})\in T_{N,e,{\mathbb{C}}}$.
Corollary 5.10 (Corollary 1.2,(ii),(iii)).
Assume $0\in A$ is a non-singular closed point of a variety of dimension $N$ over an algebraically closed field. Then the following hold:
  1. (ii)
    (Rationality of $\operatorname{lct}$}’s and their accumulation points) For a rational exponent $e$ and the multi-ideal ${\mathfrak{a}}^{e}$ on $A$, the log canonical threshold ${\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})$ is a rational number and an accumulation point of the log canonical thresholds is also a rational number.
  2. (iii)
    (ACC of $\operatorname{lct}$}) Let $D\subset{\mathbb{R}}_{>0}$ be a DCC set. Then, ACC holds for the set
    $T_{N,k}(D):=\{{\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})\mid e=(e_{1},\ldots,e_{r}),\forall e_{i}\in D\}.$
Proof.
By Corollary 5.9, for rational $e$, every value $c={\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})$ belongs to $T_{N,e,{\mathbb{C}}}\subset{\mathbb{Q}}$, and $T_{N,e,{\mathbb{C}}}$ is known to be closed ([4]), therefore (ii) follows immediately. As the set $T_{N,{\mathbb{C}}}(D)$ satisfies ACC by [3] for a special case and by [7] for the general case, its subset $T_{N,k}(D)$ satisfies ACC as well. ∎
6. A Conjecture
Theorem 1.1 is a weaker version of the following conjecture, whose form is simpler than the theorem.
Conjecture 6.1.
[…]
  1. (i)
    $k_{E}=k_{E_{\mathbb{C}}}$;
  2. (ii)
    For every non-zero ideal ${\mathfrak{a}}\subset{\mathcal{O}}_{{\mathbb{A}}_{k}^{N}}$, there exists an ideal ${\widetilde{\mathfrak{a}}}\subset{\mathcal{O}}_{{\mathbb{A}}_{\mathbb{C}}^{N}}$ which is a lifting of $\mathfrak{a}$} and satisfies
[…]
Note that if (i) and (ii) hold, then we obtain the following:
For every multi-ideal ${\mathfrak{a}}^{e}:={\mathfrak{a}}_{1}^{e_{1}}\cdots{\mathfrak{a}}_{r}^{e_{r}}$ $(e_{i}\in{\mathbb{R}}_{>0}$ on ${\mathbb{A}}_{k}^{N}$), there exists a multi-ideal ${\widetilde{\mathfrak{a}}}^{e}={\widetilde{\mathfrak{a}}}_{1}^{e_{1}}\cdots{\widetilde{\mathfrak{a}}}_{r}^{e_{r}}$ on ${\mathbb{A}}_{\mathbb{C}}^{N}$ such that
$a(E;{\mathbb{A}}_{k}^{N},{\mathfrak{a}}^{e})=a(E_{\mathbb{C}};{\mathbb{A}}_{\mathbb{C}}^{N},{\widetilde{\mathfrak{a}}}^{e}).$
The conjecture is a more powerful bridge than Theorem 1.1. Actually we can transport more statements in characteristic 0 to the positive characteristic.
Corollary 6.2.
Assume the conjecture holds true, then, for a perfect field $k$ and for $A={\mathbb{A}}_{k}^{N}$ and the origin $0\in A$, it follows:
  1. (i)
    For a fixed $e$, there is an injection:
    $\begin{array}[]{l}\left\{{\mathrm{mld}}(0;A,{\mathfrak{a}}^{e})\mid{\mathfrak{a}}^{e}\ \mbox{is \ a\ multi-ideal\ on}A\right\}\\ \ \ \ \ \ \ \ \ \ \ \ \ \hookrightarrow\left\{{\mathrm{mld}}({\tilde{0}};{\mathbb{A}}_{\mathbb{C}}^{N},{\mathfrak{b}}^{e})\mid{\mathfrak{b}}^{e}\ \mbox{is \ a\ multi-ideal\ on}\ {\mathbb{A}}_{\mathbb{C}}^{N}\right\}.\end{array}$
  2. (ii)
    If ACC Conjecture for minimal log discrepancies holds in characteristic 0, then it also holds in positive characteristic.
Proof.
Once the conjecture holds, we obtain the injection
[…]
by associating $a(E;A,{\mathfrak{a}}^{e})$ to $a(E_{\mathbb{C}},{\mathbb{A}}_{\mathbb{C}}^{N},{\widetilde{\mathfrak{a}}}^{e})$, where $E_{\mathbb{C}}$ and ${\widetilde{\mathfrak{a}}}^{e}$ are given by the Conjecture 6.1. Here, we note that this map is well defined between these sets in (42), i.e., ${\widetilde{\mathfrak{a}}}^{e}$ is log canonical by Lemma 3.12.
As the second set in (42) is discrete, the first set is also discrete, which yields a prime divisor $E$ over $(A,0)$ computing ${\mathrm{mld}}(0;A,{\mathfrak{a}}^{e})$. Applying Conjecture to $E$ and ${\mathfrak{a}}^{e}$, there exist $E_{\mathbb{C}}$ and ${\widetilde{\mathfrak{a}}}^{e}$ such that
${\mathrm{mld}}(0;A,{\mathfrak{a}}^{e})=a(E;A,{\mathfrak{a}}^{e})=a(E_{\mathbb{C}};{\mathbb{A}}_{\mathbb{C}}^{N},{\widetilde{\mathfrak{a}}}^{e})\geq{\mathrm{mld}}({\tilde{0}};{\mathbb{A}}_{\mathbb{C}}^{N},{\widetilde{\mathfrak{a}}}^{e}).$
On the other hand, by Lemma 3.12, we have the opposite inequality
${\mathrm{mld}}(0;A,{\mathfrak{a}}^{e})\leq{\mathrm{mld}}({\tilde{0}};{\mathbb{A}}_{\mathbb{C}}^{N},{\widetilde{\mathfrak{a}}}^{e})$
which gives the equality and therefore (i) holds. Now (ii) is automatic. ∎
When $N=2$ the statement (i) in Corollary 6.2 is known to hold by [11], moreover in this case the inclusion becomes equality.
There are many interesting questions based on our results.
Questions 6.3.
  1. (1)
    Is $T_{N,k}$ closed as in the case $\mathbb{C}$}?
  2. (2)
    Do the accumulation points of $T_{N,k}$ belong to $T_{N-1,k}$?
  3. (3)
    Is $T_{N,k}\subset T_{N,{\mathbb{C}}}$ equal?
  4. (4)
    Does the set $T_{N,k}$ of $\operatorname{lct}$}’s depend on the characteristic?