### v3

# Liftings of ideals in positive characteristic to those in characteristic zero: up to dimension three

Abstract.
We study log discrepancies and log canonical thresholds for multi-ideals with real exponents on smooth varieties of dimension at most three over fields of positive characteristic. We introduce a lifting construction that relates this setting to characteristic zero over the complex numbers, allowing us to compare these invariants across characteristics. Using this bridge, we prove that for fixed exponents the set of log discrepancies is discrete. We also show that the set of log canonical thresholds in positive characteristic is contained in the corresponding set over the complex numbers. Consequently, when all exponents are rational, log canonical thresholds and their accumulation points are rational. Finally, we establish the ascending chain condition for log canonical thresholds of such multi-ideals in dimension at most three in positive characteristic.
Mathematical Subject Classification 2020: 14B05,14E18, 14J17
Key words: singularities in positive characteristic, jet schemes, minimal log discrepancy, log canonical threshold
The author is partially supported by Grant-In-Aid (c) 22K03428 of JSPS and by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.
1. Introduction
There are many tools available for studying singularities in characteristic zero. These include resolutions of singularities, Bertini’s theorem (which addresses generic smoothness), and vanishing theorems of Kodaira type. However, these tools do not necessarily hold in positive characteristic, which creates difficulties when studying singularities in that setting. In order to overcome these difficulties, we construct a “bridge” between positive characteristic and characteristic zero, allowing us to transport results from characteristic zero to positive characteristic. In general, such a strategy does not work. For example, the resolution of singularities in positive characteristic does not follow from that in characteristic zero. However, for our specific problems, this approach is effective.
Our focus is on the invariants of singularities, specifically log discrepancies, minimal log discrepancies, and log canonical thresholds. Our bridge illustrates the relationships between the set of invariants in positive characteristic and the set of those in characteristic zero. As a result, the properties of these sets in characteristic zero are inherited by their counterparts in positive characteristic.
The main theorem of this paper serves as a “bridge” between characteristic 0 and positive characteristic.
Theorem 1.1.
Let $N$ be an integer $\geq 2$, and $k$ an algebraically closed field of characteristic $p>0$. Let $A={\mathbb{A}}_{k}^{N}$ be the affine space over $k$ of dimension $N$, and 0 the origin of $A$. Let $E$ be a prime divisor over $A$ with the center at 0, and ${\mathfrak{a}},{\mathfrak{a}}_{1},\ldots,{\mathfrak{a}}_{r}\subset k[x_{1},\ldots,x_{N}]$ non-zero ideals. Assume either:
  1. (1)
    $E$ is obtained by blow-ups at closed points, or
  2. (2)
    $N\leq 3$.
Then, there exists a prime divisor $F_{\mathbb{C}}$ over the affine space $A_{\mathbb{C}}={\mathbb{A}}_{\mathbb{C}}^{N}$ over $\mathbb{C}$} with the center at the origin $\tilde{0}$} satisfying the following:
  1. (i)
    $2(N-1)+k_{E}=k_{F_{\mathbb{C}}}$;
  2. (ii)
    there exists a lifting ${\widetilde{\mathfrak{a}}}\subset{\mathbb{C}}[x_{1},\ldots,x_{N}]$ of $\mathfrak{a}$} satisfying
[…]
  3. (iii)
    there exist liftings ${\widetilde{\mathfrak{a}}}_{i}\subset{\mathbb{C}}[x_{1},\ldots,x_{N}]$ of ${\mathfrak{a}}_{i}$ $(i=1,\ldots,r)$ such that
    $v_{E}({\mathfrak{a}}_{i})=v_{F_{\mathbb{C}}}({\widetilde{\mathfrak{a}}}_{i})$
    and for every multi-ideal ${\mathfrak{a}}^{e}:={\mathfrak{a}}_{1}^{e_{1}}\cdots{\mathfrak{a}}_{r}^{e_{r}}$ $(e_{i}\in{\mathbb{R}}_{>0})$ on $A$, the lifting ${\widetilde{\mathfrak{a}}}^{e}={\widetilde{\mathfrak{a}}}_{1}^{e_{1}}\cdots{\widetilde{\mathfrak{a}}}_{r}^{e_{r}}$ of ${\mathfrak{a}}^{e}$ on $A_{\mathbb{C}}$ satisfies
    $2(N-1)+a(E;A,{\mathfrak{a}}^{e})=a(F_{\mathbb{C}};A_{\mathbb{C}},{\widetilde{\mathfrak{a}}}^{e}).$
The constant term 2(N1)2(N-1) may look mysterious to the reader. This arises from two more blow-ups of the variety on which $E$ appears. The reason why we do this is briefly explained before Lemma 5.1.
We obtain the following corollaries as applications of the theorem.
Corollary 1.2.
Let $N\leq 3$ be a positive integer and $k$ an algebraically closed field of characteristic $p>0$. Then, the following hold:
  1. (i)
    (discreteness of log discrepancies) Let $0\in A$ be a smooth closed point of an $N$-dimensional algebraic variety $A$ defined over $k$. For a fixed exponent $e\in{\mathbb{R}}_{>0}^{r}$ and a real number $a>0$, the set of log discrepancies
[…]
    is finite. In particular, for fixed $N\leq 3$ and $e$, the set
    $\{{\mathrm{mld}}(0,A,{\mathfrak{a}}^{e})\mid{\mathfrak{a}}^{e}\ \mbox{non-zero\ multi-ideal}\ \}$
    is finite. Also for a fixed ${\mathfrak{a}}^{e}$, the discreteness of log discrepancies gives another proof of the existence of a prime divisor computing ${\mathrm{mld}}(0,A,{\mathfrak{a}}^{e})$.
  2. (ii)
    (Inclusion of the sets of lcts ) Let $e\in{\mathbb{R}}_{>0}^{r}$ and
    $T_{(A,0),e}:=\{{\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})\mid{\mathfrak{a}}^{e}:\mbox{non-zero\ multi-ideal\ on \ smooth }\ A\},\mbox{and}$
    $T_{({\mathbb{A}}_{\mathbb{C}}^{N},0),e}:=\{{\operatorname{lct}}({\tilde{0}};{\mathbb{A}}_{\mathbb{C}}^{e},{\mathfrak{b}}^{e})\mid{\mathfrak{b}}^{e}:\mbox{non-zero\ multi-ideal\ on}\ {\mathbb{A}}_{\mathbb{C}}^{N}\},$
    where ${\tilde{0}}\in{\mathbb{A}}_{\mathbb{C}}^{N}$ is the origin of the affine space $A_{\mathbb{C}}^{N}$ defined over $\mathbb{C}$}. Then,
    $T_{(A,0),e}\subset T_{({\mathbb{A}}_{\mathbb{C}}^{N},{\tilde{0}}),e}.$
    In particular, when the exponents $e_{i}$’s in $e=(e_{1},\ldots,e_{r})$ are all rational, then ${\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})$ is a rational number, and an accumulation point of these log canonical thresholds is also a rational number.
  3. (iii)
    (ACC of lcts) Let $D\subset{\mathbb{R}}_{>0}$ be a DCC set. Then, ACC holds for the set
    $T_{N,k}(D):=\left\{{\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})\middle|\begin{array}[]{l}(A,0):\mbox{\ smooth\ point\ of\ a\ variety\ of\ dimension}\ N\\ over\ k\ \mbox{and}\ e\in D\end{array}\right\},$
    where $e\in D$ means $e_{i}\in D$ $(i=1,\ldots,r)$ for $e=(e_{1},\ldots,e_{r})$.
Remark 1.3.
For a $2$- or $3$-dimensional variety in positive characteristic, ACC of lct is proved by Birkar [2, Theorem 1.10] in a strong form. However, our Corollary 1.2 (iii) is not contained in [2, Theorem 1.10]. First, our statement (iii) is characteristic free, while [2] assumes characteristic >5>5. Second, our (iii) is about multi-ideal ${\mathfrak{a}}^{e}$, while [2] is about ($\mathbb{R}$}-)Cartier divisor. In characteristic 0, lct of ideal is the same as lct of a general element of the ideal. But in positive characteristic case, it is not true. Because Bertini’s theorem (generic smoothness) fails in general, and log resolution of an ideal is not necessarily a log resolution of a general element of the ideal.
For $N=1$, the statements of Corollary 1.2 are trivial, because a prime divisor over $(A,0)$ is just the divisor $\{0\}\subset A$, and $v_{\{0\}}({\mathfrak{a}}^{e}):=\sum_{i}e_{i}{\mathrm{ord}}({\mathfrak{a}}_{i})$ and $k_{\{0\}}+1=1$ hold for any base field. So, we may assume $N\geq 2$ and can apply Theorem 1.1.
Note that the statements in Corollary 1.2 are known to hold in characteristic 0. More specifically, statement (i) is proven by Kawakita in [12], the statement about the rationality in (ii) is established by T. De Fernex and M. Mustaţǎ in [5], and statement (iii) for a special case is demonstrated by T. De Fernex, L. Ein and M. Mustaţǎ in [4] and for the general case by C. Hacon, J. McKernan and C. Xu in [8].
These are excellent results, and the proofs are elaborated based on resolutions of singularities, Bertini’s theorem, and cohomology vanishing of the Kodaira type, which are not available in the positive characteristic case. Therefore, the same proofs cannot be applied in positive characteristic settings.
Our proof in positive characteristic involves transporting these results from characteristic 0 into the realm of positive characteristic, utilizing the bridge established in Theorem 1.1.
This paper is organized as follows: In the second section, we introduce basic notions on liftings objects from positive characteristic to those in characteristic 0.
In the third section we introduce log discrepancies, minimal log discrepancies, and log canonical thresholds which are birational invariants to measure the nature of singularities of pairs. The description of these invariants by the arc spaces will allow us to compare these invariants in positive characteristic with those in characteristic 0.
In the fourth section, from a given prime divisor over the affine space in positive characteristic, we construct a prime divisor over the affine space defined over $\mathbb{C}$} such that the divisors have some common properties.
In the fifth section, we give the proof for the main theorem and the corollaries.
Acknowledgement. The author would like to thank Lawrence Ein, Kazuhiko Kurano, Mircea Mustaţǎ, Yuri Prokhorov, Kohsuke Shibata and Shoji Yokura for useful discussions and comments. The proof of the corollary about log canonical thresholds is largely indebted to the ideas by Mircea Mustaţǎ and his joint paper [5]. The author expresses her heartfelt thanks to János Kollár for providing critical examples and comments for the preliminary version of this paper.
Conventions
  1. (1)
    When we say just “a point” of a variety $A$, it is not necessarily a closed point. By abuse of notation, for a point $P\in A$ we write $\dim P:=\dim\overline{\{P\}}$. The codimension of $P$ is defined accordingly ; ${\mathrm{codim}}P:=\dim A-\dim P$.
  2. (2)
    A blow-up with the center by the defining ideal of the closure $\overline{\{P\}}$ is called “the blow-up at PP” and we also call $P$ the center of the blow-up.
  3. (3)
    Let $L_{1}$ be a ring, $L_{2}$ an $L_{1}$-algebra and $A$ a scheme of finite type over $L_{1}$. We denote the scheme $A\times_{{\mathrm{Spec\ }}L_{1}}{\mathrm{Spec\ }}L_{2}$ by $A\otimes_{L_{1}}L_{2}$ to avoid the bulky expression.
  4. (4)
    For an integral domain $R$, we denote the field of fractions of $R$ by $Q(R)$.
  5. (5)
    A prime ideal ${\mathfrak{p}}\subset R$ of a Noetherian integral domain is called a regular prime ideal if the local ring associated to $\mathfrak{p}$} is a regular local ring. Here, note that if $R$ is a graded ring and $\mathfrak{p}$} is a homogeneous prime ideal the associated local ring is $R_{({\mathfrak{p}})}$ by the symbol in [9, Definition, p.116].
  6. (6)
    Let ${\mathfrak{p}}\subset R$ be a regular prime ideal. A set of elements $\{f_{1},\ldots,f_{c}\}\subset{\mathfrak{p}}$ is called a regular system of parameters (RSP for short) of $\mathfrak{p}$} if it is an RSP in the local ring associated to $\mathfrak{p}$}. If $R$ is graded and $\mathfrak{p}$} is homogeneous, the set gives an RSP in $R_{({\mathfrak{p}})}$ by an appropriate dehomogenization.
  7. (7)
    Let $E_{i}$ $(i=1,\ldots,r)$ be closed subschemes of a scheme $X$ and $P\in X$ a point, and let ${\widetilde{E}}_{i}$ $(i=1,\ldots,r)$ be closed subschemes of a scheme $\widetilde{X}$ and ${\widetilde{P}}\in X$ a point. We say
    $\{E_{1},\ldots,E_{r},P\}$ has the same configuration as that of $\{{\widetilde{E}}_{1},\ldots,{\widetilde{E}}_{r},{\widetilde{P}}\}$
    if the following hold:
    $E_{i}\cap E_{j}\neq\emptyset\ \ \mbox{if\ and\ only\ if}\ \ \ {\widetilde{E}}_{i}\cap{\widetilde{E}}_{j}\neq\emptyset\ \mbox{and}$
    $P\in E_{j}\ \ \ \mbox{if\ and\ only\ if}\ \ \ {\widetilde{P}}\in{\widetilde{E}}_{j}.$
2. Basics on liftings
In this paper we construct a “lifting” which is an object on a variety over $\mathbb{C}$} corresponding to that in positive characteristic. Basic notions about liftings are discussed in [11], and here we summarize some of them, which we will use in this paper.
Definition 2.1.
Let $R$ be an integral domain which is finitely generated $k$-algebra, where $k$ is a field of characteristic $p\geq 0$. A subring $S\subset R$ is called a skeleton of $R$, if $S$ is finitely generated ${\mathbb{Z}}/(p)$}-subalgebra of $R$ such that $k[S]=R$.
Definition 2.2.
  1. (1)
    (compatible skeletons) Let $k$ be a field of characteristic $p>0$ and $K$ a field of characteristic 0. Let $R$ and $\widetilde{R}$} be finitely generated $k$-algebra and $K$-algebra, respectively. Let $S\subset R$ and ${\widetilde{S}}\subset{\widetilde{R}}$ be skeletons. If the skeletons satisfy
    $\widetilde{S}$} is $\mathbb{Z}$}-flat at $p$, and
    $S\simeq{\widetilde{S}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)},$
    we call them compatible skeletons. We sometimes say that a skeleton $\widetilde{S}$} is a lifting of the skeleton $S$. We also call the canonical surjection $\Phi_{p}:{\widetilde{S}}\to S$ compatible skeletons. In this case we write
    ${\widetilde{R}}({\operatorname{mod}\ }p)=R.$
    If schemes $\widetilde{A}$} and $A$ are defined by $\widetilde{R}$} and $R$ (eg. as $\mathrm{Spec\ }$} or $\mathrm{Proj\ }$}) which have compatible skeletons, we also write
    ${\widetilde{A}}({\operatorname{mod}\ }p)=A.$
    Here, we note that if $\widetilde{R}$} is an integral domain (therefore, so is $\widetilde{S}$}), then $\mathbb{Z}$}-flatness is automatic. In this case $S\simeq{\widetilde{S}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}$ is the only condition for $\widetilde{S}$} and $S$ to be compatible skeletons.
[…]
When we think of liftings of elements, we always use a common pair of compatible skeletons in a continuous discussion.
Example 2.3.
The following are examples of liftings. The first one is a trivial example and the second and third are more practical and will play basic roles in this paper. Assume that $k$ is a field of characteristic $p>0$.
(1) We have ${\mathbb{C}}({\operatorname{mod}\ }p)=k$ with compatible skeletons $S={{\mathbb{Z}}/(p)}$ and $\widetilde{S}$$\mathbb{Z}$}.
(2) For any finite elements $a_{1},\ldots,a_{r}\in k$ there exist liftings ${\widetilde{a}}_{1},\ldots,{\widetilde{a}}_{r}\in{\mathbb{C}}$.
Indeed compatible skeletons $\Sigma\subset k$ and $\widetilde{\Sigma}\subset{\mathbb{C}}$ giving these liftings are constructed in [11, Proposition 2.3]. However, since this statement plays a fundamental role in the paper, we restate the proof here.
Proof.
Considering the subring ${{\mathbb{Z}}/(p)}[a_{1},\ldots,a_{n}]\subset k$, we obtain canonical surjections:
$R:={\mathbb{Z}}[Y_{1},\ldots,Y_{n}]\stackrel{{\scriptstyle\psi}}{{\longrightarrow}}S:={{\mathbb{Z}}/(p)}[Y_{1},\ldots,Y_{n}]\stackrel{{\scriptstyle\varphi}}{{\longrightarrow}}{{\mathbb{Z}}/(p)}[a_{1},\ldots,a_{n}]$
with $Y_{i}\mapsto a_{i}$. Let $P:=Ker\varphi\subset S$ and $Q:=Ker\varphi\circ\psi\subset R$, then these are prime ideals in regular rings. Therefore $R_{Q}$ and $S_{P}$ are also regular local rings. Hence we obtain
  1. (1)
    $f_{1},\ldots,f_{c}\ (\in P)$ which form a regular system of parameters of $S_{P}$,
  2. (2)
    ${\tilde{f}}_{1},\ldots,{\tilde{f}}_{c},p\ (\in Q)$ which form a regular system of parameters of $R_{Q}$ and
  3. (3)
    $\psi({\tilde{f}}_{i})=f_{i}$ for $i=1,\ldots,c$.
Then, $R_{Q}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})R_{Q}$ is also a regular local ring, in particular it is an integral domain. By considering the homomorphism ${\mathbb{Z}}\to R_{Q}$ and the regular sequence $\{f_{1},\ldots,f_{c}\}\subset S_{P}=R_{Q}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}$, we obtain that $R_{Q}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})R_{Q}$ is flat over $\mathbb{Z}$} by [16, Corollary, p.177]. In particular, the homomorphism ${\mathbb{Z}}\to R_{Q}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})R_{Q}$ is injective, which implies the ring $R_{Q}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})R_{Q}$ is of characteristic 0. Note that the algebras ${\widetilde{\Sigma}}^{\prime}:=R_{Q}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})R_{Q}$ and $\Sigma^{\prime}:=S_{P}/(f_{1},\ldots,f_{c})S_{P}$ satisfies
${\widetilde{\Sigma}}^{\prime}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}=\Sigma^{\prime}={{\mathbb{Z}}/(p)}(a_{1},\ldots,a_{n}).$
However, these $\widetilde{\Sigma}$$^{\prime}$ and $\Sigma^{\prime}$ are not appropriate for the skeletons, because these are not finitely generated as $\mathbb{Z}$}-algebra and ${\mathbb{Z}}/(p)$}-algebra, respectively. So we replace $\widetilde{\Sigma}$$^{\prime}$ and $\Sigma^{\prime}$ by smaller ones.
As $PS_{P}=(f_{1},\ldots,f_{c})S_{P}$, there exists $h\in S\setminus P$ such that $PS_{h}=(f_{1},\ldots,f_{c})S_{h}$. Take ${\tilde{h}}\in R\setminus Q$ such that $\psi({\tilde{h}})=h$ and let
$\widetilde{\Sigma}:=R_{\tilde{h}}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})R_{\tilde{h}}$.
This, as well, is an integral domain and of characteristic 0. Now noting that $S/P={{\mathbb{Z}}/(p)}[a_{1},\ldots,a_{c}]$, we obtain
$\Sigma:=\widetilde{\Sigma}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}=\frac{R_{\tilde{h}}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})R_{\tilde{h}}}{p({R_{\tilde{h}}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{c})R_{\tilde{h}}})}=S_{h}/(f_{1},\ldots,f_{c})S_{h}$
$(S/P)_{h}\subset{{\mathbb{Z}}/(p)}(a_{1},\ldots,a_{c})\subset k.$
[…]
$\phi:K_{0}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\phi(K_{0})\subset{\mathbb{C}}.$
Then, we obtain $\{\widetilde{a}_{1},\ldots,\widetilde{a}_{n}\}\subset{\mathbb{C}}$ and $\{\widetilde{a}_{1},\ldots,\widetilde{a}_{n}\}({\operatorname{mod}\ }p)=\{a_{1},\ldots,a_{n}\}$. As is seen, $\widetilde{\Sigma}$$\subset{\mathbb{C}}$ and $\Sigma\otimes{{\mathbb{Z}}/(p)}$ are compatible skeletons. ∎
(3) Let $\Sigma\subset k$ and $\widetilde{\Sigma}$$\subset{\mathbb{C}}$ be compatible skeletons. Then, $S:=\Sigma[x_{1},\ldots,x_{N}]\subset k[x_{1},\ldots,x_{N}]$ and ${\widetilde{S}}:={\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]\subset{\mathbb{C}}[x_{1},\ldots,x_{N}]$ are also compatible skeletons for ${\mathbb{C}}[x_{1},\ldots,x_{N}]({\operatorname{mod}\ }p)=k[x_{1},\ldots,x_{N}]$ in the canonical way. Explicitly, a polynomial $f\in S$ has a lifting ${\tilde{f}}\in{\widetilde{S}}$ in the following form:
$f=\sum_{i\in{\mathbb{Z}}^{N}_{\geq 0}}a_{i}{\mathbb{x}}^{i}\ \ \ ,\ \ \mbox{then}\ \ \ {\tilde{f}}=\sum_{i\in{\mathbb{Z}}^{N}_{\geq 0}}{\widetilde{a}}_{i}\mathbb{x}^{i},$
where ${\widetilde{a}}_{i}$ is a lifting of $a_{i}$.
Definition 2.4.
Let $R$ and $\widetilde{R}$} be finitely generated algebras over fields $k$ and $K$ of characteristic $p>0$ and of 0, respectively.
Assume ${\widetilde{R}}({\operatorname{mod}\ }p)=R$ holds by compatible skeletons $\Phi_{p}:{\widetilde{S}}\to S$, where $S\subset R$ and ${\widetilde{S}}\subset{\widetilde{R}}$.
For ideals ${\mathfrak{a}}_{R}\subset R$ and ${\widetilde{\mathfrak{a}}}_{\widetilde{R}}\subset{\widetilde{R}}$ we write ${\widetilde{\mathfrak{a}}}_{\widetilde{R}}({\operatorname{mod}\ }p)={\mathfrak{a}}_{R}$ and say ${\widetilde{\mathfrak{a}}}_{\widetilde{R}}$ is a lifting of ${\mathfrak{a}}_{R}$ if we can take a generator system of ${\widetilde{\mathfrak{a}}}_{\widetilde{R}}$ from $\widetilde{S}$} whose image by $\Phi_{p}$ generates ${\mathfrak{a}}_{R}\subset R$. It is equivalent to the following:
  1. (1)
    There are ideals ${\mathfrak{a}}\subset S$ and ${\widetilde{\mathfrak{a}}}\subset{\widetilde{S}}$ such that ${\widetilde{\mathfrak{a}}}_{\widetilde{R}}={\widetilde{\mathfrak{a}}}{\widetilde{R}}$, ${\mathfrak{a}}_{R}={\mathfrak{a}}R$, and
  2. (2)
    $\Phi_{p}({\widetilde{\mathfrak{a}}})={\mathfrak{a}}$.
In this case, we call $\mathfrak{a}$} a skeleton of ${\mathfrak{a}}_{R}$, and $\widetilde{\mathfrak{a}}$} a skeleton of ${\widetilde{\mathfrak{a}}}_{\widetilde{R}}$.
Remark 2.5.
According to the definition above, the unit ideal $\widetilde{R}$} itself is a lifting of any ideal ${\mathfrak{a}}_{R}\subset R$ generated by elements of the skeleton $S\subset R$. Because the pull-back ${\widetilde{\mathfrak{a}}}:=\Phi_{p}^{-1}({\mathfrak{a}})$ satisfies the conditions above and contains $p\in{\mathbb{Z}}\subset{\widetilde{R}}$, a unit element of $\widetilde{R}$}.
This gives a trivial example that a lifting of a prime ideal in $R$ is not necessarily prime. For a prime ideal, we define a “prime lifting ” and “weak prime lifting” which are again a prime ideals in the skeleton rings.
Definition 2.6.
Let $S\subset R$ and ${\widetilde{S}}\subset{\widetilde{R}}$ be compatible skeletons for finitely generated algebras $R$ and $\widetilde{R}$} over fields $k$ and $K$ of characteristic $p>0$ and of 0, respectively. Let ${\mathfrak{p}}\subset S$ and ${\widetilde{{\mathfrak{p}}}}\subset{\widetilde{S}}$ be prime ideals.
  1. (i)
    we call $\widetilde{{\mathfrak{p}}}$} a prime lifting of $\mathfrak{p}$} in the skeleton level if $\Phi_{p}({\widetilde{{\mathfrak{p}}}})={\mathfrak{p}}$ and $0pt{\mathfrak{p}}=0pt{\widetilde{{\mathfrak{p}}}}$.
  2. (ii)
[…]
When a prime ideal ${\mathfrak{p}}_{R}\subset R$ is generated by a prime ideal ${\mathfrak{p}}\subset S$, a prime lifting (resp. weak prime lifting) $\widetilde{{\mathfrak{p}}}$} of $\mathfrak{p}$} is called a prime lifting (resp. weak prime lifting) of ${\mathfrak{p}}_{R}$ in the skeleton level. Sometimes we omit the term “in the skeleton level” when it is clear that we are working in the skeletons.
Remark 2.7.
The following are equivalent:
  1. (1)
    A prime ideal ${\widetilde{{\mathfrak{p}}}}\subset{\widetilde{S}}$ is a weak prime lifting of ${\mathfrak{p}}\subset S$;
  2. (2)
    There exist an element $g\in S\setminus{\mathfrak{p}}$ together with a lifting $\tilde{g}\in\tilde{S}$ of $g$ such that $\tilde{\mathfrak{p}}_{\tilde{g}}$ is a prime lifting of ${\mathfrak{p}}_{g}$;
  3. (3)
    The inclusions ${\widetilde{{\mathfrak{p}}}}\subset\Phi_{p}^{-1}({\mathfrak{p}})\subset{\widetilde{S}}$ hold with the equality $0pt{\mathfrak{p}}=0pt{\widetilde{{\mathfrak{p}}}}$. And the canonical homomorphism ${\widetilde{S}}_{\Phi_{p}^{-1}({\mathfrak{p}})}\to S_{\mathfrak{p}}$ induces the surjection
    ${\widetilde{{\mathfrak{p}}}}{\widetilde{S}}_{\Phi_{p}^{-1}({\mathfrak{p}})}\to{\mathfrak{p}}S_{\mathfrak{p}}.$
  4. (4)
    If ${\widetilde{{\mathfrak{p}}}}\subset{\widetilde{S}}$ and ${\mathfrak{p}}\subset S$ define the irreducible subschemes $\widetilde{C}$} and $C$ (e.g., as $\mathrm{Spec\ }$} or $\mathrm{Proj\ }$}), respectively, such that $\widetilde{C}$} is dominant to $\mathrm{Spec\ }$$\mathbb{Z}$}. It follows
    ${\widetilde{C}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}=C\cup C^{\prime},$
    where $C^{\prime}$ is a closed subscheme whose support does not contain $C$.
Note that $C^{\prime}=\emptyset$ in (4) if and only if $\widetilde{{\mathfrak{p}}}$} is a prime lifting of $\mathfrak{p}$}. In this case we say that $\widetilde{C}$} is a prime lifting of $C$ in $S$ and $\widetilde{S}$}.
In general, a prime lifting or a weak prime lifting in the skeleton level does not exist. However, it does exist under certain conditions. Constructing a good weak prime lifting is one of the central ingredients of our approach. The following is the simplest example for prime lifting in the skeleton level, which shows an important role in the proof of Proposition 4.9, (1).
Proposition 2.8.
Let $k$ be an algebraically closed field of characteristic $p>0$ with a skeleton $\Sigma\subset k$. Let $\widetilde{\Sigma}$$\subset{\mathbb{C}}$ be a compatible skeleton of $\Sigma$. Let $R$ be a finitely generated $k$-algebra of dimension $N\geq 2$, and $\widetilde{R}$} a finitely generated $\mathbb{C}$}-algebra. Assume
${\widetilde{R}}({\operatorname{mod}\ }p)=R$
with the compatible skeletons $S\subset R$ and ${\widetilde{S}}\subset{\widetilde{R}}$ such that these are $\Sigma$-subalgebra and $\widetilde{\Sigma}$}-subalgebra of $R$ and of $\widetilde{R}$}, respectively.
Let $X:={\mathrm{Spec\ }}R$, and $x\in X$ a regular closed point. Assume the defining ideal ${\mathfrak{m}}_{x}\subset R$ is generated by $N$ elements of $S$, then ${\mathfrak{m}}_{x}$ has a prime lifting in the skeleton level.
Proof.
By the assumption, the maximal ideal ${\mathfrak{m}}_{x}\subset R$ is generated by $N$ elements $f_{1},\ldots,f_{N}\in S$ $(N=\dim R)$. Denote ${\mathfrak{m}}={\mathfrak{m}}_{x}\cap S$. By the construction, it follows ${\mathfrak{m}}\cap\Sigma=\{0\}$. For each $1\leq i\leq N$, take a lifting, ${\tilde{f}}_{i}\in{\widetilde{S}}$. Note that these elements are not in $\widetilde{\Sigma}$}. As $\{f_{1},\ldots,f_{N}\}$ is a regular sequence in $S_{\mathfrak{m}}$, the lifting
$\{{\tilde{f}}_{1},\ldots,{\tilde{f}}_{N}\}\subset{\widetilde{S}}_{\Phi_{p}^{-1}({\mathfrak{m}})}$
forms a regular sequence and also
$W:={\widetilde{S}}_{\Phi_{p}^{-1}({\mathfrak{m}})}/({\tilde{f}}_{1},\ldots,{\tilde{f}}_{N})$
is $\mathbb{Z}$}-flat at $p$ by [16, Corollary p. 177]. As $W\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}=S_{\mathfrak{m}}/(f_{1},\ldots,f_{N})$ is an integral domain, $p\in W$ is a prime element and therefore $W$ is an integral domain. It implies that $({\tilde{f}}_{1},\ldots,{\tilde{f}}_{N})$ is prime in ${\widetilde{S}}_{\Phi_{p}^{-1}({\mathfrak{m}})}$. Define ${\widetilde{\mathfrak{m}}}:=({\tilde{f}}_{1},\ldots,{\tilde{f}}_{N}){\widetilde{S}}_{\Phi_{p}^{-1}({\mathfrak{m}})}\cap{\widetilde{S}}$, then it is a prime ideal of $\widetilde{S}$} of height $N$ and $\Phi_{p}({\widetilde{\mathfrak{m}}})={\mathfrak{m}}$. ∎
Corollary 2.9.
Under the condition of Proposition 2.8, assume $H\subset X$ is a smooth subvariety, and its defining ideal is $I_{H}$ has a prime lifting ${\widetilde{I}}\subset{\widetilde{S}}$ in the skeleton level. If a closed point $x\in X$ is contained in $H$, i.e.,
$I_{H}\subset{\mathfrak{m}}_{x},$
then there is a prime lifting ${\widetilde{\mathfrak{m}}}\subset{\widetilde{S}}$ of the defining ideal ${\mathfrak{m}}_{x}$ of $x$, such that
${\widetilde{I}}\subset{\widetilde{\mathfrak{m}}}$.
Proof.
By the assumption that $I_{H}$ has a prime lifting, we may assume that $I_{H}\subset R$ is generated by $I=I_{H}\cap S\subset S$ and ${\mathfrak{m}}_{x}$ is generated by ${\mathfrak{m}}={\mathfrak{m}}_{x}\cap S$. Then, $S^{\prime}=S/I$ is a skeleton of $R/I_{H}$ which is the affine ring of smooth variety $H$. We can regard ${\widetilde{S}}^{\prime}={\widetilde{S}}/{\widetilde{I}}$ as a compatible skeleton of $S^{\prime}$. Apply Proposition 2.8 to ${\mathfrak{m}}^{\prime}={\mathfrak{m}}/I\subset S^{\prime}$, then we obtain its prime lifting
${\widetilde{\mathfrak{m}}}^{\prime}\subset{\widetilde{S}}^{\prime}$,
[…]
Remark 2.10.
If a regular prime ideal ${\mathfrak{p}}_{R}\subset R$ is generated by ${\mathfrak{p}}:={\mathfrak{p}}_{R}\cap S$, then $\mathfrak{p}$} is again a regular prime ideal of the same height by Lemma 2.11, (ii). In this case, a weak prime lifting ${\widetilde{{\mathfrak{p}}}}\subset{\widetilde{S}}$ of $\mathfrak{p}$} is also a regular prime ideal. Because an RSP $f_{1},\ldots,f_{c}$ $(c=0pt{\mathfrak{p}})$ of $\mathfrak{p}$} admits an regular sequence ${\tilde{f}}_{1},\ldots,{\tilde{f}}_{c}$ $(c=0pt{\mathfrak{p}})$ in $\widetilde{{\mathfrak{p}}}$}. As $c=0pt{\mathfrak{p}}=0pt{\widetilde{{\mathfrak{p}}}}$, the local ring ${\widetilde{S}}_{\widetilde{{\mathfrak{p}}}}$ is regular.
In this paper, we work on algebras of the special type: i.e., a finitely generated $k$-subalgebra $R$ of a polynomial ring. On such an algebra $R$, properties on an ideal $\mathfrak{a}$} of $R$ are reduced to the properties of its skeleton ideal ${\mathfrak{a}}\cap S$ as follows:
Lemma 2.11.
Let $K$ be a field and $\Sigma$ a skeleton of $K$. Denote the field of fractions of $\Sigma$ by $Q$. Let $S\subset\Sigma[x_{1},\ldots,x_{m}]$ be a finitely generated $\Sigma$-subalgebra of the polynomial ring and ${\mathfrak{a}}\subset S$ an ideal which does not contain any nonzero element of $\Sigma$. Let
$Q[S]\subset Q[x_{1},\ldots,x_{m}]\ \ \mbox{ and}\ \ K[S]\subset K[x_{1},\ldots,x_{m}]$
be the $Q$- and $K$-subalgebras generated by $S$, respectively. Let
${\mathfrak{a}}_{Q}\subset Q[S]\ \ \mbox{ and}\ \ {\mathfrak{a}}_{K}\subset K[S]$
be the ideals generated by $\mathfrak{a}$}. Then, it follows:
  1. (i)
    $0pt{\mathfrak{a}}\leq 0pt_{+}{\mathfrak{a}}=0pt{\mathfrak{a}}_{Q}=0pt{\mathfrak{a}}_{K},$ where $0pt_{+}{\mathfrak{a}}$ is the minimal height of the associated primes $\mathfrak{p}$} such that ${\mathfrak{p}}\cap\Sigma=\{0\}$.
  2. (ii)
    If $\mathfrak{a}$} is a prime ideal, then so is ${\mathfrak{a}}_{Q}$, while ${\mathfrak{a}}_{K}$ is not necessarily prime. In this case, it follows:
    $0pt{\mathfrak{a}}=0pt{\mathfrak{a}}_{Q}=0pt{\mathfrak{a}}_{K}.$
    Moreover, if ${\mathfrak{a}}_{K}\subset K[S]$ is a regular prime ideal, then ${\mathfrak{a}}\subset S$ is regular.
  3. (iii)
    Assume ${\mathfrak{a}}\subset S$ is a prime ideal. For an element $f\in S$ which is not in $\Sigma$, the following are equivalent for each $\in{\mathbb{Z}}_{>0}$:
    1. (a)
      $f\in{\mathfrak{a}}^{n}\setminus{\mathfrak{a}}^{n+1}$ at the generic point of $\mathfrak{a}$}, which means $f\in{\mathfrak{a}}^{n}S_{\mathfrak{a}}\setminus{\mathfrak{a}}^{n+1}S_{a}$,
    2. (b)
      $f\in{\mathfrak{a}}_{Q}^{n}\setminus{\mathfrak{a}}_{Q}^{n+1}$ at the generic point of ${\mathfrak{a}}_{Q}$,
    3. (c)
      In case ${\mathfrak{a}}_{K}$ is prime, $f\in{\mathfrak{a}}_{K}^{n}\setminus{\mathfrak{a}}_{K}^{n+1}$ at the generic point of ${\mathfrak{a}}_{K}$.
  4. (iv)
    In particular, for ${\operatorname{char}}K=0$, assume ${\mathfrak{a}}_{Q}$ is a locally principal prime ideal then, the following is also equivalent to (a), (b) above:
    1. (d)
      Every minimal prime $\mathfrak{p}$} of ${\mathfrak{a}}_{K}$ is locally principal at the generic point of $\mathfrak{p}$} and $f\in{\mathfrak{p}}^{n}\setminus{\mathfrak{p}}^{n+1}$.
Proof.
As $Q\hookrightarrow K$ is faithfully flat, it follows $Q[S]\hookrightarrow K[S]=K\otimes_{Q}Q[S]$ is also faithfully flat, in particular it is flat. Then, by [9, III, Corollary 9.6] (or algebraically [16, Theorem 15.1]), we have the height of a minimal prime $\mathfrak{p}$} of ${\mathfrak{a}}_{Q}$ is equal to the height of a minimal prime of the extension ${\mathfrak{p}}_{K}$ of $\mathfrak{p}$}. Therefore it follows:
$0pt{\mathfrak{a}}_{Q}=0pt({\mathfrak{a}}_{Q}K[S])=0pt{\mathfrak{a}}_{K}.$
This gives the last equality in (i).
Since for every $g\in Q[S]$ there exists an element $r\in\Sigma$ such that $rg\in S$, we have the following bijective correspondence:
$\left\{\begin{array}[]{l}\mbox{prime ideals\ of\ }S\\ \mbox{without\ nonzero\ elements\ in\ }\Sigma\\ \end{array}\right\}$ $\leftrightarrow$ $\left\{\begin{array}[]{l}\mbox{prime\ ideals\ of\ }\\ Q[S]\\ \end{array}\right\}$
by associating ${\mathfrak{p}}\subset S$ to ${\mathfrak{p}}_{Q}\subset Q[S]$ in the forward direction and ${\mathfrak{q}}\subset Q[S]$ to ${\mathfrak{q}}\cap S\subset S$ in the backward direction. Here, the inclusion relations are preserved under the correspondence. Therefore, we have the bijective correspondence between minimal associated primes of ${\mathfrak{a}}_{Q}$ and those of $\mathfrak{a}$} without nonzero elements of $\Sigma$, which gives the middle equality of (i). The first inequality if (i) is trivial, which completes the proof of (i). By applying this discussion for the case that $\mathfrak{a}$} is a prime ideal, it follows ${\mathfrak{a}}_{Q}$ is prime and
$0pt{\mathfrak{a}}=0pt_{+}{\mathfrak{a}}=0pt{\mathfrak{a}}_{Q},$
which show the first and second statement of (ii). For the last statement of (ii), note that $S\hookrightarrow K[S]$ is flat, as it is a base change of the flat homomorphism $\Sigma\hookrightarrow K$. Then, regularity of ${\mathfrak{a}}_{K}$ yields regularity of ${\mathfrak{a}}={\mathfrak{a}}_{K}\cap S$ (see, for example, [16, Theorem 15.1,(ii)]).
As for the proof of the equivalence of (a) and (b) in (iii), it is sufficient to prove the equivalence for $f\in S$:
$f\in{\mathfrak{a}}^{n}\Leftrightarrow f\in{\mathfrak{a}}_{Q}^{n}$
at the generic point of $\mathfrak{a}$} and ${\mathfrak{a}}_{Q}$, respectively.
As $\Sigma\setminus\{0\}\subset S\setminus{\mathfrak{a}}$, it follows
$(S_{Q})_{{\mathfrak{a}}_{Q}}=S_{\mathfrak{a}}\ \ \mbox{and}\ \ \ {\mathfrak{a}}_{Q}(S_{Q})_{{\mathfrak{a}}_{Q}}={\mathfrak{a}}S_{\mathfrak{a}}.$
Additionally, commutativity of multiplication and localization yields:
${\mathfrak{a}}^{n}_{Q}(S_{Q})_{{\mathfrak{a}}_{Q}}={\mathfrak{a}}^{n}S_{\mathfrak{a}}$
for every $n\in{\mathbb{Z}}_{>0}$, which gives the equivalence of (a) and (b).
As for the equivalence (b) and (c), by the faithful flatness of $Q\hookrightarrow K$ it follows ${\mathfrak{a}}_{K}^{n}\cap Q[S]={\mathfrak{a}}_{Q}^{n}$. In case ${\mathfrak{a}}_{K}$ is prime, then (b) holds if and only if $f\in{\mathfrak{a}}_{K}^{n}\setminus{\mathfrak{a}}_{K}^{n+1}$ around ${\mathfrak{a}}_{K}$, which yields (c).
To prove (iv), first note that $Q[S]\hookrightarrow K[S]$ is flat, as it is a base change of the faithfully flat homomorphism $Q\hookrightarrow K$. Therefore, it follows $0pt{\mathfrak{a}}_{Q}=0pt{\mathfrak{p}}$ for every minimal prime $\mathfrak{p}$} of ${\mathfrak{a}}_{K}$ by [16, Theorem 15.1]. As $0pt{\mathfrak{a}}_{Q}=1$ around ${\mathfrak{a}}_{Q}$, it follows $0pt{\mathfrak{p}}=1$ around $\mathfrak{p}$}. On the other hand, $Q\hookrightarrow Q[S]$ is smooth around ${\mathfrak{a}}_{Q}$, because the local ring $Q[S]_{{\mathfrak{a}}_{Q}}$ is regular with dimension one. Then, the base change $K\hookrightarrow K[S]$ is also smooth around every minimal prime of ${\mathfrak{a}}_{K}$, in particular, every minimal prime $\mathfrak{p}$} of ${\mathfrak{a}}_{K}$ is regular and of height one, therefore $\mathfrak{p}$} is principal at the generic point of $\mathfrak{p}$}. As the natural local homomorphism
$Q[S]_{{\mathfrak{a}}_{Q}}\to K[S]_{\mathfrak{p}}$
is flat, and ${\rm char}K=0$, it follows that the maximal ideal of $K[S]_{\mathfrak{p}}$ is the extension of the maximal ideal of $Q[S]_{{\mathfrak{a}}_{Q}}$ (see for example [18, Lemma 33.5.3]). Now, we have that the generator of ${\mathfrak{a}}_{Q}Q[S]$ becomes the generator of ${\mathfrak{p}}K[S]$. Thus, $f\in{\mathfrak{a}}_{Q}^{n}$ at the generic point of ${\mathfrak{a}}_{Q}$ if and only if $f\in{\mathfrak{p}}^{n}$ at the generic point of $\mathfrak{p}$}. This shows the equivalence of (b) and (d) in case ${\rm char}K=0$. ∎
Remark 2.12.
Lemma 2.11 also holds for the case that $S\subset\Sigma[x_{1},\ldots,x_{N}]$ have a graded algebra structure and ideals are homogeneous ideals, where the degree is given by a weighted degree of $x_{i}$’s.
Lemma 2.13.
Let $k$ be a field, and $\Sigma\subset k$ a skeleton. Let $R$ be a finitely generated $k$-subalgebra of $k[x_{1},\ldots,x_{m}]$ and $S\subset R$ a skeleton such that it is a $\Sigma$-subalgebra of $\Sigma[x_{1},\ldots,x_{m}]$. Let
$f_{i}\in R\ \ (i=1,\ldots,n)$
be any finite number of elements of $R$. Then, by replacing $\Sigma$ by a larger skeleton $\Sigma^{\prime}\supset\Sigma$, and $S$ by $\Sigma^{\prime}[S]$. We obtain a new skeleton $S\subset R$ such that
$f_{i}\in S\ \ \ (i=1,\ldots,n).$
Proof.
As $S$ is finitely generated over ${\mathbb{Z}}/(p)$, where $p={\operatorname{char}}\ k$, $S$ is finitely generated also over $\Sigma$ so that we can express, $S=\Sigma[s_{1},..,s_{n}]$. Then, $R=k[s_{1},..,s_{n}]$, because $R=k[S]$. Therefore, an element $f\in R$ is expressed as
$f=F(s_{1},...,s_{n}),\ \mbox{where}\ \ F(Y_{1},\ldots,Y_{n})\in k[Y_{1},\ldots,Y_{n}],$
We call the coefficients of $F$ “coefficients of ff.” Let $\{a_{i}\}\subset k$ be the set of all coefficients of $F$. Note that it is a finite set. If these already belong to $\Sigma$, then $f\in S$. If not, let $\Sigma^{\prime}$ be $\Sigma[a_{i}]$. Then, $S^{\prime}:=\Sigma^{\prime}[S]$ becomes a larger skeleton of $R$ with $f\in S^{\prime}$. In this way, adding the coefficients of the finite number of elements $\{f_{i}\}$ of $R$ to the original $\Sigma$, we obtain a new skeleton $S$ of $R$ containing all $\{f_{i}\}$. ∎
This extension method is used repeatedly throughout the paper. When we use this method, we say “by adding the coefficients” or “by enlarging the coefficient skeleton”.
Remark 2.14.
If the original skeleton $S\subset\Sigma[x_{1},\ldots,x_{m}]$ has a lifting skeleton ${\widetilde{S}}\subset{\widetilde{\Sigma}}[x_{1},\ldots,x_{m}]$, then an enlarging $\Sigma^{\prime}\supset\Sigma$ of the coefficient skeleton induces an enlarged skeleton ${\widetilde{\Sigma}}^{\prime}\supset{\widetilde{\Sigma}}$ by Example 2.3, which yields enlarged compatible skeletons
${\widetilde{S}}^{\prime}:={\widetilde{\Sigma}}^{\prime}[{\widetilde{S}}]\to S^{\prime}:=\Sigma^{\prime}[S]$
Proposition 2.15.
Let $k$ be a field of characteristic $p>0$. Let $R\subset k[x_{1},\ldots,x_{m}]$ be a finitely generated $k$-subalgebra and ${\widetilde{R}}\subset{\mathbb{C}}[x_{1},\ldots,x_{m}]$ a finitely generated $\mathbb{C}$}-algebra. Assume ${\widetilde{R}}({\operatorname{mod}\ }p)=R$ with the compatible skeletons $S\subset R$ and ${\widetilde{S}}\subset{\widetilde{R}}$ such that these are $\Sigma$-subalgebra and $\widetilde{\Sigma}$}-subalgebra of $\Sigma[x_{1},\ldots,x_{m}]$ and ${\widetilde{\Sigma}}[x_{1},\ldots,x_{m}]$, respectively. Let ${\mathfrak{a}}\subset R$ be a non-zero proper ideal and ${\widetilde{\mathfrak{a}}}\subset{\widetilde{R}}$ its lifting with respect to these compatible skeletons $S$ and $\widetilde{S}$}. Then,
$0pt{\mathfrak{a}}\leq 0pt{\widetilde{\mathfrak{a}}}$.
Proof.
Let ${\mathfrak{p}}_{1},\ldots,{\mathfrak{p}}_{s}$ be the minimal associated primes of ${\mathfrak{a}}\subset R$. Then, $\underline{{\mathfrak{p}}_{i}}:={\mathfrak{p}}_{i}\cap S$’s are the minimal associated primes of $\underline{{\mathfrak{a}}}={\mathfrak{a}}\cap S$ and satisfies $\underline{{\mathfrak{p}}_{i}}\cap\Sigma=\{0\}$, since otherwise ${\mathfrak{p}}_{i}$ would contain a unit in $R$, a contradiction. Here, we may assume that $\Sigma$ contains all coefficients of the generators of $\underline{{\mathfrak{p}}}_{i}$’s by enlarging the coefficient skeleton $\Sigma$ as in Lemma 2.13. Therefore, ${\mathfrak{p}}_{i}\subset R$ is the extensions of $\underline{{\mathfrak{p}}}_{i}\subset S$, and by Lemma 2.11, (i), it follows
$0pt{\mathfrak{a}}=0pt\underline{{\mathfrak{a}}}.$
On the other hand, let $\underline{{\widetilde{\mathfrak{a}}}}\subset{\widetilde{S}}$ be the skeleton ideal of ${\widetilde{\mathfrak{a}}}\subset{\widetilde{R}}$, and $\widetilde{{\mathfrak{p}}}$} a minimal associated prime of ${\widetilde{\mathfrak{a}}}\subset{\widetilde{R}}$ such that $0pt{\widetilde{\mathfrak{a}}}=0pt{\widetilde{{\mathfrak{p}}}}$.
Then, as $\widetilde{R}$} is flat over $\widetilde{S}$}, it follows
(1) $0pt{\widetilde{{\mathfrak{p}}}}\geq 0pt\underline{{\widetilde{{\mathfrak{p}}}}}$,
where $\underline{{\widetilde{{\mathfrak{p}}}}}={\widetilde{{\mathfrak{p}}}}\cap{\widetilde{S}}$ (see, for example, [16, Theorem 15.1, (ii)]). As $\underline{{\widetilde{{\mathfrak{p}}}}}\cap{\widetilde{\Sigma}}=\{0\}$, a fortiori $\underline{{\widetilde{{\mathfrak{p}}}}}\cap{\mathbb{Z}}=\{0\}$ holds, which yields
[…]
where the second equality follows from the flatness of $\widetilde{S}$} over $\mathbb{Z}$}. Here, since $\Phi_{p}(\underline{{\widetilde{{\mathfrak{p}}}}})\supset\Phi_{p}(\underline{{\widetilde{\mathfrak{a}}}})=\underline{{\mathfrak{a}}}$, it follows
(3) $0pt\Phi_{p}(\underline{{\widetilde{{\mathfrak{p}}}}})\geq 0pt\underline{{\mathfrak{a}}}.$
By connecting inequalities (1), (2) and (3), we obtain $0pt{\widetilde{\mathfrak{a}}}\geq 0pt{\mathfrak{a}}$. ∎
3. log discrepancies for multi-ideals and log canonical thresholds of ideals
In this section we study a pair $(A,{\mathfrak{a}}^{e})$ consisting of a smooth affine variety $A$ of dimension $N$ defined over a field of arbitrary characteristic and a “multi-ideal” ${\mathfrak{a}}^{e}={\mathfrak{a}}_{1}^{e_{1}}\cdots{\mathfrak{a}}_{r}^{e_{r}}$ on $A$, where ${\mathfrak{a}}_{i}$’s are non-zero coherent ideal sheaves on $A$ and $e$ is a combination of the exponents $e=(e_{1},\ldots,e_{r})\in{\mathbb{R}}_{>0}^{r}$.
Definition 3.1.
Let $A$ be a normal variety defined over a field $k$ and $P\in A$ a (not necessarily closed) point.
Let $\varphi_{1}:A_{1}\to A$ be a proper birational morphism with $A_{1}$ normal and $E_{1}\subset A_{1}$ an irreducible divisor. Let $\varphi_{2}:A_{2}\to A$ be another proper birational morphism from a normal $A_{2}$ with an irreducible divisor $E_{2}\subset A_{2}$. Define
$E_{1}\sim E_{2}$
if the birational map $\varphi_{2}^{-1}\circ\varphi_{1}:A_{1}\dasharrow A_{2}$ is a local isomorphism at the generic points of $E_{1}$ and $E_{2}$. Then, $\sim$ becomes an equivalence relation, and such an equivalence class is called a prime divisor over $A$. In this case we denote $E_{1}$ and $E_{2}$ by the same symbol, say $E$, and we say “$E$ appears on $A_{1}$ and on $A_{2}$”. (Strictly speaking, we should be talking about the corresponding valuation instead.) If $P$ is the generic point of $\varphi_{1}(E)$, then we call $E$ a prime divisor over $A$ with the center at $P$ or just a prime divisor over $(A,P)$.
Definition 3.2.
Let $(A,{\mathfrak{a}}^{e})$ be a pair as in the beginning of this section. The log discrepancy of such a pair at a prime divisor $E$ over $A$ is defined as
$a(E;A,{\mathfrak{a}}^{e}):=k_{E}+1-\sum_{i=1}^{r}e_{i}\cdot v_{E}({\mathfrak{a}}_{i}),$
where $k_{E}$ is the coefficient of the relative canonical divisor $K_{{A}^{\prime}/A}$ at $E$.
Here, $\varphi:{A^{\prime}}\to A$ is a birational morphism such that $E$ appears on a normal variety $A^{\prime}$}. The valuation of an ideal ${\mathfrak{a}}_{i}$ at $E$ is defined as follows:
$v_{E}({\mathfrak{a}}_{i}):=\min\{v_{E}(f)\mid f\in{\mathfrak{a}}_{i}\}$
We say that the pair $(A,{\mathfrak{a}}^{e})$ is log canonical at a (not necessarily closed) point $P\in A$ if
(4) $a(E;A,{\mathfrak{a}}^{e})\geq 0,$
holds for every exceptional prime divisor $E$ over $A$ with the center containing $P\in A$.
Definition 3.3.
Let $A$ be a smooth variety, $P\in A$ a point and ${\mathfrak{a}}^{e}$ a non-zero multi-ideal on $A$. We define the log canonical threshold of the pair $(A,{\mathfrak{a}}^{e})$ at $P$ as follows:
${\operatorname{lct}}(P;A,{\mathfrak{a}}^{e})=\sup\{c\in{\mathbb{R}}_{>0}\mid(A,({\mathfrak{a}}^{e})^{c})\ \mbox{is\ log \ canonical\ at \ }P\}.$
Let $E$ be a prime divisor over $A$ whose center on $A$ is contained in the locus of $\mathfrak{a}$}. Define
$z(E;A,{\mathfrak{a}}^{e}):=\frac{k_{E}+1}{v_{E}({\mathfrak{a}}^{e})},$
where, we define $v_{E}({\mathfrak{a}}^{e}):=\sum_{i}e_{i}\cdot v_{E}({\mathfrak{a}}_{i})$. Then, the following is well known:
[…]
Definition 3.4.
Let $(A,{\mathfrak{a}}^{e})$ and $P\in A$ be as in Definition 3.2. Then, the minimal log discrepancy is defined as follows:
  1. (1)
    When $\dim A\geq 2$,
    ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e}):=\inf\{a(E;A,{\mathfrak{a}}^{e})\mid E:\ {\operatorname{prime\ divisor\ over}}\ (A,P)\}.$
  2. (2)
    When $\dim A=1$, define ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})$ by the same definitions as in (1) if the right hand side of the equality (1) is non-negative and otherwise define ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})=$-$\infty$.
Here, we remark that for every pair $(A,{\mathfrak{a}}^{e})$, either ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})\geq 0$ or ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})=$-$\infty$ holds.
Definition 3.5.
Let $A$, NN, and ${\mathfrak{a}}^{e}$ and $P\in A$ be as in Definition 3.2. We say that a prime divisor $E$ over $(A,P)$ computes ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})$, if
[…]
Let $A$, $P$ and ${\mathfrak{a}}^{e}$ be as in Definition 3.3. Let $E$ be a prime divisor over $A$ with the center containing $P$. We say that a prime divisor $E$ computes ${\operatorname{lct}}(P;A,{\mathfrak{a}}^{e})$ if
$z(E;A,{\mathfrak{a}}^{e})={\operatorname{lct}}(P;A,{\mathfrak{a}}^{e}).$
Remark 3.6.
Let ${\mathfrak{m}}_{P}\subset{\mathcal{O}}_{A}$ be the ideal defining $P$. If there is a log resolution of $(A,{\mathfrak{m}}_{P}\cdot{\mathfrak{a}}^{e})$ in a neighborhood of $P$, or if $e$ is a set of rational numbers, then there exists a prime divisor computing ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})$. Otherwise, the existence of such a divisor was not known in general. We will show the existence of a prime divisor computing ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})$ for a smooth point of arbitrary dimension as a corollary of our main theorem (Corollary 1.2, (i)).
Similarly, if there is a log resolution of $(A,{\mathfrak{a}}^{e})$ in a neighborhood of $P$, then there exists a prime divisor over $A$ computing ${\operatorname{lct}}(P;A,{\mathfrak{a}}^{e})$. However, even in this case a prime divisor computing ${\operatorname{lct}}(P;A,{\mathfrak{a}}^{e})$ is not necessarily a prime divisor over $(A,0)$, while its center on $A$ just contains $P$.
Definition 3.7.
Let $X$ be a scheme of finite type over a field $k$ and $k^{\prime}\supset k$ a field extension. For $m\in{\mathbb{Z}}_{\geq 0}$ a $k$-morphism ${\mathrm{Spec\ }}k^{\prime}[t]/(t^{m+1})\to X$ is called an $m$-jet of $X$ and $k$-morphism ${\mathrm{Spec\ }}k^{\prime}[[t]]\to X$ is called an arc of $X$.
Let $X_{m}$ be the space of $m$-jets of $X$. It is well known that $X_{m}$ has a scheme structure of finite type over $k$. If $X$ is an affine variety, then there exists the projective limit
$X_{\infty}:=\lim_{\overleftarrow{m}}X_{m}$
and it is called the space of arcs or the space of $\infty$-jet of $X$.
Definition 3.8.
Denote the canonical truncation morphisms induced from $k[[t]]\to k[t]/(t^{m+1})$ and $k[t]/(t^{m+1})\to k$ by $\psi_{m}:X_{\infty}\to X_{m}$ and $\pi_{m}:X_{m}\to X$, respectively. In particular we denote the morphism $\psi_{0}=\pi_{\infty}:X_{\infty}\to X$ by $\pi$.
Definition 3.9.
Let $\mathfrak{a}$} be a non-zero ideal on a variety $X$. We define the subsets “contact loci of $\mathfrak{a}$}” in the space of arcs as follows:
${\mathrm{Cont}}^{\geq m}({\mathfrak{a}})=\{\gamma\in X_{\infty}\mid{\mathrm{ord}}_{\gamma}({\mathfrak{a}}):={\mathrm{ord}}_{t}\gamma^{*}({\mathfrak{a}})\geq m\},$
where $\gamma^{*}:{\mathcal{O}}_{X}\to k^{\prime}[[t]]$ is the homomorphism of rings corresponding to $\gamma$.
Example 3.10.
Let $Z$ be a closed subscheme of affine $N$-space $A:={\mathbb{A}}_{k}^{N}\\ ={\mathrm{Spec\ }}k[x_{1},\ldots,x_{N}]$ defined over a field $k$. Let ${\mathfrak{a}}\subset k[x_{1},\ldots,x_{N}]$ be the defining ideal of $Z$ in $A$. Assume $\mathfrak{a}$} is generated by $f_{1},\ldots,f_{r}\in k[x_{1},\ldots,x_{N}]$. We define polynomials
$F_{i}^{(j)}\in k\left[x_{\ell}^{(q)}\middle|1\leq\ell\leq N,0\leq q\leq j\right]$
so that
$f_{i}\left(\sum_{q\geq 0}x_{1}^{(q)}t^{q},\sum_{q\geq 0}x_{2}^{(q)}t^{q},\ldots,\sum_{q\geq 0}x_{N}^{(q)}t^{q}\right)=F_{i}^{(0)}+F_{i}^{(1)}t+\cdots+F_{i}^{(j)}t^{j}+\cdots.$
Then, the contact locus ${\mathrm{Cont}}^{\geq m}{\mathfrak{a}}\subset A_{\infty}$ is defined by the ideal of $k[x_{\ell}^{(q)}\mid 1\leq\ell\leq N,0\leq q]$ generated by
$F_{i}^{(j)}\ \ (i=1,\ldots,r,j=0,1,\ldots,m-1).$
Here, we note that if all coefficients of $f_{i}$’s are in a subring $\Sigma\subset k$, then all coefficients of $F_{i}^{(j)}$’s are also in $\Sigma$. (This fact will be used in the proof of Lemma 3.12.)
The fiber $\pi_{m}^{-1}(0)$ of the origin $0\in{\mathbb{A}}_{k}^{N}=A$ by the truncation morphism $\pi_{m}:A_{m}\to A$ is defined by the ideal
$(x_{1}^{(0)},\ldots,x_{N}^{(0)})\subset k[x_{\ell}^{(q)}\mid 1\leq\ell\leq N,0\leq q\leq m].$
The following Proposition 3.11 and Lemma 3.12 require perfectness of the base field $k$. As our base field $k$ is algebraic closed, we can apply these formula.
Proposition 3.11.
Let $A$ be a smooth variety defined over a perfect field, $P\in A$ a closed point and $e=(e_{1},\ldots,e_{r})\in{\mathbb{R}}_{>0}^{r}$. Let ${\mathfrak{a}}^{e}$ be a multi-ideal ${\mathfrak{a}}_{1}^{e_{1}}\cdots{\mathfrak{a}}_{r}^{e_{r}}$ on $A$. Then, it follows:
(5) ${\mathrm{mld}}(P;A,{\mathfrak{a}}^{e})=\inf_{m\in{\mathbb{Z}}_{\geq 0}^{r}}\left\{{\mathrm{codim}}\left(\bigcap_{i}({\mathrm{Cont}}^{\geq m_{i}}({\mathfrak{a}}_{i})\cap\pi^{-1}(P),A_{\infty}\right)-\sum_{i=1}^{r}m_{i}e_{i}\right\}.$
(6) ${\operatorname{lct}}(P;A,{\mathfrak{a}}^{e})=\inf_{m\in{\mathbb{Z}}_{\geq 0}}\left\{\frac{{\mathrm{codim}}_{P}\left(\bigcap_{i}{\mathrm{Cont}}^{\geq m_{i}}({\mathfrak{a}}_{i}),A_{\infty}\right)}{\sum_{i}m_{i}e_{i}}\right\},$
where ${\mathrm{codim}}_{P}({\mathrm{Cont}}^{\geq m}({\mathfrak{a}}),A_{\infty})$ is the minimum of codimensions of the irreducible components $T$ of ${\mathrm{Cont}}^{\geq m}({\mathfrak{a}})$ in $A_{\infty}$ such that $\pi(T)\cap\{P\}\neq\emptyset$.
Proof.
The equality (5) is obtained by [6] for characteristic 0 and by [10] for arbitrary characteristic. The equality (6) is obtained by [17] for characteristic 0 and by [20] for arbitrary characteristic. ∎
Lemma 3.12.
[…]
(7) ${\mathrm{mld}}(0;A,{\mathfrak{a}}^{e})\leq{\mathrm{mld}}({\tilde{0}};{\widetilde{A}},{\widetilde{\mathfrak{a}}}^{e}),$
(8) ${\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})\leq{\operatorname{lct}}({\tilde{0}};{\widetilde{A}},{\widetilde{\mathfrak{a}}}^{e}).$
Proof.
By the assumption, there are compatible skeletons $\widetilde{\Sigma}\to\Sigma$ with $\widetilde{\Sigma}\subset{\mathbb{C}}$ and $\Sigma\subset k$ such that the coefficients of the generators of ${\mathfrak{a}}_{i}$ and of ${\widetilde{\mathfrak{a}}}_{i}$ are in $\Sigma$ and in $\widetilde{\Sigma}$, respectively.
First, we prove the following inequalities which are essential for (7) and (8):
[…]
Let ${\mathfrak{a}}_{i}$ be generated by $f_{ih}\in\Sigma[x_{1},\ldots,x_{N}]$ and ${\widetilde{\mathfrak{a}}}_{i}$ by their liftings ${\tilde{f}}_{ih}\in{\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]$ $(i=1,\ldots,r,h=1,\ldots,s_{i})$, then the defining ideal $I$ of
$\bigcap_{i=1}^{r}{\mathrm{Cont}}^{\geq m_{i}}({\mathfrak{a}}_{i})\cap\pi^{-1}(0)$
in $A_{\infty}$ is generated by $x_{\ell}^{(0)}(\ell=1,\ldots,N)$ and
$F_{ih}^{(j)}\in\Sigma\left[x_{\ell}^{(q)}\middle|\ell=1,\ldots,N,q\geq 0\right]\ \ (i=1,\ldots,r,h=1,\ldots,s_{i},j\leq m_{i}-1).$
Then, we can see that the liftings $x_{\ell}^{(0)}(\ell=1,\ldots,N)$ and
$\widetilde{F}_{ih}^{(j)}\in\widetilde{\Sigma}\left[x_{\ell}^{(q)}\middle|\ell=1,\ldots,N,q\geq 0\right]\ \ (i=1,\ldots,r,h=1,\ldots,s_{i},j\leq m_{i}-1)$
generate the ideal $\widetilde{I}$ defining
[…]
in $\widetilde{A}$$_{\infty}$. Here, $\widetilde{F}_{ih}^{(j)}$ is obtained from $\widetilde{f}_{ih}^{(j)}$ as in Example 3.10. So, we note that $\widetilde{F}_{ih}^{(j)}$ is a lifting of $F_{ih}^{(j)}$. By Lemma 2.11 and Proposition 2.15, it follows that $0ptI\leq 0pt\widetilde{I}$, which yields (9). This completes the proof of (7), by using (5) in Proposition 3.11. Here, we note that Proposition 2.15 states for finitely generated $k$-algebra and $\mathbb{C}$}-algebra, while our statement is for infinitely generated algebras. However as the monomials appearing in the generators and the number of generators of $I$ and $\widetilde{I}$ are finite, we can work in the same way a in Proposition 2.15. Concretely, we may assume for $m\gg 0$, the ideal $I\subset k\left[x_{\ell}^{(q)}\middle|\ell=1,\ldots,N,q\geq 0\right]$ is generated by elements of $k\left[x_{\ell}^{(q)}\middle|\ell=1,\ldots,N,0\leq q\leq m\right]$. Let
$I_{m}\subset k\left[x_{\ell}^{(q)}\middle|\ell=1,\ldots,N,0\leq q\leq m\right]$
be the ideal generated by the generators of $I$. In the same way, we obtain the ideal
${\widetilde{I}}_{m}\subset{\mathbb{C}}\left[x_{\ell}^{(q)}\middle|\ell=1,\ldots,N,0\leq q\leq m\right]$
generating $\widetilde{I}$}. As ${\widetilde{I}}_{m}$ is a lifting of $I_{m}$, applying Proposition 2.15, we have
$0ptI_{m}\leq 0pt{\widetilde{I}}_{m},$
which yields
$0ptI\leq 0pt{\widetilde{I}}$,
where we use $0ptI=0ptI_{m}$ and $0pt{\widetilde{I}}=0pt{\widetilde{I}}_{m}$. The inequality (7) follows from this using the formula in Proposition 3.11.
For (8), we need one more step. We note that
${\operatorname{lct}}(0;A,{\mathfrak{a}})=\inf_{m}\frac{{\mathrm{codim}}_{0}(\bigcap_{i}{\mathrm{Cont}}^{\geq m_{i}}({\mathfrak{a}}_{i}),A_{\infty})}{\sum_{i}m_{i}e_{i}}\leq\inf_{m}\frac{{\mathrm{codim}}(\bigcap_{i}{\mathrm{Cont}}^{\geq m_{i}}({\mathfrak{a}}_{i})\cap\pi^{-1}(0),A_{\infty})}{\sum_{i}m_{i}e_{i}},$
because an irreducible component of $\bigcap_{i}{\mathrm{Cont}}^{\geq m_{i}}({\mathfrak{a}}_{i})\cap\pi^{-1}(0)$ is contained in an irreducible component $T$ of $\bigcap_{i}{\mathrm{Cont}}^{\geq m_{i}}({\mathfrak{a}}_{i})$ with $\pi(T)\cap\{0\}\neq\emptyset$. Then, by (9), it follows
$\inf_{m}\frac{{\mathrm{codim}}(\bigcap_{i}{\mathrm{Cont}}^{\geq m_{i}}({\mathfrak{a}}_{i})\cap\pi^{-1}(0),A_{\infty})}{\sum_{i}m_{i}e_{i}}\leq\inf_{m}\frac{{\mathrm{codim}}(\bigcap_{i}{\mathrm{Cont}}^{\geq m_{i}}({\widetilde{\mathfrak{a}}}_{i})\cap\pi^{-1}({\tilde{0}}),{\widetilde{A}}_{\infty})}{\sum_{i}m_{i}e_{i}}$
On the other hand, in characteristic 0, by [5, Lemma 2.6], we have the equality
${\operatorname{lct}}({\tilde{0}};{\widetilde{A}},{\widetilde{\mathfrak{a}}}^{e})=\inf_{m}\frac{{\mathrm{codim}}(\bigcap_{i}{\mathrm{Cont}}^{\geq m_{i}}({\widetilde{\mathfrak{a}}}_{i})\cap\pi^{-1}({\tilde{0}}),{\widetilde{A}}_{\infty})}{\sum_{i}m_{i}e_{i}},$
which yields (8).
4. Construction of a weak lifting of the prime divisor
As we mentioned just after the Corollary 1.2, the corollary holds for the dimension of the variety $N=1$. So, from now on, we assume $N\geq 2$.
Proposition-Definition 4.1.
(Zariski, see e.g.,[14, VI, 1.3]) Let $E$ be a prime divisor over a variety $A^{\prime}$ defined over a field $k$ and appear on $\psi:A^{\prime\prime}\to A^{\prime}$. Let $Q\in A^{\prime\prime}$ be the generic point of $E$ and $Q_{0}\in A^{\prime}$ the image of $Q$. Then, we have a sequence of varieties $A^{\prime}_{0},A^{\prime}_{1},\ldots,A^{\prime}_{n}$ and birational maps as follows:
  1. $(0)$
    $A^{\prime}_{0}=A^{\prime}$, $\psi^{\prime}_{0}=\psi$.
  2. $\cdots$
  3. (i)
    If $\psi^{\prime}_{i}:A^{\prime\prime}\dashrightarrow A^{\prime}_{i}$ is already defined, then let $Q_{i}=\psi_{i}(Q)$. Let $A^{\prime}_{i+1}=Bl_{Q_{i}}A^{\prime}_{i}$ and $\psi^{\prime}_{i+1}:A^{\prime\prime}\dashrightarrow A^{\prime}_{i}$ be the induced map.
  4. $\cdots$
  5. (n)
    Then, ${\mathrm{codim}}(Q_{n},A^{\prime}_{n})=1$ and the final birational map $\psi^{\prime}_{n}:A^{\prime\prime}\dashrightarrow A^{\prime}_{n}$ is the normalization around $Q_{n}$.
Here, $Bl_{Q_{i}}A^{\prime}_{i}$ is the blow-up of $A^{\prime}_{i}$ at $Q_{i}$. Let $\varphi^{\prime}_{i}:A^{\prime}_{i}\to A^{\prime}_{i-1}$ be the blow-up morphism and $E_{i}\subset A^{\prime}_{i}$ be the exceptional divisor dominating $Q_{i}$. We call this sequence $A^{\prime}_{n}\to\cdots\to A^{\prime}_{1}\to A^{\prime}_{0}=A^{\prime}$ the Zariski sequence for $E$ over $A^{\prime}$.
Note that if $Q_{0}\in A^{\prime}$ is a smooth point, then $Q_{i}$’s are all smooth for $i=1,\ldots,n-1$. Indeed, this is proved inductively. As $A^{\prime}_{1}$ is the blow-up at a smooth point $Q_{0}$, it follows that $A^{\prime}_{1}$ and $E_{1}$ are smooth in a neighborhood of ${\psi^{\prime}_{1}}^{-1}(Q_{0})$. Since $Q_{1}\in{\psi^{\prime}_{1}}^{-1}(Q_{0})$, it follows that $Q_{1}$ is again a smooth point. In the same way, we obtain that $Q_{i}$’s are all smooth.
Remark 4.2 (see for example [15]).
Let $A_{n}\to A_{n-1}\to\cdots A_{1}\to A$ be a sequence of blow-ups with the center at subschemes over the projective space $A={\mathbb{P}}_{k}^{N}$. Then, each $A_{i}$ is described as $A_{i}={\mathrm{Proj\ }}R_{i}$, where $R_{i}$ is a graded ring with its degree 0 part $(R_{i})_{0}=k$ as follows:
(10) $R_{i}=\oplus_{\ell\geq 0}k[x_{0},x_{1},\ldots,x_{N},I_{1}T_{1},\ldots,I_{i}T_{i}]_{(\ell g_{0},\ell g_{1},\ldots,\ell g_{i-1},\ell g_{i})},$
$\mbox{for}\ \ g_{0}\gg g_{1}\gg\cdots\gg g_{i}\gg..\gg g_{n},$
where $x_{\iota}$ and $T_{j}$ are indeterminates for all $\iota$ and $j$, and ${I}_{1}\subset R_{0},\ldots,I_{i}\subset R_{i-1},\ldots,I_{n}\subset R_{n-1}$ are the homogeneous ideals defining the center of each blow-up. We define the multi-degree of an element $r\left(\prod_{j=0}^{N}x_{j}^{a_{j}}\right)T_{1}^{e_{1}}\cdots T_{i}^{e_{i}}\in R_{i}$ $(r\in k,\ e_{0}:=\sum_{j=0}^{N}a_{j})$ as follows:
$\deg r\left(\prod_{j=0}^{N}x_{j}^{a_{j}}\right)T_{1}^{e_{1}}\cdots T_{i}^{e_{i}}=(e_{0},\cdots,e_{i}).$
By this degree, the ring $R_{0}[I_{1}T_{1},\ldots,I_{i}T_{i}]$ becomes a multi-graded ring and $R_{0}[I_{1}T_{1},\ldots,I_{i}T_{i}]_{(\ell g_{0},\ldots,\ell g_{i-1},\ell g_{i})}$ is its homogeneous part of degree $(\ell g_{0},\ldots,\ell g_{i-1},\ell g_{i})$. Then, $R_{i}$ is a “diagonal” sub-graded ring of $R_{0}[I_{1}T_{1},\ldots,I_{i}T_{i}]$. From now on, we regard $R_{i}$ as a graded ring with respect to $\ell\in{\mathbb{Z}}_{\geq 0}$ and a homogeneous ideal of $R_{i}$ is in this grading.
Usually the Zariski sequence
$A^{\prime}_{n}\stackrel{{\scriptstyle\varphi^{\prime}_{n}}}{{\longrightarrow}}A^{\prime}_{n-1}\to\cdots\to A^{\prime}_{1}\stackrel{{\scriptstyle\varphi^{\prime}_{1}}}{{\longrightarrow}}A^{\prime}_{0}={\mathbb{A}}_{k}^{N}$
for a divisor $E$ over ${\mathbb{A}}_{k}^{N}$ does not satisfy the vanishing:
$R^{1}(\varphi^{\prime}_{1}\cdots\varphi^{\prime}_{n})_{*}{\mathcal{O}}_{A^{\prime}_{n}}=0$
because $A^{\prime}_{i}$ ($i>0$) is not necessarily normal. In the proof of our main theorem, we need this vanishing. So we modify the Zariski sequence to a sequence with normality of each variety. We now present a fundamental lemma for constructing an alternative sequence for Zariski sequence in the upcoming proposition:
Lemma 4.3.
Let $A$ be an irreducible Noetherian scheme and $Z\subset A$ a closed subscheme with locally a complete intersection in $A$. Let $\psi:A^{\prime}\to A$ be the blow-up by $Z$. Then, the following hold:
  1. (i)
    $\psi^{-1}(Z)\simeq{\mathbb{P}}(I_{Z}/I_{Z}^{2})$, where $I_{Z}$ is the defining ideal of $Z$ in $A$. In particular, the irreducible components of $\psi^{-1}(Z)$ and the irreducible components of $Z$ correspond bijectively.
  2. (ii)
    Assume $A$ is a normal Gorenstein variety over a field $k$, and for any irreducible component $Z^{\prime}$ of $Z$, the singular loci ${\mathrm{Sing}}A$ and ${\mathrm{Sing}}Z$ of $A$ and of $Z,$ respectively, satisfy:
    (11) $Z^{\prime}\not\subset({\mathrm{Sing}}A)\cup({\mathrm{Sing}}Z).$
    Then, $\psi^{-1}(Z)$ is Gorenstein and $A^{\prime}$ is normal and Gorenstein.
Proof.
As the embedding $Z\subset A$ is locally a complete intersection, by [7, IV, 16.9.4], it follows
$\bigoplus_{n\geq 0}I_{Z}^{n}/I^{n+1}_{Z}\simeq S(I_{Z}/I_{Z}^{2}),$
where $I_{Z}/I_{Z}^{2}$ is a locally free ${\mathcal{O}}_{Z}$-module of rank $c={\mathrm{codim}}(Z,A)$. The description yields $\psi^{-1}(Z)\simeq{\mathbb{P}}(I_{Z}/I_{Z}^{2})$ as required in (i). On the other hand, as $\psi^{-1}(Z)$ is a projective bundle on $Z$, its irreducible component is in the form $\psi^{-1}(Z^{\prime})$ for an irreducible component $Z^{\prime}$ of $Z$, which yields the second statement of (i).
For (ii), we use (11) to see that $A^{\prime}$ satisfies $R_{1}$ condition. On the other hand, we have that $Z$ is Gorenstein by the assumption on $Z$ and $A$. Then, it follows $\psi^{-1}(Z)$ is locally isomorphic to $Z\times{\mathbb{P}}^{c-1}$, which implies $\psi^{-1}(Z)$ is Gorenstein. As $\psi^{-1}(Z)$ is locally a hypersurface of the variety $A^{\prime}$, it follows that $A^{\prime}$ is Gorenstein. Applying the Serre’s criterion to $A^{\prime}$, we obtain that $A^{\prime}$ is normal. ∎
For later use, we prove the following vanishing theorem for a blow-up as in Lemma 4.3.
Lemma 4.4.
Let $A$ be a normal variety over a field $k$ and $Z\subset A$ a closed subscheme locally a complete intersection in $A$ with the property (11). Let $\psi:A^{\prime}\to A$ be the blow-up by the center $Z$. Then, it follows:
[…]
Proof.
The equality $\psi_{*}{\mathcal{O}}_{A^{\prime}}={\mathcal{O}}_{A}$ follows immediately from normality of $A$. For the vanishing $R^{i}\psi_{*}{\mathcal{O}}_{A^{\prime}}=0$ $(i>0)$, we remark the following facts:
  1. (1)
    $R^{i}\psi_{*}{\mathcal{O}}_{E}=0$ $(i>0)$ for $E:=\psi^{-1}(Z)={\mathbb{P}}(I_{Z}/I_{Z}^{2})$,
  2. (2)
    $R^{i}\psi_{*}(\mathcal{I}^{j}\otimes_{{\mathcal{O}}_{A^{\prime}}}{\mathcal{O}}_{E})=0$ $(i,j>0)$ for the defining ideal sheaf $\mathcal{I}$ of $E$ in $A^{\prime}$,
  3. (3)
    $R^{i}\psi_{*}\mathcal{I}^{n}=0$ for $n\gg 0$.
(1) follows from that $E\to Z$ is locally isomorphic to the canonical projection $Z\times{\mathbb{P}}^{c-1}\to Z$, where $c={\mathrm{codim}}(Z,A)$. (2) follows from the same reason and $\mathcal{I}$ is relatively very ample on $E$. (3) follows from the Serre’s vanishing for the relatively ample invertible sheaf $\mathcal{I}$ on $A^{\prime}$.
Now consider the $R^{i}\psi_{*}$ of the exact sequences:
$0\to\mathcal{I}^{j+1}\to\mathcal{I}^{j}\to\mathcal{I}^{j}\otimes{\mathcal{O}}_{E}\to 0$
for $j=0,\ldots,n-1$. Then, by (1)–(3), we obtain $R^{i}\psi_{*}{\mathcal{O}}_{A^{\prime}}=0$ $(i>0)$. ∎
Corollary 4.5.
Let $A_{0}$ be a normal Gorenstein variety over a field $k$ and $\psi$ be the composite
$A_{n}\stackrel{{\scriptstyle\varphi_{n}}}{{\longrightarrow}}A_{n-1}\to\cdots A_{1}\stackrel{{\scriptstyle\varphi_{1}}}{{\longrightarrow}}A_{0}$
of blow-ups $\varphi_{i}$ by the center of locally a complete intersection in $A_{i-1}$ satisfying the condition of (ii) of Lemma 4.3. Then, it follows
$R^{i}\psi_{*}{\mathcal{O}}_{A_{n}}=\left\{\begin{array}[]{ll}0&i>1\\ {\mathcal{O}}_{A_{0}}&i=0\\ \end{array}\right.$.
Proof.
As it holds if $i=1$ by Lemma 4.4, we prove the corollary by induction on $n$. Assume $n\geq 2$ and for $\psi^{\prime}:=\varphi_{1}\circ\cdots\circ\varphi_{n-1}$,
$R^{i}\psi^{\prime}_{*}{\mathcal{O}}_{A_{n-1}}=\left\{\begin{array}[]{ll}0&i>1\\ {\mathcal{O}}_{A_{0}}&i=0\\ \end{array}\right.$.
Consider the Leray’s spectral sequence.
$E_{2}^{p,q}=R^{p}\psi^{\prime}_{*}R^{q}{\varphi_{n}}_{*}{\mathcal{O}}_{A_{n}}\Rightarrow E^{p+q}=R^{p+q}(\psi^{\prime}\circ\varphi_{n})_{*}{\mathcal{O}}_{A_{n}}$
As we know that $A_{n-1}$ is normal Gorenstein by Lemma 4.3, it follows $(\varphi_{n})_{*}{\mathcal{O}}_{A_{n}}={\mathcal{O}}_{A_{n-1}}$. Then, we have $E_{2}^{p,q}=0$ for $(p,q)\neq(0,0)$. Therefore $E^{i}=0$ for $i>0$ and $E^{0}={\mathcal{O}}_{A_{0}}$ which give the required equalities. ∎
Proposition 4.6.
Let $k$ be an algebraically closed field of characteristic $p\geq 0$ and $E$ a prime divisor over the affine space $A={\mathbb{A}}_{k}^{N}$ over $k$ with the center at the origin $0\in A$. Then, there exists a sequence:
$A_{n}\stackrel{{\scriptstyle\varphi_{n}}}{{\longrightarrow}}A_{n-1}\to\cdots A_{1}\stackrel{{\scriptstyle\varphi_{1}}}{{\longrightarrow}}A$
of blow-ups such that
  1. (i)
    $E$ appears on $A_{n}$,
  2. (ii)
    the first blow-up $\varphi_{1}$ is at the origin of ${\mathbb{A}}_{k}^{N}$ i.e. by the maximal ideal ${\mathfrak{m}}=(x_{1},\ldots,x_{N})\subset k[x_{1},\ldots,x_{N}]$, and if the center of $E$ on $A_{i-1}$ is a closed point, $\varphi_{i}:A_{i}\to A_{i-1}$ is the blow-up at the point,
  3. (iii)
    for each $i$ such that the center of $E$ on $A_{i-1}$ is not a closed point, the morphism $\varphi_{i}:A_{i}\to A_{i-1}$ is the blow-up along a closed subscheme $Z_{i}\subset A_{i-1}$ of locally a complete intersection in $A_{i-1}$. One of the irreducible components of $Z_{i}$ is the center of $E$ on $A_{i-1}$, while the other component, if exist, are not contained in the fiber over $0\in A(=:A_{0})$ on $A_{i-1}$,
  4. (iv)
    $A_{i}$’s are all normal Gorenstein varieties and smooth at the generic point of the center of $E$ and also smooth outside of the fiber over $0\in A$.
Proof.
We need the blow-up sequence only over $A={\mathbb{A}}_{k}^{N}$, however by a technical reason we compactify it to $\overline{A}={\mathbb{P}}_{k}^{N}$ with the identification
$A\subset{\overline{A}},\ \ \ (x_{1},\ldots,x_{N})\mapsto(1:x_{1}:\cdots:x_{N})$
and will construct the sequence over $\overline{A}$}. Then, we can restrict it over $A$ to obtain the required sequence of the proposition.
Let $\varphi_{1}:{\overline{A}}_{1}\to{\overline{A}}$ be the blow-up by $(x_{1},\ldots,x_{N})\subset{\overline{R}}_{0}:=k[x_{0},x_{1},\cdots,x_{N}]$, the ideal defining $0:=(1:0:\ldots:0)\in A\subset{\overline{A}}$, where 0 is the center of $E$ on $A$. (automatically (ii) and (iii) are satisfied for $i=1$). As ${\overline{A}}_{1}$ is smooth, (iv) is satisfied for $i=1$. We will construct $\varphi_{i}$ by induction on $i$ so that (iii) and (iv) are satisfied.
Let $i\geq 2$ and assume we constructed ${\overline{A}}_{i-1}$ with the properties (iii) and (iv). If the center of $E$ on $A_{i-1}$ is a closed point, then we can blow up $A_{i-1}$ at the point to gain $A_{i}$. Then, we may assume that the center of $E$ is not a closed point in $A_{i-1}$. As ${\overline{A}}_{i-1}\to{\overline{A}}$ is projective and $\overline{A}$} is projective, we can write ${\overline{A}}_{i-1}={\mathrm{Proj\ }}{\overline{R}}_{i-1}$ and
${\overline{R}}_{i-1}=\oplus_{n=0}^{\infty}J_{n},$
where $J_{n}$ is the degree $n$ part of the graded ring ${\overline{R}}_{i-1}$ and $J_{0}=k$. By replacing ${\overline{R}}_{i-1}$ by a suitable Veronese subring, we may assume that ${\overline{R}}_{i-1}$ is generated by $J_{1}$. Let ${\mathfrak{p}}_{i}\subset{\overline{R}}_{i-1}$ be the homogeneous prime ideal defining the center of $E$ in ${\overline{A}}_{i-1}$.
Claim. We can take a subset $\{f_{1},\ldots,f_{t}\}\subset{\mathfrak{p}}_{i}$ such that
  1. (1)
    the equality of subschemes $V(f_{1},\ldots,f_{t})=V({\mathfrak{p}}_{i})$ holds and each $f_{j}$ has the common degree $d>0$ in the graded ring ${\overline{R}}_{i-1}$,
  2. (2)
    the linear system $\Lambda$ defined by the $k$-vector subspace
    $\langle f_{1},\ldots,f_{t}\rangle\subset\Gamma({\overline{A}}_{i-1},{\mathcal{O}}_{{\overline{A}}_{i-1}}(d))$
    is very ample on ${\overline{A}}_{i-1}\setminus V({\mathfrak{p}}_{i})$. In prticular, $\Lambda$ is very ample on ${\overline{A}}_{i-1}\setminus\phi_{i-1}^{-1}(0)$, where $\phi_{i-1}:{\overline{A}}_{i-1}\to{\overline{A}}$ is the composite $\varphi_{1}\circ\cdots\circ\varphi_{i-1}$ and $0\in{\overline{A}}$ is the origin of $A$. Here, note that $V({\mathfrak{p}}_{i})\subset\phi_{i-1}^{-1}(0)$.
Let $\{h_{1},\ldots,h_{s}\}$ be a generator system of $J_{1}$ over $k$, then the linear system over $k$ defined by $\{h_{1},\ldots,h_{s}\}$ is very ample on ${\overline{A}}_{i-1}$. To prove the Claim above, we start by taking any homogeneous generator system $\{f^{\prime}_{1},\ldots,f^{\prime}_{r}\}$ of ${\mathfrak{p}}_{i}\subset{\overline{R}}_{i-1}$ with the degree $d_{1}\leq\cdots\leq d_{r}$, respectively. We will construct an appropriate system of elements with the properties claimed in (1) and (2). Here, $V({\mathfrak{p}}_{i})\subset{\overline{A}}_{i-1}$ is the center of $E$ on ${\overline{A}}_{i-1}$.
Let $d>d_{r}$. Define
$f_{i\gamma}:=f^{\prime}_{i}\cdot X_{\gamma}^{(d-d_{i})},$
where
$\left\{X_{\gamma}^{(d-d_{i})}\middle|\gamma=1,\ldots,\left(\begin{array}[]{c}ms+d-d_{i}-1\\ d-d_{i}\\ \end{array}\right)\right\}$
is the set of all monomials in variables $h_{j}$’s of degree $d-d_{i}$. Then, the set of elements
$\left\{f_{i\gamma}\middle|1\leq i\leq r,1\leq\gamma\leq\left(\begin{array}[]{c}ms+d-d_{i}-1\\ d-d_{i}\\ \end{array}\right)\right\}$
gives the closed subset $V(f_{i\gamma})=V({\mathfrak{p}}_{i})$. Note that all $f_{i\gamma}$ are of degree $d$. Thus (1) in the Claim holds. Let $\Lambda$ be the linear system defined by the $k$-vector subspace $\langle f_{i\gamma}\rangle\subset\Gamma({\overline{A}}_{i-1},{\mathcal{O}}_{{\overline{A}}_{i-1}}(d))$. Then, $\Lambda$ has the base locus $V({\mathfrak{p}}_{i})$ and is very ample outside of $V({\mathfrak{p}}_{i})$. As $V({\mathfrak{p}}_{i})\subset\phi_{i-1}^{-1}(0)$, the linear system $\Lambda$ is very ample outside of $\phi_{i-1}^{-1}(0)$. Now, denote the $\{f_{i\gamma}\}$ just by $\{f_{i}\mid i=1,\ldots,t\}$ for simplicity of notation. Then, the system $\{f_{i}\mid i=1,\ldots,t\}$ satisfies (2) required in the Claim.
Now, let $c=0pt{\mathfrak{p}}_{i}$ and take $c$ general linear combinations of $f_{i}$’s over $k$:
$\sum_{i=1}^{t}a_{ji}f_{i}\ (a_{ji}\in k\ \mbox{ for}\ j=1,\ldots,c)$
and rename them as $f_{1},\ldots,f_{c}$. Then, $\{f_{1},\ldots,f_{c}\}$ becomes a regular sequence in $R_{i-1}$. Let $I_{i}\subset{\overline{R}}_{i-1}$ be the ideal generated by $\{f_{1},\ldots,f_{c}\}$ and $\mathcal{I}_{i-1}$ the ideal sheaf of ${\mathcal{O}}_{{\overline{A}}_{i-1}}$ defined by $I_{i-1}$. Similarly, let $\mathcal{P}_{i}$ be the ideal sheaf of ${\mathcal{O}}_{{\overline{A}}_{i-1}}$ defined by ${\mathfrak{p}}_{i}$. By the choice of $f_{1},\ldots,f_{c}$, these give a regular system of parameters of ${\mathcal{O}}_{{\overline{A}}_{i-1},P_{i}}$, where $P_{i}\in{\overline{A}}_{i-1}$ is the generic point of $V({\mathfrak{p}}_{i})$. Therefore,
$\mathcal{I}_{i}{\mathcal{O}}_{{\overline{A}}_{i-1},P_{i}}=\mathcal{P}_{i}{\mathcal{O}}_{{\overline{A}}_{i-1},P_{i}}.$
Therefore, one of the irreducible components of $Z_{i}:=V(I_{i})\subset{\overline{A}}_{i-1}$ is the center of $E$ on ${\overline{A}}_{i-1}$.
As $Z_{i}\subset{\overline{A}}_{i-1}$ is obtained by successive cutting ${\overline{A}}_{i-1}$ with $c$ general hypersurfaces
$H_{i}:=V(f_{i})\ \ (i=1,\ldots,c)$
from the linear system $\Lambda$, we obtain the following:
  1. (a)
    $\dim({\mathrm{Sing}}{\overline{A}}_{i-1})\cap Z_{i}<N-c$ and
  2. (b)
    $\dim(\phi_{i-1}^{-1}(0)\setminus V({\mathfrak{p}}_{i}))\cap Z_{i}<N-c$.
Indeed, first we have
$\dim({\mathrm{Sing}}{\overline{A}}_{i-1}\setminus V({\mathfrak{p}}_{i}))\cap Z_{i}<N-c.$
Because $\Lambda$ is very ample on ${\overline{A}}_{i-1}\setminus V({\mathfrak{p}}_{i})$, each cut by a general member of $\Lambda$ makes the dimension of the intersection with ${\mathrm{Sing}}{\overline{A}}_{i-1}\setminus V({\mathfrak{p}}_{i})$ exactly 1 less. Next we have
$\dim({\mathrm{Sing}}{\overline{A}}_{i-1}\cap V({\mathfrak{p}}_{i}))\cap Z_{i}<N-c,$
because $\dim V({\mathfrak{p}}_{i})=N-c$ and the generic point of $V({\mathfrak{p}}_{i})$ is smooth in ${\overline{A}}_{i-1}$. Thus (a) follows. The inequality (b) also follows in a similar way.
By (a), any irreducible component of $Z_{i}$ is not contained in ${\mathrm{Sing}}{\overline{A}}_{i-1}$, since the dimension of every irreducible component of $Z_{i}$ is NcN-c. By (b), any irreducible component of $Z_{i}$, except for $V({\mathfrak{p}}_{i})$, is not contained in the fiber $\phi_{i-1}^{-1}(0)$. Then, $Z_{i}$ satisfies the condition (iii) of the proposition.
Moreover, we obtain that
$({\overline{A}}_{i-1}\setminus\phi_{i-1}^{-1}(0))\cap Z_{i}$
is a smooth scheme by Bertini’s theorem “for very ample linear system” $\Lambda$ on a smooth quasi-projective scheme ${\overline{A}}_{i-1}\setminus\phi_{i-1}^{-1}(0)$. Note that the “very ample version” of Bertini’s theorem holds for any characteristic base field (see the proof of [9, II, Theorem 8.18]). A fortiori, every irreducible component of $Z_{i}$ is not contained in ${\mathrm{Sing}}Z_{i}$ nor in ${\mathrm{Sing}}{\overline{A}}_{i-1}$. Now the condition (11) in Lemma 4.3 is satisfied. Therefore, if $\varphi_{i}:{\overline{A}}_{i}\to{\overline{A}}_{i-1}$ is the blow-up with the center $Z_{i}\subset{\overline{A}}_{i-1}$, then, ${\overline{A}}_{i}$ is also normal and Gorenstein by the Lemma 4.3. As the blow-up $\varphi_{i}$ is the same as the blow-up at the center $V({\mathfrak{p}}_{i})\subset{\overline{A}}_{i-1}$ of $E$ around the generic point of $V({\mathfrak{p}}_{i})$, the generic point of the center of $E$ on $A_{i}$ is also smooth (see the comment below Proposition-Definition 4.1).
As ${\overline{A}}_{i-1}\setminus\phi_{i-1}^{-1}(0)$ and (${\overline{A}}_{i-1}\setminus\phi_{i-1}^{-1}(0))\cap Z_{i}$ are smooth, we also obtain the blown-up space
${\overline{A}}_{i}\setminus\phi_{i}^{-1}(0)$
of ${\overline{A}}_{i-1}\setminus\phi_{i-1}^{-1}(0)$ by the center $({\overline{A}}_{i-1}\setminus\phi_{i-1}^{-1}(0))\cap Z_{i}$ is also smooth, which completes the proof of (iv).
As the sequence of the blow-ups is the same as the Zariski sequence for $E$ around the centers of $E$ on any ${\overline{A}}_{i}$ by (iii), the prime divisor $E$ appears on ${\overline{A}}_{n}$ for a certain $n\in{\mathbb{Z}}_{\geq 0}$. Hence, (i) follows. ∎
Remark 4.7.
Under the situation in of Proposition 4.6, if Nc>0N-c>0 then, Bertini’s theorem for very ample linear system gives that the $Z_{i}$ has at most two irreducible components, $V({\mathfrak{p}}_{i})$ and $Z_{i}^{\prime}$, where $Z^{\prime}_{i}$ is not contained in $\phi_{i}^{-1}(0)$. On the other hand, if $N=c$, then $Z^{\prime}_{i}$ may consist of some closed points.
Definition 4.8.
For a prime divisor $E$ over $A$ with the center at a closed point $0\in A$, the sequence satisfying the conditions (i)–(iv) in Proposition 4.6 is called a normal Gorenstein Zariski sequence for $E$. This name makes sense, as the sequence is the same as the Zariski sequence for $E$ around the generic point of the center of $E$ on each $A_{i}$ by the properties (i) and (iii) in Proposition 4.6. Note that a normal Gorenstein Zariski sequence are not unique for a prime divisor $E$ over $(A,0)$.
Proposition 4.9.
We use the same notation as in Proposition 4.6. Let $k$ be an algebraically closed field of characteristic $p>0$, and set $A={\mathbb{A}}_{k}^{N}$. Let $E$ be a prime divisor over $A$ defined over $k$. Then, the following hold.
  1. (1)
    Assume that $E$ is obtained by a Zariski sequence
    $A_{n}\xrightarrow{\varphi_{n}}A_{n-1}\xrightarrow{\varphi_{n-1}}\cdots\xrightarrow{\varphi_{2}}A_{1}\xrightarrow{\varphi_{1}}A_{0}:=A,$
    where, for each $1\leq i\leq n$, the morphism $\varphi_{i}\colon A_{i}\to A_{i-1}$ is the blow-up of a closed point $P_{i}\in A_{i-1}$. Then, there exist a subfield $L\subset{\mathbb{C}}$ and a sequence
    ${\widetilde{A}}_{n}\xrightarrow{\widetilde{\varphi}_{n}}{\widetilde{A}}_{n-1}\xrightarrow{\widetilde{\varphi}_{n-1}}\cdots\xrightarrow{\widetilde{\varphi}_{2}}{\widetilde{A}}_{1}\xrightarrow{\widetilde{\varphi}_{1}}{\widetilde{A}}_{0}:={\widetilde{A}}={\mathbb{A}}_{L}^{N}$
    such that:
    1. (i)
      ${\widetilde{A}}_{i}(\bmod p)=A_{i}$ for all $0\leq i\leq n$;
    2. (ii)
      for each $1\leq i\leq n$, the morphism $\widetilde{\varphi}_{i}\colon{\widetilde{A}}_{i}\to{\widetilde{A}}_{i-1}$ is the blow-up of a closed point ${\widetilde{P}}_{i}\in{\widetilde{A}}_{i-1}$;
    3. (iii)
      let $E_{i}\subset A_{i}$ be the exceptional divisor of $\varphi_{i}$ and let ${\widetilde{E}}_{i}\subset{\widetilde{A}}_{i}$ be the exceptional divisor of $\widetilde{\varphi}_{i}$. Then, for every $1\leq i\leq n$,
      $k_{{\widetilde{E}}_{i}}=k_{E_{i}}.$
  2. (2)
    Assume that $N\leq 3$. Then, $E$ admits a normal Gorenstein Zariski sequence
    $A_{n}\xrightarrow{\varphi_{n}}A_{n-1}\xrightarrow{\varphi_{n-1}}\cdots\xrightarrow{\varphi_{2}}A_{1}\xrightarrow{\varphi_{1}}A_{0}:=A$
    with the center $Z_{i}\subset A_{i-1}$ for the blow-up $\varphi_{i}$. Let $P_{i}\in A_{i-1}$ be the generic point of the center of $E$ on $A_{i-1}$, and let $E_{i}\subset A_{i}$ be the exceptional divisor dominating $P_{i}$ by $\varphi_{i}$. Let $E_{i}^{\prime}\subset A_{i}$ be the other exceptional divisor of $\varphi_{i}$, if it exists. Then, there exist a subfield $L\subset{\mathbb{C}}$ and a sequence
    ${\widetilde{A}}_{n}\xrightarrow{\widetilde{\varphi}_{n}}{\widetilde{A}}_{n-1}\xrightarrow{\widetilde{\varphi}_{n-1}}\cdots\xrightarrow{\widetilde{\varphi}_{2}}{\widetilde{A}}_{1}\xrightarrow{\widetilde{\varphi}_{1}}{\widetilde{A}}_{0}:={\widetilde{A}}={\mathbb{A}}_{L}^{N}$
    satisfying the following:
    1. (i)
      ${\widetilde{A}}_{i}(\bmod p)=A_{i}$ for all $0\leq i\leq n$;
    2. (ii)
      for each $1\leq i\leq n$, the morphism $\widetilde{\varphi}_{i}\colon{\widetilde{A}}_{i}\to{\widetilde{A}}_{i-1}$ is the blow-up by locally a complete intersection center ${\widetilde{Z}}_{i}\subset{\widetilde{A}}_{i-1}$. One of the irreducible components of ${\widetilde{Z}}_{i}$ has the generic point ${\widetilde{P}}_{i}\in{\widetilde{A}}_{i-1}$ contained in the fiber of $0\in{\widetilde{A}}$;
      corresponding to a weak prime lifting of $P_{i}\in A_{i-1}$;
    3. (iii)
      let ${\widetilde{E}}_{i}\subset{\widetilde{A}}_{i}$ be the exceptional divisor mapped to ${\widetilde{P}}_{i}$ by $\widetilde{\varphi}_{i}$. Then, for every $1\leq i\leq n$,
      $k_{{\widetilde{E}}_{i}}=k_{E_{i}}.$
Proof.
We regard $A={\mathbb{A}}_{k}^{N}$ as an affine chart of ${\mathbb{P}}_{k}^{N}$ and let $0\in{\mathbb{A}}_{k}^{N}$ correspond to $(1:0:\cdots:0)\in{\mathbb{P}}_{k}^{N}.$ As in the proof of Proposition 4.6, we assume that $A={\mathbb{P}}_{k}^{N}$ and construct the sequences on ${\mathbb{P}}^{N}$, and then restrict them on ${\mathbb{A}}^{N}$. Now, denote $A={\mathrm{Proj\ }}R_{0}$, $A_{i}={\mathrm{Proj\ }}R_{i}$ $(i=1,\ldots,n)$, where $R_{0}=k[x_{0},\ldots,x_{N}]$, and
$R_{i}=\oplus_{\ell\geq 0}R_{0}[I_{1}T_{1},\ldots,I_{i}T_{i}]_{(\ell g_{0},\ldots,\ell g_{i-1},\ell g_{i})},\ \ \ \mbox{for}\ \ g_{0}\gg g_{1}\gg\cdots\gg g_{i}\gg..\gg g_{n},$
as in Remark 4.2. Here, $I_{i}\subset R_{i-1}$ is the defining ideal of $Z_{i}$, the center of the blow-up $\varphi_{i}:A_{i}\to A_{i-1}$.
Case (1). Assume $N\geq 2$ and $\varphi_{i}$’s are all blow-ups by closed points $P_{i}\in A_{i-1}$.
Step 0. For an arbitrary chosen skeleton $\Sigma$ of $k$, we can define $S_{0}=\Sigma[x_{0},x_{1},\ldots,x_{N}]$. Then, $S_{0}\subset R_{0}$ is a skeleton of $R_{0}$. (Later, we will restrict the choice of $\Sigma$ such that it contains certain necessary elements in the following steps.) By choosing a compatible skeleton $\widetilde{\Sigma}$} of $\Sigma$, we have a compatible skeleton ${\widetilde{S}}_{0}={\widetilde{\Sigma}}[x_{0},x_{1},\ldots,x_{N}]$ of $S_{0}$. Denote
[…]
then,
${\widetilde{B}}_{0}({\operatorname{mod}\ }p)=B_{0}.$
Step 1. As $I_{1}={\mathfrak{m}}=(x_{1},\ldots,x_{N})\subset k[x_{0},x_{1},\ldots,x_{N}]$, we can take
$\underline{I}_{1}=(x_{1},\ldots,x_{N})\subset S_{0}=\Sigma[x_{0},x_{1},\ldots,x_{N}]$
for arbitrary chosen skeleton $\Sigma$ of $k$. We denote the point in $B_{0}$ defined by the ideal $\underline{I}_{1}$ by $P_{1}\in B_{0}$, the same symbol of $P_{1}\in A_{0}$. If we denote
$S_{1}:=\oplus_{\ell\geq 0}S_{0}[\underline{I}_{1}T_{1}]_{(\ell g_{0},\ell g_{1})},$
then $S_{1}$ is a skeleton of $R_{1}$. Take
$\underline{{\widetilde{I}}_{1}}=(x_{1},\ldots,x_{N})\subset{\widetilde{S}}_{0}={\widetilde{\Sigma}}[x_{0},x_{1},\ldots,x_{N}]$
for a compatible skeleton $\widetilde{\Sigma}$} of $\Sigma$ and let the ideal define ${\widetilde{P}}_{1}\in{\widetilde{B}}_{0}$. Then, ${\widetilde{P}}_{1}$ is a prime lifting of $P_{1}$ and
${\widetilde{S}}_{1}:=\oplus_{\ell\geq 0}{\widetilde{S}}_{0}[\underline{{\widetilde{I}}}_{1}T_{1}]_{(\ell g_{0},\ell g_{1})}$
is a compatible skeleton of $S_{1}$. Denote
$B_{1}:={\mathrm{Proj\ }}S_{1}\ \ \mbox{and}\ \ {\widetilde{B}}_{1}:={\mathrm{Proj\ }}{\widetilde{S}}_{1},$
and denote the exceptional divisor over $P_{1}\in B_{0}$ and ${\widetilde{P}}_{1}\in{\widetilde{B}}_{0}$ by $E_{1}$ and ${\widetilde{E}}_{1}$, respectively. Then, by the definition, ${\widetilde{E}}_{1}$ is a prime lifting of $E_{1}$.
Step 2. The next center $P_{2}$ of the blow up $\varphi_{2}$ is on $E_{1}$, by the definition of Zariski sequence. We can take the lifting ${\widetilde{P}}_{2}\in{\widetilde{E}}_{1}$. Indeed, if $P_{2}\in A_{1}$ is in an affine open subset ${\mathrm{Spec\ }}(R_{1})_{(\ell_{1})}$ for a homogeneous element $\ell_{1}\in R_{1}$ such that $P_{2}$ is defined by $N$ elements of $(R_{1})_{(\ell_{1})}$, we enlarge $\Sigma$ such that $\ell_{1}\in S_{1}$, and $P_{2}\in{\mathrm{Spec\ }}(S_{1})_{(\ell_{1})}$ is defined by $N$ elements of $(S_{1})_{(\ell_{1})}$. We enlarge $\Sigma$ again if necessary so that $I_{2}\subset R_{1}$ is generated by elements of $S_{1}$ and write $\underline{I}_{2}:=I_{2}\cap S_{1}$.
Then, take a homogeneous lifting $\widetilde{\ell}_{1}$ of $\ell_{1}$ with the same degree. As the ideal of ${\widetilde{E}}_{1}$ in $({\widetilde{S}}_{1})_{(\widetilde{\ell}_{1})}$ is a prime lifting of the ideal of $E_{1}$ in $(S_{1})_{(\ell_{1})}$, we have a prime lifting ${\widetilde{P}}_{2}\in{\mathrm{Spec\ }}({\widetilde{S}}_{1})_{(\widetilde{\ell}_{1})}$ of $P_{2}$ such that ${\widetilde{P}}_{2}\in{\widetilde{E}}_{1}$ by Corollary 2.9. Hence the configurations of $\{E_{1},P_{2}\}$ on $B_{1}$ and of $\{{\widetilde{E}}_{1},{\widetilde{P}}_{2}\}$ on ${\widetilde{B}}_{1}$ are the same.
Define
$S_{2}=\oplus_{\ell\geq 0}S_{0}[\underline{I}_{1}T_{1},\underline{I}_{2}T_{2}]_{(\ell g_{0},\ell g_{1},\ell g_{2})},\ \ \ \mbox{for}\ \ g_{0}\gg g_{1}\gg g_{2},$
and $B_{2}:={\mathrm{Proj\ }}S_{2}$, then, it is a skeleton of $R_{2}$. Let $\underline{{\widetilde{I}}}_{2}\subset{\widetilde{S}}_{1}$ be the homogeneous ideal defining ${\widetilde{P}}_{2}$ and define
${\widetilde{S}}_{2}=\oplus_{\ell\geq 0}{\widetilde{S}}_{0}[\underline{{\widetilde{I}}}_{1}T_{1},\underline{{\widetilde{I}}}_{2}T_{2}]_{(\ell g_{0},\ell g_{1},\ell g_{2})},\ \ \ \mbox{for}\ \ g_{0}\gg g_{1}\gg g_{2},$
and ${\widetilde{B}}_{2}:={\mathrm{Proj\ }}{\widetilde{S}}_{2}$. Here, we note that $\underline{{\widetilde{I}}}_{2}$ does not contain nonzero elements of $\widetilde{\Sigma}$}, because $\underline{{\widetilde{I}}}$ is the homogenization of the ideal defining the point ${\widetilde{P}}_{2}$ in the affine open subset.
Step i. Assume we have a skeleton $S_{i-1}$ of $R_{i-1}$ and its compatible skeletons ${\widetilde{S}}_{i-1}$ with the same configurations $\{E_{1},\ldots,E_{i-1}\}$ and $\{{\widetilde{E}}_{1},\ldots,{\widetilde{E}}_{i-1}\}$ on $B_{i-1}$ and on ${\widetilde{B}}_{i-1}$, respectively. Let
$J:=I\left(\bigcap_{P_{i}\in E_{j}}E_{j}\right)\subset R_{i-1}$
be the homogeneous ideal defining $\bigcap_{P_{i}\in E_{j}}E_{j}$. Note that $J$ is generated by $\underline{J}:=J\cap S_{i-1}$ which has a prime lifting
$\underline{\widetilde{J}}=I\left(\bigcap_{{\widetilde{P}}_{i}\in{\widetilde{E}}_{j}}{\widetilde{E}}_{j}\right).$
Then, restricting on an affine neighborhood of $P_{i}$ and enlarging $\Sigma$ appropriately as in Step 2, we apply Corollary 2.9. Then, we obtain a prime lifting ${\widetilde{P}}_{i}\in{\widetilde{B}}_{i-1}$ of $P_{i}$ whose defining ideal contains $\underline{\widetilde{J}}$. This yields
[…]
Consequently, the configurations of $\{E_{1},\ldots,E_{i-1},P_{i}\}$ and of $\{{\widetilde{E}}_{1},\ldots,{\widetilde{E}}_{i-1},{\widetilde{P}}_{i}\}$ are the same. Define
$S_{i}=\oplus_{\ell\geq 0}S_{0}[\underline{I}_{1}T_{1},\ldots,\underline{I}_{i}T_{i}]_{(\ell g_{0},\ell g_{1},\ldots,\ell g_{i})},\ \ \ \mbox{for}\ \ g_{0}\gg g_{1}\ldots\gg g_{i},$
then, it is a skeleton of $R_{i}$. Let $\underline{\widetilde{I}}_{i}\subset{\widetilde{S}}_{i-1}$ be the homogeneous prime ideal defining ${\widetilde{P}}_{i}\in{\widetilde{S}}_{i-1}$, then, it is a prime lifting of $\underline{I}_{i}$ and
${\widetilde{S}}_{i}=\oplus_{\ell\geq 0}{\widetilde{S}}_{0}[\underline{{\widetilde{I}}}_{1}T_{1},\ldots,\underline{{\widetilde{I}}}_{i}T_{i}]_{(\ell g_{0},\ell g_{1},\ldots,\ell g_{i})},\ \ \ \mbox{for}\ \ g_{0}\gg g_{1}\ldots\gg g_{i}.$
Define $B_{i}:={\mathrm{Proj\ }}S_{i}$ and ${\widetilde{B}}_{i}:={\mathrm{Proj\ }}{\widetilde{S}}_{i}$ and denote the exceptional divisors over $P_{i}$ and ${\widetilde{P}}_{i}$ by $E_{i}\subset B_{i}$ and ${\widetilde{E}}_{i}\subset{\widetilde{B}}_{i}$, respectively. Then, the configuration of $\{E_{1},\ldots,E_{i}\}$ on $B_{i}$ and that of $\{{\widetilde{E}}_{1},\ldots,{\widetilde{E}}_{i}\}$ are the same.
By carrying out each step up to $i=n$, we obtain the sequence:
(12) ${\widetilde{B}}_{n}\to\ldots\to{\widetilde{B}}_{0}$
of blow-ups by the ideals $\underline{\widetilde{I}}_{i}\subset{\widetilde{B}}_{i-1}$ such that
(13) ${\widetilde{P}}_{i}\in{\widetilde{E}}_{j}\ \ \mbox{iff}\ \ P_{i}\in E_{j}\ \ \mbox{ for}\ \ j\leq i-1.$
On the other hand, by (ii) of Lemma 2.11, we have
(14) $N=0ptI_{i}=0pt\underline{I}_{i},\ \mbox{and}\ \ 0pt\underline{\tilde{I}}_{i}=0pt(\underline{\tilde{I}}_{i}\otimes_{\widetilde{\Sigma}}L),$
[…]
As $\underline{\tilde{I}}_{i}$ is a prime lifting of $\underline{I}_{i}$, we obtain:
(15) $0pt\underline{I}_{i}=0pt\underline{\tilde{I}}_{i}.$
By (14) and (15), it follows:
(16) $0ptI_{i}=0pt(\underline{\tilde{I}}_{i}\otimes_{\widetilde{\Sigma}}L),$
Let ${\widetilde{A}}_{i}:={\widetilde{B}}_{i}\otimes_{\widetilde{\Sigma}}L$, then ${\widetilde{A}}_{i}({\operatorname{mod}\ }p)=A_{i}$ with the compatible skeletons ${\widetilde{S}}_{i}$ and $S_{i}$, and it follows the sequence of blow-ups:
${\widetilde{A}}_{n}\stackrel{{\scriptstyle\widetilde{\varphi}_{n}}}{{\longrightarrow}}{\widetilde{A}}_{n-1}\to\cdots{\widetilde{A}}_{1}\stackrel{{\scriptstyle\widetilde{\varphi}_{1}}}{{\longrightarrow}}{\widetilde{A}}={\mathbb{A}}_{L}^{N}.$
We denote the exceptional divisor on ${\widetilde{A}}_{i}$ corresponding to ${\widetilde{E}}_{i}\subset{\widetilde{B}}_{i}$ by the same symbol ${\widetilde{E}}_{i}$. Then, (13) and (16) yields $k_{{\widetilde{E}}_{i}}=k_{E_{i}}$, because
$k_{E_{i}}=0ptI_{i}-1+\sum_{P_{i}\in E_{j}}k_{E_{j}}\ \ \mbox{and}$
$k_{{\widetilde{E}}_{i}}=0pt(\underline{{\widetilde{I}}}_{i}\otimes_{\widetilde{\Sigma}}L)-1+\sum_{{\widetilde{P}}_{i}\in{\widetilde{E}}_{j}}k_{{\widetilde{E}}_{j}}.$
Here, we note that the point ${\widetilde{P}}_{i}\in{\widetilde{A}}_{i-1}$ is smooth, as the local ring ${\mathcal{O}}_{{\widetilde{A}}_{i-1},{\widetilde{P}}_{i}}$ has the maximal ideal generated by $N$ elements, where $N$ is the height of the defining ideal of $P_{i}$ which is equal to that of ${\widetilde{P}}_{i}$.
Case (2) Assume $N\leq 3$.
If $N=2$, the Zariski sequence is the composite of blow-ups by closed points, which is included in Case 1. Therefore, we may assume that $N=3$. As the first blow-up $\varphi_{1}:A_{1}\to A$ is at the closed point $P_{1}=(1:0:0:0)\in A={\mathbb{P}}_{k}^{3}$, there is $s\geq 2$ such that $\varphi_{s}:A_{s}\to A_{s-1}$ is the first blow-up by a curve. Let $C_{s}\subset E_{s-1}\subset A_{s-1}$ be the center. As $E_{s-1}\simeq{\mathbb{P}}_{k}^{2}$, there is the following isomorphism up to the Veronese transformation.
(17) $k[z_{0},z_{1},z_{2}]\simeq R_{s-1}/I_{E_{s-1}}$
Step s1s-1. By the arguments in Case 1, we have constructed the sequences of skeletons:
$B_{s-1}\stackrel{{\scriptstyle\varphi_{s-1}}}{{\longrightarrow}}B_{s-2}\to\cdots B_{0}={\mathbb{P}}_{\Sigma}^{3}$
and their liftings
${\widetilde{B}}_{s-1}\stackrel{{\scriptstyle\widetilde{\varphi}_{s-1}}}{{\longrightarrow}}{\widetilde{B}}_{s-2}\to\cdots{\widetilde{B}}_{0}={\mathbb{P}}_{{\widetilde{\Sigma}}}^{3}.$
Also until i=s1i=s-1 the statements (i),(ii) and (iii) of the proposition hold by the arguments in Case (1). Let
[…]
be the exceptional divisors for the latest blow-ups $\varphi_{s-1}$ and $\widetilde{\varphi}_{s-1}$, then, these exceptional divisors $E_{s-1}$ and ${\widetilde{E}}_{s-1}$ are the projective planes
[…]
respectively. Note that the latter is a prime lifting of the former in $B_{s-1}$ and ${\widetilde{B}}_{s-1}$. Then, $S_{s-1}/I_{E_{s-1}}$ and ${\widetilde{S}}_{s-1}/{\widetilde{I}}_{{\widetilde{E}}_{s-1}}$ are compatible skeletons. Then, by (17),
$\Sigma[z_{0},z_{1},z_{2}]\simeq S_{s-1}/I_{E_{s-1}},\ \ \mbox{and}\ \ {\widetilde{\Sigma}}[z_{0},z_{1},z_{2}]\simeq{\widetilde{S}}_{s-1}/I_{{\widetilde{E}}_{s-1}},$
up to the Veronese transformation of the common degree. Hence, we may assume that both are real isomorphisms for the simplicity of notation. The center $C_{s}\subset A_{s-1}$ of $E$ is an irreducible curve of degree $d$ on ${\mathbb{P}}_{k}^{2}=E_{s-1}$ for some $d>0$. As $C_{s}\subset A_{s-1}$ is locally a complete intersection, we can take $\varphi_{s}:A_{s}\to A_{s-1}$ the blow-up by the defining ideal $I_{C_{s}}\subset R_{s-1}$. Enlarging $\Sigma$, we may assume that $I_{C_{s}}$ is generated by elements of $S_{s-1}$. By abuse of notation, we use the same symbols $E_{s-1}$ and $C_{s}$ for subschemes of $A_{s-1}$ and for the corresponding subschemes of its skeleton $B_{s-1}$.
Step $s$. As $C_{s}\subset A_{s-1}$ is of codimension two and contained in $E_{s-1}$, there are two possibilities on the position of $C_{s}$:
(Is) $C_{s}=E_{s-1}\cap E_{j}$ for some $j\leq s-2$ or (IIs) $C_{s}\subset E_{s-1}\setminus E_{j}$ for any $j\leq s-2$.
In case (Is), as $E_{s-1}\cap E_{j}$ has the lifting ${\widetilde{E}}_{s-1}\cap{\widetilde{E}}_{j}$, we have the canonical lifting ${\widetilde{C}}_{s}:={\widetilde{E}}_{s-1}\cap{\widetilde{E}}_{j}\subset{\widetilde{B}}_{s-1}$ of $C_{s}\subset B_{s-1}$. Then, by compatibility of the blow-ups $B_{s}\to B_{s-1}$ and ${\widetilde{B}}_{s}\to{\widetilde{B}}_{s-1}$ by $C_{s}$ and ${\widetilde{C}}_{s}$, we obtain for $L=Q(\Sigma)$ (i) ${\widetilde{A}}_{s}({\operatorname{mod}\ }p)=A_{s}$. For (ii), remark that ${\widetilde{C}}_{s}\subset{\widetilde{E}}_{s-1}\subset{\widetilde{B}}_{s-1}$ which implies that ${\widetilde{C}}_{s}\subset\widetilde{\varphi}_{s-1}^{-1}(P_{1})$, where $P_{1}=(1:0:0:0)\in{\mathrm{Proj\ }}{\widetilde{\Sigma}}[x_{0},x_{1},x_{2},x_{3}]={\widetilde{B}}_{0}$. This yields the corresponding subscheme ${\widetilde{C}}_{s}\subset{\widetilde{A}}_{s-1}$ is contained in $\phi_{s-1}^{-1}(0)$ for $0=(1:0:0:0)\in{\mathbb{P}}_{L}^{N}$. Let $P_{s}$ and ${\widetilde{P}}_{s}$ be the generic points of $C_{s}$ and ${\widetilde{C}}_{s}$. For (iii), note that the configuration of $\{E_{1},\cdots,E_{s-1},P_{s}\}$ and that of $\{{\widetilde{E}}_{1},\cdots,{\widetilde{E}}_{s-1},{\widetilde{P}}_{s}\}$ are the same. Hence, we have:
$k_{E_{s}}=k_{E_{s-1}}+k_{E_{j}}+1=k_{{\widetilde{E}}_{s-1}}+k_{{\widetilde{E}}_{j}}+1=k_{{\widetilde{E}}_{s}},$
In case (IIs), by enlarging $\Sigma$, we may assume that the coefficients of the defining equation $f=0$ of $C_{s}\subset E_{s-1}={\mathbb{P}}_{k}^{2}$ are in $\Sigma$}. Then, we can take an equation $\tilde{f}$} of degree $d$ with coefficients in a compatible skeleton $\widetilde{\Sigma}$} of $\Sigma$, which becomes a lifting of $f$. Let $\widetilde{C}_{s}$ be the curve defined by the ideal (f~)($\tilde${f}}). Then, $\widetilde{C}_{s}\subset{\widetilde{E}}_{s-1}$ is a prime lifting of $C_{s}\subset E_{s-1}$. Therefore,
${\widetilde{C}}_{s}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}=C_{s}$
which implies that ${\widetilde{C}}_{s}\subset{\widetilde{B}}_{s-1}$ is a prime lifting of $C_{s}\subset B_{s-1}$ by Remark 2.7,(4).
As $C_{s}\subset E_{s-1}\subset A_{s-1}$ is locally a complete intersection and the generic point of $C_{s}$ is a regular point of $A_{s-1}$, it follows that the blow-up $\varphi_{s}:A_{s}\to A_{s-1}$ by $C_{s}$ gives a normal and Gorenstein variety $A_{s}$ by Lemma 4.3. Let ${\underline{I}}_{s}$ and ${\underline{{\widetilde{I}}}}_{s}$ be the defining ideal of $C_{s}$ and ${\widetilde{C}}_{s}$ in $S_{s-1}$ and ${\widetilde{S}}_{s-1}$, respectively. Let
$S_{s}:=\oplus_{\ell\geq 0}S_{s-1}[{\underline{I}}_{s}T_{s}]_{(\ell g_{0},\cdots,\ell g_{s})}\ \ \mbox{for}\ \ g_{0}\gg\cdots\gg g_{s},$
and $B_{s}:={\mathrm{Proj\ }}S_{s}$ then $S_{s}$ is a skeleton of $R_{s}$ and its lifting is
${\widetilde{S}}_{s}:=\oplus_{\ell\geq 0}{\widetilde{S}}_{s-1}[\underline{{\widetilde{I}}}_{s}T_{s}]_{(\ell g_{0},\cdots,\ell g_{s})}\ \ \mbox{for}\ \ g_{0}\gg\cdots\gg g_{s}.$
[…]
$k_{E_{s}}=k_{E_{s-1}}+1=k_{{\widetilde{E}}_{s-1}}+1=k_{{\widetilde{E}}_{s}},$
because the configuration of $\{E_{1},\cdots,E_{s-1},P_{s}\}$ and that of $\{{\widetilde{E}}_{1},\cdots,{\widetilde{E}}_{s-1},{\widetilde{P}}_{s}\}$ are the same, where $P_{s}$ and ${\widetilde{P}}_{s}$ are the generic points of $C_{s}$ and ${\widetilde{C}}_{s}$.
Step $s+1$. For the next step, let $C_{s+1}\subset A_{s}$ be the center of $E$ on $A_{s}$ and assume $C_{s+1}\neq E$. Then, the center $C_{s+1}\subset E_{s}\subset A_{s}$ is a curve dominating $C_{s}$. As $C_{s+1}\subset A_{s}$ has codimension two, there are two possibilities on the position of $C_{s+1}$:
(Is+1) $C_{s+1}=E_{s}\cap E_{s-1}$ or (IIs+1) $C_{s+1}\subset E_{s}$ and $\not\subset E_{s-1}$.
In case (Is+1), we have $C_{s+1}\simeq C_{s}$ and the blow-up of $A_{s}$ at $C_{s+1}$ gives a similar situation as in (Is). Then, by the same discussion, it follows (i),(ii) and (iii).
In case (IIs+1), note that $E_{s}$ has a ${\mathbb{P}}^{1}$-bundle structure on $C_{s}$. Remove the zero locus $(\xi=0)$ from $C_{s}$ (for some homogeneous element $\xi\in R_{s-1}$) to obtain an open affine subset $U\subset C_{s}$ such that the restriction
$\varphi_{s}|_{E_{s}|_{U}}\colon E_{s}|_{U}\to U$
is trivial:
$E_{s}|_{U}\simeq U\times{\mathbb{P}}_{k}^{1}.$
By enlarging $\Sigma$ and taking its compatible skeleton $\widetilde{\Sigma}$}, we may assume that $\xi\in S_{s-1}$ and the open subset of $C_{s}\subset B_{s}$ by removing the zero locus of $\xi\in S_{s-1}$ is denoted by the same symbol $U$ by abuse of notation. Then, we have:
$E_{s}|_{U}\simeq U\times{\mathbb{P}}_{\Sigma}^{1}.$
Removing the zero locus of an appropriate lifting $\widetilde{\xi}\in{\widetilde{S}}_{s-1}$ of $\xi$, we obtain an open affine subset $\widetilde{U}\subset{\widetilde{C}}_{s}$ such that
${\widetilde{E}}_{s}|_{\widetilde{U}}\simeq\widetilde{U}\times{\mathbb{P}}_{{\widetilde{\Sigma}}}^{1}.$
Let $g\in\Gamma(U,{\mathcal{O}}_{C_{s}})[y_{0},y_{1}]$ be a generator of the defining ideal of $C_{s+1}|_{U}$, where $y_{0},y_{1}$ are homogeneous coordinates of ${\mathbb{P}}^{1}$. Then, we obtain a lifting
${\tilde{g}}\in\Gamma({\widetilde{U}},{\mathcal{O}}_{{\widetilde{C}}_{s}})[y_{0},y_{1}]$
with the same degree as $g$ with respect to $y_{0},y_{1}$. Now, take the closure ${\widetilde{C}}_{s+1}$ of the zero locus $({\tilde{g}}=0)\subset{\widetilde{E}}_{s}|_{\widetilde{U}}$ in ${\widetilde{E}}_{s}$. Then,
${\widetilde{C}}_{s+1}\subset{\widetilde{E}}_{s}$
is a weak prime lifting of $C_{s+1}\subset E_{s}$ because the generator $g$ of the defining ideal of $C_{s+1}$ at its generic point comes from the generator $\tilde{g}$} of the defining ideal of ${\widetilde{C}}_{s+1}$ at its generic point (cf. Remark 2.7, (3)). Geometrically this shows that
${\widetilde{C}}_{s+1}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}=C_{s+1}\cup C^{\prime},$
where $C^{\prime}$ is the sum of other components whose supports do not contain $C_{s+1}$ (Remark 2.7, (4)). This yields that ${\widetilde{C}}_{s+1}\subset{\widetilde{B}}_{s}$ is a weak prime lifting of $C_{s+1}\subset B_{s}$ by Lemma 2.7,(4). Then, we obtain the algebraic description for the situation: Let ${\mathfrak{p}}_{s+1}\subset R_{s}$ be a homogeneous prime ideal defining $C_{s+1}\subset A_{s}$. Then, by enlarging $\Sigma$ if necessary, we may assume that ${\mathfrak{p}}_{s+1}$ is generated by elements of $S_{s}$ and denote $\underline{{\mathfrak{p}}}_{s+1}:={\mathfrak{p}}_{s+1}\cap S_{s}$, where $\underline{{\mathfrak{p}}}_{s+1}$ defines $C_{s+1}\subset B_{s}$. Denoting the homogeneous prime ideal defining ${\widetilde{C}}_{s+1}\subset{\widetilde{B}}_{s}$ by $\underline{{\widetilde{{\mathfrak{p}}}}}_{s+1}\subset{\widetilde{S}}_{s}$, we have $\underline{{\widetilde{{\mathfrak{p}}}}}_{s+1}$ is a weak prime lifting of $\underline{{\mathfrak{p}}}_{s+1}$ (Remark 2.7, (4)). Therefore, we have the surjective map:
(18) $\underline{{\widetilde{{\mathfrak{p}}}}}_{s+1}({\widetilde{S}}_{s})_{\Phi_{p}^{-1}(\underline{{\mathfrak{p}}}_{s+1})}\to\underline{{\mathfrak{p}}}_{s+1}(S_{s})_{\underline{{\mathfrak{p}}}_{s+1}}.$
By the definition and the construction of a normal Gorenstein Zariski sequence in Proposition 4.6, there is a complete intersection subscheme $Z_{s}\subset A_{s-1}$ containing $C_{s+1}$ as an irreducible component and plays the role of the center of blow-up $\varphi_{s+1}$. By the construction we may take the generator
$f_{1},f_{2}\in R_{s}$
of the defining ideal $I_{s+1}$ of $Z_{s+1}$ such that these form RSP of ${\mathfrak{p}}_{s+1}$ and the other irreducible component $Z^{\prime}_{s+1}$ of $Z_{s+1}$ is not contained in $\phi_{s}^{-1}(0)$, where $0=(1:0:0:0)\in A={\mathbb{P}}_{k}^{3}$ and $\phi_{s}:A_{s}\to A$ is the composite $\varphi_{1}\circ\cdots\circ\varphi_{s}$. By enlarging $\Sigma$ and take a compatible skeleton $\widetilde{\Sigma}$} of $\Sigma$ accordingly, we may assume that $f_{1},f_{2}\in S_{s}$ and write the ideal $\underline{I}_{s+1}:=I_{s+1}\cap S_{s}$. Define
(19) $S_{s+1}:=\oplus_{\ell\geq 0}S_{s}[\underline{I}_{s+1}T_{s+1}]_{(\ell g_{0},\ldots,\ell g_{s+1})}\ \mbox{for}\ \ g_{0}\gg\cdots\gg g_{s+1},$
which gives the blow-up $\varphi_{s+1}:B_{s+1}\to B_{s}$ by the ideal $(f_{1},f_{2})=\underline{I}_{s+1}\subset S_{s}$. As $\underline{{\mathfrak{p}}}_{s+1}(S_{s})_{\underline{{\mathfrak{p}}}_{s+1}}=(f_{1},f_{2})(S_{s})_{\underline{{\mathfrak{p}}}_{s+1}}$, by surjectivity (18), there are liftings
${\tilde{f}}_{1},{\tilde{f}}_{2}\in\underline{{\widetilde{{\mathfrak{p}}}}}_{s+1}({\widetilde{S}}_{s+1})_{\Phi_{p}^{-1}(\underline{{\mathfrak{p}}}_{s+1})}$
of $f_{1},f_{2}$. As $\{f_{1},f_{2}\}$ is a regular sequence, its lifting $\{{\tilde{f}}_{1},{\tilde{f}}_{2}\}$ is also a regular sequence in $\underline{{\widetilde{{\mathfrak{p}}}}}_{s+1}({\widetilde{S}}_{s})_{\Phi_{p}^{-1}(\underline{{\mathfrak{p}}}_{s+1})}$. Hence, $\{{\tilde{f}}_{1},{\tilde{f}}_{2}\}$ is an RSP of $\underline{{\widetilde{{\mathfrak{p}}}}}_{s+1}$, as the height of $\underline{{\widetilde{{\mathfrak{p}}}}}_{s+1}$ is two. This shows that ${\widetilde{{\mathfrak{p}}}}_{s+1}\subset{\widetilde{S}}_{s+1}$ is a regular prime ideal. For ${\tilde{f}}_{1},{\tilde{f}}_{2}\in({\widetilde{S}}_{s})_{\Phi_{p}^{-1}(\underline{{\mathfrak{p}}}_{s+1})}$, there exists an element ${\widetilde{a}}\in{\widetilde{S}}_{s}\setminus\Phi_{p}^{-1}(\underline{{\mathfrak{p}}}_{s+1})$ such that ${\tilde{h}}_{i}:={\widetilde{a}}{\tilde{f}}_{i}\in{\widetilde{S}}_{s}$ for $i=1,2$. Define $\underline{{\widetilde{I}}}_{s+1}:=({\tilde{h}}_{1},{\tilde{h}}_{2})\subset{\widetilde{S}}_{s}$ and
(20) ${\widetilde{S}}_{s+1}:=\oplus_{\ell\geq 0}{\widetilde{S}}_{s}[\underline{{\widetilde{I}}}_{s+1}T_{s+1}]_{(\ell g_{0},\ldots,\ell g_{s+1})}\ \mbox{for}\ \ g_{0}\gg\cdots\gg g_{s+1},$
which gives the blow-up $\widetilde{\varphi}_{s+1}:{\widetilde{B}}_{s+1}\to{\widetilde{B}}_{s}$ by the ideal $\underline{{\widetilde{I}}}=({\tilde{h}}_{1},{\tilde{h}}_{2})\subset S_{s}$.
[…]
$\Phi_{p}(\underline{{\widetilde{I}}}_{s+1})=a\underline{I}=(af_{1},af_{2})\subset S_{s}$
[…]
$k_{{\widetilde{E}}_{s+1}}=k_{{\widetilde{E}}_{s}}+1=k_{E_{s}}+1=k_{E_{s+1}}.$
Step $s+2$. For the next step, let $C_{s+2}\subset A_{s+1}$ be the center of $E$ on $A_{s+1}$ and assume $C_{s+2}\neq E$. Then, by a similar consideration as in Step $s+1$, there are three cases:
(Is+2) $C_{s+2}\subset E_{s+1}\cap E_{s-1}$ or (IIs+2) $C_{s+2}\subset E_{s+1}$ and $\not\subset E_{j}$ for any $j\leq s$ or (IIIs+2) $C_{s+2}\subset E_{s+1}\cap E_{s}.$
Here, (Is+2) happens only after (Is+1). In this case, in the same way as (Is+1), it follows (i), (ii) and (iii) in our proposition.
In case (IIs+2), we can construct the center $Z_{s+2}\subset A_{s+1}$ of the blow-up $\varphi_{s+2}:A_{s+2}\to A_{s+1}$ being locally a complete intersection and having an irreducible component $C_{s+2}$ and the other component not contained in $\phi_{s+1}^{-1}(0)\subset A_{s+1}$, in the same way as (IIs+1). Note that ${\widetilde{E}}_{s+1}$ is a weak prime lifting of $E_{s+1}$ as ${\widetilde{C}}_{s+1}$ is a weak prime lifting of $C_{s+1}$. Then, we can construct a weak prime lifting ${\widetilde{C}}_{s+2}\subset{\widetilde{E}}_{s+1}$ of $C_{s+2}$ in the same way as (IIs+1). Then, we construct a complete intersection subscheme $Z_{s+2}\subset A_{s+1}$ containing $C_{s+1}$ as an irreducible component with the other component not contained in $\phi_{s+1}^{-1}(0)\subset A_{s+1}$. Also we obtain a complete intersection subscheme ${\widetilde{Z}}_{s+2}\subset{\widetilde{B}}_{s+1}$ containing ${\widetilde{C}}_{s+2}$ as an irreducible component. Let
$\varphi_{s+2}:B_{s+2}\to B_{s+1}$
be the blow-up by $Z_{s+2}$ and
$\widetilde{\varphi}_{s+2}:{\widetilde{B}}_{s+2}\to{\widetilde{B}}_{s+1}$
the blow-up by ${\widetilde{Z}}_{s+2}$. Then, these are compatible as in the proof for (IIs+1). We see the configuration of $\{E_{1},\ldots,E_{s+1},P_{s+2}\}$ is the same as that of $\{{\widetilde{E}}_{1},\ldots,{\widetilde{E}}_{s+1},{\widetilde{P}}_{s+2}\}$, where $P_{s+2}$ and ${\widetilde{P}}_{s+2}$ be the generic points of $C_{s+2}$ and ${\widetilde{C}}_{s+2}$. Therefore, for $s+2$, we obtain (i),(ii), and (iii):
$k_{{\widetilde{E}}_{s+2}}=k_{{\widetilde{E}}_{s+1}}+1=k_{E_{s+1}}+1=k_{E_{s+2}}.$
In case (IIIs+2), we can construct the center $Z_{s+2}\subset A_{s+1}$ of the blow-up $\varphi_{s+2}:A_{s+2}\to A_{s+1}$ is locally a complete intersection and has an irreducible component $C_{s+2}$ and the other component not contained in
$\phi_{s+1}^{-1}(0)\subset A_{s+1},$
in the same way as (IIs+1). Up to here, the discussions are the same as (IIs+2) given just before. The different point from (IIs+2) is that in our case, the weak prime lifting ${\widetilde{E}}_{s+1}\cap{\widetilde{E}}_{s}$ of $C_{s+2}=E_{s+1}\cap E_{s}$ is already given. So, we define ${\widetilde{C}}_{s+2}:={\widetilde{E}}_{s+1}\cap{\widetilde{E}}_{s}$ and do the same discussions to obtain the results (i), (ii), (iii) in the proposition.
Continuing these process until $C_{n}$ becomes $E$, we obtain the statements in our proposition. Here, we note that the generic point ${\widetilde{P}}_{i}$ of ${\widetilde{C}}_{i}$ is regular, because its ideal is generated by two elements at the local ring of dimension two. ∎
Remark 4.10.
In the above proposition, we focus on the exceptional divisor $E_{i}$ mapped to $P_{i}$ by $\varphi_{i}$ and ${\widetilde{E}}_{i}$ mapped to ${\widetilde{P}}_{i}$ by ${\widetilde{\varphi}}_{i}$, respectively. However, we should not forget that there may be another irreducible component of $\varphi_{i}^{-1}(Z_{i})$ and ${\widetilde{\varphi}}_{i}^{-1}({\widetilde{Z}}_{i})$.
Remark 4.11.
Under the notation of Proposition 4.9, the reader may expect that for an ideal ${\mathfrak{a}}\subset k[x_{1},\ldots,x_{N}]$, there exists a lifting ${\widetilde{\mathfrak{a}}}\subset L[x_{1},\ldots,x_{N}]$ such that
[…]
It is a reasonable and preferable expectation. In some simple cases, we can verify it (see Proposition 4.12 below). However, at present, we cannot construct a lifting $\widetilde{\mathfrak{a}}$} satisfying (21) for $E_{n}$ and ${\widetilde{E}}_{n}$. A compromised statement is our main theorem (Theorem 1.1) which will be proved in the next section. The following are simple examples that the equality (21) holds for an appropriate lifting $\widetilde{\mathfrak{a}}$} of $\mathfrak{a}$}.
Proposition 4.12.
Under the notation of the statement in Proposition 4.6, the following hold:
  1. (i)
[…]
  2. (ii)
    Let $E_{2}\subset A_{2}$ be a prime divisor obtained by the blow-up at a $k$-valued closed point $P_{2}\in E_{1}\subset A_{1}$ and ${\mathfrak{a}}\subset k[x_{1},\ldots,x_{N}]$ a nonzero ideal. Then, the following hold:
    1. (a)
[…]
    2. (b)
      By choosing compatible skeletons $\Sigma\subset k$ and $\widetilde{\Sigma}$$\subset{\mathbb{C}}$, construct the skeleton schemes $B_{i}$ and ${\widetilde{B}}_{i}$ for $i=1,2$ as in the proof of Proposition 4.6. Denote the canonical morphism $B_{2}\to{\mathbb{A}}_{\Sigma}^{N}$ and ${\widetilde{B}}_{2}\to{\mathbb{A}}_{\widetilde{\Sigma}}^{N}$ by $\psi$ and $\widetilde{\psi}$, respectively. Then, the canonical map
      $\widetilde{\psi}_{*}({\mathcal{O}}_{{\widetilde{B}}_{2}}(-m{\widetilde{E}}_{2}))\to\psi_{*}({\mathcal{O}}_{B_{2}}(-mE_{2}))$
      is surjective for every $m\in{\mathbb{Z}}_{\geq 0}$.
    3. (c)
      For every pair $m<t$ of positive integers the canonical map is surjective:
      $\widetilde{\psi}_{*}\left(({\mathcal{O}}_{{\widetilde{B}}_{2}}(-m{\widetilde{E}}_{2}))/{\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{2}}\right)\to\psi_{*}\left(({\mathcal{O}}_{B_{2}}(-mE_{2}))/{\mathfrak{m}}^{t}{\mathcal{O}}_{B_{2}}\right),$
      where ${\mathfrak{m}}=(x_{1},\ldots,x_{N})\subset\Sigma[x_{1},\ldots,x_{N}]$ and ${\widetilde{\mathfrak{m}}}=(x_{1},\ldots,x_{N})\subset{\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]$.
Proof.
[…]
For (i), take appropriate compatible skeletons $\Sigma$ and $\widetilde{\Sigma}$} so that $f\in\Sigma[x_{1},\ldots,x_{N}]$. For
$f=\sum_{i\in{\mathbb{Z}}^{N}_{\geq 0}}a_{i}{\mathbb{x}}^{i}\ \ \ ,\ \ \mbox{take}\ \ \ {\tilde{f}}=\sum_{i\in{\mathbb{Z}}^{N}_{\geq 0}}{\widetilde{a}}_{i}\mathbb{x}^{i},$
where ${\widetilde{a}}_{i}\in{\widetilde{\Sigma}}$ is a lifting of $a_{i}\in\Sigma$ such that ${\widetilde{a}}_{i}=0$ if $a_{i}=0$. Then, we can see that ${\mathrm{ord}}_{\mathbb{x}}f={\mathrm{ord}}_{\mathbb{x}}{\tilde{f}}$, which implies
[…]
For (ii), we may assume that locally the following hold as $P_{2}$ is a $k$-valued point:
$(B_{1},P_{2})=\left({\mathrm{Spec\ }}\Sigma\left[x_{1},\frac{x_{2}}{x_{1}},\ldots,\frac{x_{N}}{x_{1}}\right],\left(x_{1},\frac{x_{2}}{x_{1}},\ldots,\frac{x_{N}}{x_{1}}\right)\right)$
and
$({\widetilde{B}}_{1},{\widetilde{P}}_{2})=\left({\mathrm{Spec\ }}{\widetilde{\Sigma}}\left[x_{1},\frac{x_{2}}{x_{1}},\ldots,\frac{x_{N}}{x_{1}}\right],\left(x_{1},\frac{x_{2}}{x_{1}},\ldots,\frac{x_{N}}{x_{1}}\right)\right),$
for an appropriate compatible skeletons $\Sigma\subset k$ and $\widetilde{\Sigma}$$\subset{\mathbb{C}}$.
For $f\in\Sigma[x_{1},\ldots,x_{N}]$, if ${\mathrm{ord}}_{\mathbb{x}}f=m_{1}$, then $f$ factors as follows:
[…]
on $B_{1}$. If
$g=\sum_{e}a_{e}x_{1}^{e_{1}}\left(\frac{x_{2}}{x_{1}}\right)^{e_{2}}\ldots\left(\frac{x_{N}}{x_{1}}\right)^{e_{N}},$
where $a_{e}\in\Sigma$ and $e=(e_{1},\ldots,e_{N})$, define
${\tilde{g}}=\sum_{e}{\widetilde{a}}_{e}x_{1}^{e_{1}}\left(\frac{x_{2}}{x_{1}}\right)^{e_{2}}\ldots\left(\frac{x_{N}}{x_{1}}\right)^{e_{N}},$
by using a lifting ${\widetilde{a}}_{e}\in{\widetilde{\Sigma}}$ of $a_{e}$ such that ${\widetilde{a}}_{e}=0$ if $a_{e}=0$. Then, ${\mathrm{mult}}_{P_{2}}g={\mathrm{mult}}_{{\widetilde{P}}_{2}}{\tilde{g}}$, which we denote by $m_{2}$. Therefore, the lifting ${\tilde{f}}:=x_{1}^{m_{1}}{\tilde{g}}$ of $f$ satisfies
[…]
which yields (ii), (a).
For the proof of (ii),(b), note that for every polynomial $f\in\Sigma[x_{1},\ldots,x_{N}]$ with $v_{E_{2}}(f)=m$ ($m\in{\mathbb{Z}}_{\geq 0}$) has a lifting $\tilde{f}$} such that $v_{{\widetilde{E}}_{2}}({\tilde{f}})=m$ as is constructed above. This yields that the modulo $p$ reduction $\widetilde{\psi}_{*}({\mathcal{O}}_{{\widetilde{B}}_{2}}(-m{\widetilde{E}}_{2}))\to\psi_{*}({\mathcal{O}}_{B_{2}}(-mE_{2}))$ is surjective for $m\in{\mathbb{Z}}_{\geq 0}$.
For the proof of (ii), (c), in the construction $\tilde{f}$} from $f$, by the expression of $g$ and $\tilde{g}$}, we have
$\deg_{x}f=\max\{e_{1}+m_{1}-(e_{2}+\cdots+e_{N})\mid a_{e}\neq 0\}=\deg_{x}{\tilde{f}}$.
[…]
5. Proof of the Main Theorem and applications
We fix the notation used in this section. Let $k$ be an algebraically closed field of characteristic $p>0$. Let $A={\mathbb{A}}_{k}^{N}$ be the affine space over $k$ of dimension $N\leq 3$ and 0 the origin of $A$.
Let $E$ be a prime divisor over $(A,0)$ and
$A_{n}\stackrel{{\scriptstyle\varphi_{n}}}{{\longrightarrow}}A_{n-1}\to\cdots A_{1}\stackrel{{\scriptstyle\varphi_{1}}}{{\longrightarrow}}A$
be a normal Gorenstein Zariski sequence for $E$. Proposition 4.9 guarantees the existence of a reasonable lifting of the sequence:
${\widetilde{A}}_{n}\stackrel{{\scriptstyle\widetilde{\varphi}_{n}}}{{\longrightarrow}}{\widetilde{A}}_{n-1}\to\cdots{\widetilde{A}}_{1}\stackrel{{\scriptstyle\widetilde{\varphi}_{1}}}{{\longrightarrow}}{\widetilde{A}}={\mathbb{A}}_{L}^{N},$
For a technical reason, we will blow up two more times from $A_{n}$ and lift them to the blow-ups from ${\widetilde{A}}_{n}$, which will lead to the main statement in this paper. Here, focusing at the part of the extra two blow-ups, we see that an ideal in positive characteristic is lifted to that in characteristic 0 with the preferable property (see, Proposition 4.12, (ii)). We will reduce our problem into the two blow-ups case in Proposition 4.12, (ii). In the following discussions, we will use the notation in Proposition 4.9.
Lemma 5.1.
Let $E$ be a prime divisor over $(A,0)$, and
$A_{n}\stackrel{{\scriptstyle\varphi_{n}}}{{\longrightarrow}}A_{n-1}\to\cdots A_{1}\stackrel{{\scriptstyle\varphi_{1}}}{{\longrightarrow}}A$
be a normal Gorenstein Zariski sequence for $E$, We use the same notation as in Proposition 4.9. Let
$A_{n+2}\stackrel{{\scriptstyle\varphi_{n+2}}}{{\longrightarrow}}A_{n+1}\stackrel{{\scriptstyle\varphi_{n+1}}}{{\longrightarrow}}A_{n}$
be the composite of two blow-ups at general closed points $P_{n+1}\in E_{n}\subset A_{n}$ and $P_{n+2}\in F_{1}\subset A_{n+1}$, where $F_{1}$ is the exceptional divisor for $\varphi_{n+1}$. Denote the exceptional divisor for $\varphi_{n+2}$ by $F_{2}$.
Then, by taking a skeletons $\Sigma\subset k$ sufficiently large, and letting $\widetilde{\Sigma}$$\subset{\mathbb{C}}$ be its lifting, one obtains liftings ${\widetilde{P}}_{n+1}\in{\widetilde{A}}_{n}$ and ${\widetilde{P}}_{n+2}\in{\widetilde{A}}_{n+1}$ of $P_{n}$ and $P_{n+1}$, respectively, and the sequence:
${\widetilde{A}}_{n+2}\stackrel{{\scriptstyle\widetilde{\varphi}_{n+2}}}{{\longrightarrow}}{\widetilde{A}}_{n+1}\stackrel{{\scriptstyle\widetilde{\varphi}_{n+1}}}{{\longrightarrow}}{\widetilde{A}}_{n},$
the composite of two blow-ups at ${\widetilde{P}}_{n+1}\in{\widetilde{E}}_{n}\subset{\widetilde{A}}_{n}$ and ${\widetilde{P}}_{n+2}\in{\widetilde{F}}_{1}\subset{\widetilde{A}}_{n+1}$, where ${\widetilde{F}}_{1}$ and ${\widetilde{F}}_{2}$ are the exceptional divisor of $\widetilde{\varphi}_{n+1}$ and $\widetilde{\varphi}_{n+2}$, respectively.
Then, we obtain
$k_{F_{1}}=k_{{\widetilde{F}}_{1}}=N-1+k_{E_{n}}=N-1+k_{{\widetilde{E}}_{n}}\ \mbox{and}$
$k_{F_{2}}=k_{{\widetilde{F}}_{2}}=2(N-1)+k_{E_{n}}=2(N-1)+k_{{\widetilde{E}}_{n}}.$
Proof.
Take a general closed point $P_{n+1}\in E_{n}\subset A_{n}$ such that
$P_{n+1}\not\in\bigcup_{j<n}(E_{j}\cup E^{\prime}_{j}),$
where we denote the proper transform of $E_{j}$ and $E^{\prime}_{j}$ on $A_{n}$ by the same symbols as $E_{j}$ and $E^{\prime}_{j}\subset A_{j}$. Here, we denote the fiber $\varphi_{j}^{-1}(Z_{j})=E_{j}+E^{\prime}_{j}$ as Cartier divisors on $A_{j}$.
Next, take a general closed point $P_{n+2}\in F_{1}\subset A_{n+1}$ such that
$P_{n+2}\not\in E_{n},$
where $E_{n}$ denotes the proper transform of $E_{n}$ on $A_{n+1}$ as well. Here, we note that $P_{n+1}$ and $P_{n+2}$ are $k$-valued point as $k$ is algebraically closed. So, we can realize their defining ideals in the skeletons $S_{n}$ and $S_{n+1}$ for sufficiently large $\Sigma\subset k$. Then, in the same way as in Proposition 4.6, we can construct a sequence
${\widetilde{A}}_{n+2}\stackrel{{\scriptstyle\widetilde{\varphi}_{n+2}}}{{\longrightarrow}}{\widetilde{A}}_{n+1}\stackrel{{\scriptstyle\widetilde{\varphi}_{n+1}}}{{\longrightarrow}}{\widetilde{A}}_{n},$
of blow-ups with the centers ${\widetilde{P}}_{n+1}\in{\widetilde{E}}_{n}$ and ${\widetilde{P}}_{n+2}\in{\widetilde{F}}_{1}$ of liftings of $P_{n+1}$ and $P_{n+2}$, respectively. Hence, by the choice of the points of blow-ups, we obtain the required equalities. ∎
Lemma 5.2.
Under the notation as in the beginning of this section and in Lemma 5.1, consider the ideals
$\underline{\widetilde{I}}_{1}={\widetilde{\mathfrak{m}}}=(x_{1},\ldots,x_{N})\subset{\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]={\widetilde{S}}_{0}\ \ \mbox{and}$
$\underline{I}_{1}={\mathfrak{m}}=(x_{1},\ldots,x_{N})\subset\Sigma[x_{1},\ldots,x_{N}]=S_{0}.$
Then, the following hold for $1\leq i\leq n+2$:
  1. (i)
    ${\mathcal{O}}_{{\widetilde{B}}_{i}}/{\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{i}}$ is $\mathbb{Z}$}-flat at $p\in{\mathbb{Z}}$ for every $t\in{\mathbb{N}}$, and
  2. (ii)
    ${\mathcal{O}}_{{\widetilde{B}}_{i}}/{\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{i}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}={\mathcal{O}}_{B_{i}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{B_{i}}$.
Proof.
For $1\leq i\leq n+2$, denote the proper transforms on $B_{i}$ of the prime divisors $E_{j}\subset B_{j}$ and ${\widetilde{E}}_{j}\subset{\widetilde{B}}_{j}$ $(i<j)$ by the same symbols $E_{j}$ and ${\widetilde{E}}_{j}$, as we did in the previous discussions. (For consistency of the notation, only in this proof, we denote $F_{i}$ and ${\widetilde{F}}_{i}$ $(i=1,2)$ by $E_{n+i}$ and ${\widetilde{E}}_{n+i}$, respectively.) As $Z_{i}$ ($i\geq 2$) is not necessarily irreducible, the irreducible component of the exceptional divisor of $\varphi_{i}:B_{i}\to B_{i-1}$ is not only $E_{i}$.
As $\widetilde{\varphi}_{1}$ is the blow-up by $\widetilde{\mathfrak{m}}=(x_{1},\ldots,x_{N})\subset{\widetilde{S}}_{0}$, the pull-back ${\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{1}}$ is an invertible ideal sheaf on ${\widetilde{B}}_{1}$. Then, for every $i\geq 1$ its pull-back ${\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{i}}$ on ${\widetilde{B}}_{i}$ is also an invertible ideal sheaf on ${\widetilde{B}}_{i}$. For a similar reason, ${\mathfrak{m}}^{t}{\mathcal{O}}_{B_{i}}$ is also an invertible ideal sheaf for every $i\geq 1$.
For the proof of (i), let ${\widetilde{D}}_{i}\subset{\widetilde{B}}_{i}$ be the subscheme defined by ${\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{i}}$, and $D_{i}\subset B_{i}$ the subscheme defined by ${\mathfrak{m}}^{t}{\mathcal{O}}_{B_{i}}$ for $1\leq i\leq n+2$. Then, ${\widetilde{D}}_{i}\subset{\widetilde{B}}_{i}$ is locally a hypersurface and its support is contained in the exceptional divisor of $\widetilde{\phi}_{i}:{\widetilde{B}}_{i}\to{\widetilde{B}}_{0}$. Since the exceptional divisor of $\widetilde{\phi}_{i}$ does not contain $B_{i}$, the closed subscheme ${\widetilde{D}}_{i}$ does not contain the fiber $B_{i}$ over $p$, which yields that ${\widetilde{D}}_{i}$ is flat over $\mathbb{Z}$} at $p$ (see, [16, Corollary, p.177]). This completes the proof of (i).
For the proof of (ii), we first claim that the supports of $D_{i}$ and ${\widetilde{D}}_{i}$ are expressed as follows:
(22) ${\mathrm{supp}}D_{i}=\bigcup_{j\leq i}E_{j},\ \mbox{and}\ \ {\mathrm{supp}}{\widetilde{D}}_{i}=\bigcup_{j\leq i}{\widetilde{E}}_{j}\ \ \mbox{in\ a\ neighborhood\ of}\ p.$
(23) ${\mathrm{supp}}\left({\widetilde{D}}_{i}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}\right)=\bigcup_{j\leq i}E_{j}.$
When $E_{j}$’s are all obtained by closed-point-blow-ups, these three equalities are obvious. Otherwise, we remark again that the irreducible components of the exceptional set of $\widetilde{\phi}_{i}$ are not only ${\widetilde{E}}_{1},\ldots,{\widetilde{E}}_{i}$, and the irreducible components of the exceptional set of $\phi$ are not only $E_{1},\ldots,E_{i}$, because each center of the blow-up ${\widetilde{\varphi}}_{j}$ and $\varphi_{j}$ $(j\leq i)$ is not necessarily irreducible. We will prove the equalities in (22) by induction on $i$.
Let $i>1$ be the smallest number that $\varphi_{i}:A_{i}\to A_{i-1}$ is not a blow-up by a closed point. We may assume that (22) hold for $D_{i-1}\subset B_{i-1}$ and ${\widetilde{D}}_{i-1}\subset{\widetilde{B}}_{i-1}$. Now, assume there exists an irreducible component $C$ of $D_{i}$ such that
$C\not\subset\bigcup_{j\leq i}E_{j}.$
Then, obviously $\varphi_{i}(C)\subset D_{i-1}$. Therefore, $\varphi_{i}(C)\subset Z_{i}$, because otherwise, $\varphi_{i}$ is isomorphic at the generic point of $C$ and by the hypothesis of the induction, we have $\varphi_{i}(C)\subset E_{j}\subset B_{i-1}$ for some $j<i$, which gives $C\subset E_{j}\subset B_{i}$, a contradiction to the assumption. Hence, we obtain
$\varphi_{i}(C)\subset Z_{i}\cap D_{i-1}\subset B_{i-1}\ \mbox{ and}\ \ \varphi_{i}(C)\not\subset V({\mathfrak{p}}_{i}).$
By the construction of $Z_{i}$ (see, Proposition 4.6), it follows
${\mathrm{codim}}(Z_{i}\cap D_{i-1}\setminus V(\underline{{\mathfrak{p}}}_{i}),B_{i-1})=c_{i}+1:=0pt{\mathfrak{p}}_{i}+1,$
which yields
${\mathrm{codim}}(C,B_{i})\geq{\mathrm{codim}}(\varphi_{i}^{-1}(\varphi_{i}(C)),B_{i})\geq{\mathrm{codim}}(\varphi_{i}(C),B_{i-1})-(c_{i}-1)$
$\geq{\mathrm{codim}}(Z_{i}\cap D_{i-1}\setminus V(\underline{{\mathfrak{p}}}_{i}),B_{i-1})-(c_{i}-1)=2,$
a contradiction to that $C$ is an irreducible component of a closed subscheme defined by a locally principal ideal on an irreducible reduced scheme $B_{i}$. Therefore we obtain
${\mathrm{supp}}D_{i}=\bigcup_{j\leq i}E_{j}.$
For the second equality of (22), we use
${\mathrm{codim}}({\widetilde{Z}}_{i}\cap{\widetilde{D}}_{i-1}\setminus V({\underline{\widetilde{\mathfrak{p}}}}_{i}),{\widetilde{B}}_{i-1})\geq{\mathrm{codim}}(Z_{i}\cap D_{i-1}\setminus V(\underline{{\mathfrak{p}}}_{i}),B_{i-1})=c_{i}+1$
in a neighborhood of $p\in{\mathbb{Z}}$. Then, by a similar discussion as above, we obtain
${\mathrm{supp}}{\widetilde{D}}_{i}=\bigcup_{j\leq i}{\widetilde{E}}_{i}.$
[…]
${\mathrm{supp}}({\widetilde{D}}_{i}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)})\subset{\mathrm{supp}}\phi_{i}^{-1}(V({\mathfrak{m}}))=\bigcup_{j\leq i}E_{j}\ \ \mbox{as\ sets}.$
On the other hand, for each $j\leq i$,
${\mathrm{supp}}({\widetilde{E}}_{j}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)})\supset{\mathrm{supp}}E_{j},$
because ${\widetilde{E}}_{j}$ is a weak prime lifting of $E_{j}$ by Proposition 4.9. Therefore, we obtain the equality of the sets
${\mathrm{supp}}({\widetilde{D}}_{i}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)})=\bigcup_{j\leq i}E_{j}.$
In order to prove (ii), it is sufficient to show the equality
${\mathfrak{m}}^{t}{\mathcal{O}}_{B_{i}}={\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{i}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}$
[…]
${\mathfrak{m}}^{t}{\mathcal{O}}_{B_{iQ}}={\mathcal{O}}_{B_{iQ}}\left(-\sum_{j\leq i}a_{j}E_{j}\right),$
${\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{A}}_{i}}={\mathcal{O}}_{{\widetilde{A}}_{i}}\left(-\sum_{j\leq i}{\widetilde{a}}_{j}{\widetilde{E}}_{j}\right).$
As $\left({\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{i}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)}\right)\otimes_{\Sigma}Q(\Sigma)={\mathcal{O}}_{B_{iQ}}(-\sum_{j\leq i}{\widetilde{a}}_{j}E_{j})$, for the proof of (ii), it is sufficient to show $a_{j}={\widetilde{a}}_{j}$ ($j\leq i$). We will prove this by induction on $i$.
For $i=1$, it is clear that $a_{1}={\widetilde{a}}_{1}$. Assume $i\geq 2$ and $a_{j}={\widetilde{a}}_{j}$ holds for $j\leq i-1$. By the construction, $P_{i}\in B_{i-1Q}$ and its weak prime lifting ${\widetilde{P}}_{i}\in{\widetilde{A}}_{i-1}$ satisfy
$P_{i}\in E_{j}$ if and only if ${\widetilde{P}}_{i}\in{\widetilde{E}}_{j}$ for $j\leq i$.
$a_{i}=\sum_{P_{i}\in E_{j}}a_{j}=\sum_{{\widetilde{P}}_{i}\in{\widetilde{E}}_{j}}{\widetilde{a}}_{j}={\widetilde{a}}_{i},$
where the first and the last equalities follows from the fact that $P_{i}\in A_{i-1}$ and ${\widetilde{P}}_{i}\in{\widetilde{A}}_{i-1}$ are smooth (see the last part of the proof of Proposition 4.9). Hence, the statement holds for $i$. ∎
Theorem 5.3 (Theorem 1.1, (i), (ii)).
Let $N$ be an integer $\geq 2$, and $k$ an algebraically closed field of characteristic $p>0$. Let $A={\mathbb{A}}_{k}^{N}$ be the affine space over $k$ of dimension $N$, and 0 the origin of $A$. Let $E$ be a prime divisor over $A$ with the center at 0, and ${\mathfrak{a}}\subset k[x_{1},\ldots,x_{N}]$ a non-zero ideal. Assume either:
  1. (1)
    $E$ is obtained by blow-ups at closed points, or
  2. (2)
    $N\leq 3$.
Then, there exists a prime divisor $F_{\mathbb{C}}$ over the affine space $A_{\mathbb{C}}={\mathbb{A}}_{\mathbb{C}}^{N}$ over $\mathbb{C}$} with the center at the origin $\tilde{0}$} satisfying the following:
  1. (i)
    $2(N-1)+k_{E}=k_{F_{\mathbb{C}}}$;
  2. (ii)
    there exists a lifting ${\widetilde{\mathfrak{a}}}\subset{\mathcal{O}}_{A_{\mathbb{C}}}$ of $\mathfrak{a}$} satisfying
[…]
Proof.
First, note that although we assume (1) and (2) for the prime divisor $E$, our proof here works for any $E$ such that the statement of Proposition 4.9 and Lemma 5.2 hold.
[…]
Let $F_{\mathbb{C}}$ be one of the irreducible components of ${\widetilde{F}}_{2}\times_{{\mathbb{A}}_{L}^{N}}{\mathbb{C}}$ on ${\widetilde{A}}_{n+2}\times_{{\mathbb{A}}_{L}^{N}}{\mathbb{C}}$. Then, as $k_{{\widetilde{F}}_{2}}=k_{F_{\mathbb{C}}}$, the statement (i) follows from Lemma 5.1.
We prove the statements (ii) for $N\geq 3$ first, and then, applying Lemma 5.4, we will obtain the results for $N=2$.
Now, let us assume $N\geq 3$. First remember the definition of the blow-up centers. In Lemma 5.1, we took the closed point $P_{n+1}\in A_{n}$ general. We may assume
$P_{n+1}\not\in Z({\mathfrak{a}}_{A_{n}})\cup\bigcup_{j<n}E_{j},$
where ${\mathfrak{a}}_{A_{n}}$ is the weak transform of the ideal ${\mathfrak{a}}\subset R_{0}$ on $A_{n}$ and $E_{j}$ $(j<n)$ denotes the proper transform of $E_{j}$ on $A_{n}$. (For the definition of weak transform, see, e.g., [13]. The important property we use is only that the locus of the weak transform does not contain the whole $E_{n}$.) We take $P_{n+2}\in A_{n+1}$ as in Lemma 5.1.
By the choice of $P_{n+1}$ and $P_{n+2}$, it follows
$v_{E_{n}}({\mathfrak{a}})=v_{F_{2}}({\mathfrak{a}}).$
Denote the divisors on $B_{n}$ and on $B_{n+j}$ $(j=1,2)$ corresponding to $E_{n}$ and $F_{j}$ by the same symbols, and the ideal in the skeleton $S_{0}$ generating $\mathfrak{a}$} is denoted by $\underline{{\mathfrak{a}}}$. Then, we have
$v_{E_{n}}(\underline{{\mathfrak{a}}})=v_{F_{2}}(\underline{{\mathfrak{a}}}).$
Therefore, it is sufficient to show that for a nonzero ideal ${\mathfrak{a}}\subset k[x_{1},\ldots,x_{N}]$ there exist compatible skeletons $\Sigma\subset k$ and $\widetilde{\Sigma}$$\subset{\mathbb{C}}$ such that $\mathfrak{a}$} is generated by $\underline{{\mathfrak{a}}}\subset\Sigma[x_{1},\ldots,x_{N}]$ and a lifting ${\widetilde{\mathfrak{a}}}\subset{\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]$ of $\underline{{\mathfrak{a}}}$ satisfies
[…]
We will prove (24) for the case $N\geq 3$ first, and then will prove for the case $N=2$.
[Proof of (ii) for the case $N\geq 3$]
Assume $N\geq 3$. Consider the commutative diagram
$\begin{array}[]{ccc}B_{n+2}&\stackrel{{\scriptstyle\varphi}}{{\longrightarrow}}&B_{0}\\ \downarrow&&\downarrow\\ {\widetilde{B}}_{n+2}&\stackrel{{\scriptstyle{\widetilde{\varphi}}}}{{\longrightarrow}}&{\widetilde{B}}_{0},\end{array}$
where the vertical arrows are the canonical closed immersions induced from the surjection ${\widetilde{\Sigma}}\to\Sigma$.
Note that $v_{F_{2}}(\underline{{\mathfrak{a}}})\geq m$ if and only if $\underline{{\mathfrak{a}}}\subset\varphi_{*}({\mathcal{O}}_{B_{n+2}}(-mF_{2}))$. Then, to prove (24), it is sufficient to prove that there exists a lifting ${\widetilde{\mathfrak{a}}}\subset{\widetilde{S}}_{0}={\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]$ of $\underline{{\mathfrak{a}}}$ such that
${\widetilde{\mathfrak{a}}}\subset{\widetilde{\varphi}}_{*}({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2}))\ \ \mbox{for}\ \ m=v_{F_{2}}(\underline{{\mathfrak{a}}}),$
because $v_{{\widetilde{F}}_{2}}({\widetilde{\mathfrak{a}}})\leq v_{F_{2}}(\underline{{\mathfrak{a}}})$ is clear for any lifting $\widetilde{\mathfrak{a}}$} of $\underline{{\mathfrak{a}}}$. Therefore, we have only to prove the surjectivity of the canonical map:
(25) ${\widetilde{\varphi}}_{*}({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2}))\to\varphi_{*}({\mathcal{O}}_{B_{n+2}}(-mF_{2}))$
for every $m\in{\mathbb{N}}$. As the monomials of degree >m>m do not affect the valuation $m=v_{F_{2}}(f)$ of the polynomial $f\in\Sigma[x_{1},\ldots,x_{N}]$, we can neglect higher degree term of the polynomial. Hence, it is sufficient to show the surjectivity of
(26) ${\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2}))/{\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{n+2}}\right)\stackrel{{\scriptstyle\alpha}}{{\longrightarrow}}\varphi_{*}\left({\mathcal{O}}_{B_{n+2}}(-mF_{2}))/{\mathfrak{m}}^{t}{\mathcal{O}}_{B_{n+2}}\right),$
for $t>m$. Now we reduce the problem into the problem about coherent sheaves with the support on the exceptional divisors. Here, denote ${\widetilde{\mathfrak{m}}}=(x_{1},\ldots,x_{N})\subset{\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]$ and ${\mathfrak{m}}=(x_{1},\ldots,x_{N})\subset\Sigma[x_{1},\ldots,x_{N}]$. For simplicity of the notation, we denote
${\mathcal{F}}:={\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2}))/{\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{n+2}},$
then, $\mathcal{F}$} is $\mathbb{Z}$}-flat and the following holds
${\mathcal{O}}_{B_{n+2}}(-mF_{2}))/{\mathfrak{m}}^{t}{\mathcal{O}}_{B_{n+2}}={\mathcal{F}}\otimes_{\mathbb{Z}}{{\mathbb{Z}}/(p)},$
by Lemma 5.2.
As the map $\alpha$ in (26) is a part of the following long exact sequence:
[…]
it is sufficient to show that the last map is injective, i.e.,
[…]
Fix an integer $t$ such that $t>m$ and factor the locally principal ideals ${\mathfrak{m}}^{t}{\mathcal{O}}_{B_{n+2}}$ and ${\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{n+2}}$ into the following forms:
${\mathfrak{m}}^{t}{\mathcal{O}}_{B_{n+2}}=I\cdot I_{1}^{r_{1}}\cdot I_{2}^{r_{2}},$
${\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{n+2}}={\widetilde{I}}\cdot{\widetilde{I}}_{1}^{r_{1}}\cdot{\widetilde{I}}_{2}^{r_{2}},$
where for $i=1,2$, the ideals $I_{i}$ and ${\widetilde{I}}_{i}$ are the defining ideals of the proper transforms of the exceptional divisors $F_{i}$ and ${\widetilde{F}}_{i}$ on $B_{n+2}$ and ${\widetilde{B}}_{n+2}$, respectively. We denote the remaining factors by $I$ and $\widetilde{I}$}, respectively. Here, by the construction of blow-ups, ${\widetilde{I}},{\widetilde{I}}_{1}$ and ${\widetilde{I}}_{2}$ are liftings of $I,I_{1}$ and $I_{2}$, and
$r_{1}=r_{2}=r,$
where $r$ is the exponent of the defining ideal of $E_{n}$ in $I$, which is the same as that of ${\widetilde{E}}_{n}$ in $\widetilde{I}$} by Lemma 5.2. Consider the following exact sequence of Mayer-Vietoris type:
0
$\downarrow$
${\mathcal{F}}={\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/({\widetilde{I}}\cdot{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r})$
(27) $\downarrow$
$({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}})\oplus({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r})$
$\downarrow$
${\mathcal{O}}_{{\widetilde{B}}_{n+2}}/({\widetilde{I}}+{\widetilde{I}}_{1}^{r})$
$\downarrow$
$0.$
Here, we note that $\left({\mathrm{supp}}\ {\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\right)\cap{\widetilde{F}}_{2}=\emptyset$, which yields
${\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/{\widetilde{I}}={\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}$.
From this exact sequence, we obtain the long exact sequence of direct image sheaves. First consider the case $m=0$ as follows:
0
$\downarrow$
${\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\cdot{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)$
$\iota_{0}\downarrow$
${\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\right)\oplus{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)$
$\pi_{0}\downarrow$
${\widetilde{\varphi}}_{*}({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/({\widetilde{I}}+{\widetilde{I}}_{1}^{r}))$
(28) $\downarrow$
$R^{1}{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\cdot{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)\$
$\downarrow$
$R^{1}{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\right)\oplus R^{1}{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)$
$\downarrow$
$R^{1}{\widetilde{\varphi}}_{*}{\mathcal{O}}_{{\widetilde{B}}_{n+2}}/({\widetilde{I}}+I_{1}^{r}).$
Here, we will show the following under the condition $N\geq 3$:
  1. (a)
[…]
  2. (b)
    $R^{1}{\widetilde{\varphi}}_{*}({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\cdot{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r})=0.$
  3. (c)
    $R^{1}{\widetilde{\varphi}}_{*}({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}+{\widetilde{I}}_{1}^{r})$ is $\mathbb{Z}$}-flat at $p$.
  4. (d)
[…]
[Proof of (a)] We will prove the following canonical map is isomorphic
[…]
We study of ${\widetilde{B}}_{n+1}$ instead of ${\widetilde{B}}_{n+2}$. Denote ${\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{n+1}}={{{{\widetilde{J}}}_{1}}}^{r}\cdot{\widetilde{I}}$, where ${\widetilde{J}}_{1}$ is the defining ideal of the latest exceptional divisor ${\widetilde{F}}_{1}\subset{\widetilde{B}}_{n+1}$. By the definition of the center ${\widetilde{P}}_{n+2}\in{\widetilde{B}}_{n+1}$ of blow-up ${\widetilde{\varphi}}_{n+2}$, we have the isomorphism
${\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\simeq{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{I}}$.
Here, we use the same symbol $\widetilde{I}$} to describe the defining ideals of the image of $Z({\widetilde{I}})\subset{\widetilde{B}}_{n+2}$ in ${\widetilde{B}}_{n+1}$. As
${\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\cdot{{\widetilde{I}}_{1}}^{r}\cdot{{\widetilde{I}}_{2}}^{r}\right)={\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]/(x_{1},\ldots,x_{N})^{t}={\widetilde{\psi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{I}}\cdot{{\widetilde{J}}_{1}}^{r}\right),$
it is sufficient to show
[…]
where ${\widetilde{\psi}}:{\widetilde{B}}_{n+1}\to{\widetilde{B}}_{0}$ is the composite ${\widetilde{\varphi}}_{1}\circ\cdots\circ{\widetilde{\varphi}}_{n+1}$. Considering the exact sequence:
$0\to{\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{{\widetilde{J}}_{1}}^{r}\to{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{I}}\cdot{{\widetilde{J}}_{1}}^{r}\to{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{I}}\to 0,$
we will see the following vanishing for $i=0,1$
(31) $R^{i}{\widetilde{\psi}}_{*}({\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{{\widetilde{J}}_{1}}^{r})=0,$
which yields the isomorphism (30).
Indeed, for $j\leq r-1$, noting that
(32) ${\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{F}}_{1}}={\widetilde{I}}_{{\widetilde{E}}_{n}}^{r}\otimes{\mathcal{O}}_{{\widetilde{F}}_{1}}={\mathcal{O}}_{{\mathbb{P}}_{\widetilde{\Sigma}}^{N-1}}(-r)\ \ \mbox{and}\ \ {\widetilde{I}}\otimes{\widetilde{J}}_{1}^{j}\otimes{\mathcal{O}}_{{\widetilde{F}}_{1}}={\mathcal{O}}_{{\mathbb{P}}_{\widetilde{\Sigma}}^{N-1}}(j-r),$
[…]
(33) $0\to{\widetilde{I}}\otimes{\widetilde{J}}^{j}_{1}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}\to{\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/({\widetilde{J}}_{1}^{j+1})\to{\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/({\widetilde{J}}_{1}^{j})\to 0,$
we obtain the long exact sequence of the relative cohomologies
(34) $\to R^{i}{\widetilde{\psi}}_{*}({\widetilde{I}}\otimes{\widetilde{J}}^{j}_{1}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1})\to R^{i}{\widetilde{\psi}}_{*}({\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/({\widetilde{J}}_{1}^{j+1}))\to R^{i}{\widetilde{\psi}}_{*}({\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/({\widetilde{J}}_{1}^{j}))\to.$
Assume i<N1i<N-1, apply the vanishings of $R^{i}{\widetilde{\psi}}$ for the two sheaves in (32) with $j=1$ first, then by induction on $j$ up to j=r1j=r-1, we obtain (31) :
$R^{i}{\widetilde{\psi}}_{*}({\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{{\widetilde{J}}_{1}}^{r})=0,$
in particular, it holds for $i=0,1$ as we assume $N\geq 3$. This completes the proof of (30), and therefore the proof of $(a)$ as well.
[Proof of (b)]First we show for every $t\in{\mathbb{N}}$
(35) $R^{1}\varphi_{*}({\mathcal{O}}_{A_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{n+2}})=0,$
where we denote the maximal ideal $(x_{1},\ldots,x_{N})\subset k[x_{1},\ldots,x_{N}]$ also by $\mathfrak{m}$}. Factor the morphism $\varphi:A_{n+2}\to A$ by $\psi:A_{n+2}\to A_{1}$ and $\varphi_{1}:A_{1}\to A$. By Corollary 4.5 it follows:
(36) $R^{i}\psi_{*}{\mathcal{O}}_{A_{n+2}}=\left\{\begin{array}[]{cc}0&i>0\\ {\mathcal{O}}_{A_{1}}&i=0.\end{array}\right$.
Therefore, we have
(37) $R^{i}\psi_{*}{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{n+2}}=R^{i}\psi_{*}(\psi^{*}({\mathfrak{m}}^{t}{\mathcal{O}}_{A_{1}}))=(R^{i}\psi_{*}{\mathcal{O}}_{A_{n+2}})\otimes{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{1}}=\left\{\begin{array}[]{cc}0&i>0\\ {\mathfrak{m}}^{t}{\mathcal{O}}_{A_{1}}&i=0,\end{array}\right$.
where, as ${\mathfrak{m}}^{t}{\mathcal{O}}_{A_{1}}$ is invertible, we can apply the projection formula to obtain the second equality. Considering the exact sequence
$0\to{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{n+2}}\to{\mathcal{O}}_{A_{n+2}}\to{\mathcal{O}}_{A_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{n+2}}\to 0,$
by using (36) and (37) we obtain
(38) $R^{i}\psi_{*}({\mathcal{O}}_{A_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{n+2}})=\left\{\begin{array}[]{cc}0&i>0\\ {\mathcal{O}}_{A_{1}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{1}}&i=0.\end{array}\right$.
Now, take the Leray’s spectral sequence:
$E_{2}^{p,q}=R^{p}\varphi_{1*}R^{q}\psi_{*}({\mathcal{O}}_{A_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{n+2}})\Rightarrow E^{p+q}=R^{p+q}\varphi_{*}({\mathcal{O}}_{A_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{n+2}}).$
As $E_{2}^{1,0}=R^{1}\varphi_{1*}({\mathcal{O}}_{A_{1}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{1}})=0$ and $E_{2}^{0,1}=\varphi_{1*}R^{1}\psi_{*}({\mathcal{O}}_{A_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{n+2}})=0$, the canonical exact sequence $0\to E_{2}^{1,0}\to E^{1}\to E_{2}^{0,1}$ gives $E^{1}=0$, i.e.
$R^{1}\varphi_{*}({\mathcal{O}}_{A_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{A_{n+2}})=0,$
which is first required. This cohomology is the faithfully flat base change of
$R^{1}\varphi_{*}(({\mathcal{O}}_{B_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{B_{n+2}})\otimes_{\Sigma}Q(\Sigma))$
by the field extension $Q(\Sigma)\hookrightarrow k$, where $Q(\Sigma)$ is the field of fractions of $\Sigma$. Therefore,
(39) $R^{1}\varphi_{*}(({\mathcal{O}}_{B_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{B_{n+2}})\otimes_{\Sigma}Q(\Sigma))=0.$
Now, we denote
${\mathcal{G}}:={\mathcal{O}}_{B_{n+2}}/{\mathfrak{m}}^{t}{\mathcal{O}}_{B_{n+2}}\ \mbox{and}$
${\widetilde{{\mathcal{G}}}}:={\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{\mathfrak{m}}}^{t}{\mathcal{O}}_{{\widetilde{B}}_{n+2}}.$
[…]
(40) $(R^{1}\varphi_{*}{\mathcal{G}})\otimes_{\Sigma}Q(\Sigma)\simeq R^{1}\varphi_{*}({\mathcal{G}}\otimes_{\Sigma}Q(\Sigma))=0,$
where the vanishing follows from (39). On the other hand, the following exact sequence
$\to R^{1}{\widetilde{\varphi}}_{*}{\widetilde{{\mathcal{G}}}}\stackrel{{\scriptstyle\times p}}{{\longrightarrow}}R^{1}{\widetilde{\varphi}}_{*}{\widetilde{{\mathcal{G}}}}\to R^{1}\varphi_{*}{\mathcal{G}}\to\cdots$
gives the inclusion $R^{1}{\widetilde{\varphi}}_{*}{\widetilde{{\mathcal{G}}}}\otimes_{\widetilde{\Sigma}}\Sigma\hookrightarrow R^{1}\varphi_{*}{\mathcal{G}}$ from the cokernel of the map $\times p$. By applying the tensor $\otimes_{\Sigma}Q(\Sigma)$ on the inclusion, we obtain
$R^{1}{\widetilde{\varphi}}_{*}{\widetilde{{\mathcal{G}}}}\otimes_{\widetilde{\Sigma}}Q(\Sigma)\hookrightarrow R^{1}\varphi_{*}{\mathcal{G}}\otimes_{\Sigma}Q(\Sigma).$
By the vanishing (40), we obtain $R^{1}{\widetilde{\varphi}}_{*}{\widetilde{{\mathcal{G}}}}\otimes_{\widetilde{\Sigma}}Q(\Sigma)=0$, which yields
$R^{1}{\widetilde{\varphi}}_{*}{\widetilde{{\mathcal{G}}}}=0\ \ \mbox{around}\ \ \ p{\widetilde{\Sigma}}\in{\mathrm{Spec\ }}{\widetilde{\Sigma}}$.
[Proof of (c)] The support of ${\mathcal{O}}_{{\widetilde{B}}_{n+2}}/({\widetilde{I}}+{\widetilde{I}}_{1}^{r})$ is $C:={\widetilde{E}}_{n}\cap{\widetilde{F}}_{1}\simeq{\mathbb{P}}_{\widetilde{\Sigma}}^{N-2}$. As ${\widetilde{B}}_{n+2}$ is isomorphic to ${\widetilde{B}}_{n+1}$ around $C$, we will study ${\widetilde{B}}_{n+1}$ instead of ${\widetilde{B}}_{n+2}$ as we did in the proof of (a). We use the same symbols and discussion as in the proof of (a). Consider the following exact sequence:
$0\to{\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}^{r}\to{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}^{r}\to{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/({\widetilde{I}}+{\widetilde{J}}_{1}^{r})\to 0.$
Here, for $N\geq 4$, we have the following vanishing:
(41) $R^{2}{\widetilde{\psi}}_{*}({\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}^{r})=0.$
For $N=3$, we have
(42) $R^{2}{\widetilde{\psi}}_{*}({\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}^{r})\simeq\oplus_{j=0}^{r-1}H^{2}({\mathbb{P}}_{\widetilde{\Sigma}}^{2},{\mathcal{O}}_{{\mathbb{P}}_{\widetilde{\Sigma}}^{2}}(j-r)),$
which is torsion free, therefore it is $\mathbb{Z}$}-flat at $p$.
On the other hand, we can prove
(43) $R^{1}{\widetilde{\psi}}_{*}{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}^{r}=0,$
by using $R^{1}{\widetilde{\psi}}_{*}({\widetilde{J}}_{1}^{j}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1})=H^{1}\left({\mathbb{P}}^{N-1}_{\widetilde{\Sigma}},{\mathcal{O}}_{{\mathbb{P}}^{N-1}_{\widetilde{\Sigma}}}(j)\right)=0$ $(j\geq 0)$ and the exact sequences for $j>0$:
$0\to{\widetilde{J}}_{1}^{j}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}\to{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}^{j+1}\to{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}^{j}\to 0.$
By (43) we have exact sequence:
$0\to R^{1}{\widetilde{\psi}}{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/({\widetilde{I}}+{\widetilde{J}}_{1}^{r})\to R^{2}{\widetilde{\psi}}_{*}({\widetilde{I}}\otimes{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/{\widetilde{J}}_{1}^{r}).$
Now, by (41) and (42) the submodule $R^{1}{\widetilde{\psi}}{\mathcal{O}}_{{\widetilde{B}}_{n+1}}/({\widetilde{I}}+{\widetilde{J}}_{1}^{r})$ is also $\mathbb{Z}$}-flat at $p$ as required in (c).
[Proof of (d)] Due to the sequence (28), the statements (d) follows automatically from (b) and (c).
Now, we are on the last step to show $\mathbb{Z}$}-flatness of $R^{1}{\widetilde{\varphi}}_{*}{\mathcal{F}}$ at $p$ for $N\geq 4$. We can use (a)–(d). For $m>0$, by taking direct image sheaves of the exact sequence (27), we obtain the following long exact sequence:
0
$\downarrow$
${\widetilde{\varphi}}_{*}{\mathcal{F}}={\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/{\widetilde{I}}\cdot{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)$
$\iota_{m}\downarrow$
${\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\right)\oplus{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)$
$\pi_{m}\downarrow$
${\widetilde{\varphi}}_{*}({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/({\widetilde{I}}+{\widetilde{I}}_{1}^{r}))$
(44) $\downarrow$
$R^{1}{\widetilde{\varphi}}_{*}{\mathcal{F}}=R^{1}{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/{\widetilde{I}}\cdot{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)$
$\downarrow$
$R^{1}{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\right)\oplus R^{1}{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)$
$\downarrow$
$R^{1}{\widetilde{\varphi}}_{*}{\mathcal{O}}_{{\widetilde{B}}_{n+2}}/({\widetilde{I}}+{\widetilde{I}}_{1}^{r}).$
Applying (a) – (d) to the exact sequence (44), we obtain
$R^{1}{\widetilde{\varphi}}_{*}{\mathcal{O}}_{{\widetilde{B}}_{n+2}}/({\widetilde{I}}+{\widetilde{I}}_{1}^{r})\ \mbox{is}\ {\mathbb{Z}}\mbox{-flat}\ \mbox{at}\ p$
$R^{1}{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\right)\ \mbox{is}\ {\mathbb{Z}}\mbox{-flat}\ \mbox{at}\ p.$
Then, we have the exact sequence:
$0\to{\rm Coker\ }\pi_{m}\to R^{1}{\widetilde{\varphi}}_{*}{\mathcal{F}}\stackrel{{\scriptstyle\mu}}{{\longrightarrow}}R^{1}{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\right)\oplus R^{1}{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right),$
where the right factor of the last term is $\mathbb{Z}$}-flat at $p$, because our last two blow-ups are the same as the ones in Proposition 4.12, (ii) and the surjection in (c) of Proposition 4.12, (ii) gives the $\mathbb{Z}$}-flatness of the last term at $p$. Now, we obtain ${\rm Coker\ }\mu$ is $\mathbb{Z}$}-flat at $p$, because it is a submodule of $\mathbb{Z}$}-flat at $p$. Hence, in order to prove that $R^{1}{\widetilde{\varphi}}_{*}{\mathcal{F}}$ is $\mathbb{Z}$}-flat at $p$, it is sufficient to prove that ${\rm Coker\ }\pi_{m}$ is $\mathbb{Z}$}-flat at $p$.
[…]
(45) ${\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)\simeq{\widetilde{\varphi}}_{*}({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/({\widetilde{I}}+{\widetilde{I}}_{1}^{r})).$
Under the identification by (45), describe
$\pi_{m}:{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\right)\oplus{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)\to{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)$
by $\pi_{m}=(\gamma,\beta)$, where
$\gamma:{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\right)\to{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)$
$\beta:{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}(-m{\widetilde{F}}_{2})/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right)\to{\widetilde{\varphi}}_{*}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}_{1}^{r}\cdot{\widetilde{I}}_{2}^{r}\right).$
As the map $\beta$ and $\gamma$ appear in the following diagram with the first line exact:
[…]
Here, as ${\rm{supp}}\left({\mathcal{O}}_{{\widetilde{B}}_{n+2}}/{\widetilde{I}}\right)\cap{\widetilde{F}}_{2}=\emptyset$, the map $\delta$ is the zero map, which yields ${\rm Im}\gamma\subset{\rm Im}\beta$. Therefore, ${\rm Im}\pi_{m}={\rm Im}\gamma+{\rm Im}\beta={\rm Im}\beta,$ and ${\rm Coker\ }\pi_{m}={\rm Coker\ }\beta.$ By exactness of the first line of (46), it follows ${\rm Coker\ }\beta$ is contained in the $\mathbb{Z}$}-flat module ${\widetilde{\varphi}}_{*}{\mathcal{O}}_{m{\widetilde{F}}_{2}}$. Therefore, ${\rm Coker\ }\beta$ is $\mathbb{Z}$}-flat at $p$, and so is ${\rm Coker\ }\pi_{m}$ as required. [QED for $N\geq 3$]
[Proof of (ii) for the case $N=2$]
As we proved (ii) of Theorem 5.3 for $N\geq 3$, the following lemma provides the proof of (ii) for $N=2$. ∎
Lemma 5.4.
Let $N\geq 2$ be the dimension of the affine space $A={\mathbb{A}}_{k}^{N}$. If the Theorem 5.3 is proved for a prime divisor obtained by composite of blow-ups by closed points for the $(N+1)$-dimensional affine space by the method of the proof as above, then the theorem also holds for $N$-dimensional affine space.
Proof.
We use the notation in the proof of Theorem 5.3 for $N\geq 3$ shown above. Let $E$ be a prime divisor over $(A,0)$ obtained by composite of blow-ups by closed points for $A={\mathbb{A}}_{k}^{N}$, and $0\in A$ the origin. In the proof of Theorem 5.3 for case (2), we constructed sequence of schemes of compatible skeletons for any dimension $N$:
[…]
where the prime divisor $E=E_{n}$ appears on $A_{n}$ whose skeleton is $B_{n}$. Let $P_{i}\in B_{i-1}$ be the center of $E$ and also give the center of the blow-up $B_{i}\to B_{i-1}$. On the other hand, ${\widetilde{P}}_{i}\in{\widetilde{B}}_{i-1}$ is the prime lifting of $P_{i}$ and the center of the blow-up ${\widetilde{B}}_{i}\to{\widetilde{B}}_{i-1}$ with the appropriate conditions. Here, ${\widetilde{Z}}_{i}$ is a lifting of $Z_{i}$. Then, we already proved the statement (i) for arbitrary dimension $N$. Therefore, the problem is (ii). In order to reduce the discussion to the $(N+1)$-dimensional case, embed our affine spaces into higher dimensional affine spaces as follows:
$B_{0}={\mathbb{A}}_{\Sigma}^{N}={\mathrm{Spec\ }}\Sigma[x_{1},\ldots,x_{N}]\subset\overline{B}_{0}:={\mathbb{A}}_{\Sigma}^{N+1}={\mathrm{Spec\ }}\Sigma[x_{1},\ldots,x_{N},x_{N+1}]$
${\widetilde{B}}_{0}={\mathbb{A}}_{\widetilde{\Sigma}}^{N}={\mathrm{Spec\ }}{\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]\subset\overline{{\widetilde{B}}}_{0}:={\mathbb{A}}_{\widetilde{\Sigma}}^{N+1}={\mathrm{Spec\ }}{\widetilde{\Sigma}}[x_{1},\ldots,x_{N},x_{N+1}],$
as the hyperplanes defined by $x_{N+1}=0$. For $i=1,2,\ldots,n+2$, take the blow-ups
$\overline{B}_{i}\to\overline{B}_{i-1}\ \mbox{ with\ the\ center\ by\ }P_{i}\subset B_{i-1}\subset\overline{B}_{i-1}\ \mbox{and}$
$\overline{{\widetilde{B}}}_{i}\to\overline{{\widetilde{B}}}_{i-1}\ \mbox{ with\ the\ center\ by\ }{\widetilde{P}}_{i}\subset{\widetilde{B}}_{i-1}\subset\overline{{\widetilde{B}}}_{i-1}.$
Here, we note that $B_{i}$ and ${\widetilde{B}}_{i}$ are isomorphic to the proper transforms of $B_{0}$ and ${\widetilde{B}}_{0}$ in $\overline{B}_{i}$ and in $\overline{{\widetilde{B}}}_{i}$, respectively. So we identify $B_{i}$ and ${\widetilde{B}}_{i}$ with the proper transforms of $B_{0}$ and ${\widetilde{B}}_{0}$ in $\overline{B}_{i}$ and in $\overline{{\widetilde{B}}}_{i}$, respectively. As $\overline{{\widetilde{B}}}_{0}$ and $\overline{B}_{0}$ are compatible skeletons, and ${\widetilde{P}}_{1}\in{\widetilde{B}}_{0}\subset\overline{{\widetilde{B}}}_{0}$ is a lifting of $P_{1}\in B_{0}\subset\overline{B}_{0}$, it follows that $\overline{{\widetilde{B}}}_{1}$ and $\overline{B}_{1}$ are compatible skeletons.
Here, we note that
$\overline{B}_{n+2}\to\overline{B}_{n+1}\to\overline{B}_{n}\to\ldots\to\overline{B}_{1}\to\overline{B}_{0}={\mathbb{A}}_{\Sigma}^{N+1}$
is a normal Gorenstein Zariski sequence, i.e. it satisfies (i)–(iv) of Proposition 4.6. And
$\overline{{\widetilde{B}}}_{n+2}\to\overline{{\widetilde{B}}}_{n+1}\to\overline{{\widetilde{B}}}_{n}\to\ldots\to\overline{{\widetilde{B}}}_{1}\to\overline{{\widetilde{B}}}_{0}={\mathbb{A}}_{\widetilde{\Sigma}}^{N+1}$
is its lifting.
Let $E_{i}\subset B_{i}$ and $\overline{E}_{i}\subset\overline{B}_{i}$ be the exceptional divisors dominating $P_{i}\in B_{i-1}$ and $P_{i}\in\overline{B}_{i-1}$, respectively. Let ${\widetilde{E}}_{i}\subset{\widetilde{B}}_{i}$ and $\overline{{\widetilde{E}}}_{i}\subset\overline{{\widetilde{B}}}_{i}$ be the exceptional divisors dominating ${\widetilde{P}}_{i}$. (For $i=n+1,n+2$, we use character $F$ instead of $E$ as we did before.) As we chose ${\widetilde{P}}_{i}\in{\widetilde{B}}_{i-1}$ as follows:
$P_{i}\in E_{j}$ if and only if ${\widetilde{P}}_{i}\in{\widetilde{E}}_{j},\ \ (j\leq i-1)$
Therefore, it follows
${P}_{i}\in\overline{E}_{j}$ if and only if ${{\widetilde{P}}}_{i}\in\overline{{\widetilde{E}}}_{j}\ \ (j\leq i-1)$
Let ${\mathfrak{a}}\subset k[x_{1},\ldots,x_{N}]$ be a non-zero ideal. We may assume that $\mathfrak{a}$} is generated by elements of $\Sigma[x_{1},\ldots,x_{N}]$ by taking $\Sigma\subset k$ sufficiently large. Denote the ideal ${\mathfrak{a}}\cap\Sigma[x_{1},\ldots,x_{N}]$ by the same symbol $\mathfrak{a}$}. Let
$\overline{{\mathfrak{a}}}:={\mathfrak{a}}[x_{N+1}]\subset\Sigma[x_{1},\ldots,x_{N},x_{N+1}].$
Then, applying the theorem to $\overline{{\widetilde{E}}}_{n}$ and $\overline{{\mathfrak{a}}}$, there exists a lifting $\widetilde{\overline{{\mathfrak{a}}}}\subset{\widetilde{\Sigma}}[x_{1},\ldots,x_{N},x_{N+1}]$ of $\overline{{\mathfrak{a}}}$ such that
(47) $v_{\overline{{\widetilde{F}}}_{2}}(\widetilde{\overline{{\mathfrak{a}}}})=v_{\overline{F}_{2}}({\overline{{\mathfrak{a}}}})=v_{\overline{E}_{n}}(\overline{{\mathfrak{a}}}).$
Define ${\mathfrak{a}}^{\prime}\subset{\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]$ to be the image of $\widetilde{\overline{{\mathfrak{a}}}}\subset{\widetilde{\Sigma}}[x_{1},\ldots,x_{N},x_{N+1}]$ by the canonical projection. Then, $\mathfrak{a}$$^{\prime}$ is a lifting of $\mathfrak{a}$}. By the definition of the ideals it follows
(48) $v_{{\widetilde{F}}_{2}}({\mathfrak{a}}^{\prime})=v_{\overline{{\widetilde{F}}}_{2}}({\mathfrak{a}}^{\prime}[x_{N+1}])\geq v_{\overline{{\widetilde{F}}}_{2}}(\widetilde{\overline{{\mathfrak{a}}}})=v_{\overline{F}_{2}}({\overline{{\mathfrak{a}}}})=v_{F_{2}}({\mathfrak{a}}),$
where the inequality in the middle follows from the minimality of ${\mathfrak{a}}^{\prime}[x_{N+1}]$ among the ${\widetilde{\Sigma}}[x_{1},\ldots,x_{N+1}]$-ideals whose images in ${\widetilde{\Sigma}}[x_{1},\ldots,x_{N}]$ coincide with $\mathfrak{a}$$^{\prime}$. The inequality (48) gives $v_{{\widetilde{F}}_{2}}({\mathfrak{a}}^{\prime})\geq v_{F_{2}}({\mathfrak{a}}),$ while the opposite inequality holds in general. Now, we obtain that the lifting $\mathfrak{a}$$^{\prime}$ of $\mathfrak{a}$} satisfies
$v_{{\widetilde{F}}_{2}}({\mathfrak{a}}^{\prime})=v_{F_{2}}({\mathfrak{a}})=v_{E_{n}}({\mathfrak{a}}),$
as desired. ∎
Corollary 5.5 (Theorem 1.1, (iii)).
Let $k$ be an algebraically closed field of characteristic $p>0$, $A={\mathbb{A}}_{k}^{N}$ the affine space over $k$ of dimension $N\geq 2$ and 0 the origin of $A$. Let $E$ be a prime divisor over $(A,0)$ and ${\mathfrak{a}}_{1},\ldots,{\mathfrak{a}}_{r}\subset k[x_{1},\ldots,x_{N}]$ non-zero ideals. Then, there exists a prime divisor $F_{\mathbb{C}}$ over the affine space $A_{\mathbb{C}}={\mathbb{A}}_{\mathbb{C}}^{N}$ over $\mathbb{C}$} with the center at the origin $\tilde{0}$} satisfying the following:
There exist liftings ${\widetilde{\mathfrak{a}}}_{i}\subset{\mathcal{O}}_{A_{\mathbb{C}}}$ of ${\mathfrak{a}}_{i}$ $(i=1,\ldots,r)$ such that for every multi-ideal ${\mathfrak{a}}^{e}:={\mathfrak{a}}_{1}^{e_{1}}\cdots{\mathfrak{a}}_{r}^{e_{r}}$ $(e_{i}\in{\mathbb{R}}_{>0})$ on $A$, the lifting ${\widetilde{\mathfrak{a}}}^{e}={\widetilde{\mathfrak{a}}}_{1}^{e_{1}}\cdots{\widetilde{\mathfrak{a}}}_{r}^{e_{r}}$ of ${\mathfrak{a}}^{e}$ on $A_{\mathbb{C}}$ satisfies
$2(N-1)+a(E;A,{\mathfrak{a}}^{e})=a(F_{\mathbb{C}};A_{\mathbb{C}},{\widetilde{\mathfrak{a}}}^{e}).$
Proof.
By (i) and (ii) of the theorem, the equality
[…]
is obvious for an ideal $\mathfrak{a}$}. For multi-ideal case, as the number of ideals is finite in the multi-ideal, we can take common centers $P_{n+1}$ and $P_{n+2}$ general for all ideals. Then, we obtain a lifting ${\widetilde{\mathfrak{a}}}_{i}$ of ${\mathfrak{a}}_{i}$ for each $i$ such that $v_{E}({\mathfrak{a}}_{i})=v_{F_{\mathbb{C}}}({\widetilde{\mathfrak{a}}}_{i})$, which completes the proof of our corollary. ∎
Until now, we assume the base space $A$ to be the affine space. However, we can apply the theorem also to a pair consisting of a smooth variety and a multi-ideal with a real exponent. In such a case, we cannot define a lifting of a regular function around the point 0, but passing through the completion we can reduce the discussion to the case of the affine space.
Corollary 5.6 (Corollary 1.2, (i)).
Let $k$ be an algebraically closed field of characteristic $p>0$. Let $0\in A$ be a smooth closed point of a variety $A$ of dimension $N\leq 3$. Then, for fixed $N$ and an exponent $e\in{\mathbb{R}}_{>0}^{r}$, the set of log discrepancies
$\Lambda_{(A,0),e}:=\left\{a(E;A,{\mathfrak{a}}^{e})\middle|\begin{array}[]{l}E:\mbox{prime\ divisor\ over\ }(A,0)\\ (A,{\mathfrak{a}}^{e}):\mbox{log\ canonical\ at\ }0\\ \end{array}\right\}$
is discrete in the sense of Corollary 1.2, (i). In particular, for a fixed $e$, the set $\{{\mathrm{mld}}(0,A,{\mathfrak{a}}^{e})\mid{\mathfrak{a}}^{e}:\ \mbox{non-zero\ multi-ideal}\ \}$ is finite.
Proof.
[…]
$v_{E}({\mathfrak{a}}^{e})=v_{E}(\widehat{\mathfrak{a}}^{e})\ \ \ \mbox{for\ any \ prime\ divisor\ }E\ \mbox{over}\ (A,0)\ \mbox{and}$
(49) ${\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})={\operatorname{lct}}(\widehat{0};\widehat{A},\widehat{\mathfrak{a}}^{e}).$
Then, for a given prime divisor $E$ over $(A,0)$ there exists $d\in{\mathbb{N}}$ such that $v_{E}(\widehat{\mathfrak{a}}_{i})<v_{E}(\widehat{\mathfrak{m}}^{d})$ for all $i$, where $\widehat{\mathfrak{m}}$ is the maximal ideal of $\widehat{\mathcal{O}}_{A,0}=k[[x_{1},\ldots,x_{N}]]$. Here, noting that $\widehat{\mathfrak{a}}_{i}+\widehat{\mathfrak{m}}^{d}$ is the extension of an ideal $\underline{{\mathfrak{a}}}_{id}\subset k[x_{1},\ldots,x_{N}]$, we obtain
$v_{E}(\widehat{\mathfrak{a}}^{e})=v_{E}((\widehat{\mathfrak{a}}+\widehat{\mathfrak{m}}^{d})^{e})=v_{E}(\underline{{\mathfrak{a}}}_{d}^{e})\ \mbox{and}$
(50) ${\operatorname{lct}}(0;{\mathbb{A}}_{k}^{N},\underline{{\mathfrak{a}}}_{d}^{e})\geq{\operatorname{lct}}(\widehat{0};\widehat{A},\widehat{\mathfrak{a}}^{e}),$
where the notation $(\widehat{\mathfrak{a}}+\widehat{\mathfrak{m}}^{d})^{e}$ means the formal product
$\prod_{i=1}^{r}(\widehat{\mathfrak{a}}_{i}+\widehat{\mathfrak{m}}^{d})^{e_{i}}.$
Here, the above inequality (50) of lcts follows from $\underline{{\mathfrak{a}}}_{id}k[[x_{1},\ldots,x_{N}]]=\widehat{\mathfrak{a}}_{i}+\widehat{\mathfrak{m}}^{d}\supset\widehat{\mathfrak{a}}_{i}$. Therefore, the log discrepancy $a(E;A,{\mathfrak{a}}^{e})=a(E;\widehat{A},(\widehat{\mathfrak{a}}+\widehat{\mathfrak{m}}^{d})^{e})=a(E;{\mathbb{A}}_{k}^{N},\underline{{\mathfrak{a}}}_{d}^{e})$ belongs to the set
$\Lambda_{({\mathbb{A}}_{k}^{N},0),e}:=\left\{a(E;{\mathbb{A}}_{k}^{N},{\mathfrak{a}}^{e})\middle|\begin{array}[]{l}E:\mbox{prime\ divisor\ over\ }({\mathbb{A}}_{k}^{N},0)\\ ({\mathbb{A}}_{k}^{N},{\mathfrak{a}}^{e}):\mbox{log\ canonical\ at\ }0\\ \end{array}\right\}.$
Now, we can reduce the discussion to the case over $({\mathbb{A}}_{k}^{N},0)$. By Theorem 1.1, (iii), for a fixed exponent $e\in{\mathbb{R}}_{>0}^{r}$, the set of numbers of the following form:
$\{2(N-1)+a(E;{\mathbb{A}}_{k}^{N},{\mathfrak{a}}^{e})\mid({\mathbb{A}}_{k}^{N},{\mathfrak{a}}^{e})\ \mbox{is\ log\ canonical\ at\ }0,\ E:\ \mbox{prime\ divisor \ over \ }({\mathbb{A}}_{k}^{N},0)\}$
is contained in
$\{a(E^{\prime};{\mathbb{A}}_{\mathbb{C}}^{N},{\mathfrak{b}}^{e})\mid(A_{\mathbb{C}},{\mathfrak{b}}^{e})\ \mbox{is\ log\ canonical\ at\ }{\tilde{0}},\ E^{\prime}:\mbox{prime\ divisor \ over \ }({\mathbb{A}}_{\mathbb{C}}^{N},0)\}$
which is known to be discrete by [12]. Here, we use that a lifting ${\widetilde{\mathfrak{a}}}^{e}$ of log canonical multi-ideal ${\mathfrak{a}}^{e}$ is also log canonical by Lemma 3.12. Hence, we obtain that the set of log discrepancies of log canonical multi-ideals with a fixed exponent in positive characteristic is also discrete.
As the minimal log discrepancy is either -$\infty$ or in the interval $[0,N]$, the discreteness of log discrepancies implies the finiteness of the set of the minimal log discrepancies. ∎
Definition 5.7.
Let $k$ be a field and $0\in A$ a smooth $k$-valued point of a variety $A$ of dimension $N\geq 2$. Fix an exponent $e\in{\mathbb{R}}_{>0}^{r}$. We define
$T_{(A,0),e}:=\{{\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})\mid{\mathfrak{a}}^{e}\ \mbox{ a\ multi-ideal\ with\ the\ exponent\ }e\mbox{ on}\ A\}$
the set of log canonical thresholds.
Corollary 5.8.
Let $k$ be an algebraically closed field $k$ of positive characteristic and $0\in A$ a smooth closed point in a variety $A$ of dimension $N\leq 3$ defined over $k$. Fix an exponent $e\in{\mathbb{R}}_{>0}^{r}$. Then,
$T_{(A,0),e}\subset T_{({\mathbb{A}}_{\mathbb{C}}^{N},{\tilde{0}}),e}.$
Proof.
In the corollary, we state the result only for $N\leq 3$; however, the corollary follows once Theorem 1.1 holds. Thus, for future use, we give the proof for every $N\geq 2$, assuming that Theorem 1.1 holds. For simplicity, we denote $\sum e_{i}v_{E}(\mathfrak{a}_{i})$ by $v_{E}(\mathfrak{a}^{e})$. We use the notation from the proof of Corollary 5.6. By (49), it is sufficient to prove
$T_{(\widehat{A},\widehat{0}),e}\subset T_{({\mathbb{A}}_{\mathbb{C}}^{N},0),e}.$
When the characteristic of $k$ is zero, we know the following for a multi-ideal $\widehat{\mathfrak{a}}^{e}$ on $\widehat{A}$ by [5, Corollary 2.10] for every $d\in{\mathbb{N}}$:
(51) $\left|{\operatorname{lct}}(\widehat{0};\widehat{A},(\widehat{\mathfrak{a}}+\widehat{\mathfrak{m}}^{d})^{e})-{\operatorname{lct}}(0;\widehat{A},(\widehat{\mathfrak{a}}^{e}))\right|\leq N/d.$
This inequality was proven using the method of jet schemes. We now know that the same method applies based on the description of lct in Proposition 3.11, (6) in characteristic $p>0$, as well as by [20, Corollary 3.7]. Therefore, (51) holds in positive characteristic too.
Now, assume ${\operatorname{char}}k=p>0$ and let $\widehat{\mathfrak{a}}^{e}$ be a multi-ideal on $\widehat{A}$. Noting that $(\widehat{\mathfrak{a}}+\widehat{\mathfrak{m}}^{d})^{e}$ is the extension of a multi-ideal $\underline{{\mathfrak{a}}}_{d}^{e}$ on ${\mathbb{A}}_{k}^{N}$, we obtain ${\operatorname{lct}}(\widehat{0};\widehat{A},(\widehat{\mathfrak{a}}+\widehat{\mathfrak{m}}^{d})^{e})={\operatorname{lct}}(0;{\mathbb{A}}_{k}^{N},\underline{{\mathfrak{a}}}_{d}^{e})$. Therefore, we have
(52) $\left|{\operatorname{lct}}(0;{\mathbb{A}}_{k}^{N},\underline{{\mathfrak{a}}}_{d}^{e})-{\operatorname{lct}}(\widehat{0};\widehat{A},(\widehat{\mathfrak{a}}^{e}))\right|\leq N/d.$
In [5], they prove that $T_{{\mathbb{A}}_{\mathbb{C}}^{N-1},e}\subset T_{{\mathbb{A}}_{\mathbb{C}}^{N},e}$ and the set of accumulation points of $T_{{\mathbb{A}}_{\mathbb{C}}^{N},e}$ coincides with $T_{{\mathbb{A}}_{\mathbb{C}}^{N-1},e}$. Actually, they stated it only for $e=(1)\in{\mathbb{R}}_{>0}$. However, their proof also works for every fixed $e\in{\mathbb{R}}^{r}_{>0}$.
Now, we look at a positive characteristic pair $({\mathbb{A}}_{k}^{N},\underline{{\mathfrak{a}}}_{d}^{e})$ and let $c:={\operatorname{lct}}(\widehat{0};\widehat{A},\widehat{\mathfrak{a}}^{e})$. Then, by (52) we have:
(53) $\lim_{d\to\infty}{\operatorname{lct}}(0;{\mathbb{A}}_{k}^{N},\underline{{\mathfrak{a}}}_{d}^{e})=c$
If $c\in T_{{\mathbb{A}}_{\mathbb{C}}^{N-1},e}$, then $c\in T_{{\mathbb{A}}_{\mathbb{C}}^{N},e}$, which completes the proof. Assume $c\not\in T_{{\mathbb{A}}_{\mathbb{C}}^{N-1},e}$, then by the result of [5], $c$ is not an accumulation point of $T_{({\mathbb{A}}_{\mathbb{C}}^{N},{\tilde{0}}),e}$. Therefore, we obtain $\epsilon>0$ such that:
$(c,c+\epsilon)\cap T_{N,{\mathbb{C}}}=\emptyset.$
We claim that there is a prime divisor $E$ over $({\mathbb{A}}_{k}^{N},0)$ and an integer $d\gg 0$ such that
(54) $c\leq z(E;{\mathbb{A}}_{k}^{N},\underline{{\mathfrak{a}}}_{d}^{e})=\frac{k_{E}+1}{v_{E}(\underline{{\mathfrak{a}}}_{d}^{e})}<c+\frac{\epsilon}{2},\ \ \mbox{and}$
(55) $\frac{2(N-1)}{v_{E}(\underline{{\mathfrak{a}}}_{d}^{e})}<\frac{\epsilon}{2}.$
Indeed, (53) follows from (52). Regarding inequality (54), based on the definition of $\operatorname{lct}$} and the inequality (53), it is evident that there exists a prime divisor $E^{\prime}$ over $A$ with its center “containing 0” that satisfies the inequality (54):
$z(E^{\prime};{\mathbb{A}}_{k}^{N},\underline{{\mathfrak{a}}}_{d}^{e})=\frac{k_{E^{\prime}}+1}{v_{E^{\prime}}(\underline{{\mathfrak{a}}}_{d}^{e})}<c+\frac{\epsilon}{2}.$
We need to find a prime divisor “with the center at 0” satisfying the inequality (54). If the center of $E^{\prime}$ is not $\{0\}$, then by suitable blow-ups, we obtain a normal variety $\overline{A}$ on which $E^{\prime}$ appears, and there is an exceptional divisor $E^{\prime}_{1}$ on $\overline{A}$ with the center at 0 such that
$E^{\prime}\cap E^{\prime}_{1}\neq\emptyset.$
We may assume that $\overline{A}$ is smooth around $E^{\prime}\cap E^{\prime}_{1}$ outside a closed subset of codimension $4$, and $E^{\prime}$ and $E^{\prime}_{1}$ intersect normally there. Because even over a field of positive characteristic, there is a log resolution of singularities for dimension $\leq 3$ ([1],[3]).
If $z(E^{\prime}_{1};A,\underline{{\mathfrak{a}}}_{d}^{e})\geq c+{\epsilon}/{2}$, then take the blow-up at the intersection $E^{\prime}\cap E^{\prime}_{1}$ to obtain a new prime divisor $E^{\prime}_{2}$ dominating $E^{\prime}\cap E^{\prime}_{1}$ such that
$z(E^{\prime};{\mathbb{A}}_{k}^{N},\underline{{\mathfrak{a}}}_{d}^{e})<z(E^{\prime}_{2};{\mathbb{A}}_{k}^{N},\underline{{\mathfrak{a}}}_{d}^{e})<z(E^{\prime}_{1};{\mathbb{A}}_{k}^{N},\underline{{\mathfrak{a}}}_{d}^{e}).$
Next, blow-up at $E^{\prime}\cap E^{\prime}_{2}$ to obtain $E^{\prime}_{3}$. Continue this process, then we have a sequence of prime divisors $\{E^{\prime}_{i}\}$ over $({\mathbb{A}}_{k}^{N},0)$ such that
$\lim_{i\to\infty}z(E^{\prime}_{i};{\mathbb{A}}_{k}^{N},\underline{{\mathfrak{a}}}_{d}^{e})=z(E^{\prime};{\mathbb{A}}_{k}^{N},\underline{{\mathfrak{a}}}_{d}^{e}).$
This gives a prime divisor $E^{\prime}_{n}$ over $({\mathbb{A}}_{k}^{N},0)$ satisfying (54). The inequality (55) in the claim holds when one takes sufficiently many times of blow-ups.
For $E$ and $\underline{a}_{d}^{e}$, choose a prime divisor $F_{\mathbb{C}}$ and a lifting $\tilde{a}_{d}^{e}$ on $\mathbb{A}_{\mathbb{C}}^{N}$ as in Theorem 1.1. Then, applying Lemma 3.12, we obtain the following inequalities:
(56) $c+\epsilon>\frac{k_{E}+1+2(N-1)}{v_{E}(\underline{a}_{d}^{e})}=\frac{k_{F_{\mathbb{C}}}+1}{v_{F_{\mathbb{C}}}(\tilde{a}_{d}^{e})}\geq{\operatorname{lct}}(\tilde{0};\mathbb{A}_{\mathbb{C}}^{N},\tilde{a}_{d}^{e})\geq{\operatorname{lct}}(0;\mathbb{A}_{k}^{N},\underline{a}_{d}^{e}).$
If ${\operatorname{lct}}(0;\mathbb{A}_{\mathbb{C}}^{N},\underline{a}_{d}^{e})=c$ for some $d>0$, then we are done. Otherwise, by the definition of $\epsilon$, we have
${\operatorname{lct}}(0;\mathbb{A}_{\mathbb{C}}^{N},\underline{a}_{d}^{e})<c$
for every $d>0$. On the other hand, by (53),
$c={\operatorname{lct}}(\widehat{0};\widehat{A},\widehat{a}^{e})=\lim_{d\to\infty}{\operatorname{lct}}(0;\mathbb{A}_{k}^{N},\underline{a}_{d}^{e}).$
Therefore, the inequalities (56) yield an ascending chain
$\cdots<{\operatorname{lct}}(\tilde{0};\mathbb{A}_{\mathbb{C}}^{N},\tilde{a}_{d_{i}}^{e})<{\operatorname{lct}}(\tilde{0};\mathbb{A}_{\mathbb{C}}^{N},\tilde{a}_{d_{i+1}}^{e})<\cdots<c,$
whose limit is $c$, while none of the terms is equal to $c$. This contradicts the fact that the ACC holds for the set of log canonical thresholds in characteristic 0 [4, 8]. ∎
Corollary 5.9 (Corollary 1.2,(ii),(iii)).
Assume $k$ is an algebraically closed field and $N\leq 3$ is an integer. Let $e=(e_{1},\ldots,e_{r})\in{\mathbb{R}}_{>0}^{r}$. Define
[…]
Then, the following hold:
  1. (ii)
    (Rationality of lcts and their accumulation points) For a rational exponent $e$ and the multi-ideal ${\mathfrak{a}}^{e}$ on $A$, the following inclusion holds:
    $T_{N,k}(e)\subset{\mathbb{Q}}_{\geq 0}.$
    And an accumulation point of $T_{N,k}(e)$ is also a rational number.
  2. (iii)
    (ACC of lcts) Let $D\subset{\mathbb{R}}_{>0}$ be a DCC set. Then, ACC holds for the set
    $T_{N,k}(D):=\bigcup_{e_{i}\in D}T_{N,k}(e).$
Proof.
By Corollary 5.8, for rational $e$, every value $c={\operatorname{lct}}(0;A,{\mathfrak{a}}^{e})$ belongs to $T_{N,e,{\mathbb{C}}}\subset{\mathbb{Q}}$, and $T_{N,e,{\mathbb{C}}}$ is known to be closed ([5]), therefore (ii) follows immediately. As the set $T_{N,{\mathbb{C}}}(D)$ satisfies ACC by [4] for a special case, and by [8] for the general case, its subset $T_{N,k}(D)$ satisfies ACC as well. ∎